A Complete Bibliography of The Journal of the American Statistical Association: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

05 October 2024 Version 1.10

Title word cross-reference

#4059 [CMS87].

2^{31} − 1 [FM82]. 2 × 2 [AG83, GR81, Hab80, HAL82, HAJ83, MPG85, MPG86, SS80a]. 2 × 2 × 2 [CGM83]. 2 × C [Gro80]. A [Jac89b]. AR(1) [CR89]. ARIMA(p, 1, q) [SD85b]. χ^{2} [Lin81b, Stu80]. χ_{r}^{2} [JH82]. D [CT80, Stu82]. ED50 [JJW84a]. F [Arn80, Dye82, HC84, JH84, Lin81b, TW80, WHH88]. G [MG84]. I [DL88]. K [Boo86, Flu84, MW81, SS87a, WB81, Wei81a, BC82c, Cha81, HN87, MW82, Wal80a]. L [ER86, KP87, PS82]. L_{p} [Spo82]. M [Boo80b, HRR81, Cha81, MZ89]. N [CLNL87, CL85a, FD85, BD81b]. P [BS87a, BS87b, Bor87, CB87b, Dic87, Goo87b, Hin87b, Mor87d, Pra87b, SO83, Var87, TW89]. P (X_{1} + X_{2} + ⋯ + X_{k} < z) [YM85, YM86]. p_{1} − p_{2} [SS80a]. p_{1}/p_{2} [SS80a]. R [JJW84a, JJW84b, Hen79, Hen80, Zac86]. r × c [MP83, PH81]. S [WG81]. t [Ben83a, Ben83b, Hor83, KT88, LLT89, Lin81b, PMF88, Som80].

-estimates [MZ89]. -estimation [KP87]. -Estimator [JJW84b, JJW84a]. -estimators [HRR81]. -Optimal [CT80, Stu82]. -optimality [Jac89b]. -out-of- [Cha81]. -Sample [BC82c, Wal80a, HN87, MW82, MW81, SS87a]. -Spectral [MG84]. -statistics [PS82]. -Tail [Som80]. -Values [TW89].

1980 [EKT89a]. 1981 [And84a]. 1982 [And84a].

2nd [Blo84]. 2SAE [Kar81].

62009a [Bor87].

71 [CMS87]. 73 [FH81c]. 74 [Krz81]. 75 [JTB81, RC82, WM81]. 76 [BC82b, DD82, FH81a, KP82, Law82, MW82, MT82, Pra82, Sch82b, de 83a]. 77 [HG83]. 78 [NS86, Rin85a].

80 [And84a, BK86, MPG86, YM86]. 80k [FH81c]. 81 [Mac87, Sim87b]. 81a [Krz81]. 81e [WM81]. 81f [JTB81]. 82 [Bor87, Cla89, Kot89]. 82c [FH81a]. 82d [MW82, RC82]. 82h [KP82, Law82, MT82, Sch82b]. 83a [de 83a]. 83b [BC82b]. 83e [DD82]. 83h [Pra82]. 84 [Fab94, Whi92]. 84d [HG83]. 84i [NS86]. 85a [Rin85a]. 87b [YM86]. 87c [BK86]. 87j [Mac87]. 88h [Kot89]. 88k [Cla89]. 89a [Bor87].

92e [Fab94, Whi92].

A. [Oak85]. Abraham [Ber84a, MS84a, MS84b]. Absolute [Sie83, Sch87g]. accelerated [Sch88b, SS82a]. Acceptance [SS80b, Shi85, KP81c, KP82]. acceptance-rejection [KP81c, KP82]. Acceptance/rejection [SS80b]. Accident [Tom88]. Account [SO83]. Accounting [AR89, Zie87]. Accounts [TW84b, KY86]. Accuracy [Hod87a, JJS83, LH86, Lin81b, LP84, Sto83b]. Accurate [TK86]. Across [DL80]. Activity [Zie87]. Actual [Say80]. Adaptation [WH86]. Adaptive [Ruk84, CS87, HB88]. Addition [WM82]. Additional [Kiv80]. Additive [Bac89, EHK84, HT87a]. Additive-Plus-Multiplicative [EHK84]. Additivity [Tib88, MJ82]. Adequacy [FMT89, Jen86a]. Adjust [BS89]. Adjusted [BW85]. Adjusting [EKT89a, EKT89b, Sel86, SO83]. Adjustment [Bel89b, CD82, CT82, HW85, PM87, Roc89, Ros87c, HT82, Kar87b, O’G82, Sch81b, Sch82b, Wat87]. Adjustments [Bel89b, CD89b, CD89c, Cle89, Gew89, Pie89, Wat89, WY89]. Administrative [SM84c]. Admissible [MV85]. Adrian [DH83b]. Advances [Hil84, Jar89]. Adversary [Fis86]. After [Tay86, FH78, FH81c, Ked80]. Again [BCF^{+}81]. Against [Gro80, Mee81a, Stu80, BHS81, HM88a, iCL87, ZC88]. Age [And84a, AH86, Smi89b, TBLM89, SS82b]. Age-Specific [TBLM89]. Aged [MSV86]. Aggregate [BP86, HW89, KD89, KK84b]. Aggregation [Sto85a]. Aging [HPP86]. Agreement [TY85]. agreements [SG83]. AIDS [BG88, KL89]. Aircraft [Ber84a, MS84a, MS84b]. Akaike [Nef82]. Albert [Cha84a]. Alcohol [SSR^{+}88]. algebraic [SB81]. Algorithm [BM88, Dyk83, EO85, GN82, Rub87, Tri84a, JK89, LLS87, MP83, PH81, Spi82, TZ81]. Algorithms [LB88, Vel80, SS87b]. Aliasing [HW88]. Aligned [Shi85]. Allais [Sha84a]. Allocated [Sha81a]. Allocation [Sha83, Sha85, BM85]. Alternative [BMR88, DLST86, HH89a, HH89b, Hol89, MSV86, Mof89, MK80, Rub87, SM84b, PH86]. Alternatives [AG82b, AG82a, Fra81a, Gro80, MW81, Shi85, Stu80, BD88, BS86a, HN87, KR84, MW82, Nai87c, ZC88]. Aman [Hil84]. Amer [BK86, Bor87, BC82b, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, MW82, Mac87, MPG86, MT82, NS86, Pra82, Rin85a, RC82, Sch82b, Sim87b, WM81, Whi92, YM86, de 83a]. Ames [MKR89]. Among [HH89a, HH89b, Hol89, MSV86, Mof89, TY85, KD87, Thu85, Uus85]. Amos [Sha84b]. analog [SM81]. Analogies [HC81]. analogs [Rot83, Yan84]. Analogy [TW84b, SW83]. Analyses [DD83, Ole85]. Analysis [Ala84, All85, BL88a, BL88b, Ban89, Bas80a, Bas80b, BC89, Bec89, BJ81, Bel89b, BS88a, Ber84c, BD81a, Bic84, BM89, BC82a, BS88b, BLML83, Bro83b, BB80, CD89b, CD89c, CHK89, CR84a, Cau80, CD80, Cle89, Con89a, DFH89, Dok84, Fat83, FMW85, Flu85, Fri89, FB83, Gas85, GCC89, Gew89, GN82, Goo81, GWL87, GL88, Hab82, HMR80, HSG80, HSG82, HD80, HW88, Hin80, Hin84, Hol85, JK87, JJ88, Joh84a, Kad81, Kad83, KL89, Kee82, Kem80, LWZ86a, LWZ86b, LO81, Lan80, Lar84a, Lee81, LMR89, Lin80b, Lit81, Lon81, Mac86b, McD81a, McD81b, Muk84, Oga88, Par84, PN81, Pie89, Pre86b, Pro84, RS81, Rav89, RT87, Rog86a, Rub80a, Rub84a, SH89, Say80, SJ89, Sel86, Sha81a, SB87, She84a, Sie83, SH87, Sto89, SSR^{+}88, SR86]. Analysis [Tak83, TW84a, Tay84, TL88, Tuk85c, WB85, Wal81a, WLW89a, Wat89, Wer80, Whi86b, WABA89, WM85, ZB86, BOP88, BM85, BD88, BB89a, BH88, Boi86, BC82b, CGG85, CM84a, Chr87, Cla87b, Cla89, Clo82, CG84, DWL87, Efr88, FH86, GH86, GH88, GH87b, HR84, HVS84, Hol87, IHC84, IS87, JG83, JK89, KL85, LS86, LV88, MB88c, NS83, NS86, Pud82, RW85, Sim87c, SHS86, Tsi82, TH82, TV85, WLW89b, Won82a, Dic84, McK84, O’B84]. analysis-of-variance [BD88]. Analyzing [DL80, Tsu88, Was80a]. Ancillarity [God82a]. Ancillary [BN82, Bue82a, Cox82, Hin82b, Bue82b]. and/or [Gui88]. Anderson [SS87a, SS88b]. Animal [Din86, MV88, MC84]. Annual [And84a, Zie87]. ANOCOVA [BS86a]. ANOVA [Boi86, KR86, MS80, MO83, MB89, Tak83]. Answers [Rub86]. apparent [Efr86b]. applicability [Hur85b]. Application [BS88b, CM84b, EK82, GH83b, GAF83, JK87, KD89, LM85b, LMR89, O’N80, RA89a, Rea81, Sch86d, Sco85, Sha81a, Sto88, Tai89, Yan81, ZJ80, Abr88, GG87, LLS87, Ole85, SS88c, VW87, Wij85]. Applications [Ao86, Ban86b, BG80, Ber83b, Bra84b, CCE85, CZ85, Dem83, DK81, Fra81a, GK83, HT87a, Hin83, LH89, Leo83, Lin83a, McC82, Mor83a, NJ88, Sch88a, SS84, Sti83b, Sto83a, TW84a, TB81, AG85, Boi86, BJM81, CS87, ER86, Hwa86, JTB80, JTB81, KSC89, Mor83b, Yan84, YD84, Rip84]. Applied [CF87a, HD80, Jar89, Ran84, Dic84, Lin84a, O’B84]. Apportionment [Spe85b]. Appraisal [CV89, Coh89, Sin89a, Tur89b, dGJ89]. Approach [Ala80, BS86c, fC81, GE80, GCC89, IPRB85, Kaf82, KP81a, KP81b, LM85b, LC85a, LC85b, MRB86, MH89, OS87a, Rav89, Sch84d, Tam88b, Wei86, BD88, BJM81, CS87, Fin83, Gil84, Gil86, GM86b, HT82, JSG88, Mor86, Oma81, Roy86, Sin86, TJW88, WA83]. Approaches [BS89]. Approximate [Atw84, Bly86, Boo80b, Cha81, Cla87a, Fel80, KS89, MB89, PFV89, Swe84]. Approximating [SRMS81, TW80]. Approximation [Lut80, Mar84b, MT81, MT82]. Approximations [DKMO89, Gla89, Hod87a, KH84, Lin81b, LP84, Nau82, SS88b, WK88, DT87b, ER86, PMF88, Sha89a, Sha89b, TK86, TKK89]. Arbitrary [Van80a, Lut80, Wei80a, Wes87a]. Arcsine [AG80]. Area [Som80]. Areas [BHF88]. ARIMA [HT82, KA86, Krä86, Mar86, SD85a, Sol84]. ARIMA-model-based [HT82]. Arising [BHR83]. Arithmetic [Nan80]. ARMA [HM89, BY86, GW86, HM88b, Has80, Kiv80, TT84, WG81]. ARMAX [Sto86]. Array [WG81]. Article [Kiv80]. Ascertainment [KL89]. Aspects [Bic83, Car83, CW83, Hub83a, KW83, Mal83, Spe85b, GKR88, Hub83b, Sen84a]. Assay [MKR89]. Assessing [Arj88, AM84, DSW83, FG84, Has84, Hod87a, LPS84a, LPS84b, Rub84b]. Assessment [BBL89, CK89, Sin88]. assessments [Agn85]. assessors [VM83]. Assignment [BR83b, Ros84c, HP88]. Assoc [BK86, Bor87, BC82b, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, MW82, Mac87, MPG86, MT82, NS86, Pra82, Rin85a, RC82, Sch82b, Sim87b, WM81, Whi92, YM86, de 83a]. associated [Boo82, YD84]. Association [Agr81, BC89, Bec89, CHK89, DH80, Goo81, Ros84c, ACK87, Clo82, Gil86, GH88, GKR88]. assumption [CM85, KL85, Kru88]. Assumptions [Att80, Stu80, FJV89]. Asymmetric [Dag82, SB80b, Wec81, Wec82, AP87, SG83, Zel86a]. Asymmetrical [FR81a]. Asymmetry [RFPW80, Boo82]. Asymptotic [Arn80, BG80, Cha80a, Cro85, Has80, HI88, Kun80, LW85, Mor81, PP88, Rea81, SW81, SL87, Sha86b, Shi85, WB81, Wei81a, Whi89, Ber88b, Fab94, Gro81b, Whi92]. Asymptotically [LaR86a, MZ89, RFPW80, IHC84, KR86, Sin86, WL84]. Asymptotics [Bha85]. attained [But84]. attenuation [FH78, FH81c]. attributes [Tam81]. Augmentation [Dem87, Hab87, Mor87b, O’H87b, Rub87, TW87a, TW87b]. Augmented [HMM85]. Australian [Kad81, Lee81, Lit81, Lon81, McD81a, McD81b, Wal81a]. Author [Ano80e, Bue82b, Daw82b, Efr87c, Gew82b, God82c, Hol86b, Kit87, Sha82c]. authors [BS87a, BH88, BF85b, CB87a, DH83b, HHM88b, HL81b, LL87b, LF82b, LV88, MB88b, PT82b, TW87b]. Autocorrelated [Krä80, ZB86, Kob85]. Autocorrelation [DP84, WM82, Ali89, BHS81, DP89, KD87, TT84]. Autocorrelations [Has80]. Automatic [OYR86, Wah80]. Automatically [HHM88a, Joh88, Sco88, HHM88b]. Autoregressive [Ano87d, Ked80, SB83, Bha81, BHS81, BA83, FH81a, FH81b, Har84, KY85, Mac86a, Mac87, Spl83, Sti87, TT84, Tsa89b]. Auxiliary [Isa83, Smo82, Wri83]. Average [Orr89, Osb82, Pie85, HS89a, Spl83]. Average-Width [Pie85]. Averages [Ano87d, Mac86a, Mac87].

B [Daw82b, Kas85, Moo84a, Mye84, Par84, Sim87b, MR82]. Back [Ano80a, Ano80b, Ano80c, Ano80d, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano88a, Ano88b, Ano88c, Ano89a, Ano89b, Ano89c, Ano89d]. Backcalculation [Wei86]. Background [Mar80a]. Backward [All85, BB88, Sti88]. Bahadur [Rog88]. Balanced [Ber88a, Bro83b, LL89, MO83, CS89, HS89c, MS88a, PH86]. Bands [CS80, KSY85, Pie85, Pie87, BK89, HB88, Kan88, Nai87a, TT86, Uus83]. Bandwidth [Sta89c]. Bank [KD89]. Barndorff [Bue82b]. Barndorff-Nielsen [Bue82b]. Barrie [How84]. Bartlett [DK80, Riv86, Roc89]. Based [Ao86, Ban86b, Bha85, BR80, Cha80a, Der87, DW83, DvdM81, Efr87a, Fel80, GS80d, Hil87a, KL89, KE86, LL87a, Mor87c, Ros87c, SS88a, SF81a, SF81b, Sin88, TW84a, Atw84, BS86b, BHT88, CSS89, CS87, HM88a, HM83, LL87b, LQ85, LL81, MH82, MC84, OW82, Ran89, Ros84a, Sch86b, SW82, Sta89b, Ste81, Tar87, TH82, Wat87, HT82]. Basic [Zah84a]. Basu [Leh81a]. Bathtub [Gla80]. Bayes [Ber83b, BC82d, CH83, Cre89, DL81, Del89, Dem80, Dem83, Der87, DH83b, DH83a, Efr87a, GM86a, GL87, Gol83b, Hil87a, Hin83, HB83, JK87, Kas83, KS89, Kay80, Kre83, LL87b, LL87a, Leo83, Lin83a, Lou84, MWS^{+}89, MV85, Mor83b, Mor83a, Mor87c, NM83a, O’H87a, Rei85, Rog80, Rub80b, Rub80c, Sco80b, Smi83a, Sti83b, Umb80, Wei80c, WF82]. Bayes-like [WF82]. Bayesian [AG83, AG85, Alb88b, AP89a, Ber82, BD88, Ber88c, BB89a, BH88, Ber84c, CB87a, CS86, Dav85b, Daw82b, Daw82a, DC85, DJK87, FT85, Goe83, GP85, GL88, Hsu82a, JW88, Kad82, KS89, Len88, LHT89, Lin82b, Los85, MS85a, Mal88b, Mee81a, Mey85, MB88b, MB88a, Nyc88, RN80, Sha89b, SW80a, SW80b, Sha83, SS82b, Smo82, Smo84, Tai89, Tam88b, VM83, WHM85b, WHM85a, WH86, Win85, Won87, Zel86a]. Be [And84b, WB85, BF83, BHT88]. Beef [Eld81]. Behavior [HJ80, SB89, Sch84e, Nai84, SM84a]. Behrens [FP81, JW88]. Being [GRZ82, NM83a]. Belief [Gol86b]. Benchmarking [HT87b]. Benjamin [MZ80]. Berger [Bor87, MB88b]. Bernoulli [MO85, Tam85]. Best [AP87, EH83, Esc86, HV86, HE83, SH81, Sam85c, Tam88b, But84, EH84, Hsu82b, Wil84]. Best-Choice [Sam85c, Tam88b]. best-fitting [But84].

Best-Linear-Unbiased [Esc86]. bet [SG83]. Beta [Pre86a, SS80b]. Beta-Binomial [Pre86a]. Better [Ber87, Bic87, Cox87a, DT87a, Efr87c, Efr87b, KS80, LW87, PF87, Sch87b, Sch87f, Smi89b, Ruk87, Tam85, Wil84]. Betting [Mee81a]. Between [AS85b, AG81, Van80a, EGRW86, Fou80, Gew82a, Krz79, Krz81, NS82, Par82, Pie82, Wei82, WG81, Zel82, CR82, Gew82b, Gew84b, HS89c, Kob86, dAJ86]. Between-Groups [Krz79, Krz81]. Beyond [Bai85, EK85a, EK85b, Fay85a, Fel85, Han85a, Hau85, Pas85, Pre85, Rol85, Sed85, Tuk85b, Tuk85a]. Bias [AN81, BBL89, MMP86, RR84b, CSC86, MZ89]. Biased [ST87, Efr86b]. Biases [Sha84b]. biasing [Oma81]. Bickel [BC82b]. bidec [KT88]. Bilinear [BSB80]. Binary [AP83, Kee82, LMR89, Pre86a, Wu85, APS81, BS84b, LZ89, Man88, Min87, Ros89b, TH82]. binary-type [APS81]. Binomial [BGT80, BS83a, Bly86, Cas86b, LL86, OPZ81, Pre86a, RA89a, SM83b, WY89, AG85, BD81b, CL85a, De 83b, HPE81]. Binomially [Sam80]. bioassay [MS88b, Sch84c]. Biochemistry [And84a]. Biosciences [CZ85]. Birth [Ao86, HW89, Sti86]. Births [CL86, Kad81, Lee81, Lit81, Lon81, McD80, McD81a, McD81b, Wal81a]. Bispectral [HW88]. Bivariate [AS88, Die82, GH80, JT81, Oak89, ON89, Sar87, Sch84d, Sto80a, TH80, Cha84d, KR82, MO85]. Biweight [Hor88, Kaf82]. Block [HI88, KE86, YBN85, HS89c, Hsu82b, Jac89b, KR84]. Blockmodels [WW87a]. Blocks [Shi85, CS89, IHC84, Vos86]. Blood [WLW89a]. Blood-Lead [WLW89a]. Blue [DG85, SSdM85a, SSdM85b]. Bonferroni [BCF^{+}81]. Book [Abb83, Abe83, Abe89, Abr81, Abr84, Adk84, Adl84, Agr83a, Agr88, Aic81, Alb85, Ald87, Ald88, Ale80, All87, Alw88, Ame83, Ame85, Amm81, And84a, And83, Ang83, Ano80k, Ano80l, Ano80m, Ano80n, Ano81f, Ano81g, Ano81h, Ano81i, Ano82f, Ano82g, Ano82h, Ano82i, Ano83i, Ano83j, Ano83k, Ano83l, Ano84j, Ano84k, Ano84l, Ano84m, Ano85k, Ano85l, Ano85m, Ano85n, Ano86i, Ano86j, Ano86k, Ano86l, Ano87i, Ano87j, Ano87k, Ano88l, Ano88m, Ano88n, Ano89k, Ano89l, Ano89m, Ano89n, Ans85, Ans80b, Arm82, Arn81, Arn86, Arn87, Arn89a, Arn89b, Art82, Arv86, Asc84, Ash82a, Ash82b, Ass83, Ath83, Ath87, Atw81, Aue86, Aus89, Bab88, Bai82, Bai89, Bal86, Ban81, Ban82, Ban86a, Ban88, Bar86a, Bar86b, Bar89, BN84]. Book [Bar82b, Bas83, Bas85a, Bas85b, Bas87a, Bas87b, BK84a, Bea80, Bea83a, Bea83b, Bec80, Bec84, Bel89a, Bel81, Ben82, Ber83a, Ber84b, Ber85a, Ber83c, Ber83d, Ber85b, Ber85c, Ber86, Bha87, Bha82, Bie83a, Bil81, Bil83c, Bil83a, Bil83b, Bil84a, Bil84b, Bil85, Bin85, Bis88, Blo84, Boc89, Bod85, Boe83, Bon87, Boo81b, Bow83, Bow86, Boz85, Bra81, Bra84a, Bra85a, Bra85b, Bra87, Bra80, Bra84b, Bra88, Bra86, Bra85c, Bre86, Bri83, Bro80a, Bro80b, Bro80c, Bro81a, Bro82a, Bro82b, Bro83a, Bro84, Bro87, Bro80d, Bro83d, Bro88, Bry84, Bub80, Bub83, Bub84, Buc89, Bue81, Bue83a, Bue83b, Bue84, Bus87, Bus84, But85, BH85, Cal87, Cal85, Cam86, Can82, Car80, Car86a, Car86b, Car87, Cas86a, Cas87, Cas88, Cas89, Cha84a, Cha86a, Cha87]. Book [Cha89, Cha84b, Cha84c, Cha80b, Cha84e, Cha85a, Cha85c, Cha85b, Cha86c, Che89, Chi85, Chr85, Chr88a, Chr88b, Cla87c, Clo80a, Clo80b, Clo83a, Clo83b, Cob82, Cob84a, Cob84b, CG86, Cob87b, Cob87a, Coh82, Con89b, Coo83, Cor88a, Cor88b, Cor88c, CW89, Cor89a, Cor89b, Cor89c, Cos82a, Cos82b, Cos83, Cov86, Cov83, Cov85, Cow82, Cox80b, Cox85, Cre84, Dal83a, Dal84, Dar80, Dav84b, Dav84a, Dav84d, Dav84c, Dav84e, Dav85a, DS87, DeG83, Dea89a, Dea89b, Dea89c, Den84, Dev84, Dev80, Dic81, Dic82a, Dic82b, Dic83a, Dic84, Die89, Dom88, Don87, Dra87, Dub84, Duf82, Dun81, Dun85a, Dun85b, Dun87, Dyk88, Dyk89a, Dyk89b, Eas80, Eat82, Eat83, Ecc85, Edd82, Edd83, Edd84, ES81, ES82a, Eli83, Eme81, Emi83, Eri81b, Esp84, Esp87, Eub86a, Ewb85]. Book [Fah81, Fai85, Fai86, Far83a, Far88, Far87, Fay80, Fei82, Fei88, Fel86, Fel88, Fen89, Fie87, Fin86, Fin87, Fin88, Fir88, Fis80, Fla86, Fle84, Fli86, Fol85, Fol88, Fom86, Fon85, Fow88, Fra84a, Fra86a, Fra88a, Fra82, Fra88b, Fre83, Fri86, Fu87, Gar84, Gar83b, Gar83a, Gas83, Gat89, Gav89, GJ89, Gel86a, Gel86b, Geo84, Geo87, GG83, Gho83, Gho88, Gib88, Gif87, Gla88, Gle86, Gle85a, God85, Goe81, GW83a, GW83b, Gol83a, Gol85, Gol86a, Gol88, Goo80a, Goo80b, Goo83, Goo85, Goo87a, Gra84, Gra82, Gre81a, Gre81b, Gre81c, Gre83, Gre86a, Gre84, Gre86b, Gre85b, Gre88, Gre89, Gri80, Gri89, Gro81a, Gro87, Gui85, Gup81, Gup84, Haa88, Hai81, Han81, Han83, Han85b, Han84, Har82, Har85a, Har86, Har89, Hau82, Haw89, Hel83]. Book [Hen83, Her80, Her85, Het87, Hey82a, Hey82b, Hil89, Hil80, Hil83, Hil84, Hil85, HM85b, Hil86a, Hil86b, Hil86c, Hil87b, Hin82a, Hin87a, Hir81, Hir82, Hir83, Ho84, Hod87b, Hog85, Hol83, Hol85, Hoo87, Hor84, Hor89, Hos86, How84, How86, How89, Hsi80, Hsu80, Hsu85, Hsu89, Iac82a, Iac82b, Isa81, Ive86a, Ive86b, Ive87, Ive88a, Ive88b, Ive89a, Ive89b, Ive85, Iye89, Ize88, Jac87, Jac85, Jac84, Jac89a, Jar86a, Jar86b, Jay82, Jay81, Jeb82, Joh81, Joh80, Joh84b, Joh85, Joh87c, Joh87a, Joh87b, Joi82, Kad84, Kaf88, Kaf89, Kar85a, Kar80b, Kar87a, Kas85, Kat81, Kat86, Kay81, Kay89, Kel86, Kem82, Kem83, Kem88, Ken86, Khu89, Kie86, Kie88, Kim89a, Kim89b, Kin81, Kin83, Kin87, Kin89, Kir84, Kle84, Klu87, Klu88, Kme80, Kme82, Kme83a]. Book [Kme83b, Kme84, Kme85a, Kme85b, Kme85c, Kme88a, Kme88b, Kme89, Kno83, Koc89, Koe81, Koe82b, Koe82a, Koh88, Koo88a, Koo88b, Kos85b, Kou82, Kra89, Kri87, Kru85, Ksh88, KBD89, Kue87, Kuk87, Kul84, Kul86, Lac89, Lad84, Lah89, Lam86a, Lam82a, Lam86b, Lam88, Lam89, Lan87, Lan85, Lan84, Lar84a, Lar84b, Lar85, Lar86b, Lar86c, Lar87, Lar89a, Lar89b, Lar80, Las84, Lea83, Led81a, Led81b, Leh81b, Leh82, Leh84, Leh89, Len86, Leo89, Lep85, Leu85, Lev87, Lii80, Lin87, Lin89a, Lin80c, Lin81a, Lin82c, Lin83b, Lin83c, Lin84a, Lin84b, Lin86, Lit86, Loh89, Lor86, LH85, Low82, Low83a, Low83b, Luc82, Lun84, Lur82, Lus85, Mac89, Mad85, Mad89, Maj88, Mal88a, Mal85, Man84, Mar84a, Mar87b, Mas85, Mas88, Mat89, Max88, May83b, May83c, May83a]. Book [May84, McC83, McC81a, McC87, McC84, McC88, McD89, McK88, McK84, McL80, McL85, Mee81b, Mee81c, Mee84, Meh82, Men83, Men80, Met84, Met85, Mey86, Mey87b, Mil81a, Mil81b, Mil85, Mil86a, Mil86b, Mil87, Mis87, Mit84, Mit88, Mit86, Mon85, Mon88, Moo80, Moo81a, Moo82a, Moo82b, Moo83a, Moo83b, Moo84b, Moo84a, Moo87, Moo89, Mor80, Mor87a, Mor84, Mos85, Mos80, Mos86, Mue80, Muk84, Mul83, Mye84, NN86, Nam83, Nam84, Nam85a, Nam85b, Nam86, Nam87a, Nam87b, Nam87c, Nam88, Nat88, Nel89, Ner86, Neu81a, Neu81b, Neu83, New84, New81a, New81b, Noa81, Nob81, Noe80, Nor86, Nor85, Not89, Nyc89, O’B80a, O’B80b, O’B81a, O’B81b, O’B82, O’B83, O’B84, O’F80, O’F84, O’L85, O’R88, O’S87b, Obe86, Ode81, Ord85, Ore83a, Ore83b, Pac84, Pal86, Pan87, Par80a]. Book [Par85a, Par85b, Par85c, Par88, Par89, Par80b, Par83, Par84, Par85d, Pas80, Pat87, Pea87, Peb81, Pec80, Ped88, Ped89, Pek80, Pey89, Pfe87, Pic89, Pie81a, Pie81b, Pie86, Pij86, Pla83, Pla80, Pli87, Pli88, Plo87, Pol84b, Pol87, Pol85, Poo80, Por86, Pot83, Pra87a, Pra86, Pre81b, Pre89a, Pre89b, Pro82, Pro83a, Pro83b, Pro84, Pru88, Pru84, Pru85, Psa86, Pul85, Pur88, Pur89, Put80, Put82, Put85, Qua80, Qui82, Rag89, Ram85a, Ram85b, Ran86, Rao80a, Rao84, Red86, Reg80, Rei80, Rei81, Ren85, fSMR84, Res85, Ric88a, Ric84, Ric88b, Ric85, Rid82, Rin85b, Rip84, Rob89a, Rob89b, Roc85, Rod82a, Rod82b, Rod85, Rod86, Rog83, Roh83a, Roh83b, Roh83c, Roh83d, Ros87a, Ros88a, Ros87b, Ros88b, Ros84d, Ros82, Ros84e, Ros84f, Ros89c]. Book [Ros89d, Row87, Rus82, Rus89, Rya88, SB80a, Sam85b, Sam88, Sam89, San89, Sau83a, Sau83b, Sau85, Sav80, Sav81, Sav86, Sca87, Sch87a, Sch81a, Sch85a, Sch82a, Sch83a, Sch84a, Sch89a, Sch89b, Sch82c, Sch82d, Sch84b, Sch83c, Sch87c, Sch88c, Sch87d, Sch89c, Sch86c, Sch80b, Sch87e, Sch85e, Sch88d, Sco82, Sco83, Sea86, Sea82, Sea89, Sed84, Sen80, Sen81a, Sen81b, Sen81c, Sen82, Sen83a, Sen83b, Sen83c, Sen84a, Sen84b, Sen84c, Ser86, Sha84b, Sha84a, She80a, She85, She80b, She82a, She83a, She84a, She81, She82b, She82c, She83b, She84b, She89, Shi84a, Shr84, Shr87, Sil86, Sim80, Sim85b, Sim85c, Sim85a, Sim87a, Sin82, Sin89b, Sin80, Sin89c, Sir86, Ska89, Ska83, Ski87, Sko88, Slo88, Slo87, Sly88, Smi80, Smi84, Smi85b, Smi85a, Smi89a, Sne86, Sok81]. Book [Sok82, Sok83, Sok84a, Sok84b, Sol85, Sol88, Sol87, Spa88, Spe86, Spi86, Spi83, Spu87a, Spu87b, Sta85, Sta89a, Ste87a, Ste88, Ste82, SK84, Ste89, Sti83a, Sti84, Sti82, Sti83c, Sto85c, Sto85b, Sud83, Sud87a, Sud87b, Sud89, Suk87, Sus86, Sut85, Swa81, Swa84, Swi83, Swi85, Syl81, Tae81, TCD^{+}89, Tat88, Tav84, Tav85, Tay80a, Tay80b, Tay84, Tay87a, Tay87b, Tel86, Thi89, Tho82a, Tho89, Tho82b, Tho82c, Tho83b, Tho83a, Tho84c, Tho84d, Tho84a, Tho84e, Tho84b, Tho85b, Tho85c, Tho85d, Tho85e, Tho85a, Tho85f, Tho86b, Tho86a, Tho87, Tho85g, Tho83c, Thu82, Thu87, Thu88, Tie82, Tie84, Tie85, Tie81a, Tie81b, Tit85, Tol82, Tor83, Tor82, Tra83, Tsa89a, Tsu85, Tuf85, Tur87, Tur89a, Tve84, Ull88, Utt87, Van86a]. Book [Van81, Van80b, Van80c, Van86b, Van86c, Van86d, Var85, Var80, Var82a, Var82b, Var86, Var89, Vau82, Vel81b, Vel81a, Ver89a, Ver89b, VJ83, Vre80, Wai83, Wak85, Wal84a, Wal86a, Wal81b, War80, Was80b, Was85, Was89, Wat80, Wat84a, Wat85a, Wat88, Wat84b, Wat85b, Web89, Weg80, Weg89, Weh80, Wei81b, Wei83b, Wei84b, Wei85, Wei89, Wel84, Wel86, Wel89, Whi83a, Whi83b, Whi86a, Wic80, Wic81, Wid83a, Wid83b, Wij83, Wij86, Wil85a, Wil88a, Wil85b, Wil86, Win80, Win83, Win84a, Win84b, Win89, Wit84, Wol89, Wol80, Wol82, Wol83, Wol85, Wol87b, Won89, Woo87, Wra83, Yan89, Yan83, Yao87, Yas87, You83, Zac89a, Zac89b, Zah84a, Zah84b, Zah89, Zal82, Zel87a, Zel80a, Zel84, Zel85, Zel86b, Zil80, Zin89, dC81, dL86a]. Bookie [Mee81a]. Books [Kos85b]. Boostrap [Bic87]. Bootstrap [Ber87, BY88, Cox87a, Der87, DT87a, DT87b, Efr81, Efr87b, Efr87a, FH89, FH90, Gon86, Hil87a, LL87a, LW87, Mor87c, PF87, Roc89, Sch85b, Sch87b, Sch87f, Sti85, Ber88b, BK89, Efr87c, LL87b, Rom88]. Bootstrapping [Akr86, FP84, HB88, Sim86]. Borrowing [KD89]. Both [Bro82c, Bro83c, GAF83, CM86]. Bound [LNPS81, BE85, Kim86, Oma81]. Boundaries [Net86]. boundary [Cha88b]. Bounded [Bic83, Car83, CW83, Hub83a, KD89, KW82, KW83, Mal83, Sam85a, AP87, Hub83b, KSC89, Rup85, dJdWW88]. Bounded-Influence [Bic83, Car83, CW83, Hub83a, KW82, KW83, Mal83, Sam85a, Hub83b, KSC89, Rup85]. bounded-influence-regression [dJdWW88]. Bounds [Ano88e, BG88, KK84a, Kar81, Lea81, SL81, Sch86d, Som80, Del89, Gla89, HHO88, Kob86, Nai83, Nai87b, YM85, YM86]. Box [Kru80a, Spi82, Tay85, WG81]. Box-and-Cox [Spi82]. Branching [SW87]. breakdown [Sim89, YZ88]. breakdown-point [YZ88]. Breast [MRB86]. Breast-Feeding [MRB86]. Breiger [HL81b]. Breiman [LV88]. Brock [Bro84]. Brown [Rot83]. Brown-Forsythe [Rot83]. Bulletin [And84a]. Bump [Ask80, Edd80, GG80a, GG80b, Par80c, Poi80, Sco80a, Sil80, Tar80]. Bump-Hunting [Ask80, Edd80, GG80a, GG80b, Par80c, Poi80, Sco80a, Sil80, Tar80]. Bureau [NE84]. Bureaucracies [Mar81a]. Business [Zah84a].

C [Bow87, Pro84, Sim87b]. C. [God82c, MB88b]. C.-M. [God82c]. Cahoon [Bow87]. Calculating [BW85, Kap83]. Calculation [Dem87, Hab87, Mor87b, O’H87b, Rub87, TW87a, TW87b]. Calendar [BH83b, CD80, CD82]. Calibrated [Daw82a, Kad82, Lin82b, Daw82b]. Calibrating [Daw85, Loh87, Sch85d, Oak85]. Calibration [VM83]. call [Ros84a]. Campbell [Mar80b]. can [BHT88]. Canadian [Sta86]. Cancer [DH83a, Kas83, Kre83, MWS^{+}89, Smi83a, DH83b]. Canonical [GH86, Goo81, Gil86, GH88]. Capital [KJ87, Orr89]. Capitalizing [BR83b]. Capture [CM86, MSW^{+}83]. Capture-recapture [CM86, MSW^{+}83]. carcinogenesis [MC84]. Carcinogenicity [AR89, MV88]. Care [LTAP88, NRMM80]. Carl [Tay84]. Carlo [JTB81, Hur85a, JTB80, JV85a, Lat81, LC85a, LC85b, NJ88, NS82, Raj80, SF81a, SF81b, Sha89b, Tuk85c, TV85]. Carryover [LM85b]. Carryover-Effect [LM85b]. Case [Bel89b, CD89b, CD89c, Cha81, Cle89, Dun80a, Dun80b, Gew89, HH89a, HH89b, Hol89, Lee80, Mof89, Pie89, SM84b, SW80c, Wat89, Akr88, Gui83, Hoe84, Pic87, Wil84]. Cases [Kad83]. Cassel [God82c]. Castles [Ege81, Jac81, KH81a, KH81b, Wai81]. casual [Kru88]. Catastrophe [CZ85]. Categorical [BL88a, BL88b, DJK87, EO85, Nor84, RS81, Sel86, Wal84b, Con89a]. Categories [Goo81, Clo82]. Categorized [HM85c, HM85d]. Categorizing [ME86]. Cauchy [Fra81a]. Causal [Cox86a, Fay86, Gly86, Gra86, Hol86a, Rub86, Zie87, Hol86b]. Causation [Ros84c, Lin82c]. Cause [AR89, Din86]. Cause-of-Death [AR89, Din86]. Causes [fC81]. Causing [Voo81]. cell [BOP88, Hab88]. Cells [FK80, MHMEH82]. Censored [BR83a, Bha85, BC82c, DJK87, Efr81, Esc86, KW85, LaR86a, LO81, Lat81, MB82, SS88a, SJ89, ZB86, AA87, Alb88a, Gui88, Pad86, SM83a, SM84a, Tsi82, VJ88, WG83, Wei84a, WP85, Zac86]. censored-data [VJ88]. Censoring [BM81, BM83, HWW80, Lin89b, SY84, SW80c]. Censorship [Wei80a]. Census [Bai85, Cre89, EK85a, EK85b, EKT89a, EKT89b, Fay85a, Fel85, Han85a, Hau85, Jar89, Pas85, Pre85, Rol85, Sed85, Tuk85b, Wol86, Tuk85a]. centenary [Sti88]. Center [And84a]. Central [BB85a, Mar81b, ZC88]. central-mixture [ZC88]. Century [NE84, PH84]. Certain [Ger80, Rea81, Stu80, SM84a]. Chain [Ped80]. Challenger [DFH89]. Change [Sto85a, HRR81, SW86, WG83, Pro84]. change-of-variance [HRR81]. characteristic [YD84]. Characterization [Ger80, SS88c, JG83]. Characterizations [Gla80]. characterize [Sto82]. Characterizing [Umb80]. Charles [Zah84b]. Chemistry [BHR83]. Chi [Fay85b, GK83, Hab80, LaR86a, iCL87, Nai87c, RS81, Hab88, KOS85a, Ole85]. Chi-Square [GK83]. Chi-Squared [Fay85b, Hab80, LaR86a, RS81, iCL87, Hab88, KOS85a, Ole85]. Chi-squared-type [Nai87c]. Child [Mau83, MRB86, PH84]. Chlorofluoromethanes [RT87]. Choice [Bro81b, SA86, Sam85c, Tam88b, Car82, GC82, Hur85b, Sil84]. Choosing [Cam84, ES82b, ES84, HH89a, HH89b, Hol89, Mof89, Oma81]. Chosen [HHM88a, Joh88, Sco88, HHM88b]. Chow [Thu85]. Christ [Sti86]. Christian [BH88]. Circle [CHM82]. Circular [Moo84b]. Civilian [LH86]. Class [Cau80, DM88b, Fli81, Gro80, HPP86, Shi80, UH80, Voo81, LZ89, NS83, NS86, OW82, Ste81, Tsi82]. classes [KOS85a, Wal86b]. Classification [BS89, BC82d, CD89a, Cha80a, Moo84b, O’N80, Ruk84, Sch84e, Sel86, Sne82, Kor81, LV88, Sin86]. Classifications [BM89, Goo81, Agr83b, Clo82, Gro81b]. Classified [Tom88, WMH85]. Cleary [McK84]. Clevenson [Tsu84, Tsu86]. Clevenson-Zidek-type [Tsu84, Tsu86]. Cliff [Rip84]. Clinical [BBL89, BHR83, BB89b, SW89b, TGG89]. closed [Sch87g]. Closeness [GS89]. Closure [BG81]. Cluster [FH82, PN81, Pro85, BM85, MS85a, Roy86, TR87]. clustered [CM86, Ros89b]. Clustering [DL80, LNPS81, Won82b]. Clusterings [FM83b, FM83c, JM83, Van83, Tur83, Wal83, Won83a]. clusters [Har81a, PFV89, Won82b]. Coarsely [HSG80, HSG82]. Cochran [Hol85, Riv86]. Coefficient [GH87a, Sto80a, Cha84d, Mui85, Sen86, SG82]. Coefficients [AS85b, GS80b, Hen79, Hen80, Kiv80, Loh87, SL81, Hod88, Kob86, NM83b].

Cognitive [WLW89a]. coherent [Ger82]. Cohort [BLML83, dB85]. Coincidences [DM89]. Collaborators [Ano85f]. Collapsed [GK83]. collecting [APS81]. Combinations [GW80, See80, Flu85]. Combined [AA87, Alb88a, BM81, MH89, TG80]. Combining [DSW83, DH83a, Gas85, Geo86, HM85a, Kas83, Kre83, LL86, Sch86a, Smi83a, DH83b, Kin85, LQ85, LL81]. Comment [Ans80a, Ask80, AM84, Bai85, BN82, Bec83, Bel89b, Ber87, Ber83b, Ber84a, Ber88c, Bic83, Bic84, Bic87, Boa83, Bor87, BC82b, Bre81, Bri82, BK85b, Cah87, Car83, CR84a, CL85b, Cas82, Cle82, Cle89, Coh89, CW83, Cox82, Cox86b, Cox86a, Cox87a, CS83, Dal83b, Dav85b, Daw84, Daw85, DeG82, Dem80, Dem82, Dem83, DC85, Dem87, Der87, DT87a, Dic87, DG85, Dok84, Edd80, Edd81, Efr87a, Ege81, Eri82, Fab94, Far83b, Fay85a, Fel85, FW81, Fie82, FG84, FT85, FK85, Fra81b, Fra86b, Ful81, Gai83, Gas86, Gay83, Gen81, Gew84a, Gew89, Gly86, Goo82a, Goo82b, Goo87b, Gra86, Gun80, GP85, Hab81, Hab87, Han85a, Har81b, Has84, Hau85, HCGL85, Hil82, Hil87a, Hin80, Hin82b, Hin83, Hin87b, Hol89, Jac81, Joh84a, Joh88, JM83, Jon82, Kad81]. Comment [Kad82, Kas82, Kas83, Kau81, Kay80, Kea85, Kem80, Kna82, Kot89, KW83, Kre83, Kru80b, LM85a, Lan80, Lee81, Leo82, Leo83, Lin80a, Lin80b, Lin82a, Lin82b, Lin83a, Lit81, Lit83b, LW87, Lon81, Los85, Mac86b, Mal81, Mal83, Mal88b, Mar80b, Mar80c, May81, Mei82, Mey85, Mof89, Mor87b, Mor87c, Mor87d, Van83, O’H87b, Oak83, Obe80, Ord87, Par80c, Par82, Pas85, PR80, PF87, Pie82, Pie89, Poi80, Pra87b, Pre81a, PV85, Pre86b, Pre85, Rob86, Rog86a, Rog80, Rol85, RR84a, Rot82, Roy83, Rub80a, Rub83, Rub84a, Rub84b, Rub85, Rub86, Rub87, Sär81, Sch87b, Sch85c, Sch85d, Sch87f, Sch81c, Sco80a, Sco80b, Sco88, Sed85, Sil80, Sim87b, Sin89a, Smi81, Smi83a, Smi83b, Sti83b, Sto80b, Sug82, SR86, TW83, Tar80, Tha85, Thi80, Tho81, Tuk85b, Tuk85c]. Comment [TV85, Tum82, Tur83, Tur89b, Van80d, Var87, Vin80, Wai81, Wal83, Wal81a, Wat89, Wei82, Wei80b, Wei80c, Whi86b, Win85, Won83a, Wre81, Zel80b, Zel82, dGJ89, JV85b, LF82b, LV88]. Comments [Kiv80, Tuk85a, BS87a, BH88, Bow87, Bue82b, CB87a, Daw82b, DH83b, Efr87c, God82c, HHM88b, HL81b, Kit87, MB88b, Oak85, PT82b]. Common [FR83, Flu84, Fou80, HAL82, HAJ83, JH84, Rao80b, SW88, GG87, Mat83, MPG85, MPG86, Noe87]. Comparative [DFG87, BT81]. Comparing [Boo86, Efr84, FM83b, FM83c, JM83, Nai83, Van83, O’B88a, Tur83, Wal83, WB81, Won83a, CMST88, Jac89b]. Comparison [BS89, BB80, CGG85, Gew81, Hab80, IHC84, KJ87, Kob85, Krz79, LS83, MV88, ME86, MRW87, OPZ81, Sch80a, SM84c, Vel80, WTB80, CS85, CR82, Krz81, Wes87a, WPSL84]. Comparisons [BG81, Bof85, Dun80a, Dun80b, EH83, Hay89, HWH82, HE83, LP84, NM83b, PH86, PG83, Rin83b, Sha81b, BB88, EH84, Fli84, HVS84, Rea84, SI85, TR87, Uus85, Wit88]. Compatible [AP89b]. Complete [HI88, IHC84, KR84]. Completely [Lit88]. Completeness [Kre86, Leh81a]. Complex [Fay85b, RS81, RW88, BS85a]. Complexity [Sha81b]. Component [AC81, SM84c, Wan80b, Ger82, HV86, Pud82, SHS86, Wes87a].

Components [AH81, BHF88, Bro81b, DS82, DRT81, DGK81, Flu84, Kap83, Krz79, LC85a, LC85b, See80, Stu80, ELR87, Khu81, KR86, Krz81, Loh86, MS88a, Wes87b]. Composite [BB85a, Kuo89]. Composition [Smi89b, SS82b]. Compound [Col87, AY82]. Computable [Wes87b]. Computation [Sch88a, Wan80b]. Computational [Alb88b, Ano80f, MMP87]. Computationally [GH83a]. Computations [Hur85a, LLS87]. computer [KP81c, KP82, Mar84b]. Computing [HMP87, MPG85, MPG86]. Concavity [Pra81, Pra82]. Concentration [MP85]. Concentrations [WLW89a]. Concept [Goo82a, PT82a, Sug82, PT82b]. Concerning [GB80, BS84a, Ber89]. Concomitant [BNT86, DM88b]. Concomitant-Variable [DM88b]. Concomitants [Yan81]. Concurrent [PM87]. Condensed [SM83b]. Condition [Ber81, Cas85, Kar80a]. Conditional [AP89b, HWW80, McC86, Ros84b, RC85, SH89, Tsa87, Val87a, Coh81, CGM83, Con89a, FW88, Gew84b, HHO88, SC85]. Conditionally [KS89, KSC89]. Conditionals [AS88]. Conditioning [Kot89, Rob87]. conditions [Bie83b, SL87]. cone [DL88]. Cones [DG85, SSdM85a, SSdM85b]. Confidence [Ber87, Bic87, BK89, BS83a, Bly86, Boo80b, CS80, Cox87a, Der87, DT87a, Efr87b, Efr87a, FH89, Ger82, GW80, Hil87a, KSY85, LL87a, LL81, LW87, Loh87, MB82, Mor87c, PF87, Pie85, Pie87, RC85, SS80a, SL81, Sch85b, Sch87b, Sch86d, Sch87f, See80, TT86, Ber88a, CH83, Cla87a, DT87b, Efr87c, FMD82, FH90, FS87, HB88, IS87, Kan88, Khu81, Kim86, LL87b, Mal82, MSW^{+}83, MPG85, MPG86, Mey87a, Mor86, Nai83, Nai87a, Nai87b, Nyc88, Oma81, PFV89, Ric82, SI85, Uus83, Uus85, YM85, YM86]. configural [Mor86]. confounding [SB81]. Congruential [FM82]. Conjugate [GN82]. Conjugate-Gradient [GN82]. Consensus [Ber81]. Consequences [Whi81a]. Conservative [Ben83a, Ben83b]. conservativeness [Wei81a]. considerations [VM83]. Consistency [Har81a, Bie83b, Sha89a]. Consistent [Sär84, TT84, Wit88]. Constancy [Nyb89]. constant [Nai83]. constant-width [Nai83]. Constrained [EK82, KP81a, KP81b, Cha84d, Sch84c, VW87]. Constraints [HB86, DL88, Nai87b, Wol87a]. Constructing [Boo80b, VY80, Ric82]. construction [CJ88, SB81, Vos86, Wat87]. Constructive [Cox87b]. Consumption [SSR^{+}88]. contagious [Hol87]. Containing [HS81]. Containment [Far83b, Liu83]. Contaminated [BHR83, LWZ86a, LWZ86b, Mac86b, Pre86b, Rog86a, Shi85, SR86, Whi86b]. Contemporary [SK84]. Content [She84a]. Context [Ano81e, DS80, BM85, Gut82]. Contingency [BC89, Bec89, FK80, Fuc82, Gas85, GK83, Gro80, Koe86, SS80a, ACK87, AG83, BOP88, CG84, Coh81, CGM83, CG82, Dal82, GH86, Gil86, GH88, GKR88, Gro81b, GSS89, MP83, MPG85, MPG86, Nai87c, PH81, Sar89]. contingent [SG83]. Continuation [Mar80a]. Continuity [Hab80]. Continuous [McC89a, ME86, Sar87, Sto88, TH82]. Continuous-Time [Sto88]. contour [Har87]. contrast [JK89]. Contrasts [Rub89, MRW87]. Control [BR83b, Cox80a, HC81, CMST88, Fli84, Jac89b, MRW87, Spu88]. controllable [HVY83]. Controlled [Cox87b, VY80]. Controls [DSW83, HY86]. convergence [CH88]. convex [Har87]. Convolution [Dye82, Spr83]. Coordinate [Lea81]. Coordinate-Free [Lea81]. Cornell [Lar84a]. correct [FM83a, Kim86]. correcting [FH78, FH81c]. Correction [Ano85e, BL88a, BK86, BS88b, CMS87, Cla89, FH90, FH81a, FH81c, HM89, HM85c, JTB81, Krz81, LC85a, Mac87, MPG86, NS86, WM81, Whi81b, YM86, KBD89, Sch88c]. Corrections [Ano87d, Ano88e, Ano88d, Ano88f, CSS86, Hab80, Rin85a]. Correlated [GCC89, Hay89, LMR89, WB81, Wal80b]. Correlation [Arn80, BF85a, BK85b, FK85, GS80a, GH87a, Goo81, Hen79, Hen80, PV85, Pre86a, SM84c, Sto80a, Sto83a, Bin82, BF85b, Gil86, KY86, Mui85, SG82, VJ88, dAJ86, Lin82c]. correspondence [GH87b]. Corrigenda [AG82b, Ano80e, Ano80f, Ano81e, Ano84e, Bro83c, EH84, Fli81, GE80, HSG82, HD80, HAJ83, Hen80, HG83, Lin81b, SF81a, Wec82, de 83a, May83a]. Corrigendum [DD82, KP82, Law82, MW82, MT82, Pra82, RC82, Sch82b]. Cost [DD81, EK82, Far83b, Liu83, Sam85c, Tay88, DD82]. Cost-Containment [Far83b, Liu83]. costs [FE88]. Count [Lee80, TL88, Sim87c]. Counting [GM86b, CSC86]. counts [BOP88, Hab88]. County [BHF88]. Covariance [BK84b, BB80, DRT81, Gew81, JF80, LO81, LL89, Rea81, Wer80, CGG85, HVS84, Kar81, Lee88, MB88c, Rei82, SB87, TB86]. Covariance-Stationary [Gew81]. Covariate [ME86, Pre86a, SD88, WK88]. Covariates [BS89, SJ89, Kar87b]. Coverage [RC85, Wol86]. coverages [Yan84]. Cox [Lin84a, Arj88, Bue82b, Gil84, LW89, Spi82, SC85, Tay85]. CPS [DLST86]. Craig [Sti86]. Cramer [Koz80]. Creating [Rub87]. Crime [BR83b, Cha84a, WTB80]. Crime-Control [BR83b]. Criminal [Cha84a]. Criteria [Osb82]. Criterion [Sha81b, Nef82]. Critical [DK80, Har81b, Kau81, May81, Sch81c, Sha81c, Wio81a, Wio81b]. Critique [Gun80, Lin80a, Mar80c, Obe80, PR80, SC80a, SC80b, Thi80, Van80d, Vin80]. Crop [BHF88]. Cross [Gon86, Goo81, HM85c, HM85d, NWGP84, PC84, Sto85a, Agr83b, Clo82, Efr83, GA82, Hur85b, KY86, ST87, Sil84, Sto82]. Cross- [Goo81]. cross-classifications [Agr83b, Clo82]. cross-correlation [KY86]. Cross-Section [Sto85a, GA82, Sto82]. Cross-Sectional [HM85c, HM85d]. Cross-validated [NWGP84]. Cross-Validation [Gon86, PC84, Efr83, Hur85b, ST87, Sil84]. crossings [Cha88b]. Crossover [KF81, LM85b]. Curvature [CW85]. curvatures [Cla87b, Cla89]. Curve [Ano88f, LL89, MS85b, ZJ80, Efr88, HRR81, Mül87, Rei82]. Curves [HS89b, Sch86d, SH86, SW87, Abr88, CC83, Lee88]. Cycling [SSR^{+}88].

D [Hil84, Lin84a, Rip84, Sim87b]. D. [Bue82b, Daw82b]. Damned [Ano85e, Mei86]. Daniel [Sha84b, DH83b]. Darling [SS87a, SS88b]. Data [AP83, AM81a, ALF83a, ALF83b, Ao86, And84a, AH86, Ask80, Ban89, Bas80a, Bas80b, BHF88, Bec83, Bec81, BR83a, Bha85, Bic84, Boa83, BM88, BC82c, Bro83b, BS86c, CHK89, CR84a, CDP82, Cox86b, Cre86, CS83, Dem87, DJK87, DFG87, Din86, DW83, Dok84, DL86b, DL86c, DR81, Edd80, Efr81, Esc86, EO85, EPG89, FH89, FR81a, Fra86b, Fuc82, Gai83, Gas86, Gay83, GCC89, GG80a, GG80b, Hab87, HSG80, HSG82, Hin80, Hin84, HC84, HM85c, HM85d, Hur85b, IPRB85, Joh84a, KL89, Kee82, KW85, KK84b, Kem80, KE86, KS80, LO81, LH89, Lan80, Lat81, Law87, Lee80, Lin80b, LB88, LS87, Lit88, McC86, McK84, MFS84, MP85, Mor87b, Nor84, O’H87b, Oak83, Par80c, PN81, Poi80, RS81, RW88, RT87, Rob86, Rub80a, Rub84a, Rub87, SS88a]. Data [SJ89, SH86, Sco80a, SF81a, SF81b, Sel86, Sil80, Sin88, Spe85a, Sta86, TW83, TW84a, TW87a, Tar80, Tay84, TL88, Var88, Vel80, Wan89, WLW89a, Wol86, Wu85, ZB86, ZJ80, Zie87, dB85, Ala84, AA87, Alb88a, AG85, AP89a, BS84b, BP86, CR84b, CR89, Chr87, FH90, GA82, Gui88, HW86, HVY83, HR84, Hwa86, KL85, KA86, Min87, O’N84, Pad86, Ros89b, SM83a, SS86, Sim87c, SG82, Sto86, Sto82, TW87b, VJ88, WP85, WLW89b, Wit88, Lar84a]. Data-Based [SF81a, SF81b, TW84a]. Data-driven [Hur85b]. David [Lin82c, Pro84, Bue82b, HHM88b]. Davis [Cha84c]. Dawid [Oak85]. Dayal [Lun84]. Dean [O’B84]. Death [AR89, Din86, Kad83, Sti86]. Decennial [Cre89]. Decision [BCF^{+}81, Muk84, RN80, Tai89, Sha89b, WF82]. Decisions [Mar80a]. Decomposability [GK83]. Decomposition [BNT86, Har85b, Tak83]. Decompositions [JV85a]. Deconvolution [MR82]. deconvolving [CH88]. decreasing [SM83a]. Deductibles [NRMM80]. Definition [Vel80]. Deformation [Sto88]. DeGroot [Ber81]. Delay [Fou80]. deletion [Vos86]. Demand [DFG87, HM85c, HM85d, NRMM80]. Democratic [Ord87]. Demographic [IPRB85, Kad81, Lee81, Lit81, Lon81, McD81a, McD81b, TW84b, Wal81a]. Demographic-Economic [IPRB85]. Dennis [Kas85]. Denominator [JH84]. Densities [IS88, Len88, TK86]. Density [Ask80, Dav81, Edd80, FSS84, GG80a, GG80b, LF82a, Leo82, OS87a, Par80c, Poi80, Sco80a, SF81a, SF81b, Sil80, Tar80, TS85, WABA89, CH88, DH86, DM88c, Har84, Har81a, Har87, Jon89, LF82b, ST87, SW83, Won82b, Won83b, Yak85]. Density-Dependent [WABA89]. Department [And84a, Hwa86]. Dependence [Gew82a, JH82, Joe89, JH84, Par82, Ped80, Pie82, WM82, Wei82, Zel82, BS84b, Gew82b, Gew84b, GJ84, HM88a, Har84]. Dependent [BC82d, BMR88, Dal83b, HMT83a, HMT83b, Lit83b, Roy83, Rub83, Sär84, Smi83b, WABA89, Agn85, Wei81a, Zeh88, Sch83d, Sch84b]. Depends [GRZ82]. Derivatives [McC82, HS89a]. derived [ELR87, Rot83, Sch81b, Sch82b]. Design [AP83, BS88a, LG86, Sär84, Shi85, TGG89, BHS81, BS85a, BHT88, CSS89, IF82, O’G82, SHS86, Tam85]. design-based [BHT88]. Design-Consistent [Sär84]. Designing [Pro85]. Designs [BB85a, Bel80, CT80, DG86, HI88, KF81, Lar84a, LM85b, Sha81a, WG88, Won84, Wu85, YBN85, CJ88, CMST88, CS89, HS89c, Hsu82b, Jac89b, KR84, KR88, Min87, SB81, Stu82, Vos86, Wit88, dAJ86, Mye84]. destructively [dAJ86]. Detecting [CDP82, DL89, FK80, KR82, KR88].

Detection [CD80, Hsu82a, JG83, KP83, Whi81a]. Determination [DP84, DK80, Noe87]. Determining [Ash83, NP83, Cre88]. deterministic [Nyb86]. Development [CM84b, WLW89a]. developments [Tsu86]. deviance [Sim89]. Deviation [AG81]. Diagnosis [MT80]. Diagnostic [EPG89, LM85c, SS86, SC85]. Diagnostics [Law88, Pol84a]. Diagnostis [TW84a]. dialing [Pot87]. Diaries [Gro86]. Diary [JK87]. Dichotomies [SO83]. Difference [SW89b, Hay86, WP85]. Differences [BE82, SH86, MC84]. Differencing [Sol84]. Different [Rao80b]. Differential [Bra84b]. digit [Pot87]. Digital [Blo84]. Dimensional [Hub87, Abr88, Dal82, NWGP84]. Dimensions [Ege81, Jac81, KH81a, KH81b, Wai81, Har87]. diminishing [BJM81]. Direct [Ros87c, BJM81]. Directed [Bre81, FW81, Fra81b, Hab81, HL81a, WW87a, Won87, HL81b]. Directional [FH89, FH90]. directly [Eri81a]. Directory [WMH85]. Dirichlet [AG83]. Discipline [Bra82]. Disclosure [Cox80a, Cox86b, DL86b, DL86c, Fra86b, Gas86, Rob86]. Disclosure-Limited [Cox86b, DL86b, DL86c, Fra86b, Gas86, Rob86]. Discontinuity [BR83b]. discontinuous [Gle85b]. Discoveries [Sor89]. discovering [CF87b]. Discovery [Daw84, Gew84a, PS84a, PS84b, RR84a]. Discrete [McC86, Sam80, Wil88b, BS83b, Pic87]. Discretized [Jon89]. discriminability [GC82]. Discriminant [AC81, BJ81, Fat83, Fri89, Rav89, Say80, TH82, CM84a, FH86, LV88]. Discriminating [LS89]. Discrimination [Bri82, BK85c, Fie82, Jon82, Kay82a, Kay82b, O’N80]. Discussion [Pre86a, BF85b, Gew82b, HR84, Hol86b, Hub83b, LL87b, Mor83b, Sha82c, TW87b, VSK85b, WHM85b]. Dispersion [DGK81, Hab82, LC85a, LC85b, Mat83]. Dissemination [Cox86b, DL86b, DL86c, Fra86b, Gas86, Rob86]. Distance [LW85, PS80, Boo81a, ELR87, Rom88, Sim87c, TB86, WPSL84]. Distributed [HW84, LaR86a, Lut80, Sam80, SB80b, CM85, Ger82]. Distribution [AM81a, AM81b, AG82b, AG82a, AG80, BE82, Ben83a, CW80, Dyk80, Esc86, Fli81, FK80, GH80, GB80, HWW80, HSG80, HSG82, Has80, Hen80, JT81, Kor80, Lee80, LP84, McC86, Mor81, Nan80, O’N80, OPZ81, Oma85, Pre86a, RFPW80, Rin83a, RS82, Sar87, SM83b, SW81, SS88b, SRMS81, SW80c, Wal80b, WP85, YM85, YM86, AY82, Ben83b, BK89, BS86b, CL85a, CH83, Cro85, De 83b, FE88, Fli84, Gla89, IHC84, IS83, KR84, LLT89, Len88, MO85, MT81, MT82, OW82, Ran89, Rin85a, Riv86, Sam81, SZJ87, SS88c, Tad80, VJ88, Wal86b, WL84, Zac86]. Distribution-Free [Fli81, HWW80, RFPW80, Rin83a, WP85, YM85, YM86, Fli84, IHC84, KR84, OW82, Ran89, Rin85a, WL84]. Distributional [BSB80, Stu80]. Distributions [AS88, AP89b, BGT80, BM81, Bre81, BHR83, CSS81, CSS86, Dem87, Dye82, FR83, FW81, Fra81b, GR81, GW84, Ger80, GD80, Hab81, Hab87, Hen79, HL81a, HPP86, Lax85, McC89a, MB82, Mor87b, NJ88, NS83, NS86, Nan80, Nau82, O’H87b, Rub87, Say80, Sha81c, Shi85, TW80, TW87a, Wal80a, Wil88b, AG83, CKC83, GJ84, Gle85b, HMP87, HL81b, Jen86b, JTB80, JTB81, Kun80, MO85, MO88, NWGP84, PT83, SW89a, SM83a, SS82a, SB87, Suz85, TW87b, Tay85, WLW89b, Wes87a, Zel87b, vd87]. Disturbance [WM82, Kob86]. Disturbances [Raj80]. divergence [Rea84]. Diversity [Goo82a, PT82b, PT82a, Sug82, Zah84b]. Do [Daw85, Sch85d, Oak85]. Doksum [BC82b]. Domain [LS83, SH89, Tan83, Zeg85, ZH87]. Domains [Sär84]. domination [VM83]. Don’t [SG83]. Dose [BB89b, Rub89, Sch86d, CC83]. Dose-Response [Sch86d, CC83]. Double [Efr86a]. Doubly [EO85]. Dowdy [Tho84c]. Drilling [EK82]. driven [Hur85b]. Drivers [Tom88]. Dual [LG86, NM83a]. Duality [DL88]. Duration [RSW80]. Durbin [WM82]. Dynamic [AH81, Dav85b, DC85, FT85, Ghy88, GP85, Los85, Mey85, MFS84, WHM85b, WHM85a, Win85]. Dynamics [Rog86b].

Each [BNT86]. Early [Hur85a, Kes85]. Earthquake [GAF83, Oga88]. Easy [Hor83]. ecological [BP86, Zah84b]. Econometricians [Fis86]. Econometrics [Sch84b]. Economic [Ghy88, HP84, IPRB85, HT87b, Nef82]. Economics [Bro84, Zah84a]. Economy [For80a, For80b, Sto80b, Zel80b]. ed [Blo84]. ED50 [JJW84b]. Edgington [Gre81a]. Editing [LS87]. Editions [Ano89o, Ano89p]. Editor [Ano87h, Ano88j, Ano88k, Ano89i, Ano89j, SST88, SST89, SST87]. Editorial [Ano85f]. Editors [KM86]. Effect [Ber82, JH84, LL89, LM85b, MT80, ME86, NRMM80, Ped80, SH82, Sor89, TW80, BK85a, BK86, Bin82, CWR88, Hod88, SHS86, WHH88]. Effect-Size [Sor89]. Effective [Rub89]. Effects [Ano81e, Bha81, BCF^{+}81, Cah87, CK89, CD80, CD82, CW85, Dag82, DS80, DL80, Fat83, GH83a, KH84, KF81, KE86, Kiv80, LWZ86a, LWZ86b, LB88, Mac86b, MS80, Mis85, MRB86, Ord87, Pre86b, RKC81, Rei85, Rog86a, Sik85, SSR^{+}88, SR86, Whi86b, Zie87, Bow87, CR89, Dav82, Eri81a, KR88, MS88a, Rei82, Rei84, YD84]. Efficacies [TA89]. Efficiencies [Ano87d, HI88, Mac86a, Mac87, Rao81, Kob85]. Efficiency [BB81, DP89, Gas85, HHO88, Har84, JH82, Krä80, Rea81, Tar87, Bin82, CD89a, Gro81b, HD88, KD87, Rog88, Sha89a, Sim89, Wei81a]. Efficient [JJW84b, JK80, JV85a, KW82, Wu85, JJW84a, Sil84, SS87b, Spi82, Vos86, YZ88]. Eisenhart [SM84b]. Elasticity [HM85c, HM85d]. Electricity [EGRW86, HM85d]. Elementary [How84]. Elicitation [KDW^{+}80, KW88]. eliminant [Pud82]. elimination [BB88, Mar88]. elliptical [SB87]. Emission [CL85b, HCGL85, Kea85, LM85a, Rub85, Tha85, VSK85a, VSK85b]. Empirical [BK82, Bel89b, Ber83b, BS89, CD89b, CD89c, CH83, Cle89, Cre89, DL81, Dem80, Dem83, Der87, Efr87a, Ful81, Gew89, GM86a, Hil87a, Hin83, HB83, KS89, Kay80, KL80, LL87b, LL87a, Leo83, Lin83a, MWS^{+}89, Mor83a, Mor87c, Pie89, Pro85, Rei85, Rog80, RC81a, RC81b, Rub80b, Rub80c, Sär81, Sco80b, Smi81, Sti83b, Tho81, Wat89, Wei80c, Wre81, FH82, FP84, GL87, Lou84, Mor83b, RC81c]. employment [GG87]. Encyclopedia [Moo84b, Moo84b]. Endpoints [TGG89]. Energy [MH89, Hwa86]. entered [BF83]. Entropy [DvdM81, Joe89, SC89]. Entropy-Based [DvdM81]. Epidemic [BG88]. Equal [GS80b, GG87]. Equality [AS85b, BM81, Kar80a, Kob86, Sin86, Wol87a]. Equation [BG80, HMR80, Raj80, Bin82, BF83, FP84]. Equations [Wer80, Kun80, Bra84b, Kas85]. Equitable [Spe85b]. Equivalence [DM88c, Esc86, BB88, PP88]. Ericson [God82c]. Error [AH81, AN81, BHF88, Van80a, CF87a, EPG89, GAF83, Gon86, Har85b, LG86, O’N80, RKC81, Rei85, Say80, Tar86, Wol86, Bha81, Efr83, Efr86b, HV86, Hay86, Kor81, Law81, Law82, Sam85a, Sch86a, WK88, Moo84b]. Error-Components [BHF88]. Errors [BK82, Bof85, Bro82c, Bro83c, BS86c, CF80, Cha81, HW85, Jus86, KH84, Kiv80, Krä80, Lin85, MT80, Pre86a, Sel86, SB83, Tsa84, CGG85, CR89, CSC86, Hwa86, Kob85, KD87, Sch86a]. Errors-in-Variables [CF80, CGG85, Hwa86, Sch86a]. Essche [Ger80]. Essential [Bai88]. Estimate [Dav81, Fer82, MK80, Dua83, Eri81a, GA82, Hur85b, Sta89b, Moo84b]. Estimated [KK84b]. Estimates [AN81, Blo80, Boo80b, BY86, EGRW86, GL83, Hec88, LH86, PH84, SM84b, Sta89c, TS85, UH80, Wan80b, Atw84, Jon89, KR86, Kot89, Lea84, MZ89, MS88b, Rob87, TT84, Wes87b, Wes87a, YZ88]. Estimating [Bai85, BR83a, BB89a, BD81b, BM88, BPN87, BS86b, BF85b, BF85a, BK85b, Cha88b, Cle82, Efr83, EK85a, EK85b, Fay85a, Fel85, FK85, GG87, Han85a, Har81b, HP84, Hau85, HH89a, HH89b, HW84, Hol89, HM85c, HM85d, JR85, KJ87, Kau81, KF81, Kna82, Lit83a, Lou84, May81, Mei82, MP85, Mof89, Mui85, Pas85, PV85, Pre85, Pud82, Rao80b, Rol85, Rot82, RS82, Sch81c, Sed85, ST82b, Smi89c, Smi89b, Spr83, Sta86, Sti87, Sto88, Tib88, Tuk85a, Tuk85b, Tum82, WT82a, WT82b, WS89, Wil80, Wio81a, Wio81b, Zac86, CWR88, CC83, MPS88, Sha85, SS88c, WP85, dAJ86]. Estimation [AH81, AK86, Ask80, BM85, Bet89, BB81, Bro82c, Bro83c, BHR83, BS86c, CF80, Cas82, Cha86b, fC81, CKC83, CW80, Cre89, DD88, DS82, Dav81, DRT81, DW87, DGK81, DR81, DT80, Edd80, Eri82, EPG89, Fou80, FG80, FSS84, Fuc82, God82c, God82b, Gon86, GG80a, GG80b, HM88b, Har87, HC81, HMM85, HB83, Isa83, JK80, JF80, KD89, Ked80, KA86, Kou80, KW82, Kuo89, LL89, Leh83, Mau83, MV85, MT80, Mis85, Muk84, OS87a, Osb82, PS80, Par80c, Poi80, RN80, Rei84, Ruk86, RC80, ST82a, Sam80, SW89a, Sär84, SH89, SB89, Sco80a, SW80b, Sha83, Shi80, Sil80, SW87, SY84, SW89b, SM84c, Smi86, Sor89, Sto86, Sto85a, Tan83, TW84a, Tar80, Tha84, TH87, Wan80a, WM82, Yan81, Zin80, de 81, de 83a, AG83, AP87]. estimation [AP89a, BS83b, BHT88, CF87b, DC87, DB83, DM88c, Dyk82, DL88, FD85, Fin83, Fry84, FH78, FH81c, GM86a, GL87, GSS89, HM89, HW86, IS83, JSG88, KP87, Kre86, KSC89, Lee88, Mat83, NWGP84, Roy86, RS86, RC82, SH81, SM83a, Sch84c, ST87, SW80a, Sim87c, SW83, SS82b, Smo82, Spi82, Spl83, SG82, Ste87b, SZJ87, Suz85, Tak81, Tam88a, TB86, Val85, VW87, WG83, Wij85, Won83b, WPSL84, Yak85, Zel86a, vd87, Moo84a]. estimations [MSW^{+}83]. Estimator [Biy80, DW87, Ful81, GW86, GS80d, JJW84b, Mor81, MG84, RC81a, RC81b, Sär81, Smi81, Tho81, Wre81, Akr86, Boo82, Cha84d, CHL82, Cro85, DH86, FM83a, Har84, JJW84a, Kar81, Pad86, RC81c, Rup85, Yan85]. Estimators [Ban86b, Bha85, BB85b, Cas86b, Cau80, Cha81, Din86, Esc86, Ful81, Gal80, Geo86, Gib81, GHP87, GWL87, HAL82, HAJ83, HS81, KH84, Kap83, Kor80, LW85, Lax85, OPZ81, Raj80, Rao81, RKC81, Rei85, RC81a, RC81b, Sär81, SW81, SF81a, SF81b, Smi81, Tho81, Umb80, Wre81, BLE88, Bie83b, Boo81a, BT81, CL85a, Cas85, DWL87, DB83, Gol83b, HRR81, HV86, Hoe84, HT81, Kob85, Kun80, Law81, Law82, Loh86, Moo81b, Nai84, O’H87a, PMF88, RC81c, Ruk87, SM84a, SL87, Sha89a, SW83, Spo82, Tsu84, Tsu86, Val87a, Wol84]. Eugene [Gre81a]. Evaluating [Far83b, Liu83]. Evaluation [BB85a, BR83b, BS85b, Dal83b, FM82, GH83b, HMT83a, HMT83b, Lit83b, Roy83, Rub83, Smi83b, Smi86, GC82, Cob87b]. Event [MT80, KP83]. Events [SM83b, TL88]. Evidence [BS87b, Bri82, CB87b, Dic87, Fie82, Goo87b, Hin87b, Jon82, Kay82a, Kay82b, KM83, Mor87d, Pra87b, Var87, BS87a, Bor87, CB87a]. Exact [CS89, Gle85b, Gui88, Hen79, Hen80, Nan80, See80, SI85, Tri84a, Uus83, Uus85, HMP87, Mar84b, MP83, MPG85, MPG86, MPS88, PH81, Wol87a]. exact-approximation [Mar84b]. Example [Bai87, Nor84, Sta86]. examples [Sim89, Lin84a]. Excess [Gon86]. Exchangeable [Gol86b]. Exemplified [Ask80, Edd80, GG80a, GG80b, Par80c, Poi80, Sco80a, Sil80, Tar80]. Exist [Daw85, Sch85d, Oak85]. existence [KR86]. exogenous [Spl83]. Expansions [Mor81, Kun80]. Expectation [DT80, KS80]. Expectations [GD80, TKK89]. Expected [HB86, Hab88, Sha84a]. Expenditure [Gro86, KJ87]. Experience [Cha84c]. Experiences [Hub87]. Experiment [Ano81e, DS80, Far83b, Liu83]. Experimental [AP83, Bas80a, Bas80b, BS88a, Hin80, Hoo89, Kem80, Lan80, Lin80b, Rub80a, dAJ86]. Experimentation [CM84b]. Experiments [Din86, GH83b, HM85a, LMR89, Sch86d, BS85a, FB83, MV88, MC84, Lar84a]. Expert [Mei86, Sin88, LS86]. experts [Agn85]. Explicit [Gal80]. Exploratory [Ban89, ES82b, Fri87, TW84a, Gil86]. Exploring [Zeg85]. Exponential [AS88, BM81, Bre81, Cha81, FW81, Fra81b, FG80, GW88, Hab81, HL81a, MB82, RS82, Sar87, SW80c, BLE88, BK85a, BK86, CKC83, Efr86a, FMD82, HL81b, MS85c, SW80a, Tad80, Tak81, TKK89, Zac86]. Exponentiality [SS88a]. exponentially [Ger82]. Exposure [Dan89]. Extended [Pre86a, TT84]. extending [Hur85b]. Extension [Lut80, SZJ87]. Extensions [O’B88a]. Extraction [Mar86, WA83]. Extrapolation [BY88, Blo80, Sch86d]. Extreme [BMR88, CSS81, CSS86, Flu85]. Extreme-Value [CSS86]. Eye [Moo84b].

F [Dic84, DH83b, Pro84, Pro85]. Faces [CM84c, FR81a]. Factor [Dag82, EW81, GS80c, MS80, FB83]. Factorial [BCF^{+}81, CJ88, SB81, Vos86, Mye84]. Factors [MO83, Tom88, Del89]. Failure [DFH89, FG80, Gla80, TH80, Sch88b, SW83, WLW89b]. failure-rate [SW83]. failure-time [Sch88b]. failures [Suz85]. Families [Ger80, MO88, BLE88, EH87, Efr86a, Loh85, SW80a]. Family [BS88a, Bre81, FW81, Fra81b, GW88, Hab81, HL81a, JT81, LW85, McC89a, NJ88, Wil88b, CKC83, CC83, HL81b, JTB80, JTB81, MO85]. familywise [Hay86]. Far [HHM88a, Joh88, Sco88, HHM88b]. fast [EH87, Sil84, Spi82, Spl83]. fault [LS86]. Faults [JR85]. Favorable [Eth82]. Feasible [GH83a]. Features [VY80, HVY83]. Federal [KD89]. Federer [Mye84]. Feedback [Gew82a, Par82, Pie82, Wei82, Zel82, Gew82b, Gew84b]. Feedforward [Whi89, Fab94, Whi92]. Feeding [MRB86]. Fertility [Bec81, BS85b, CL86, CV89, Coh89, HW89, Sin89a, TBLM89, Tur89b, dGJ89]. Few [Rub87]. fields [Pic87]. Fienberg [Cha84a, HL81b]. Filter [MS89, Zeh88]. Filters [Dag82, Wat87, Blo84]. Finance [Bro84]. Finding [BB89b, GL88, PH81]. Findley [Dic84]. Finite [AG81, Cha86b, DD81, DT80, HMR80, HAL82, HAJ83, Krä80, Lax85, Lit83a, RC85, Smo84, Wri83, AK86, Car82, DD82, GM86a, Lin85, MS85a, Mey87a, RC81c, Smo82, Tam88a, Val85]. finite-population [RC81c]. Finite-Sample [HAL82, HAJ83, Lax85]. First [Vos86, WM82, BHS81, HM88a]. First-Order [WM82, Vos86, BHS81, HM88a]. Fish [Wei86, SS82b]. Fisher [Kru80a, AC81, Bas80a, Bas80b, FP81, Hay86, Hin80, JW88, Kem80, Kru80a, Lan80, Lin80b, MP83, Rub80a]. Fit [Arj88, CSS81, CSS86, DKMO89, Fel80, Gre85a, KL80, Koe86, McC86, MS85b, RS81, Sch85c, Sim85d, SRMS81, Akr88, Boo81a, CM85, Gle85b, KOS85a, O’N84, Ole85, Rea84, Sim86, TR87, Wei84a, Zel87b]. Fits [EHK84]. Fitted [EO85, Tay86, Tuk85c, TV85]. Fitting [Ano88f, Pro85, But84, CR84b, Mül87]. Five [Kos85b]. Fixed [KH84, MS88a]. Flexible [DG86, MH89]. Florida [Jar89]. Flows [CF87a, Sta86]. follow [Suz85]. follow-ups [Suz85]. Following [SW89b, CS87]. follows [Krä86]. Football [Har80b]. Force [Sta86]. Forecast [AN81, Kad81, Krz87, Lee81, Lit81, Lon81, McD81a, McD81b, Wal81a]. Forecaster [McK84]. Forecasting [Ao86, AS85a, Dav85b, DC85, FT85, GP85, HW89, IPRB85, Los85, McK84, Mey85, MW84, Smi88, TB85, WHM85a, WH86, Win85, WHM85b]. Forecasts [CL86, Lan86, Sti87]. form [Sch87g]. Forming [BS89]. Forms [BSB80, HVY83]. Forsythe [Rot83]. Forward [Gon86]. Fossil [Har81b, Kau81, May81, Sch81c, Wio81a, Wio81b]. Found [WB85]. four [Uus85]. Fourier [FM81, SSR^{+}88]. Fractions [Rub87]. Frailties [Oak89]. Frame [Ban86b, LG86]. Framework [Cob87b]. Frank [HL81b]. Frederick [Zah84b]. Free [Fli81, HWW80, Lea81, RFPW80, Rin83a, YBN85, CJ88, Fli84, IHC84, KR84, OW82, Ran89, Rin85a, WL84, WP85, YM85, YM86]. Frequencies [HB86, Tom88]. Frequency [KK84b, Sco85, Tan83, Zeg85, ZH87, Hab88, IS87]. Frequency-Domain [ZH87]. Frequentist [Cau80, CB87a]. frequentistic [Fin83]. Friedman [Wit88, JH82, LV88, MS80]. Friedman-type [Wit88, MS80]. Front [Ano80g, Ano80h, Ano80i, Ano80j, Ano82e, Ano83e, Ano83f, Ano83g, Ano83h, Ano84f, Ano84g, Ano84h, Ano84i, Ano85g, Ano85h, Ano85i, Ano85j, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano87g, Ano88g, Ano88h, Ano88i, Ano89e, Ano89f, Ano89g, Ano89h]. Fuel [Har81b, Kau81, May81, Sch81c, Wio81a, Wio81b]. Fully [TKK89]. Function [AC81, BK82, DP84, LF82a, Leo82, Shi80, SS88b, TW84a, Yan81, BK89, Bie83b, BS86b, Cla87a, DC87, FW88, GA82, KY86, LF82b, LP86, Mui85, Pad86, SW83, Sta89b, TT84, Yan85, de 81, de 83a]. Functional [MK80, MH89, Kun80]. Functionally [JF80]. Functions [Gal80, Kad81, KSY85, Lam81, Lee81, Lit81, Lon81, McD81a, McD81b, Osb82, Wal81a, Yan81, AP87, BK85c, De 83b, Dic83b, Dyk82, FD85, Khu81, Nai87a, OYR86, PS82, SS88c, TKK89, Val87b, Zel86a]. Further [Tsu86]. Future [Bra82, SM83b]. Fuzzy [Net86].

G [Bro84, Hol85, How84, Sch84b, Zah84a, Zah84b]. Games [Har80b]. Gamma [GB80, Nan80, Nel85, SL80, GR81]. Gamma-Poisson [Nel85]. Gas [EK82]. Gauss [HM85a, Kar85b, KP81a, KP81b]. Gaussian [MT82, DZ89, FB83, IS83, Kit87, Kor80, MT81]. Gehan [Wei80a]. Geisser [BH88]. General [DP84, ER86, Gol83b, KP81a, KP81b, O’N80, Rei85, Fin83, Gui83, HVY83, KSC89, MB88c, Pol84a, Raf87, Ric82, Ste81, SC85, Tsi82]. generalization [Zeh88]. Generalizations [Spu88, Pot87]. Generalized [Dav85b, DC85, Dye82, FT85, GN82, GWL87, GP85, HT87a, Los85, Mey85, PS86, Stu88, TW89, Umb80, Val87b, Wei80a, WHM85a, Win85, Alb88b, Bie83b, CR82, DD88, DWL87, Efr86a, FHO87, KSC89, Law81, Law82, Lea84, LV88, OYR86, PP88, Pud82, Rei84, WHM85b]. Generalizing [FL89, Tay88]. Generally [Hay89]. generated [MO85]. Generating [HVY83, SL80, Mar84b]. generation [KP81c, KP82, SS80b]. Generators [Col87, FM82]. Genest [BH88]. Genetic [Nor84, Tai89]. Geographic [Tsu88]. Geographically [WMH85]. Geometric [Nan80, GH87b, HM83]. George [PT82b]. Gilbert [Wei80a]. Gini [Sen86]. Glioma [MHMEH82]. Global [FMT89, Lea84]. GNP [Sto88]. Goals [And84b]. Going [And84b]. Good [PT82b]. Goodman [GR81]. Goodness [Arj88, Fel80, Gre85a, KL80, Koe86, McC86, MS85b, RS81, Sch85c, Sim85d, SRMS81, Zel87b, Akr88, Boo81a, CM85, Gle85b, KOS85a, O’N84, Ole85, Rea84, Sim86, TR87, Wei84a]. Goodness-of-Fit [Gre85a, KL80, Koe86, McC86, Sch85c, Sim85d, Zel87b, Akr88, Gle85b, KOS85a, O’N84, Ole85, Rea84, Sim86, TR87]. Govindarajulu [Muk84]. Gradient [GN82, JJ88]. Granularity [HD88]. Graph [CMM88, SH87]. Graphical [Arj88, AM84, CD80, CT82, CM84b, FG84, FR81a, Has84, HWH82, LPS84a, LPS84b, Rub84b, Cre88]. Graphics [Cle87, Edd81, FR81b, FR81c, Gen81, Mal81, Pre81a, Kos85b]. Graphs [Bre81, FW81, Fra81b, FS86, Hab81, HL81a, WW87a, Won87, FH82, HL81b]. Grassle [Zah84b]. Greenberg [Pro84]. Gross [Sta86]. Group [Fou80, Mis85, TGG89, HPE81, Ans80a, Kru80b, Wal80c, Wal80d]. group-testing [HPE81]. Grouped [HSG80, HSG82, Ala84, SG82]. Grouping [AM81a, Eld81, Gil86]. Groups [Flu84, Krz79, Tuk85c, Krz81, TV85]. Growth [BB85a, LL89, MHMEH82, SH86, ZJ80, Abr88, BPN87, Lee88, Rei82]. Growth-Curve [LL89]. Grubbs [BB85b, Sch80a]. Grundy [Biy80]. Guessers [SZ80].

H [Fab94, LV88, SK84]. H. [Pro85]. Haberman [HL81b]. Haenszel [WW87b]. Hagen [Sha84a]. Hamming [Blo84]. Handle [BMR88]. Hans [McK84]. Hard [Ked80, Sav85]. Hard-Soft [Sav85]. Hartley [Riv86]. Hausman [Thu85]. Having [Goo81, Sto83a, Agr83b, Clo82]. Hazard [BNT86, TW84a, Tha84, SM83a]. Hazards [Arj88, LW89, Wei84a]. Health [And84a, LWZ86a, LWZ86b, Mac86b, MRB86, Pre86b, Rog86a, SR86, Whi86b, And84a]. Hearing [Kad83]. Heart [ALF83a, ALF83b, Bec83, Boa83, CS83, Gai83, Gay83, Oak83, TW83]. Hedayat [Mye84]. Heights [Cle82, Kna82, Mei82, Rot82, Tum82, WT82a, WT82b]. Helge [Sch84a]. Hellinger [Sim87c, Sim89, TB86]. Hematologial [And84a]. Herfindahl [MP85]. Hersh [Cha84c]. Heterogeneity [MSV86]. Heteroscedastic [Tsa87, CR82, FH86, Kar81]. Heteroscedasticity [AS85b]. Heuristics [Sha84b]. Hidden [Whi89, Fab94, Whi92]. hidden-layer [Fab94, Whi92]. Hierarchical [FM83b, FM83c, GD81, JM83, KS89, Kor81, Mor87b, Van83, Tur83, Wal83, Won83a, WM85, Alb88b, Goe83, HPE81]. High [BB81, YZ88, Abr88, Har81a, Won82b]. high-density [Har81a, Won82b]. high-dimensional [Abr88]. High-Efficiency [BB81]. higher [Mar81b]. highly [Ger82]. Hildreth [Cro85]. Hildreth-Houck [Cro85]. Hinkley [Bue82b]. Histograms [MR82]. Historical [Cle82, DSW83, Kna82, Mei82, Rot82, Tum82, WT82a, WT82b, HY86]. history [Sti86]. Hoaglin [Sim87b, HIT87]. Hodges [Boo82]. Hodges-Lehmann [Boo82]. Hoel [Zah84a]. holding [FE88]. Homogeneity [ZC88]. Homogeneous [Dun80a]. Homoscedasticity [RA89b]. Hong [Gro86]. Hope [Hog89]. Hospital [Far83b, Liu83]. Houck [Cro85]. Household [Gro86]. Housing [EKT89a, EKT89b, SM84c, Smi86]. Howland [MZ80]. Hrishikesh [Hil84]. Hultquist [SB81]. Human [And84a, Ban89, Kos85b]. Human-Rights [Ban89]. Humans [DH83a, Kas83, Kre83, Smi83a, DH83b]. Hunter [Sto83a]. Hunting [Ask80, Edd80, GG80a, GG80b, Par80c, Poi80, Sco80a, Sil80, Tar80]. hybrid [FMD82, Won82b]. Hypergeometric [LP84, Dic83b, Wal86b]. Hyperparamters [GD81]. Hypotheses [FR83, GV86, HD80, Sha84a, TW89, BS84a, Ber89, Del89, HT81, Loh85, Rog88, SS82a, WR84]. Hypothesis [AG82b, AG82a, BS87b, CB87b, CGM83, Dic87, Goo87b, Hin87b, KK84b, Mor87d, Muk84, Pra87b, SD85a, SD85b, Var87, BS87a, Bor87].

Hypothetical [DFG87].

I. [PT82b, TV85]. Iain [HHM88b]. Identical [AG82b, AG82a]. identically [CM85]. identifiability [Hod88]. Identification [GH80, Man88, WG81, Sto86]. Identifying [PB87, Rub89, Won82b]. Ifs [Rub86]. Iglewicz [Sim87b, HIT87]. Ignorable [Ros84c, RS86]. II [BM81, BM83, Bha85, BHS81, Esc86, LaR86a, MB82, SM84b, SM84c]. iid [BK82]. III [Bof85]. Ill [MMP87]. Ill-Posed [MMP87]. Impact [HH89a, HH89b, Hol89, Mof89, RT87]. imperfect [WS89]. Implementation [LNPS81]. implicit [FJV89]. implied [HHO88]. Importance [Mar87a, MPS88, Rub87]. impossible [Gut82]. Improved [GSS89, Kiv80, Loh86, MO83, Mor81, Ruk86, Sim85d]. Improvement [Fin84, Efr83]. Improving [LH86, MK80]. Imputation [DLST86, GRZ82, LS87, RS86]. Imputations [Rub87]. Imputed [GRZ82]. imputing [CS86]. Inadequate [Leh83]. Inadmissibility [Biy80, Moo81b]. Incidence [EPG89, MC84]. Incidents [BS88a]. Includes [WM82]. Income [AM81a, Ash83, Att80, DLST86, HM85c, HM85d]. Income-Tested [Ash83]. Incomplete [BS88a, Din86, KS80, Wei83a, CS89, MJ82, WL84, WLW89b]. Inconsistent [Fer82]. Incorporating [BNT86, HY86, Sin88]. incorrect [Mat83]. Increasing [BEW85, BGT80]. Independence [Fel80, Ger80, Gew81, Gro80, RS81, Coh81, CGM83, Gro81b, GSS89, KY86, Kru88, Sar89]. Independent [Cha81, KS89, Krä86, SS85]. independently [HVY83]. Index [Ano80e, MP85]. indexes [Sen86]. Indicators [Ano81e, fC81, DS80, LM83, Cha84a]. indices [FH82]. Induced [Oak89, Pre86a]. Inequalities [Dyk80, Ber89, GJ84]. Inequality [VW87, Wij85, Wol87a]. Inequality-constrained [VW87]. Inference [AY82, Ber83b, BA83, CSS89, Coh81, Cox86a, Dem83, FM81, GH83a, Gly86, Gra86, Hin83, Hol86a, Jen86a, KL89, KS89, Leo83, LHT89, Lin83a, MKR89, MH82, Mor83a, Nor84, Pic87, Raf87, RW85, RW88, Rub86, Sch83b, Sch88a, Smo84, Sti83b, Sto83a, Tai89, BS84b, Cha88a, CS87, FJV89, FH82, Hol86b, Hoo89, Hsu82b, LW89, MS85a, MPS88, Mor83b, Swe84, Tay84]. Inferences [Dal83b, GB80, HMT83a, HMT83b, Lit83b, Roy83, Rub83, Smi83b, Sto80a, HM83]. Influence [Bic83, Car83, CW83, DWL87, GWL87, Hub83a, KD89, KW82, KW83, Lam81, Law88, Mal83, McC89b, MS85b, FJV89, Hub83b, KSC89, Rup85, Sam85a, dJdWW88]. influential [JG83]. Information [Ano80s, Ano81n, Ano82n, Ano83q, Ano84r, Ano85s, Ano86p, Ano88s, Ano89u, BNT86, For80a, For80b, GD81, Goe83, Isa83, Kos85b, Leh83, Mor81, Rub87, Sch84a, SH87, Sto80b, Zel80b, CS89, Sch86a, SG83, Smo82, Wri83]. Information-Processing [SH87]. Informative [Lin89b]. Inner [SL81, Mey87a]. Insensitive [DH80]. inspection [Lin85]. Institutional [LH86]. Instrumental [CF80, BT81]. Integer [LR87, Lin85]. integrals [Wij85]. Integrated [Has80, Lun84, Tar86, PP88]. Intensity [BEW85, DM88c]. Intensive [LTAP88]. Interaction [Sne82, Boi86, Dal82, MJ82, TA89]. Interactions [Bai88]. Interactive [KDW^{+}80, Sin88]. interblock [CS89]. Intercensal [LH86]. Intercepted [Var88]. Interchangeability [Wei83a]. interdirections [Ran89]. interpolated [Jon89]. interpolation [KA86, Sch88b, dJ89]. Interpretability [Sha81b]. Interpretation [Wan80b, GH87b, HM83, Leh81a]. Interpretations [McD80, Dic83b]. Interpreting [GS80c]. interpretive [GKR88]. intersection [BS86a]. Interstate [IPRB85]. Interval [Hor88, Muk84, TL88, Del89, FMD82, MPG85, MPG86, RS86]. Intervals [Ber87, Bic87, BS83a, Boo80b, BR80, Cox87a, Der87, DT87a, Efr87b, Efr87a, GW80, Hil87a, LL87a, LW87, MB82, Mor87c, PF87, RC85, SS80a, SL81, Sch85b, Sch87b, Sch87f, See80, Sti85, Atw84, DT87b, Efr87c, FS87, Gui83, Khu81, LL87b, LT88, Mal82, Mey87a, Mor86, Nyc88, PFV89, Ric82, SI85, Uus83, Uus85]. Interview [SS84]. interviewer [BS85a]. Interviewing [WTB80]. Introduction [Bun89, Edd88, Sed86, She84a, SK84, Tay84]. Introductory [Cob87b]. Invariant [Fra81a, Tak81, HV86, Moo81b]. Inverse [BK84b, Kor80, FB83, IS83]. Inverting [Tri84b]. Investigating [HS89a]. Investigation [KL80, NS82, Wol84]. Involving [Dyk80]. Irreconcilability [BS87b, CB87b, Dic87, Goo87b, Hin87b, Mor87d, Pra87b, Var87, BS87a, Bor87]. Irregular [EPG89]. Isotonic [Bac89]. Issues [MKR89].

J [BK86, Bor87, BC82b, Cha84a, CMS87, Cha84c, Cla89, DD82, Fab94, FH81a, FH81c, HL81b, HG83, JTB81, Kot89, KP82, Krz81, Law82, Lin84a, MW82, Mac87, MPG86, MT82, NS86, Oak85, PT82b, Pra82, Rin85a, Rip84, RC82, Sch82b, Sen84a, Sim87b, WM81, Whi92, YM86, Zah84a, de 83a]. J. [Bor87]. Jack [BH88]. Jackknife [Gon86, Sha89a]. Jackknifed [Fay85b]. Jackknifing [PS82, Sim86]. James [McK84, MB88b]. JASA [Sti88]. Jenkins [WG81]. Jerome [LV88]. Jessen [Zah84a]. Joan [Kru80a]. John [Kas85, Kru80a, Lar84a, Tay84, JV85b, Sti86]. Johnson [Moo84b, O’B84, Moo84b]. Johnston [Tuk85c]. Johnstone [HHM88b, TV85]. Joint [CL86, EPG89, Rog88]. Jointly [MB82]. Joseph [Daw82b]. Jr [Kas85]. Jr. [Cha84a]. Judging [Jen86a]. Judgment [Sha84b]. Judgments [CHM82]. Juries [Kad83]. Justice [Cha84a].

Kadane [Daw82b]. Kahneman [Sha84b]. Kalbfleisch [BH88]. Kale [Moo84a]. Kalman [MS89, Zeh88]. Kaplan [Akr86, CHL82, Efr88]. Kass [DH83b]. Kendall [LQ85]. Kenny [Lin82c]. Kernel [Ano88f, Dav81, HW86, MS88b, SW83, Sta89c, Bie83b, Har84, Jon89, Mül87, Pad86, Sta89b]. kernel-type [Pad86]. Kessler [Pro84]. Keuls [BG81]. Kinetics [MHMEH82]. Klaus [She84a]. knockout [Isr81]. knowing [Bha81]. Known [Sto83a, FE88, vd87]. Kolmogorov [CSS81, CSS86, Gle85b, Gui88]. Kolmogorov-type [Gui88]. Kong [Gro86]. Koopmans [SK84]. Kotz [Moo84b]. Krasker [Rup85]. Krewski [DH83b]. Kriging [Cre86]. Krippendorff [She84a]. Kruskal [GR81, Wei81a]. Kruskal-Wallis [Wei81a].

L [HL81b, MB88b, Moo84b, Mye84]. labeling [HIT86, Sim87b]. Labor [CF87a, NE84]. Labour [Sta86]. Lacking [MB89]. Lag [Lut80]. lagged [Eri81a]. Lags [HW84]. Lambert [SK84]. Laplace [Sti86, TKK89]. Large [HJ80, HS81, Sto83a, Wei80b, BOP88, BE85, CLNL87, EH87, SM84a, VJ88, WM80, WM81, Zel87b]. Large-Sample [HJ80, Wei80b, VJ88, WM80, WM81]. Larger [Blo80, HG81]. largest [Dav82]. Latent [BS86c, fC81, CG84, DM88b, Mis85, Pud82]. Latent-Class [DM88b]. Law [Dem80, Kay80, Pro85, Rog80, Rub80b, Rub80c, Sco80b, Wei80c]. Lawrence [Bow87]. Laws [BB81, Kou80, Sch84e]. Layer [Whi89, Fab94, Whi92]. Layout [HC84]. Layouts [Rin83a, Rin85a]. Lead [WLW89a]. League [Har80b]. Learning [ST82a, Whi89, Fab94, Whi92]. Least [Chr84, Dyk83, EHK84, GWL87, JF80, KP81b, Krä80, Krä86, Rou84, RC80, Sch87g, SH82, Sie83, Smo84, CR82, CGG85, CWR88, Chr87, DWL87, DP89, Hay86, Lea84, PP88, RC82, SS86, SS87b]. Least-absolute-value [Sch87g]. left [Gui88]. left-truncated [Gui88]. Legal [Kad83]. Lehmann [Boo82]. Length [Wei86]. Leo [LV88]. Leonard [LF82b]. Letters [Ano87h, Ano88j, Ano88k, Ano89i, Ano89j]. level [CJ88, Ros89b, TR87]. levels [PH81]. Levenbach [McK84]. Leverage [EHK84]. Liars [Mei86]. liberal [Riv86]. Lieberman [Sch80a]. lieu [APS81]. Life [BG80, HPP86, Kru80a, LH89, FMD82, Ger82, Kar87b, MS85c, O’N84, SS82a, Tak81, Moo84a]. Life- [BG80]. life-testing [MS85c]. lifetime [Suz85]. like [WF82]. Likelihood [Ask80, BG80, Bha85, Edd80, Fer82, Fra84b, FG80, Fuc82, GW88, GS80c, GG80a, GG80b, HB86, KW85, KM83, Mau83, MK80, Mor81, Par80c, Poi80, Pra81, Sam80, Sco80a, Sil80, SRMS81, Smi88, SW86, Tar80, TH87, Voo81, Atw84, BS83b, CR82, Cha84d, Con89a, De 83b, Dyk82, DL88, FD85, Fry84, GH86, LLS87, Lin85, Pra82, SM83a, SM84a, Sch84c, SL87, Spr83, Sta89b, TZ81, WR84, Wij85, Won82a, Won83b, WPSL84]. likelihood-based [Sta89b]. likelihoods [SC85]. Limited [Cox86b, DL86b, DL86c, Fra86b, Gas86, IPRB85, Mor81, Rob86, Sch84b]. Limited-Dependent [Sch84b]. Limiting [Ked80]. Limits [Bly86, MO83, Cla87a, Ger82]. Lindley [Daw82b, DeG82, Dem82, Goo82b, Hil82, Kas82, Lin82a, Sha82a, Sha82c, Sha82b]. Line [Bro82c, Bro83c, Tuk85c, JV85b, TV85]. Linear [Ala80, AS85b, Arn80, AC81, BK82, Bel80, Ber88c, Blo80, Boo86, CS80, CM84a, Dav85b, DC85, Efr84, Esc86, EO85, FT85, Gew82a, GD80, GW80, GP85, HB86, Har80b, KH84, KDW^{+}80, KK84a, Kar80a, Koe86, KP81a, KP81b, Krä80, LO81, LB88, Los85, Mal88b, Mat83, Mcc81b, Mey85, MB88a, MK80, Par82, Pie85, Pie82, PS86, Rav89, Rei85, RC80, Sam85c, Say80, Sch84a, SJ89, Sch88b, Sch84d, Sch87h, See80, Wei82, Wei80b, Wei86, Wer80, WHM85a, Win85, Yan81, Zel82, Alb88b, AP87, BLE88, BS84a, Ber89, CR82, Cha84d, CSS89, DP89, DL88, Efr86a, FW88, Fin83, Flu85, Gew82b, Gew84b, Gol83b, HVY83, HM83, Kan88, Kob85, Kob86, KP87, Kor81, KD87, Kun80, KSC89, LQ85, LP86, LT88, MB88b, NM83b, Nyb86, O’G82]. linear [O’H87a, OYR86, Ole85, Pol84a, Rei84, RC81c, RC82, SH81, TT86, Uus83, WHM85b, WM80, WM81, Wol87a, Pro84]. Linear-Model [Har80b]. linearly [Moo81b]. Lines [Sch84d, Tuk85c, Sch87g, SB81, TV85]. Linkage [Jar89, Tai89, Har81a]. List [Ano80k, Ano80l, Ano80m, Ano80n, Ano81f, Ano81g, Ano81h, Ano81i, Ano82f, Ano82g, Ano82h, Ano82i, Ano83i, Ano83j, Ano83k, Ano83l, Ano84j, Ano84k, Ano84l, Ano84m, Ano85k, Ano85l, Ano85m, Ano85n, Ano86i, Ano86j, Ano86k, Ano86l, Ano87i, Ano87j, Ano87k, Ano88l, Ano88m, Ano88n, Ano89k, Ano89l, Ano89m, Ano89n]. Lists [WMH85]. Livestock [Kuo89]. Local [Ano88f, Law88, McC89b, Nef85, Sta89c, TH87, HB88, Mül87, Sam85a]. Locally [BM83, GV86]. Location [AG82b, AG82a, Esc86, GL88, UH80, Boo81a, CC83, Die82, MH82, Mor86, Swe84, TB86, Tay85, WP85]. Location- [AG82b, AG82a]. location-scale [CC83, Swe84]. Log [EO85, HB86, Koe86, LO81, O’B88a, Pra81, SJ89, Wah80, DL88, FW88, FHO87, Kor81, Ole85, Pra82]. Log-Linear [EO85, HB86, Koe86, LO81, SJ89, DL88, FW88, Kor81, Ole85]. Log-Rank [O’B88a, FHO87]. Logistic [AM84, Efr88, FG84, FL89, Gon86, Has84, Jen86a, LPS84a, LPS84b, O’N80, Rub84b, Stu88, Tay88, Tri84a, WM85, CD89a, HMP87, Jen86b, Len88, LZ89, Whi81b, Whi81c]. Logit [HD80, MRB86]. Lognormal [CW80, HSG80, HSG82, Ruk86, SS85]. Long [Ben83a, Lax85, Ben83b, Riv86]. Long-Tailed [Ben83a, Lax85, Ben83b, Riv86]. Longitudinal [CHK89, EPG89, HB83, WLW89a, CR89]. look [Sti88]. Loss [Shi80, AP87, Zel86a]. Lottery [SC89]. Low [Sch86d, Sie83]. Low-Dose [Sch86d]. Lower [BG88, Del89, Kim86, WF82].

M [DH83b, Par84]. M. [God82c]. Macaque [DG85, SSdM85a, SSdM85b]. macrofunctions [Sto82]. Maddala [Sch84b]. Magazine [Dan89]. Magnitude [SB89, AK86]. Magnitudes [GAF83]. Mail [SS84]. Main [MS80]. making [WF82]. Malaysia [MRB86]. Malliaris [Bro84]. Mallows [dJdWW88, MB88b]. Mallows-type [dJdWW88]. Man [Sch80a]. Manpower [HH89a, HH89b, Hol89, Mof89]. Mantel [WW87b]. Manufacturing [Orr89]. Many [CM84c, Ege81, Jac81, KH81a, KH81b, Wai81, BF83]. Maps [CHM82, MWS^{+}89]. Marginal [Cla87b, Cla89, LHT89, Coh81, CGM83, GA82, KR82, TK86, WLW89b]. Marginalization [Mcc81b]. Marginally [CT80]. Marginals [JT81]. Marital [CL86]. Mark [Oak85]. Market [AG82b, AG82a, DFG87, Hsu82a]. marketing [Eri81a]. Markov [BS86c, Dan89, FS86, HM85a, KL85, Kar85b, KK84b, KP81a, KP81b, Ped80, Pic87, Sch83d, Yak85]. Markov-dependent [Sch83d]. Markovian [Krz87]. Markowitz [JK80]. Marriages [CL86]. martingale [Gil84]. Masking [BK85a, BK86]. Massachusetts [LWZ86a, LWZ86b, Mac86b, Pre86b, Rog86a, SR86, Whi86b]. Matched [WG88, Tam85]. Matched-Pairs [WG88]. Matching [Jar89, Ros89a, Wal84b]. Maternal [SSR^{+}88]. Mathematical [Cha84c]. Matrices [DGK81, LC85a, LC85b, Rea81, Kar81]. Matrix [BK84b, JF80, McC82, Boi86, CLNL87, Mat83]. Matter [Ano80a, Ano80b, Ano80c, Ano80d, Ano80g, Ano80h, Ano80i, Ano80j, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano84a, Ano84b, Ano84c, Ano84d, Ano84f, Ano84g, Ano84h, Ano84i, Ano85a, Ano85b, Ano85c, Ano85d, Ano85g, Ano85h, Ano85i, Ano85j, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano87a, Ano87b, Ano87c, Ano87e, Ano87f, Ano87g, Ano88a, Ano88b, Ano88c, Ano88g, Ano88h, Ano88i, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h]. Maurice [Sha84a]. max [MZ89]. Maxima [Nef85]. Maximal [AG81]. Maximin [Gas85]. Maximum [Bha85, BS83b, Dyk82, FD85, Fer82, FMT89, FG80, Fry84, Fuc82, GW88, HB86, KW85, LLS87, Mau83, MK80, Mor81, Sam80, SM83a, Smi88, SC89, Atw84, CR82, Cha84d, DL88, GH86, Hay86, Lin85, SM84a, Sch84c, SL87, Spr83, TZ81, Wij85, Won82a, WPSL84]. Mean [AN81, Cha81, GB80, Law81, Law82, LG86, Lit83a, Nan80, Rao80b, Rao81, Rei85, Shi80, Tar86, Tay86, Wan80a, Bha81, CH83, HV86, Kar87b, Sam85a, SW86, Tak81, Uus85]. Means [AG81, Cre82, FR81a, Hay89, LS83, Sch86d, BD88, BB89a, Ber89, BA83, Coh81, DB83, GL87, HG81, HT81, Mal82, SI85, Tsu86, YZ88, dJdWW88]. measure [ELR87, SS85]. Measurement [Van80a, Gew82b, Gew82a, Goo82a, Jus86, Kes85, MT80, Par82, PT82a, Pie82, Pre86a, Sel86, Sha81a, Sug82, Wei82, Zel82, PT82b, Rei82, Sch86a, WK88]. measurements [Con89a, HW86]. Measures [Agr81, Bec81, CHK89, FMT89, Gew84b, Joe89, LB88, MB89, Tay85, Abr88, Goe83, KR88, LLS87]. Median [BC82c, GS89, Rou84, Sie83, Wil80, ZH87, AP87, SS87b]. median-unbiased [AP87]. medians [Gui83]. Medical [BS88b, NRMM80]. Meetings [Hur85a]. Meier [Akr86, CHL82, Efr88]. Memory [Sik85]. Metaphysics [Gly86]. Meteorite [Ask80, Edd80, GG80a, GG80b, Par80c, Poi80, Sco80a, Sil80, Tar80]. Meteorology [MW84]. Method [Ano88d, Arj88, Ask80, Ber81, BGT80, Boo80b, BG88, CMS76, Edd80, FM83b, FM83c, GM84, GG80a, GG80b, HW84, JM83, Kap83, Van83, NP83, Par80c, Poi80, SH86, Sco80a, Sil80, Smi86, Sto83a, Tar80, Tur83, VY80, Wal83, Won83a, WG81, CMS87, Dua83, HS89a, KY86, KP81c, KP82, Loh85, Mar84b, Pud82, Ric82, Sch81b, Sch82b, Sil84, Spl83, Won82b]. Methodological [NE84]. Methodology [CV89, Coh89, Cox80a, Har80b, JK87, Jar89, Sin89a, Smi89c, Tur89b, dGJ89, BD88, She84a]. Methods [Ano88f, AM84, BOP88, Bro84, Chr84, CD80, CT82, CM84b, DLST86, DM89, DJK87, DH83a, ES82b, FM81, FG84, GH83a, Gun80, HB86, Has84, HH89a, HH89b, Hol89, JJ88, Kas83, Kas85, Kre83, Lan86, LH89, LPS84a, LPS84b, Law87, Lin80a, Mar80c, Mof89, Obe80, PR80, Rub84b, SL80, Sch80a, SH82, SC80a, SC80b, Smi83a, SM84c, TL88, Thi80, Tuk85c, Van80d, Vin80, WLW89a, AG83, AG85, Alb88b, CC83, Con89a, DH83b, Hur85b, JV85b, KR88, Kin85, LL81, Lou84, Mül87, NWGP84, O’G82, RW85, Ros89b, SS80b, TV85, WP85, Hil84, How84]. Metropolitan [EW81]. Microdynamic [HW89]. Microfluorometric [MR82]. Microsome [MKR89]. Migration [IPRB85]. min [MZ89]. min-max [MZ89]. Minimax [Bic83, Car83, CW83, Geo86, Hec88, Hub83b, Hub83a, KW83, Mal83, Nai87a, Sam81, Sch87h, vd87, Cas85, Sam85a, Wat87]. minimization [YZ88]. Minimum [Bet89, Boo81a, FG80, GH80, Kor80, LW85, PS80, Rub89, Sim87c, Ste87b, TB86, IS83, WPSL84, de 81, de 83a]. Minimum-Type [FG80]. MINQUE [Kap83, Wan80b, Wes87b]. MINQUE-type [Wes87b]. Miracles [Kru88]. misallocation [CM84a]. Misclassification [AR89]. Mises [Koz80]. Mises-Type [Koz80]. Missing [DR81, Fuc82, GRZ82, HP84, HW84, Lit88, Nor84, Rub87, SH86, CS86, KA86, SS86, Sto86, Wit88]. Misspecification [HMR80]. misspecified [KY85, Whi81a]. Misweighted [Cre82]. Mitofsky [Pot87]. Mixed [Ala80, AC81, Bro83b, KH84, LG86, LB88, Sha81a, ST82b, SM84b, Tsu88, WTB80, SH81, SB81]. Mixed-Effects [LB88]. Mixed-Mode [WTB80]. Mixing [FL89, vd87]. Mixture [Dan89, Lar84a, WPSL84, ZC88]. Mixtures [FR83, IS88, Lar84a, Sch84d, AG83]. MLE [GS80b]. Mode [LG86, WMH85, WTB80]. Model [Ala80, Arj88, Arn80, Att80, BHF88, BS86c, CL85b, CF80, fC81, CF87a, Dal83b, Dan89, DZ89, EW81, Fat83, FG80, Fuc82, GE80, GCC89, GHP87, HMT83a, HMT83b, Har80b, Hec88, HW89, HCGL85, KDW^{+}80, Kar80a, Kea85, KP81a, KP81b, Krä80, LM85a, LG86, Lin89b, Lit83b, Mau83, McC89b, MO83, Pie87, RKC81, Rei85, Ros87c, Roy83, Rub83, Rub85, RC80, Sär84, Sha81a, Smi83b, SM84b, SB83, SS84, TW84a, Tha85, TH80, Tsa86, Tsu88, VSK85a, Wan89, Wei80b, Wei86, WM85, WG81, Zie87, dB85, Alb88b, Bin82, BS83b, BHT88, BA83, CR82, CSS89, CS87, DD88, Dal82, EH87, FW88, FJV89, Fin83, FB83, Fry84, Gil84, GC82, Har84, HM83, HT82, IF82, Kar81, KD87, LP86, LT88, LW89, MJ82, MH82, MSW^{+}83, NM83b, PH86]. model [Raf87, Rei84, RC82, Sch86a, Sch88b, Spl83, TT86, VSK85b, Wat87, Wei84a, WM80, WM81, Wol87a, YD84, dJ89]. Model-Based [Ros87c, BHT88, Wat87]. Model-Dependent [Dal83b, HMT83a, HMT83b, Lit83b, Roy83, Rub83, Sär84, Smi83b]. Modeling [BH83b, BS84b, CD82, CZ85, CC89, For80a, For80b, Kad81, KG84, Lee81, Lit81, Lon81, McD81a, McD81b, MH89, Sto80b, TY85, TB81, Wal81a, Zel80b, Agr83b, Kit87, LLT89, Tsa89b, WLW89b]. Models [AH81, AC81, AM84, Bac89, BB85a, BK82, BC89, Bec89, Bel80, BS85b, BLML83, BY86, CL86, Che83, CR89, Dav85b, DM88b, DL89, DRT81, DC85, DR81, Efr84, EO85, Fay86, FG84, FT85, GW88, GS80c, GD81, Goo81, GP85, HB86, HMR80, Has84, HT87a, HM85a, KH84, KS89, KK84b, KE86, Koe86, LPS84a, LPS84b, LL89, Lar84a, LM85b, LR87, LB88, Lit82, Los85, MSV86, McD80, MS89, Mey85, Mor87b, Nel85, Oak89, Oga88, PC84, PS86, Pro84, Rip84, Rub84b, Ruk86, SD85a, ST82a, SD88, Sik85, Smo84, Sol84, Stu88, TW80, Tsa84, Tsa87, Var88, WHM85a, WH86, WY89, Whi89, Win85, Wol86, Won87, Abr88, ACK87, BK85a, BT81, CR84b, Cla87b, Cla89, Clo82, CM86, DP89, DL88, EH87, Fab94, FV88, GH88, Goe83, Hwa86, Khu81, KR86]. models [KP87, KA86, Kor81, KY85, KSC89, LZ89, Man88, Mat83, MS88a, Nef82, O’H87a, OYR86, Pol84a, Rei82, SD85b, SH81, Sha86b, Spi82, Sta89b, Sto86, Swe84, TA89, TT84, WHM85b, Whi81a, Whi92, Zeh88, Lun84, Sch84a]. Moderate [SSR^{+}88, Dav82, GR81]. Modes [FG80]. Modest [Rub87]. Modification [GH83b]. modifications [Gol83b]. Modified [Lut80, Sha86a, HPE81, SW82, Won83b]. Modulus [FM82]. moment [CKC83, HHO88]. Moments [DT80, SRMS81, Mar81b, TK86]. Monitoring [Ber84c, WH86]. Monotone [DH80, DK81, DR83, RA89a]. Monotonicity [GK83]. Monte [JTB81, Hur85a, JTB80, JV85a, Lat81, LC85a, LC85b, NJ88, NS82, Raj80, SF81a, SF81b, Sha89b, Tuk85c, TV85]. Monthly [CD80, CD82]. Months [And84a]. Morris [Tay84]. Mortality [Din86, MSV86, MWS^{+}89, Mau83, PH84, Tsu88]. Most [BM83, Fra81a]. mover [Fry84]. mover-stayer [Fry84]. Moving [Osb82, Spl83]. MR [BK86, Bor87, BC82b, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, MW82, Mac87, MT82, NS86, Pra82, Rin85a, RC82, Sch82b, WM81, Whi92, YM86, de 83a]. much [Ruk87]. Muirhead [Sen84a]. Multi [Pro85, WMH85]. Multi-Mode [WMH85]. Multi-Stage [Pro85]. multidimensional [CG84, Yan84]. multifactor [JK89]. Multilevel [WM85]. Multimodal [IS88, CKC83]. Multinomial [CF87a, DKMO89, FK80, Gre85a, Hab82, Lee80, LNPS81, Sch85c, SRMS81, AM81b, Ala84, CD89a, FS87, Hoe84, iCL87, Zel87b]. Multinomials [KL80, Sim85d, CG82, Sim86]. Multinormal [Wan80a]. Multiparameter [GV86]. multiperiod [HHO88]. Multiple [Agn85, Arn80, Ban86b, BG81, Bof85, BCF^{+}81, BF85a, BC82d, BB80, BB89b, BK85b, fC81, Dic83b, DD83, Dun80a, Dun80b, EH83, EH84, FMW85, Fou80, FK85, Gew82a, Gro86, HAL82, HAJ83, HWH82, HE83, LMR89, Par82, PG83, Pie82, PV85, Rin83b, RS86, Sha81b, Sha86a, Sha81c, TGG89, TB81, Wei82, WY89, Zel82, Ali89, BB88, BF85b, CS85, CSS89, Fli84, Gew82b, HS89a, IS87, Mui85, MRW87, Nai87a, Nai87b, Tam81, Uus85, Wal86b, Wit88, Wol87a]. multiple-design [CSS89]. Multiplicative [BLML83, EHK84, FM82, Hwa86, MJ82]. multiplicity [CSC86]. Multistage [Fel80, Fin84, FV88, MS85a]. Multistate [Rog86b]. Multitude [TW81]. Multivariate [Ala80, Bec89, BS88b, DK81, Dyk80, Edd81, EW81, FR81a, FR81b, FR81c, Gen81, Hos80, Joe89, Lit88, Mal81, MFS84, NJ88, Nel85, O’B84, Pre81a, Rei82, Rei85, Ros89b, Sch86b, Sen84a, Shi80, TBLM89, UH80, WY89, BS83b, BS84b, BA83, CH83, CSS89, Dav82, GH88, GJ84, HVY83, JW88, LZ89, MO88, NS83, NS86, Ran89, Rei84, SS85, SW89a, Spu88, SW86, SZJ87, SM81, TB86, VW87, WL84, WLW89b, Wri83]. Multiwave [BS86c]. Multiway [BM89, Clo82]. Mutagenicity [MKR89].

N [Tay84, Cas86b, OPZ81]. Nationa [WTB80]. National [And84a, For80a, For80b, Har80b, Lan86, Sto80b, Zel80b]. Natural [Umb80]. Nature [Daw84, Gew84a, PS84a, PS84b, RR84a]. nearest [Sch86b]. Nearly [YBN85, BHT88, Cha88a, Vos86]. Necessary [Ber81]. need [Isr81]. Negative [BGT80, Mar88, SM83b, SM84b, GA82]. neighbors [Sch86b]. Neoclassical [HW89]. Neonatal [SSR^{+}88]. Nested [Efr84, MB88c, WR84]. nesting [Ros89b]. Network [Whi89, Fab94, MP83, Whi92]. Networks [Was80a]. Neumann [Bar82a]. Newman [BG81]. Newton [LB88]. Neyman [BGT80]. Nielsen [Bue82b]. Nine [Rao81]. no [BK86, Bor87, BC82b, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, MW82, Mac87, MPG86, MT82, NS86, Pra82, Rin85a, RC82, Sch82b, Sim87b, TA89, WM81, Whi92, YM86, de 83a]. Noise [Sam80]. Nominal [Agr81]. Nominal-Ordinal [Agr81]. Nomination [Wil80, BS86b]. Non [DZ89, Efr84, Kit87]. Non-Gaussian [DZ89, Kit87]. Non-Nested [Efr84]. Nonadditive [BM89]. Nonadditivity [Sne82]. nonanticipative [Gut82]. Noncentral [Hen79, Hen80]. Noncentrality [Ole85]. Nonexperimental [HH89a, HH89b, Hol89, Mof89]. Nonhomogeneous [Sne82]. Nonignorable [BL88a, BL88b]. Noninstitutional [LH86]. Noniterative [Rub87]. Nonlinear [Eub86b, Gal80, GW88, GCC89, KD89, Kar85b, Kas85, ST82a, Shi84b, Val85, Vel80, BT81, Cla87a, Cla87b, Cla89, RW85, WA83, Whi81a]. Nonmetric [Rav89]. Nonnegative [GW80, KR86]. Nonnormal [Raj80, TW80, MT81, MT82]. Nonnormality [Wei80b, Dav82, WM80, WM81]. Nonnull [TW80]. Nonorthogonal [GN82, JJ88]. Nonorthogonality [JH84]. Nonparametric [Ano88f, BS88b, CS85, CF87b, DK81, Din86, FL89, Gui83, HP88, KR88, SF81a, SF81b, SW87, Sto89, Suz85, TW84a, TS85, Tha84, TL88, WG83, Yak85, Yan81, BH88, Die82, Dua83, HB88, JSG88, LQ85, Len88, Mül87, Noe87, Rom88, Sch88b, Ste81, TJW88, Yan85]. nonpositive [TKK89]. Nonrandom [BR83b]. Nonrandomly [Nor84]. Nonresponse [BL88a, BL88b, Fay86, Lit82, Sta86, AG85, RS86, Smo82]. nonstandard [SL87]. nonstationarity [Cre88]. Nonstationary [Cre86, BA83, Cha88a, Kit87, TT84]. nontrivial [Pic87]. norm [Ste87b]. Normal [AC81, BHR83, DR85, Dyk80, Fra81a, GS80b, GH80, GD80, Hod87a, KDW^{+}80, Raj80, Rao80b, Sch86d, Shi80, Sto80a, Ber89, CH83, DB83, HG81, Hoo89, Len88, LP86, Mal82, Ruk87, Sin86, SI85, SW86]. normal-theory [Hoo89]. Normality [HJ80, LaR86a, VJ88]. Normalized [Nan80]. Norman [Moo84b]. Note [BH83a, Chr84, CW85, DH80, Eub86b, HG82, Hur85a, Kar80a, Nef85, Smo84, CL85a, Fli84, HG83, Mar81b, Won83b]. Nuisance [TW89]. Null [BS87b, CB87b, Dic87, Goo87b, Hin87b, Mor87d, Pra87b, TW80, Var87, BS87a, Bor87, Del89]. Number [Col87, FM82, JR85, KOS85a, PFV89, Wal86b]. Numbers [Eth82, Cas85]. Numerical [FMT89, BHS81, Kas85]. Nuptiality [CL86]. Nutritional [And84a].

O [Bor87, GS80b]. O. [Bue82b]. oblique [WR84]. Observation [SH86, Zeh88]. Observational [Hol85, Ros84c, RR84b, Ros89a, Ros84b]. Observations [HP84, HW84, HMM85, Sha81a, Wei80a, Wei83a, APS81, CM85, CM86, JG83, NWGP84, Tak81, WG83, Wei84a, WL84]. Observed [SJ89, Wal86b]. obtained [Tay85]. Obtaining [BG88, PT83]. occasion [Sti88]. Occupation [Ped80]. Occurrences [Oga88]. Odds [HAL82, HAJ83, DD88, MPG85, MPG86]. odds-rate [DD88]. Oil [EK82]. Old [HAL82, HAJ83]. Ole [Sha84a]. Olkin [FM83a]. Olkin-Sobel-Tong [FM83a]. OLS [KD87]. One [BS88b, EW81, Gre85a, Kaf82, KK84a, LL86, MO83, RKC81, Sch85c, Wal80a, Won84, CB87a, DM88a, Die82, O’N84, PH86, Ros89b, SW80a, Sin86, Ste81]. One-Factor [EW81]. one-parameter [SW80a]. One-Sample [Kaf82, Die82, O’N84, Ste81]. One-Sided [BS88b, Gre85a, KK84a, LL86, MO83, Sch85c, Wal80a, CB87a]. One-Way [MO83, RKC81, PH86, Sin86]. Only [GD80]. Opinion [KDW^{+}80, Mcc81b, Ord87, Sin88]. opinions [LS86]. Oppression [Ban89]. Optimal [AM81a, BD89, BS85a, BF85a, Bro81b, BK85b, CH88, CMST88, CT80, DD81, DD82, FE88, FK85, GV86, LM85b, Min87, PV85, Ros84a, Ros89a, Sha83, Spe85a, SZ80, BF85b, HRR81, HPE81, LP86, Sin86, Stu82]. Optimality [BHS81, Cas82, Eri82, God82b, Pie85, God82c, Jac89b, Tar87]. Optimization [Kas85, SS82b]. Optimum [HHM88a, Joh88, MS88a, Sco88, HHM88b]. optional [Mar88]. options [Ros84a]. Ord [Rip84]. Order [ACK87, Ao86, Ano88e, BB85a, DP84, HG82, Sch86d, Sol84, WM82, Yan81, Bha81, BHS81, HM88a, HG83, iCL87, Raf87, RW85, Vos86, WR84, Zac86, de 81, de 83a]. Order-restricted [ACK87, WR84]. Ordered [GH80, Goo81, Gro80, BS86a, Clo82, Dyk82, FD85, Gro81b, KR84, Nai87c, Tak81]. Ordering [Fra84b, SW89a]. orders [CJ88]. Ordeshook [Bow87]. Ordinal [Agr81, CHK89, GKR88, Agr83b, CD89a, Sar89]. Ordinary [Arn80, Chr84, Krä80, SH82, Chr87, DP89, PP88]. Oriented [Sch83b]. Orthogonal [BB85a, BK85c, DH86, Vos86]. Orthogonality [Wil88b]. Other [DH83a, Hen79, Hen80, Kas83, Kre83, LP84, Smi83a, DH83b]. Outcome [BL88a, BL88b, LTAP88]. Outcomes [LMR89]. Outer [Mey87a, SL81]. Outgoing [KM86]. Outlier [Cha86b, HIT86, Sim87b]. Outliers [CDP82, DH80, GH87a, Jen86b, Tsa86, BK85a, BK86, Mar88]. Outlying [FK80]. Ove [HL81b]. Overdispersion [DL89]. overparameterized [Sha86b]. Oversmoothed [TS85]. Ozone [RT87, Zeg85].

P [BC82b, McK84, Oak85, Zah84b, SD85a, WY89]. P. [TV85]. Paired [Far83b, Liu83, Wei80a, Wei83a, Alb88a, WP85]. Pairing [JGM83, dJGM83]. Pairs [WG88]. Pairwise [Dun80a, Dun80b, Hay89, HS89c, HVS84, SI85]. Panel [BS86c, Pro84, Sta86, KL85]. Paradox [DeG82, Dem82, God82a, Goo82b, Hil82, Kas82, Lin82a, Sha82a, Sha82b, Gut82, Sha82c, Sha84a].

Parallel [Bro81b, Sch88a]. Parameter [CMM88, CW80, CW85, Dav81, JF80, LR87, MK80, SW80c, SZJ87, Umb80, Wil88b, BD81b, Cha88a, Cla87a, CC83, DM88c, FMD82, Lin85, Lou84, Moo81b, Mor86, Oma81, SM83a, SW80a, Sil84, Zac86]. Parameter-Effects [CW85]. Parameterization [LH86]. Parameterized [Rog86b, TBLM89]. Parameters [HHM88a, Joh88, Ked80, Kou80, LL89, Nyb89, Sco88, ST82b, SY84, TW89, ACK87, HHM88b, iCL87, MS85a, Ole85, Sha85, Spi82, Spr83, TT84]. Parametric [Ber83b, Dem83, Gal80, Hin83, KS89, Leo83, Lin83a, Mor83b, Mor83a, Sti83b, FJV89]. Parent [Ben83a, Ben83b]. Pareto [AP89a]. part [Mat83]. Partial [BJ81, DP84]. Partially [SJ89, Zin80]. Participation [Ash83]. particular [Mui85]. partitioning [APS81]. Partly [Sel86]. Path [Wer80]. Patients [LTAP88]. Patil [Zah84b]. Pattern [DG85, SSdM85a, SSdM85b, KY86]. Patterned [BK84b, HN87]. Patterns [Ban89, Bec81, Fay86]. Paul [Sha84b, Zah84a]. PC [Sin88]. PC-Based [Sin88]. Pearson [Akr88, ELR87, GK83]. Pearson-type [Akr88]. Peirce [MZ80]. Peizer [LP84]. Penalized [Ask80, Edd80, GG80a, GG80b, Par80c, Poi80, Sco80a, Sil80, Tar80]. Penalty [Kad83]. People [WB85]. Percentage [JJS83]. percentiles [TT86]. Perception [CM84b, SH87]. Performance [HIT86, Lax85, Sim87b, VY80, WM82]. performing [MP83]. Period [HW89]. Period-Specific [HW89]. Periodicity [Sie80]. Periodogram [Wah80]. Periods [NP83]. Permanent [Att80]. Permanent-Income [Att80]. Permutation [Tri84b, WG88, PT83, Ros84b]. permutational [MPS88]. Persons [And84a]. Peter [Bow87]. Pharmaceutical [DK81]. Phase [Rao81, Smi89b]. Phases [BNT86]. phi [ELR87]. phi-squared [ELR87]. Philatelic [IS88]. Philip [Cha84c]. piecewise [TZ81]. Pitman [GS89]. placements [OW82]. Planned [Fin84]. Planning [Hol85]. plans [SB81]. Plant [BB85a]. plates [Cha88b]. players [Isr81]. Playing [Cre82]. Plot [Stu87, MB88c]. Plots [DR85, Nai84, SS86]. PLU [RN80]. Plus [EHK84]. Point [BS87b, CB87b, Dic87, Goo87b, Hin87b, Mor87d, Muk84, Oga88, Pra87b, RN80, Var87, BS87a, Bor87, MSW^{+}83, YZ88]. Points [Ege81, Jac81, JJS83, KH81a, KH81b, Nef85, RSW80, Wai81, Sim89]. Poisson [AY82, BEW85, CM85, DL89, HT81, Law87, Nel85]. Polarization [AM81b]. Police [BS88a]. Policy [Sto89]. Polls [Ord87]. polychotomous [BK85c]. Polygons [Sco85]. Polynomial [BC82d, Lut80, PT83, Stu82]. Pooling [BH83a]. Pools [Mcc81b]. Population [AS85a, AG81, Bai85, Cha86b, DT80, EK85a, EK85b, EKT89b, Fay85a, Fel85, GS80b, Gre85a, Han85a, Hau85, HS81, LH86, Lan86, Lit83a, Pas85, Pre85, Rog86b, Rol85, RC85, Sch85c, Sed85, SM84c, Smi86, Smo84, Sto83b, Tuk85b, Wan80a, Wil80, AK86, CSC86, GM86a, KR82, Lin85, Lou84, MS85a, Mey87a, Raf87, RC81c, SS82b, Smo82, Tam88a, Tuk85a, Val85, Wil84, Wri83, EKT89a]. Populations [AG81, Boo86, Rao80b, Sto80a, Ala84, MT81, MT82, SS85, Sin86].

Portfolios [JK80]. Portmanteau [Hos80]. Posed [MMP87]. Positive [BB81, See80, BS84b, GA82, MZ89, Sha85]. positivity [BJM81]. Positron [CL85b, HCGL85, Kea85, LM85a, Rub85, Tha85, VSK85a, VSK85b]. Possibly [Rao80b]. Postcensal [LH86]. Posterior [Dem87, GD80, Hab87, Mor87b, O’H87b, Rub87, TW87a, OB88b, TW87b, TK86, VW87]. Postwar [Sto88]. potential [KY86]. poverty [Sen86]. Power [BEW85, BGT80, Cam84, CR84b, DKMO89, ES82b, ES84, GH83b, MB89, Tay86, WB81, Car82, FW88, Gle85b, Gro81b, Rea84, Tad80, TR87, WW87b]. Powerful [BM83, Fra81a, Ber89]. powers [CG82]. pp. [Kru80a]. Practical [Bai87]. Practice [Bai88, Mar81a, Zah84b]. pragmatic [Sin86]. Pre [DFH89]. Pre-Challenger [DFH89]. Predict [Nef85]. Predicting [LTAP88, TW81, Tom88]. Prediction [BHF88, Car82, DFH89, GAF83, Har85b, Hor88, Lee88, SM83b, Sti85, AP89a, Bha81, BHT88, Efr83, Efr86b, Gui83, KA86, LP86, Raf87, Rei84, Roy86, Sch87g, Val85, Yak85, Yan84, Zel86a]. Predictions [BD89, Har80b, SB83]. Predictive [BR80, GE80, LF82a, Leo82, NS82, JG83, LF82b, Len88]. Predictor [CS80, BK85c, Nai87b]. predictors [FH81a, FH81b, KY85, PH86]. Preliminary [GS80d]. Prepivoting [Ber88b]. Presence [Cah87, DR81, HWW80, Ord87, Tsa86, TW89, Bow87, NM83b, Smo82, Sto86]. preserved [Kor81]. preserving [BLE88, Hoe84]. Prevalence [Din86, LH86]. Prevision [Gol83c]. Prices [AG82b, AG82a]. Priestley [Par84]. Primary [LC85a, LC85b]. Principal [DGK81, Flu84, HS89b, Krz79, LC85a, LC85b, Krz81, SHS86]. Principle [God82a, Sch81b, Sch82b]. Principles [Lin84a]. Prior [Oma85, Sch84a]. priori [HT81]. Priors [Daw85, Eub86b, KG84, KM83, Sch85d, Shi84b, BB89a, OB88b, Oak85]. Priors-State [KG84]. Probabilistic [FMT89, MMP87, Dic83b]. Probabilities [Ao86, HMM85, Lin81b, BE85, CGM83, MPS88, OB88b, Sch83d, VW87, WF82]. Probability [Ano80f, Biy80, Bre81, Van80a, Dal83b, DZ82, FW81, Fra81b, GJ84, GRZ82, Hab81, HMT83a, HMT83b, HL81a, KW88, Lit83a, Lit83b, MT80, MW84, Roy83, Rub83, Sch84e, SF81a, SF81b, Smi83b, Agn85, CF87b, FM83a, HL81b, JTB80, JTB81, KP83, Kim86, Kin85, Nai84, Sti86, VM83, Whi81c]. Probability-Sampling [Dal83b, HMT83a, HMT83b, Lit83b, Roy83, Rub83, Smi83b]. Probable [Moo84b]. Probing [VY80]. Probit [GHP87, MS88b]. Problem [Edd81, Eld81, FP81, FR81b, FR81c, Gen81, HWW80, Kaf82, Mal81, Pre81a, RSW80, Sam85c, Tam88b, AA87, AG85, CB87a, Die82, GC82, HPE81, JW88, TJW88]. Problems [Kru81, MMP87, Oma85, Sav85, CCE85, HN87]. Procedure [BJ81, BG81, BHR83, Cox87b, HI88, Kad83, O’N80, Sin88, Tai89, CS86, Cre88, EH87, HPE81]. Procedures [BC82d, BB80, BB89b, FP81, HM88b, HW85, HWH82, MWS^{+}89, Ruk84, Sha86a, WTB80, CS85, HM89, IHC84, MS85c, Ole85, Tam85, Wat87]. Proceedings [Fis86]. Process [BEW85, Has80, KK84b, Law87, McK84, Bha81, BHS81, Hol87, Krä86]. Processes [Ano87d, Ano88e, DP84, Ked80, Krz87, Oga88, Ord87, Osb82, Rip84, Sto88, Was80a, Mac86a, Mac87, Tsa89b]. Processing [Kos85b, SH87]. Product [DT80, Dyk80, Wal80b, BB89a, Sha85]. production [GA82]. Productive [WB85]. Profession [Har80a]. Professional [McK84]. Program [BR83b, Far83b, Liu83]. programming [O’G82]. Programs [Ash83, HH89a, HH89b, Hol89, Mof89]. Progress [NE84]. Projection [FS81, FSS84, Fri87, LC85a, LC85b]. Projection-Pursuit [LC85a, LC85b]. Projections [BG88, LH86, Rog86b, Sto83b, TBLM89, DL88]. Propensity [RR84b, Ros84b]. Properties [AG80, BN82, BG80, Bue82a, Cha80a, Cox82, FH81a, FH81b, GS80a, GHP87, HAL82, HAJ83, HW89, Hin82b, KY85, LW85, McC89a, Raj80, Rei85, RC85, SB83, AK86, Bue82b, FM83a, Law81, Law82, SL87, Sti87, Val87a]. Property [Sch87h]. proportion [WPSL84, vd87]. Proportional [Arj88, AK86, LW89, Wei84a]. Proportions [BM88, DSW83, TG80, WB81, FS87, HY86, Val85]. provided [Wil84]. Psychometric [RA89a]. Public [Ord87]. Publication [BBL89, Cah87, Ord87, Bow87]. Publications [Ano80o, Ano80p, Ano80q, Ano80r, Ano81j, Ano81k, Ano81l, Ano81m, Ano82j, Ano82k, Ano82l, Ano82m, Ano83m, Ano83n, Ano83o, Ano83p, Ano84e, Ano84n, Ano84o, Ano84p, Ano84q, Ano85o, Ano85p, Ano85q, Ano85r, Ano86m, Ano86n, Ano86o, Ano87l, Ano87m, Ano87n, Ano88o, Ano88p, Ano88q, Ano89q, Ano89r, Ano89s, Ano89t, CG86]. Published [BBL89]. Pulses [DZ89]. Purely [Sel86]. Purposive [KJ87]. Pursuit [FS81, FSS84, Fri87, LC85a, LC85b].

q [SD85a]. Quadratic [Shi80, KR86, Ste87b]. Qualitative [Lam82b, Sch84b]. Quality [AH86, Eld81, Spe85a]. Qualms [Sch85b]. quantal [CC83, MS88b, Sch84c]. Quantile [BK82, LW85, RS82, SL81, Mey87a, Pad86, Yan85]. quantiles [OB88b]. Quantitative [Cha84a, CK89, LS87, Pro84]. Quartile [BE82]. Quasi [GW88, Sar89]. Quasi-independence [Sar89]. Quasi-Likelihood [GW88]. quasiunimodal [OB88b]. Question [Ano81e, DS80]. Questionnaire [SS84]. questions [APS81]. Quick [FS87]. quota [Kin85, Kre86]. Quotes [Ano85e].

R [Blo84, Bue82b, Lin84a]. R. [Kru80a]. Raktoe [Mye84]. Ram [Lun84]. Ramsey [Thu85]. Random [Ano88d, CMS76, Col87, Fat83, FM82, HS81, Hor88, JGM83, KH84, KJ87, LL89, Lit88, MO83, MRB86, RKC81, Rei85, SM83b, SH86, Tad80, TW80, TL88, CMS87, CR89, Dyk82, FD85, Khu81, Kot89, KP81c, KP82, Mar84b, MZ89, NM83b, PH86, Pot87, Rei82, Rei84, Rob87, RS86, YD84, dJGM83]. Random-Effect [TW80]. Random-Effects [Fat83, MRB86, Rei82]. Random-Interval [TL88]. Randomization [Bai87, Bas80a, Bas80b, BS88a, Hin80, Kem80, Lan80, Lin80b, Rub80a, Hoo89, Gre81a]. Randomized [Bel80, BM88, DSW83, HI88, SD88, Shi85, SS84, Tam81, CGG85, KR84, O’H87a]. Randomized-Blocks [Shi85]. Randomly [HW84, Wan89, AA87, Alb88a, SM84a]. Randomness [Bar82a, Ali89, HM88a]. Range [Ano88e, Hoe84, SM81]. Ranges [OB88b]. Rank [Bar82a, BM83, Boo86, CHK89, FP81, GH87a, HM88a, HC84, HI88, KW85, Lat81, MS80, MW81, O’B88a, Rin83a, Shi85, Won84, AA87, Alb88a, Boi86, BS86a, FHO87, KR84, MW82, MH82, Rin85a, TA89]. Rank-based [HM88a]. Rank-Censored [KW85]. Rank-One [Won84]. Rank-Score [HC84]. Rank-Sum [O’B88a]. rank-transform [TA89]. Rank-Transformation [HI88]. Ranked [Sto80a, SS88c]. ranked-set [SS88c]. Ranking [BJ81, BD88, FV88]. rankings [Tar87]. ranks [CSS89, HM83, Ros84a]. Raphson [LB88]. rare [CSC86, KP83]. Rate [Gla80, O’N80, DD88, Efr83, Efr86b, Hay86, SM83a, SW83]. Raters [TY85]. Rates [EW81, EPG89, HW89, HB83, LH86, MWS^{+}89, Say80, TBLM89, Tsu88, BM85, BPN87, CH88]. Ratio [Bar82a, Fra84b, Ful81, GS80c, HAL82, HAJ83, MP85, Nan80, Rao81, RC81a, RC81b, Sär81, SW81, SRMS81, Smi81, SM84c, TW80, Tha84, Tho81, Wre81, Kot89, MPG85, MPG86, Rob87, SL87, SW86, WR84]. Rational [Som80]. Ratios [JH84, KM83, Flu85, Loh86, Mal82, Wij85]. Raymond [Zah84a]. Reaching [Ber81]. Really [Ben83a, Smi89b, Ben83b]. Reanalysis [ALF83a, ALF83b, Bec83, Boa83, CS83, Gai83, Gay83, Oak83, TW83]. Rebutted [BC82a]. recapture [CM86, MSW^{+}83]. Received [Ano80o, Ano80p, Ano80q, Ano80r, Ano81j, Ano81k, Ano81l, Ano81m, Ano82j, Ano82k, Ano82l, Ano82m, Ano83m, Ano83n, Ano83o, Ano83p, Ano84e, Ano84n, Ano84o, Ano84p, Ano84q, Ano85o, Ano85p, Ano85q, Ano85r, Ano86m, Ano86n, Ano86o, Ano87l, Ano87m, Ano87n, Ano88o, Ano88p, Ano88q, Ano89q, Ano89r, Ano89s, Ano89t, CG86]. Reciprocal [Kor80]. Recommended [Sto83a]. reconciled [BHT88]. reconciliations [Sen86]. Reconciling [CB87a]. Record [Jar89, Suz85]. Record-Linkage [Jar89]. Records [JK87, SM84c, Smi88, TB85]. recover [CS89]. Recurrence [All85]. Recurrences [Sha81c]. Recurrent [TL88]. recursion [CKC83]. Recursive [Wer80, Abr88, SH81]. redescending [HRR81]. Reducing [RR84b]. reduction [Kan88]. Reference [And84a, BB89a, CGG85]. Refined [GM84]. refinements [Ber88b]. regarding [WR84]. Regional [CC89]. Regions [DG86, FH89, FH90, IS87, LL81]. Regression [Ano88f, AM84, BL88a, BL88b, Ber88c, BR83b, Bic83, BF85a, BK85b, Car83, CS80, Cau80, Che83, CW83, DL89, DW87, DD83, Dyk83, EK82, Esc86, Eub86b, FG84, FL89, FK85, FS81, FH78, FH81c, GH83a, Gal80, GW88, GL83, GD80, Gon86, GS80d, Gun80, HM88b, HM89, HHM88a, Has84, Hil84, Hub83a, Jen86a, Joh88, KK84a, KW85, KE86, Kiv80, KSY85, Kou80, Krä80, KW82, KW83, LPS84a, LPS84b, Law87, Law88, Lea81, Lin80a, Mal83, Mal88b, Mar80c, MB88a, O’N80, Obe80, Oma85, PR80, PN81, PC84, Pie85, Pol84a, PV85, Pre86a, RA89a, Roc89, Rou84, Rub84b, Sch87h, SW81, Sco88, Shi84b, SC80a, SC80b, SB83, Sti85, Tay88, Thi80, Tib88, Tri84a, Tsa84, Tuk85c, Van80d, VY80, Vin80, Wan80b, WLW89b, Wei80b, WM85, Yan81]. Regression [ZB86, Abr88, AP87, Bie83b, Bin82, BF83, BF85b, BP86, CD89a, CGG85, Cas85, Cha84d, Cla87a, Cla87b, Cla89, Efr86a, Efr88, Eri81a, FP84, Gil84, HB88, HHM88b, HS89a, HW86, HMP87, Hod88, Hub83b, IF82, JSG88, Jen86b, JG83, JV85b, Kan88, Krä86, KD87, KSC89, LQ85, LZ89, LT88, MB88b, Mül87, Nai83, Nai87a, Nai87b, OYR86, Oma81, RC81c, Rup85, Sam85a, Sil84, SS86, SS87b, Spo82, Sta89b, SC85, Stu82, Swe84, TT86, TZ81, TJW88, TV85, Uus83, WA83, WM80, Whi81a, WM81, Whi81b, Whi81c, WK88, Wol87a, Yak85, YZ88, dJdWW88]. Regression-Discontinuity [BR83b]. Regression-Simple [Esc86]. Regression-Type [GS80d, Kou80]. Regressions [AS85b, VY80, Kob85, Kob86, PP88]. regressor [But84]. regressors [PP88]. regret [Mar88, Nai87a]. Regularity [DG85, SSdM85a, SSdM85b]. Regularized [Fri89]. Reiss [Cha84a]. rejection [KP81c, KP82, SS80b]. Rejective [Sha86a]. Rejoinder [ALF83b, Bas80b, BS87b, BF85a, Bue82a, CD89c, CB87b, Daw82a, DH83a, DL86c, Efr87b, ES84, EK85b, For80b, FM83c, FR81c, Gew82a, God82b, GG80b, HMT83b, HHM88a, HH89b, Hin84, HL81a, Hol86a, Hub83a, Kay82b, KH81b, LWZ86b, LL87a, LPS84b, LF82a, MS84b, McD81b, MB88a, Mor83a, PT82a, PS84b, RC81b, Rub80c, Sha82b, SSdM85b, SC80b, TW87a, VSK85a, WT82b, Wal80d, WHM85a, Wio81b, BS87a, Bue82b, CB87a, Daw82b, Efr87c, LF82b, PT82b]. Related [AH86, Bha85, Gla80, Hen79, Hen80, JF80, KL89, Nef85, Tuk85c, BD88, JV85b, TV85]. Relation [Van80a, EGRW86, HS89c]. Relations [FMW85, CKC83]. Relationship [WG81, Kun80, Thu85]. Relationships [Kad81, Lee81, Lit81, Lon81, McD81a, McD81b, MK80, NS82, Wal81a]. Relative [BR83a, HI88, Joe89, CD89a, Cha88b, GG87, KD87, Ros84a]. released [Hwa86]. Reliability [Cha81, LS86, MMP86, Moo84a, Sin88, WS89]. reliable [Ger82]. Religious [Ano81e, DS80]. Remaining [Har81b, Kau81, May81, Sch81c, Wio81a, Wio81b]. Remarks [See80]. renewable [Ros84a]. renewal [GM86b]. repair [WS89]. Repeated [Bow87, Cah87, LB88, MB89, Ord87, Sha81a, Tsi82, Abr88, Con89a, HW86, KR88, LLS87, Rei82, SW82]. repeated-measurement [Rei82]. Repeated-Measures [LB88, MB89, KR88]. replacement [AK86]. Reply [HIT87, BH88, Bor87, BF85b, DH83b, Gew82b, God82c, HHM88b, HL81b, Hol86b, Kit87, LL87b, LV88, MB88b, Sha82c, Sim87b, TW87b]. Report [KM86, SST87, SST88, SST89]. Reporting [BD89, CSC86]. Representation [FR81a, MFS84, Rea81]. Representing [Ege81, Jac81, KH81a, KH81b, Wai81]. Reproductive [Mau83]. Required [Wal84b]. Rerandomization [GH83b, GH83a, PG83]. Resampling [RW88, Rub87]. Research [Ans80a, Bai88, Cle87, Kru80b, Wal80c, Wal80d, Tho84c]. Reserve [KD89]. Residents [Smi89c]. Residual [Oga88, Sie83, Jen86b]. Residuals [BC82d, PS86, Ran84]. Resistant [GH87a, Tuk85c, HIT86, JV85b, Sim87b, TV85]. Resources [Har81b, Kau81, May81, Sch81c, Wio81a, Wio81b]. respect [Hsu82b]. Respondent [AH86, SB89]. Response [Bel80, BM88, BS86c, CF87a, DG86, GCC89, GRZ82, HMM85, Jus86, MMP86, Mau83, SD88, Sch86d, SS84, CC83, Man88, O’H87a, Tam81, Whi81c]. Responses [BS88a, Hab82]. Restricted [Bai87, CS80, CT80, Dyk83, Kar87b, Stu80, ACK87, Car82, GH88, HP88, WR84]. restriction [iCL87]. Restrictions [Sch86d, HHO88]. Result [BSB80]. Results [BB85b, DH83a, Kas83, Kiv80, Kre83, Smi83a, Tuk85c, Whi89, BHS81, CHL82, DH83b, Fab94, FP84, Lea84, Tam88a, Tar87, TV85, Whi92]. Retina [DG85, SSdM85a, SSdM85b]. Retirement [Kes85]. retransformation [Dua83]. Retransformed [Tay86]. Retrospective [HMM85, KL89]. Returns [FR83, Hsu82a, SB80b, Zie87]. Reuben [Cha84c]. Reuse [BR80]. Review [Abb83, Abe83, Abe89, Abr81, Abr84, Adk84, Adl84, Agr83a, Agr88, Aic81, Alb85, Ald87, Ald88, Ale80, All87, Alw88, Ame83, Ame85, Amm81, And84a, And83, Ang83, Ans85, Ans80b, Arm82, Arn81, Arn86, Arn87, Arn89a, Arn89b, Art82, Arv86, Asc84, Ash82a, Ash82b, Ass83, Ath83, Ath87, Atw81, Aue86, Aus89, Bab88, Bai82, Bai89, Bal86, Ban81, Ban82, Ban86a, Ban88, Bar86a, Bar86b, Bar89, BN84, Bar82b, Bas83, Bas85a, Bas85b, Bas87a, Bas87b, BK84a, Bea80, Bea83a, Bea83b, Bec80, Bec84, Bel89a, Bel81, Ben82, Ber83a, Ber84b, Ber85a, Ber83c, Ber83d, Ber85b, Ber85c, Ber86, Bha87, Bha82, Bie83a, Bil81, Bil83c, Bil83a, Bil83b, Bil84a, Bil84b, Bil85, Bin85, Bis88, Blo84, Boc89, Bod85, Boe83, Bon87, Boo80a, Boo81b, Bow83, Bow86]. Review [Boz85, Bra81, Bra84a, Bra85a, Bra85b, Bra87, Bra80, Bra84b, Bra88, Bra86, Bra85c, Bre86, Bri83, Bro80a, Bro80b, Bro80c, Bro81a, Bro82a, Bro82b, Bro83a, Bro84, Bro87, Bro80d, Bro83d, Bro88, Bry84, Bub80, Bub83, Bub84, Buc89, Bue81, Bue83a, Bue83b, Bue84, Bus87, Bus84, But85, BH85, Cal87, Cal85, Cam86, Can82, Car80, Car86a, Car86b, Car87, Cas86a, Cas87, Cas88, Cas89, Cha84a, Cha86a, Cha87, Cha89, Cha84b, Cha84c, Cha80b, Cha84e, Cha85a, Cha85c, Cha85b, Cha86c, Che89, Chi85, Chr85, Chr88a, Chr88b, Cla87c, Clo80a, Clo80b, Clo83a, Clo83b, Cob82, Cob84a, Cob84b, CG86, Cob87b, Cob87a, Coh82, Con89b, Coo83, Cor88a, Cor88b, Cor88c, CW89, Cor89a, Cor89b, Cor89c, Cos82a, Cos82b, Cos83, Cov86, Cov83, Cov85, Cow82, Cox80b]. Review [Cox85, Cre84, Dal83a, Dal84, Dar80, Dav84b, Dav84a, Dav84d, Dav84c, Dav84e, Dav85a, DS87, DeG83, Dea89a, Dea89b, Dea89c, Den84, Dev84, Dev80, Dic81, Dic82a, Dic82b, Dic83a, Dic84, Die89, Dom88, Don87, Dra87, Dub84, Duf82, Dun81, Dun85a, Dun85b, Dun87, Dyk88, Dyk89a, Dyk89b, Eas80, Eat82, Eat83, Ecc85, Edd82, Edd83, Edd84, ES81, ES82a, Eli83, Eme81, Emi83, Eri81b, Esp84, Esp87, Eub86a, Ewb85, Fah81, Fai85, Fai86, Far83a, Far88, Far87, Fay80, Fei82, Fei88, Fel86, Fel88, Fen89, Fie87, Fin86, Fin87, Fin88, Fir88, Fis80, Fla86, Fle84, Fli86, Fol85, Fol88, Fom86, Fon85, Fow88, Fra84a, Fra86a, Fra88a, Fra82, Fra88b, Fre83, Fri86, Fu87, Gar84, Gar83b, Gar83a, Gas83, Gat89, Gav89, GJ89, Gel86a, Gel86b]. Review [Geo84, Geo87, GG83, Gho83, Gho88, Gib88, Gif87, Gla88, Gle86, Gle85a, God85, Goe81, GW83a, GW83b, Gol83a, Gol85, Gol86a, Gol88, Goo80a, Goo80b, Goo83, Goo85, Goo87a, Gra84, Gra82, Gre81a, Gre81b, Gre81c, Gre83, Gre86a, Gre84, Gre86b, Gre85b, Gre88, Gre89, Gri80, Gri89, Gro81a, Gro87, Gui85, Gup81, Gup84, Haa88, Hai81, Han81, Han83, Han85b, Han84, Har82, Har85a, Har86, Har89, Har81b, Hau82, Haw89, Hel83, Hen83, Her80, Her85, Het87, Hey82a, Hey82b, Hil89, Hil80, Hil83, Hil84, Hil85, HM85b, Hil86a, Hil86b, Hil86c, Hil87b, Hin82a, Hin87a, Hir81, Hir82, Hir83, Ho84, Hod87b, Hog85, Hol83, Hol85, Hoo87, Hor84, Hor89, Hos86, How84, How86, How89, Hsi80, Hsu80, Hsu85, Hsu89, Iac82a, Iac82b, Isa81, Ive86a, Ive86b, Ive87]. Review [Ive88a, Ive88b, Ive89a, Ive89b, Ive85, Iye89, Ize88, Jac87, Jac85, Jac84, Jac89a, Jar86a, Jar86b, Jay82, Jay81, Jeb82, Joh81, Joh80, Joh84b, Joh85, Joh87c, Joh87a, Joh87b, Joi82, Kad84, Kaf88, Kaf89, Kar85a, Kar80b, Kar87a, Kas85, Kat81, Kat86, Kau81, Kay81, Kay89, Kel86, Kem82, Kem83, Kem88, Ken86, Khu89, Kie86, Kie88, Kim89a, Kim89b, Kin81, Kin83, Kin87, Kin89, Kir84, Kle84, Klu87, Klu88, Kme80, Kme82, Kme83a, Kme83b, Kme84, Kme85a, Kme85b, Kme85c, Kme88a, Kme88b, Kme89, Kno83, Koc89, Koe81, Koe82b, Koe82a, Koh88, Koo88a, Koo88b, Kos85b, Kou82, Kra89, Kri87, Kru80a, Kru85, Ksh88, KBD89, Kue87, Kuk87, Kul84, Kul86, Lac89, Lad84, Lah89, Lam86a, Lam82a, Lam86b, Lam88, Lam89, Lan86, Lan87, Lan85, Lan84]. Review [Lar84a, Lar84b, Lar85, Lar86b, Lar86c, Lar87, Lar89a, Lar89b, Lar80, Las84, Lea83, Led81a, Led81b, Leh81b, Leh82, Leh84, Leh89, Len86, Leo89, Lep85, Leu85, Lev87, Lii80, Lin87, Lin89a, Lin80c, Lin81a, Lin82c, Lin83b, Lin83c, Lin84a, Lin84b, Lin86, Lit86, Loh89, Lor86, LH85, Low82, Low83a, Low83b, Luc82, Lun84, Lur82, Lus85, Mac89, Mad85, Mad89, Maj88, Mal88a, Mal85, Man84, Mar84a, Mar87b, Mas85, Mas88, Mat89, Max88, May81, May83b, May83c, May83a, May84, McC83, McC81a, McC87, McC84, McC88, McD89, McK88, McK84, McL80, McL85, Mee81b, Mee81c, Mee84, Meh82, Men83, Men80, Met84, Met85, Mey86, Mey87b, Mil81a, Mil81b, Mil85, Mil86a, Mil86b, Mil87, Mis87, Mit84, Mit88, Mit86, Mon85, Mon88, Moo80, Moo81a, Moo82a, Moo82b, Moo83a]. Review [Moo83b, Moo84b, Moo84a, Moo87, Moo89, Mor80, Mor87a, Mor84, Mos85, Mos80, Mos86, Mue80, Muk84, Mul83, Mye84, NN86, Nam83, Nam84, Nam85a, Nam85b, Nam86, Nam87a, Nam87b, Nam87c, Nam88, Nat88, Nel89, Ner86, Neu81a, Neu81b, Neu83, New84, New81a, New81b, Noa81, Nob81, Noe80, Nor86, Nor85, Not89, Nyc89, O’B80a, O’B80b, O’B81a, O’B81b, O’B82, O’B83, O’B84, O’F80, O’F84, O’L85, O’R88, O’S87b, Obe86, Ode81, Ord85, Ore83a, Ore83b, Pac84, Pal86, Pan87, Par80a, Par85a, Par85b, Par85c, Par88, Par89, Par80b, Par83, Par84, Par85d, Pas80, Pat87, Pea87, Peb81, Pec80, Ped88, Ped89, Pek80, Pey89, Pfe87, Pic89, Pie81a, Pie81b, Pie86, Pij86, Pla83, Pla80, Pli87, Pli88, Plo87, Pol84b, Pol87, Pol85, Poo80, Por86, Pot83, Pra87a, Pra86, Pre81b]. Review [Pre89a, Pre89b, Pro82, Pro83a, Pro83b, Pro84, Pru88, Pru84, Pru85, Psa86, Pul85, Pur88, Pur89, Put80, Put82, Put85, Qua80, Qui82, Rag89, Ram85a, Ram85b, Ran86, Rao80a, Rao84, Red86, Reg80, Rei80, Rei81, Ren85, fSMR84, Res85, Ric88a, Ric84, Ric88b, Ric85, Rid82, Rin85b, Rip84, Rob89a, Rob89b, Roc85, Rod82a, Rod82b, Rod85, Rod86, Rog83, Roh83a, Roh83b, Roh83c, Roh83d, Ros87a, Ros88a, Ros87b, Ros88b, Ros84d, Ros82, Ros84e, Ros84f, Ros89c, Ros89d, Row87, Rus82, Rus89, Rya88, SB80a, Sam85b, Sam88, Sam89, San89, Sau83a, Sau83b, Sau85, Sav80, Sav81, Sav86, Sca87, Sch87a, Sch81a, Sch85a, Sch82a, Sch83a, Sch84a, Sch89a, Sch89b, Sch82c, Sch82d, Sch84b, Sch83c, Sch87c, Sch88c, Sch87d, Sch89c, Sch86c, Sch80b]. Review [Sch87e, Sch85e, Sch81c, Sch88d, Sco82, Sco83, Sea86, Sea82, Sea89, Sed84, Sen80, Sen81a, Sen81b, Sen81c, Sen82, Sen83a, Sen83b, Sen83c, Sen84a, Sen84b, Sen84c, Ser86, Sha84b, Sha84a, She80a, She85, She80b, She82a, She83a, She84a, She81, She82b, She82c, She83b, She84b, She89, Shi84a, Shr84, Shr87, Sil86, Sim80, Sim85b, Sim85c, Sim85a, Sim87a, Sin82, Sin89b, Sin80, Sin89c, Sir86, Ska89, Ska83, Ski87, Sko88, Slo88, Slo87, Sly88, Smi80, Smi84, Smi85b, Smi85a, Smi86, Smi89a, Sne86, Sok81, Sok82, Sok83, Sok84a, Sok84b, Sol85, Sol88, Sol87, Spa88, Spe86, Spi86, Spi83, Spu87a, Spu87b, Sta85, Sta89a, Ste87a, Ste88, Ste82, SK84, Ste89, Sti83a, Sti84, Sti82, Sti83c, Sto85c, Sto85b, Sud83, Sud87a, Sud87b, Sud89, Suk87, Sus86, Sut85, Swa81]. Review [Swa84, Swi83, Swi85, Syl81, Tae81, TCD^{+}89, Tat88, Tav84, Tav85, Tay80a, Tay80b, Tay84, Tay87a, Tay87b, Tel86, Thi89, Tho82a, Tho89, Tho82b, Tho82c, Tho83b, Tho83a, Tho84c, Tho84d, Tho84a, Tho84e, Tho84b, Tho85b, Tho85c, Tho85d, Tho85e, Tho85a, Tho85f, Tho86b, Tho86a, Tho87, Tho85g, Tho83c, Thu82, Thu87, Thu88, Tie82, Tie84, Tie85, Tie81a, Tie81b, Tit85, Tol82, Tor83, Tor82, Tra83, Tsa89a, Tsu85, Tuf85, Tur87, Tur89a, Tve84, Ull88, Utt87, Van86a, Van81, Van80b, Van80c, Van86b, Van86c, Van86d, Var85, Var80, Var82a, Var82b, Var86, Var89, Vau82, Vel81b, Vel81a, Ver89a, Ver89b, VJ83, Vre80, Wai83, Wak85, Wal84a, Wal86a, Wal81b, War80, Was80b, Was85, Was89, Wat80, Wat84a, Wat85a, Wat88, Wat84b]. Review [Wat85b, Web89, Weg80, Weg89, Weh80, Wei81b, Wei83b, Wei84b, Wei85, Wei89, Wel84, Wel86, Wel89, Whi83a, Whi83b, Whi86a, Wic80, Wic81, Wid83a, Wid83b, Wij83, Wij86, Wil85a, Wil88a, Wil85b, Wil86, Win80, Win83, Win84a, Win84b, Win89, Wio81a, Wio81b, Wit84, Wol89, Wol80, Wol82, Wol83, Wol85, Wol87b, Won89, Woo87, Wra83, Yan89, Yan83, Yao87, Yas87, You83, Zac89a, Zac89b, Zah84a, Zah84b, Zah89, Zal82, Zel87a, Zel80a, Zel84, Zel85, Zel86b, Zil80, Zin89, dC81, dL86a]. Reviews [Ano80k, Ano80l, Ano80m, Ano80n, Ano81f, Ano81g, Ano81h, Ano81i, Ano82f, Ano82g, Ano82h, Ano82i, Ano83i, Ano83j, Ano83k, Ano83l, Ano84j, Ano84k, Ano84l, Ano84m, Ano85k, Ano85l, Ano85m, Ano85n, Ano86i, Ano86j, Ano86k, Ano86l, Ano87i, Ano87j, Ano87k, Ano88l, Ano88m, Ano88n, Ano89k, Ano89l, Ano89m, Ano89n, KBD89, Sch88c]. Revisions [Dag82, Mar86]. Revisited [BD81a, BC82a, BC82b]. revival [Rom88]. rho [GR81]. Richard [O’B84]. Richardson [BY88]. Ridge [Cau80, Gib81, Gun80, Lea81, Lin80a, Mar80c, Obe80, PR80, SC80a, SC80b, Thi80, Van80d, Vin80, Cas85, Eri81a, Law81, Law82, Oma81]. Right [HWW80, Wei80a, Gui88, Pad86]. right-censored [Gui88, Pad86]. Rights [Ban89]. Risk [BR83a, CK89, DFH89, Hsu82a, GG87]. Risks [MSV86, GA82]. Robb [Sen84a]. Robert [DH83b]. robin [Won82a, Won84]. Robinson [Kot89]. Robust [Bro82c, Bro83c, BY86, Cha86b, DGK81, EK82, FP81, Gas85, GM84, GL87, Hsu82a, LLT89, Lax85, LC85a, LC85b, Mor86, PS80, RN80, Rin83b, Rot83, UH80, Vel80, WB85, WG88, LW89, MZ89, PMF88, Roy86, Stu82, Tam88a]. robust-type [Stu82]. Robustification [MS89]. Robustness [AC81, Ber82, BHS81, Cas82, Eri82, GS80a, God82b, Lam82b, MS85c, Pie87, TG80, Tsu84, God82c, Tsu86]. Rodent [LMR89]. Role [Ros84c]. Rolph [Tay84]. Ronald [Pro84, HL81b]. root [Dav82]. Roots [DHF84]. Ross [Sch80a]. rotation [Cha88b]. Roulette [Eth82]. Round [Won82a, Won84]. Rounding [Cox87b]. Roy [Dav82]. rule [Efr83, Efr86b]. Rules [BCF^{+}81, Cha80a, CSC86, HIT86, HP88, Sam81, Sim87b]. Run [HM85c, HM85d, Sch83d, CJ88]. Runs [Sha81c].

S [Bow87, Gre81a, Moo84a, Sch84b]. saddlepoint [ER86]. Safe [Cre82]. Sales [EGRW86]. Salmonella [MKR89]. Salmonella/Microsome [MKR89]. Sample [AG81, BBL89, Blo80, BC82c, BR80, Dal83b, DD83, Dun80a, Edd81, Fli81, FR81b, FR81c, Gen81, GS80c, GHP87, HMR80, HWW80, HMT83a, HMT83b, Has80, HAL82, HAJ83, HJ80, Kaf82, Kar80a, Krä80, Lat81, Lax85, LS83, Lit82, Lit83b, MW81, Mal81, Noe87, PN81, Pre81a, Pro85, Raj80, RS81, Roy83, Rub83, SS80a, SS87a, Smi83b, SB83, Wal80a, Wal84b, Wei80b, Whi81b, Whi81c, Yan84, AA87, APS81, BE85, CHL82, CS86, CC83, DD88, Die82, FM83a, GR81, GC82, Gui83, HN87, MW82, MT81, MT82, O’N84, OW82, Rea84, Sch86b, SHS86, SW82, Ste81, SS88c, TR87, TT84, VJ88, WL84, WM80, WM81, Zac86]. sample-size [GC82]. Sampled [EO85]. Samples [Ban86b, Der87, DD83, Efr87a, Fay85b, Fel80, GS80b, HS81, Hil87a, Hor88, LaR86a, LL87a, Lit83a, MB82, Mor87c, Nan80, O’B88a, ST82b, Sto80a, BK86, CM84a, GL87, LL87b, RW85, RS86, SM84a, Tam85, Wei81a]. Sampling [Bet89, Biy80, Cah87, Cas82, Chr84, Dal83b, DD81, DT80, Eri82, God82b, HMT83a, HMT83b, KJ87, Lit83b, Ord87, Roy83, Rub83, Rub87, SA86, SH82, Smi83b, Smi89b, Tar86, Wil80, Zin80, AK86, Bow87, BS86b, Chr87, CSC86, DD82, GM86a, God82c, Kin85, Kot89, Kre86, MS85a, MPS88, Rob87, Roy86, Tad80, Tam88a, TR87, Val87a, Val87b, Wol84, Wri83, WHH88]. Sampling/Importance [Rub87]. Samuel [Moo84b]. Satellite [BHF88]. Scale [AG82b, AG82a, Esc86, Fli81, Lax85, CC83, MZ89, Swe84, WG83, YZ88, Zac86]. scale-change [WG83]. Scale-Shift [AG82b, AG82a]. scales [KR82]. Scaling [Dav81]. Scan [GW84, Nau82, BE85, Gla89]. Scattering [Ask80, Edd80, GG80a, GG80b, Par80c, Poi80, Sco80a, Sil80, Tar80].

Scatterplot [CLNL87, CM84c]. Scatterplots [Hub87, LS89]. Scheffé [Nai83]. Scheffé-type [Nai83]. Schemes [KJ87, Sha85]. Schervish [Oak85]. Schnabel [Kas85]. School [Dem80, Kay80, LH89, Mar80a, Rog80, Rub80b, Rub80c, Sco80b, Wei80c]. Schuss [Bra84b]. Science [Har80a]. Sciences [Moo84b]. Scientific [Hur85a]. Scientist [Kru80a]. Score [Che83, HC84, LM85c, RR84b, ACK87, MH82, Ros84b]. Scores [JV85a]. Scott [HHM88b]. Search [FMT89]. Seasonal [Bec81, Bel89b, CD89b, CD89c, CT82, Cle89, Dag82, DHF84, Gew89, HW85, PM87, Pie89, Wat89, BPN87, HT82, O’G82, Sch81b, Sch82b, Wat87]. Seasonality [Ghy88, KG84]. Seasonally [BW85]. Second [SZ80, RW85]. second-order [RW85]. Section [Sto85a, GA82, Sto82]. Sectional [HM85c, HM85d]. sections [NWGP84]. Security [And84a, SB80b]. Seemingly [Roc89, Bin82]. Selecting [Wil84, Tam85]. Selection [Ber88c, DD81, Fin84, Fuc82, GE80, LS83, Mal88b, MB88a, SO83, Sta89c, Wer80, WABA89, YD84, BD88, CS87, DD82, DH86, EH87, FM83a, FH86, FE88, GC82, HG81, Kim86, MB88b, Ros84a]. selectors [DM88c]. Self [Daw85, Oak85, Sch85d]. Self-Calibrating [Daw85, Sch85d, Oak85]. Sellke [Bor87, Bor87]. Semiparametric [EGRW86, Hec88, OS87a, Wan89]. Sensitive [MMP86, Tam81]. Sensitivity [BB85a, Bel89b, CD89b, CD89c, Cle89, Gew89, Pie89, Wat89, Lea84]. Separate [BNT86, Loh85]. Separating [KW88]. Sequence [DD81, DD82]. Sequences [MK80, Sch83d, Yak85]. Sequential [HG81, HE83, Muk84, SW80b, Sha83, SW89b, TGG89, Wan80a, Wu85, FE88, Fin83, KP83, MS85c, Ros84a, SW80a, Tak81, Tam85]. Sequentially [Sha86a]. Serial [GS80a, WM82, HM88a]. Series [BH83b, BW85, CL86, CDP82, CD80, CD82, DS82, Dic84, DHF84, DZ89, DR81, EW81, Fou80, Gew81, Gew82a, Ghy88, HP84, HW89, Kee82, KG84, LM83, McD80, MFS84, Nef85, NS82, NP83, Par82, PB87, Pie82, Sie80, TW81, TBLM89, TB81, Tsa84, Tsa86, Tsa87, Wec81, Wec82, Wei82, Zeg85, ZH87, Zel82, ZJ80, dB85, Ali89, BPN87, Cha88a, Cre88, DH86, FH81a, FH81b, Gew82b, Gew84b, GA82, HT87b, Hol87, IS87, Kit87, KY86, KY85, LZ89, Nef82, Nyb86, Par84]. Services [NRMM80, And84a]. Set [GL83, MV85, Sto80a, Car82, CG84, SS88c]. Sets [AS85b, BC89, Ber88a, CH83, Kob86, MSW^{+}83]. Several [Ban86b, Kob85, MPG85, MPG86, MRW87]. Seymour [BH88]. Shape [CMM88, SS85]. Shapiro [SS88a]. Shareowner [TW84b]. Sharing [JH84]. Sharp [Net86]. Shelby [HL81b]. Shift [AG82b, AG82a, GH83a, Hsu82a, KP83]. Shirley [Tho84c]. shock [BS83b]. Short [BG88, HM85c, HM85d]. Short-Run [HM85c, HM85d]. Short-Term [BG88]. Should [And84b, BF83, DM88a]. Shrinkage [Geo86, DB83]. Shuttle [DFH89]. Sichel [AY82, SZJ87]. Sided [BS88b, Gre85a, KK84a, LL86, MO83, Sch85c, Wal80a, CB87a, Fli84]. Sigmoidally [Sch84c]. Sign [BLE88, ON89, LQ85, Ran89, Spu88]. Sign-preserving [BLE88]. Signal [GL88, Mar86, WA83]. Signals [Sam80].

Significance [GS80d, Kru80a, TW89, But84, PH81, SW82, Tsi82]. significant [Hay86]. Simple [BCF^{+}81, CF80, Esc86, HS81, LF82a, Leo82, Pie85, Dal82, Kot89, LF82b, MH82, NM83b, PMF88, Rob87, RS86, Sin86, TR87]. simplest [Pic87]. Simplifying [PB87]. Simulated [JJS83, Rot83]. Simulating [Ano88d, CMS76, CMS87]. Simulation [Ano80f, BB85a, Gib81, Mor87b, Rao80b]. Simultaneous [HMR80, Hsu82b, HT81, IS87, KK84a, Khu81, LT88, Mal82, Nai87b, Pie87, Raj80, SW80b, Sha83, Sto83a, Ber88a, BT81, DB83, FS87, HVS84, Kun80, Nai87a, Ric82, SI85, Uus85]. Single [Whi89, Fab94, Har81a, Whi92]. Sinha [Moo84a]. SIR [Rub87]. Size [BG88, Dun80a, GS80c, SA86, Sor89, Wal84b, BE85, GC82, Noe87, NWGP84, Raf87, SS85, Whi81b, Whi81c]. Sizes [Blo80, CHM82, LS83, GR81]. skew [Tay85]. Sleep [SSR^{+}88]. Sleep-State [SSR^{+}88]. Slopes [Tuk85c, TV85]. Slovic [Sha84b]. Small [CHL82, CC83, GS80c, GHP87, Raj80, Rea84, SS80a, Sär84, SH89, SB83, TR87, BOP88, FM83a, GR81, Hab88, Vos86, Whi81c]. Small-Sample [Raj80, SS80a, SB83, CHL82, CC83, Rea84, TR87, FM83a]. Smearing [Dua83]. Smirnov [Wal80a]. Smirnov-Type [Wal80a]. Smith [Zah84b, DH83b, Mar80b, Pro85]. smooth [HS89a, PS82, Yan85]. Smoothing [GM84, HHM88a, Joh88, Sco88, Vel80, Wah80, dJ89, DM88c, HB88, HHM88b, Nyc88, OYR86, Sil84, VW87, WA83]. Smoothness [Eub86b, KG84, Shi84b]. Snell [Lin84a]. Sobel [FM83a]. Social [Ash83, HH89a, HH89b, Hol89, Mar80a, Mof89, Was80a, And84a]. Socialist [Ber84c]. Society [Kru81]. Sociometric [FMW85]. Soft [Sav85]. software [DM88a]. Solution [BG80, MMP87, JW88]. Some [AG80, BN82, BB85b, Bue82b, Bue82a, Clo82, Cox82, Fab94, FM81, Gib81, Gun80, Hab80, HT87a, HAL82, HAJ83, HS81, Hin82b, Hor83, Lat81, Lin80a, Lin81b, LP84, Mar80c, McC89a, Obe80, PR80, Pot87, See80, SC80a, SC80b, Stu82, Tam88a, Tam85, Thi80, Van80d, Vin80, Wal80a, Wei80b, WM80, WM81, Whi89, Whi92, Wol86, FP84, GKR88, HIT86, Nai84, Noe87, Rog88, Rom88, Sen86, Sim87b, Val87a, Wol84]. Sons [Kru80a]. Sources [Bel89b, For80a, For80b, Sto80b, Zel80b, CM86]. Space [KG84, Kit87, SS87b, Sto86, dJ89, DFH89]. space-efficient [SS87b]. space-time [Sto86]. spaces [Moo81b]. Spanish [Ber84c]. Sparse [KL80, Koe86, Sim85d, WB85, Sim86, Zel87b]. Spatial [CC89, DG85, KD87, SSdM85a, SSdM85b, Zeg85, Ste87b, Rip84]. Spearman [GR81, Hen79, Hen80]. Special [BSB80, CGG85, Lee88]. Species [DH83a, Kas83, Kre83, Smi83a, CF87b, DH83b]. Specific [HW89, TBLM89]. Specification [Nef82, Tsa86, Pud82, Thu85]. Specified [JT81, OB88b]. Specifying [Oma85]. Spectral [GW86, MG84, Par84]. Spectrum [CD80, Hur85b]. Sphericity [MB89]. spline [NWGP84, Sil84, WA83]. Splines [MR82, RA89a, WW83, Nyc88, VW87]. split [MB88c]. split-plot [MB88c]. SPRT [Mar81b]. Square [GK83]. Squared [AN81, Cha81, Fay85b, Hab80, LaR86a, LG86, RS81, Rei85, Tar86, Bha81, ELR87, Hab88, HV86, KOS85a, Law81, Law82, iCL87, Nai87c, Ole85, Sam85a]. Squares [Chr84, Dyk83, EHK84, GWL87, JF80, KP81b, Krä80, Rou84, RC80, SH82, Smo84, CR82, CGG85, CWR88, Chr87, DWL87, DP89, Krä86, Lea84, PP88, RC82, SS86, SS87b]. Squeeze [SL80]. Stabilization [Tib88]. Stabilizing [Cas86b, MWS^{+}89]. Stable [Ano88d, BB81, CMS76, FR83, Fra81a, Kou80, SB80b, CMS87]. Stage [Chr84, Pro85, SA86, SH82, Chr87, Roy86, Val87b, WHH88]. Standard [KH84, Wil84]. Standards [Eld81]. Standford [TW83]. Stanford [ALF83a, ALF83b, Bec83, Boa83, CS83, Gai83, Gay83, Oak83]. Stanley [Tho84c, HL81b]. State [KG84, Ped80, SSR^{+}88, Kit87, Zeh88, dJ89]. state-dependent [Zeh88]. state-space [Kit87, dJ89]. States [And84a, PH84]. Stationar [Ped80]. Stationary [Ano87d, Gew81, Ked80, Mac86a, Mac87, TT84]. Statist [BK86, Bor87, BC82b, CMS87, Cla89, DD82, Fab94, FH81a, FH81c, HG83, JTB81, Kot89, KP82, Krz81, Law82, MW82, Mac87, MPG86, MT82, NS86, Pra82, Rin85a, RC82, Sch82b, Sim87b, WM81, Whi92, YM86, de 83a]. Statistic [Ano88e, GS80c, HC84, Hos80, Koz80, LM85c, Sim85d, SS88b, SRMS81, BE85, Gla89, Raf87, SW82, WHH88]. Statistical [And84a, Bai88, Bri82, CL85b, CHM82, Cle87, CZ85, Cox80a, DW83, Fie82, FMW85, FM82, God82a, HCGL85, How84, Jon82, Kad83, Kay82a, Kay82b, Kea85, LM85a, Mar81a, McC89a, Moo84b, O’B84, Oga88, RT87, Rub85, Sch88a, Sen84a, Sor89, Spe85b, Tay84, Tha85, VSK85a, Var88, BJM81, CCE85, Dic83b, LLT89, VSK85b, Ans80a, Kru80b, Wal80c, Wal80d]. Statisticians [Fis86, Mar87a]. Statistics [And84a, BN82, Boo86, Bra82, Bue82a, Cha80a, CSS81, CSS86, Cox82, Cox86a, GW84, Gly86, Gra86, Har80a, Hin82b, Hog89, Hol86b, Hol86a, Hor83, HG82, KL80, Kru81, Lin84a, McC86, Nau82, Net86, Rub86, Wal80a, WW83, Yan81, Zah84a, Ber88b, Bue82b, ER86, ELR87, FHO87, Hab88, HG83, Kru88, LL81, MH82, NS83, NS86, PS82, RW85, Rea84, Sim86, Ste81, TR87, TA89, Tsi82, VJ88, Wit88, Zac86, de 81, de 83a, NE84, SK84, Tho84c]. Statistics-Sharp [Net86]. stayer [Fry84]. Steel [Spu88]. Stein [Ber82, Gut82]. step [CMST88, MH82]. step-type [CMST88]. Stephen [Cha84a, HL81b]. stepwise [CS85]. Stochastic [Bra84b, Fra84b, WW87a, Was80a, SW89a, Bro84]. stochastically [Dyk82, FD85]. Stock [AG82b, AG82a, FR83, Hsu82a, Zie87]. stop [DM88a]. Stopping [Sha83, Mar81b, Mar88, Sam81]. Strata [BS89, LS89, LS83]. Strategies [SZ80, Agr83b]. Stratified [Ban86b, Bet89, DD83, Fel80, GL87, RW85, Val87a, Val87b]. Stratospheric [RT87]. Stream [BNT86]. strengths [CG82]. Strong [HG82, HG83]. Stronger [Isr81]. Strongly [Ros84c]. Structural [McD80, MH89, Sto85a, Sha86b]. Structure [Daw84, DG85, Gew84a, Hsu82a, LL89, PB87, PS84a, PS84b, RR84a, SSdM85a, SSdM85b, CG84, GSS89, Rei82]. Structured [Oma85, LV88]. Structures [Gol86b, Lut80, Sto83a, GJ84, Lee88, SB87, SB81]. Student [PMF88]. Studentized [SM81]. Studies [Bel89b, BB89b, CD89b, CD89c, CHK89, Cle89, Dem80, DH83a, Gew89, HB83, Kas83, Kay80, Kre83, NJ88, Pie89, Rog80, Ros84c, RR84b, Ros89a, Rub80b, Rub80c, Sco80b, Smi83a, Wat89, Wei80c, DH83b, JTB80, JTB81, Ros84b, Cha84a, Hol85]. Study [BEW85, DFG87, Ful81, Ghy88, Gib81, JK87, Lat81, Lin81b, Nor84, Raj80, Rao80b, RC81a, RC81b, Sär81, SF81a, SF81b, Smi81, Tho81, Wre81, Zie87, BT81, RC81c]. Studying [DM89]. Sub [Bro83b]. Sub-Balanced [Bro83b]. Subclassification [RR84b]. Subject [Ano80e, BL88a, BL88b, GAF83, Sto88, Wei80a, WS89]. Subjected [HB86]. Subjective [DZ82]. Subsample [SA86]. subsamples [Ste81]. Subset [Van80a]. subspaces [BS84a]. Substitution [MH89]. Successive [SH86]. sufficiency [VM83]. Sufficient [Ber81]. Sugihara [PT82b]. Sum [O’B88a, Kre86]. sum-quota [Kre86]. superpopulation [IF82]. Supplement [And84a]. Suppression [Cox80a]. Supremum [FHO87]. Surface [DG86, GCC89]. Survey [AH86, BHF88, Cas82, CV89, Coh89, DD83, Eri82, God82b, Gro86, IF82, KE86, LG86, LS87, RW88, Sin89a, Tur89b, WTB80, dGJ89, Agr83b, CSC86, God82c, Sta86]. Surveys [Ban86b, Dal83b, HMT83a, HMT83b, Kin85, Kuo89, Lit82, Lit83b, MMP86, Pro85, RS81, Roy83, Rub83, Smi83b, WMH85, BS85a, CS86]. Survivability [Ber84a, MS84a, MS84b]. Survival [All85, BS85b, Gas85, LO81, MRB86, ME86, Oak89, SJ89, Sin88, BH88, Dyk82, Efr88, FD85, Tsi82]. Swindles [JV85a]. Symmetric [Lax85, McC82]. Symmetrizing [Cam84, ES84]. Symmetry [AKZ82, Koz80, RFPW80, RA89b]. Syndrome [Nor84]. System [JR85, Lun84]. Systematic [Zin80, Wol84]. Systems [Bro81b, Cha81, Ger82, Pud82, WS89].

T [Bor87, Mye84, Zah84b, O’B88a, Wal80b]. Table [Gro80, CGM83, GSS89, KT88, O’N84]. Tables [BC89, Bec89, EHK84, FK80, Fuc82, Gas85, GK83, Hab80, HAL82, HAJ83, Koe86, LH89, RS81, SS80a, Sie83, VJ88, WB85, ACK87, AG83, BOP88, CG84, Coh81, CG82, Dal82, GR81, GH86, Gil86, GH88, GKR88, Gro81b, Hab88, MP83, MPG85, MPG86, Nai87c, PH81, Sar89]. Tail [Lin81b, Sch84e, Som80, BE85]. Tailed [Ben83a, Lax85, Ben83b, Riv86]. Taillie [Zah84b]. Tampa [Jar89]. tau [LQ85]. technique [Pot87]. Techniques [Dem80, Kay80, LO81, Rog80, Rub80b, Rub80c, Sco80b, Wei80c, ZJ80, CLNL87, O’G82, Tam81]. tectonic [Cha88b]. Telephone [SS84, WMH85]. Temporary [Smi89c]. Tendencies [DR83]. Tensor [Tak83]. Teratological [CK89]. Teratology [LMR89]. Term [BG88, WM82]. terms [DH86]. Test [AM81b, AR89, BE82, Bar82a, Bas80a, Bas80b, Ben83a, BGT80, BC82c, BB89b, DH80, DK81, DK80, ELR87, Fay85b, Fra81a, Fra84b, FK80, Gre85a, GS80d, Gro80, Hab80, HD80, HW88, Hin80, JH82, Kem80, Lan80, Lee80, Lin80b, Lit88, LM85c, MS80, RFPW80, Rub80a, SS88a, Sch85c, Sha86a, SS88b, SW80c, WM82, Wei80a, Ben83b, Ber88b, Boo82, CMST88, Dav82, FMD82, Hay86, HVY83, Jac89b, KR84, Kob86, LQ85, LL81, MJ82, MP83, O’N84, Ran89, Sin86, Spu88, Ste81, SM81, WR84, Wei81a, WW87b, Wol87a]. Tested [Ash83, dAJ86]. Testing [AKZ82, Att80, BG80, BS84a, BS87a, BS87b, BM81, Boi86, Bor87, CB87b, CM85, DHF84, Dic87, DR83, Eth82, Goo87b, Hin87b, JGM83, KE86, KS80, Koz80, Lam81, Moo84a, Mor87d, Muk84, Nyb86, Nyb89, Pra87b, SD85a, SS82a, Sie80, SW88, Stu80, Tsa89b, TW89, Var87, Wei84a, dJGM83, CB87a, CGM83, DD88, DM88a, HY86, HPE81, KY86, Loh85, MS85c, SD85b, Tsi82, VJ88]. Tests [AS85b, Ali89, Arn80, BEW85, BM83, BS88b, CSS81, CSS86, Che83, DL89, DW83, DKMO89, DvdM81, Fel80, Fli81, GH83b, GS80a, Gas85, Gew81, GV86, HWW80, HN87, KK84b, Kiv80, Koe86, LaR86a, Lam82b, Lat81, LL86, MW81, MSW^{+}83, MC84, NS82, O’B88a, ON89, PG83, Ran84, RS81, Rin83a, Ros84c, SS87a, Shi85, TG80, Tri84b, UH80, WB81, Wei83a, Wei80b, WG88, WY89, Akr88, AA87, Alb88a, BK85a, BK86, Ber89, BS86a, CS89, CG82, Die82, FW88, Gle85b, Gro81b, Gui88, HM88a, HP88, KOS85a, iCL87, MW82, MS88a, Nai87c, Noe87, NM83b, OW82, Rin85a, Riv86, Rog88, Rom88, Ros84b, Rot83, Sch86b, SL87, SS82a, Sim89, SW82, SW86, Tar87, Thu85, WL84, WM80, WM81, Zel87b, ZC88, Gre81a]. Textbooks [Cob87b]. th [Zac86]. th-order [Zac86]. Their [BN82, Bue82a, Cox82, HHM88a, Hin82b, Joh88, Sco88, Bue82b, Cla87b, Cla89, Efr86a, HHM88b]. Theorem [Kar85b, Leh81a]. theoretical [BHS81, CR84b]. Theory [Bec81, Ber83b, CM84b, CZ85, Dem83, Ghy88, Hin83, Leo83, LC85a, LC85b, Lin83a, Mor83a, RN80, Sco85, Sti83b, Zah84b, CCE85, GM86b, Hoo89, Mor83b, Sha89b, Sha86b, Val85, VJ88, Muk84, Sen84a, Bra84b]. Third [BB85a]. Three [BCF^{+}81, CW80, Hub87, SF81a, SF81b, WTB80, NWGP84, PMF88, RW85, SI85, Uus85]. Three-Decision [BCF^{+}81]. Three-Dimensional [Hub87, NWGP84]. Three-Parameter [CW80]. threshold [Tsa89b]. Time [BH83b, BNT86, CL86, CDP82, CD80, CD82, DS82, Dic84, DHF84, DZ89, DL80, DR81, EW81, Fou80, Gew81, Gew82a, Ghy88, HP84, HW89, Hol87, Jus86, Kee82, KG84, LM83, McD80, MFS84, ME86, NS82, NP83, Nyb89, Par82, Par84, Ped80, PB87, Pie82, Sie80, SS87b, Sto88, TW81, TBLM89, TB81, Tsa84, Tsa86, Tsa87, Wec81, Wec82, Wei82, Zeg85, ZH87, Zel82, ZJ80, dB85, Ali89, BHS81, BPN87, Cha88a, Cre88, FH81a, FH81b, Gew82b, Gew84b, GA82, HT87b, IS87, Kit87, KY86, KY85, LZ89, Mar81b, Nef82, Nyb86, PT83, Sch88b, Sto86, WLW89b, Zac86]. Time- [SS87b]. time-censored [Zac86]. time-series [GA82]. Time-Varying [BNT86]. Times [All85, SH86]. Tolerance [MO83, Atw84, LT88]. Tom [LF82b]. Tomography [CL85b, HCGL85, Kea85, LM85a, Rub85, Tha85, VSK85a, VSK85b]. Tong [FM83a]. Topic [MMP86]. Total [HS81, BJM81, Smo82]. Totals [Kuo89]. tournaments [Isr81]. Toutenburg [Sch84a]. Toxicology [Sch86d]. Trading [Zie87]. Training [HH89a, HH89b, Hol89, Mof89, CM84a]. Transactions [DFG87]. transform [TA89]. Transformation [Cam84, ES84, HI88, Law88, Tay86, CS87, Spi82]. transformation-based [CS87]. Transformations [BD81a, BC82a, BF85a, BK85b, BMR88, ES82b, FK85, HJ80, PV85, RA89b, Tib88, Voo81, BC82b, BF85b, BJM81, Car82, CR84b, Tay85]. Transformed [Bic84, CR84a, DW83, Dok84, Hin84, HC84, Joh84a, Rub84a, HR84]. Transforms [Ger80]. Transfusion [KL89]. Transfusion-Related [KL89]. Transients [Tan83]. transitions [BP86]. transitivity [FH82]. Transplant [ALF83a, ALF83b, Bec83, Boa83, CS83, Gai83, Gay83, Oak83, TW83]. transportation [CCE85]. Travel [Sam85c]. Treatment [BM83, EH83, KF81, LM85b, LR87, Ros84c, SW89b, EH84, Fli84, HP88, Hsu82b, KR88, Tam85]. treatment-assignment [HP88]. Treatments [BR83b, CMST88, Jac89b, MV88, MRW87, Spu88]. treatments-versus-control [Spu88]. Tree [LV88, LS86]. Tree-structured [LV88]. Trees [BS89, Ege81, Jac81, KH81a, KH81b, Wai81]. Trend [BEW85, DK81, KG84, Lee80, Sch84d, YBN85, BPN87, CJ88, HY86, Nyb86, Rot83]. Trend-Free [YBN85, CJ88]. Trends [BC82d, DSW83, RT87, SW88, TG80]. Trial [SW89b]. Trials [BBL89, TGG89, Sch83d]. triangular [Sar89]. Trigonometric [DS82]. Trimmed [RC80, RC82, dJdWW88]. True [LM83]. Truncated [Wan89, Cha84d, Gui88, Moo81b, TJW88]. truncation [DB83, Hod88]. Tukey [Sim87b, HIT87, JV85b, Sim87b]. tumor [MC84, NWGP84]. Turn [PH84]. Turner [Nor84]. Turning [Nef85]. Tversky [Sha84b]. Two [AS85b, AC81, BC89, BM81, Chr84, CMM88, Edd81, EHK84, Fli81, Fou80, FM83b, FM83c, FR81b, FR81c, FK80, GS80a, Gen81, GS80b, Gew81, HWW80, HC84, JM83, KF81, LM85b, Lat81, MS80, Mal81, MB82, Van83, O’B88a, Ped80, Pre81a, RS81, Rao81, Rin83a, SA86, SH82, Sie83, SW82, Smi89b, Sne82, Tom88, Tur83, WB85, Wal83, Wal80b, WL84, Won83a, AA87, Bin82, Cha88b, CJ88, Chr87, DD88, Dal82, Fli84, FB83, Har87, HG81, Kob86, KY86, MV88, NWGP84, OW82, Rin85a, Roy86, SM83a, Sch86b, TA89, Val87b, WR84, WHH88, dAJ86, vd87]. Two-Component [AC81]. two-dimensional [Dal82, NWGP84]. two-equation [Bin82]. Two-Factor [MS80, FB83]. two-level [CJ88]. two-parameter [SM83a]. Two-Phase [Rao81, Smi89b]. Two-Sample [Edd81, Fli81, FR81b, FR81c, Gen81, HWW80, Lat81, Mal81, Pre81a, SW82, WL84, AA87, DD88, OW82, Sch86b]. two-sided [Fli84]. Two-Stage [Chr84, SA86, SH82, Chr87, Roy86, Val87b, WHH88]. Two-State [Ped80]. Two-Treatment [KF81, LM85b]. Two-Variable [CMM88]. Two-Way [BC89, EHK84, FK80, HC84, RS81, Rin83a, Sie83, Sne82, WB85, Rin85a, TA89]. Type [BGT80, Cha80a, Dav81, FG80, GS80d, Kou80, Koz80, MS80, MB82, Wal80a, Akr88, APS81, CMST88, Gle85b, Gui88, Nai87c, Pad86, Stu82, Wit88, dJdWW88, BM81, BM83, Bha85, Bof85, Esc86, LaR86a, SY84, SW80c, Nai83, Tsu84, Tsu86, Wes87b]. Type-II [MB82].

U.S [And84a]. U.S. [Ao86, CL86, Hwa86, LH89, MWS^{+}89, NE84, Sto88]. Ullah [Hil84]. Umbrella [MW81, MW82]. Unbalanced [TW80, JK89, Wit88]. Unbiased [Cox87b, Esc86, Kor80, AP87, BLE88, HV86, Hoe84, IS83, KR86, Kre86, KSC89, ST87, Spo82]. Unbiasedness [GS89, NM83a]. Unbounded [Voo81]. Uncertain [AS85a, Pud82]. Uncertainty [Bel89b, Sha84b, Wat87]. Unconstrained [Kas85]. Undercount [Cre89]. underlying [Riv86, Sam81, Wes87a]. Understanding [Gil84]. Unequal [Biy80, Dun80a, Dun80b, Lit83a, RKC81, KR82, Kob86, Wil84]. Unformulated [Kru81]. Unified [BM83, LR87]. Uniform [Bie83b, Sam81]. Uniformity [DvdM81]. Uniformly [Ber89, IS83, Kor80]. unifying [JSG88]. Unimodality [De 83b, HG82, HG83]. Union [BS86a]. Union-intersection [BS86a]. unions [BS84a]. Unit [DHF84, LTAP88, SM84c, Smi86]. United [And84a, PH84]. Units [HS81]. Univariate [McC89a, Akr88]. Universes [TW80]. Unknown [GS80b, Raf87]. Unrelated [Roc89, Bin82]. Unsolved [Kru81]. untitled [Abb83, Abe83, Abe89, Abr81, Abr84, Adk84, Adl84, Agr83a, Agr88, Aic81, Alb85, Ald87, Ald88, Ale80, All87, Alw88, Ame83, Ame85, Amm81, And83, Ang83, Ano88r, Ans85, Ans80b, Arm82, Arn81, Arn86, Arn87, Arn89a, Arn89b, Art82, Arv86, Asc84, Ash82a, Ash82b, Ass83, Ath83, Ath87, Atw81, Aue86, Aus89, Bab88, Bai82, Bai89, Bal86, Ban81, Ban82, Ban86a, Ban88, Bar86a, Bar86b, Bar89, BN84, Bar82b, Bas83, Bas85a, Bas85b, Bas87a, Bas87b, BK84a, Bea80, Bea83a, Bea83b, Bec80, Bec84, Bel89a, Bel81, Ben82, Ber83a, Ber84b, Ber85a, Ber83c, Ber83d, Ber85b, Ber85c, Ber86, Bha87, Bha82, Bie83a, Bil81, Bil83c, Bil83a, Bil83b, Bil84a, Bil84b, Bil85, Bin85, Bis88, Boc89, Bod85, Boe83, Bon87, Boo80a, Boo81b, Bow83, Bow86, Boz85]. untitled [Bra81, Bra84a, Bra85a, Bra85b, Bra87, Bra80, Bra88, Bra86, Bra85c, Bre86, Bri83, Bro80a, Bro80b, Bro80c, Bro81a, Bro82a, Bro82b, Bro83a, Bro87, Bro80d, Bro83d, Bro88, Bry84, Bub80, Bub83, Bub84, Buc89, Bue81, Bue83a, Bue83b, Bue84, Bus87, Bus84, But85, BH85, Cal87, Cal85, Cam86, Can82, Car80, Car86a, Car86b, Car87, Cas86a, Cas87, Cas88, Cas89, Cha86a, Cha87, Cha89, Cha84b, Cha80b, Cha84e, Cha85a, Cha85c, Cha85b, Cha86c, Che89, Chi85, Chr85, Chr88a, Chr88b, Cla87c, Clo80a, Clo80b, Clo83a, Clo83b, Cob82, Cob84a, Cob84b, Cob87a, Coh82, Con89b, Coo83, Cor88a, Cor88b, Cor88c, CW89, Cor89a, Cor89b, Cor89c, Cos82a, Cos82b, Cos83, Cov86, Cov83, Cov85, Cow82, Cox80b, Cox85, Cre84, Dal83a, Dal84, Dar80, Dav84b, Dav84a]. untitled [Dav84d, Dav84c, Dav84e, Dav85a, DS87, DeG83, Dea89a, Dea89b, Dea89c, Den84, Dev84, Dev80, Dic81, Dic82a, Dic82b, Dic83a, Die89, Dom88, Don87, Dra87, Dub84, Duf82, Dun81, Dun85a, Dun85b, Dun87, Dyk88, Dyk89a, Dyk89b, Eas80, Eat82, Eat83, Ecc85, Edd82, Edd83, Edd84, ES81, ES82a, Eli83, Eme81, Emi83, Eri81b, Esp84, Esp87, Eub86a, Ewb85, Fah81, Fai85, Fai86, Far83a, Far88, Far87, Fay80, Fei82, Fei88, Fel86, Fel88, Fen89, Fie87, Fin86, Fin87, Fin88, Fir88, Fis80, Fla86, Fle84, Fli86, Fol85, Fol88, Fom86, Fon85, Fow88, Fra84a, Fra86a, Fra88a, Fra82, Fra88b, Fre83, Fri86, Fu87, Gar84, Gar83b, Gar83a, Gas83, Gat89, Gav89, GJ89, Gel86a, Gel86b, Geo84, Geo87, GG83, Gho83, Gho88, Gib88, Gif87, Gla88]. untitled [Gle86, Gle85a, God85, Goe81, GW83a, GW83b, Gol83a, Gol85, Gol86a, Gol88, Goo80a, Goo80b, Goo83, Goo85, Goo87a, Gra84, Gra82, Gre81b, Gre81c, Gre83, Gre86a, Gre84, Gre86b, Gre85b, Gre88, Gre89, Gri80, Gri89, Gro81a, Gro87, Gui85, Gup81, Gup84, Haa88, Hai81, Han81, Han83, Han85b, Han84, Har82, Har85a, Har86, Har89, Hau82, Haw89, Hel83, Hen83, Her80, Her85, Het87, Hey82a, Hey82b, Hil89, Hil80, Hil83, Hil85, HM85b, Hil86a, Hil86b, Hil86c, Hil87b, Hin82a, Hin87a, Hir81, Hir82, Hir83, Ho84, Hod87b, Hog85, Hol83, Hoo87, Hor84, Hor89, Hos86, How86, How89, Hsi80, Hsu80, Hsu85, Hsu89, Iac82a, Iac82b, Isa81, Ive86a, Ive86b, Ive87, Ive88a, Ive88b, Ive89a, Ive89b, Ive85, Iye89, Ize88, Jac87, Jac85, Jac84, Jac89a, Jar86a, Jar86b]. untitled [Jay82, Jay81, Jeb82, Joh81, Joh80, Joh84b, Joh85, Joh87c, Joh87a, Joh87b, Joi82, Kad84, Kaf88, Kaf89, Kar85a, Kar80b, Kar87a, Kat81, Kat86, Kay81, Kay89, Kel86, Kem82, Kem83, Kem88, Ken86, Khu89, Kie86, Kie88, Kim89a, Kim89b, Kin81, Kin83, Kin87, Kin89, Kir84, Kle84, Klu87, Klu88, Kme80, Kme82, Kme83a, Kme83b, Kme84, Kme85a, Kme85b, Kme85c, Kme88a, Kme88b, Kme89, Kno83, Koc89, Koe81, Koe82b, Koe82a, Koh88, Koo88a, Koo88b, Kou82, Kra89, Kri87, Kru85, Ksh88, Kue87, Kuk87, Kul84, Kul86, Lac89, Lad84, Lah89, Lam86a, Lam82a, Lam86b, Lam88, Lam89, Lan87, Lan85, Lan84, Lar84b, Lar85, Lar86b, Lar86c, Lar87, Lar89a, Lar89b, Lar80, Las84, Lea83, Led81a, Led81b, Leh81b, Leh82, Leh84, Leh89, Len86, Leo89, Lep85]. untitled [Leu85, Lev87, Lii80, Lin87, Lin89a, Lin80c, Lin81a, Lin83b, Lin83c, Lin84b, Lin86, Lit86, Loh89, Lor86, LH85, Low82, Low83a, Low83b, Luc82, Lur82, Lus85, Mac89, Mad85, Mad89, Maj88, Mal88a, Mal85, Man84, Mar84a, Mar87b, Mas85, Mas88, Mat89, Max88, May83b, May83c, May84, McC83, McC81a, McC87, McC84, McC88, McD89, McK88, McL80, McL85, Mee81b, Mee81c, Mee84, Meh82, Men83, Men80, Met84, Met85, Mey86, Mey87b, Mil81a, Mil81b, Mil85, Mil86a, Mil86b, Mil87, Mis87, Mit84, Mit88, Mit86, Mon85, Mon88, Moo80, Moo81a, Moo82a, Moo82b, Moo83a, Moo83b, Moo87, Moo89, Mor80, Mor87a, Mor84, Mos85, Mos80, Mos86, Mue80, Mul83, NN86, Nam83, Nam84, Nam85a, Nam85b, Nam86, Nam87a, Nam87b, Nam87c, Nam88, Nat88, Nel89, Ner86, Neu81a, Neu81b, Neu83]. untitled [New84, New81a, New81b, Noa81, Nob81, Noe80, Nor86, Nor85, Not89, Nyc89, O’B80a, O’B80b, O’B81a, O’B81b, O’B82, O’B83, O’F80, O’F84, O’L85, O’R88, O’S87b, Obe86, Ode81, Ord85, Ore83a, Ore83b, Pac84, Pal86, Pan87, Par80a, Par85a, Par85b, Par85c, Par88, Par89, Par80b, Par83, Par85d, Pas80, Pat87, Pea87, Peb81, Pec80, Ped88, Ped89, Pek80, Pey89, Pfe87, Pic89, Pie81a, Pie81b, Pie86, Pij86, Pla83, Pla80, Pli87, Pli88, Plo87, Pol84b, Pol87, Pol85, Poo80, Por86, Pot83, Pra87a, Pra86, Pre81b, Pre89a, Pre89b, Pro82, Pro83a, Pro83b, Pru88, Pru84, Pru85, Psa86, Pul85, Pur88, Pur89, Put80, Put82, Put85, Qua80, Qui82, Rag89, Ram85a, Ram85b, Ran86, Rao80a, Rao84, Red86, Reg80, Rei80, Rei81, Ren85, fSMR84, Res85, Ric88a]. untitled [Ric84, Ric88b, Ric85, Rid82, Rin85b, Rob89a, Rob89b, Roc85, Rod82a, Rod82b, Rod85, Rod86, Rog83, Roh83a, Roh83b, Roh83c, Roh83d, Ros87a, Ros88a, Ros87b, Ros88b, Ros84d, Ros82, Ros84e, Ros84f, Ros89c, Ros89d, Row87, Rus82, Rus89, Rya88, SB80a, Sam85b, Sam88, Sam89, San89, Sau83a, Sau83b, Sau85, Sav80, Sav81, Sav86, Sca87, Sch87a, Sch81a, Sch85a, Sch82a, Sch83a, Sch89a, Sch89b, Sch82c, Sch82d, Sch83c, Sch87c, Sch87d, Sch89c, Sch86c, Sch80b, Sch87e, Sch85e, Sch88d, Sco82, Sco83, Sea86, Sea82, Sea89, Sed84, Sen80, Sen81a, Sen81b, Sen81c, Sen82, Sen83a, Sen83b, Sen83c, Sen84b, Sen84c, Ser86, She80a, She85, She80b, She82a, She83a, She81, She82b, She82c, She83b, She84b, She89, Shi84a, Shr84, Shr87, Sil86, Sim80, Sim85b, Sim85c]. untitled [Sim85a, Sim87a, Sin82, Sin89b, Sin80, Sin89c, Sir86, Ska89, Ska83, Ski87, Sko88, Slo88, Slo87, Sly88, Smi80, Smi84, Smi85b, Smi85a, Smi89a, Sne86, Sok81, Sok82, Sok83, Sok84a, Sok84b, Sol85, Sol88, Sol87, Spa88, Spe86, Spi86, Spi83, Spu87a, Spu87b, Sta85, Sta89a, Ste87a, Ste88, Ste82, Ste89, Sti83a, Sti84, Sti82, Sti83c, Sto85c, Sto85b, Sud83, Sud87a, Sud87b, Sud89, Suk87, Sus86, Sut85, Swa81, Swa84, Swi83, Swi85, Syl81, Tae81, TCD^{+}89, Tat88, Tav84, Tav85, Tay80a, Tay80b, Tay87a, Tay87b, Tel86, Thi89, Tho82a, Tho89, Tho82b, Tho82c, Tho83b, Tho83a, Tho84d, Tho84a, Tho84e, Tho84b, Tho85b, Tho85c, Tho85d, Tho85e, Tho85a, Tho85f, Tho86b, Tho86a, Tho87, Tho85g, Tho83c, Thu82, Thu87, Thu88, Tie82, Tie84]. untitled [Tie85, Tie81a, Tie81b, Tit85, Tol82, Tor83, Tor82, Tra83, Tsa89a, Tsu85, Tuf85, Tur87, Tur89a, Tve84, Ull88, Utt87, Van86a, Van81, Van80b, Van80c, Van86b, Van86c, Van86d, Var85, Var80, Var82a, Var82b, Var86, Var89, Vau82, Vel81b, Vel81a, Ver89a, Ver89b, VJ83, Vre80, Wai83, Wak85, Wal84a, Wal86a, Wal81b, War80, Was80b, Was85, Was89, Wat80, Wat84a, Wat85a, Wat88, Wat84b, Wat85b, Web89, Weg80, Weg89, Weh80, Wei81b, Wei83b, Wei84b, Wei85, Wei89, Wel84, Wel86, Wel89, Whi83a, Whi83b, Whi86a, Wic80, Wic81, Wid83a, Wid83b, Wij83, Wij86, Wil85a, Wil88a, Wil85b, Wil86, Win80, Win83, Win84a, Win84b, Win89, Wit84, Wol89, Wol80, Wol82, Wol83, Wol85, Wol87b, Won89, Woo87, Wra83, Yan89, Yan83, Yao87, Yas87, You83]. untitled [Zac89a, Zac89b, Zah89, Zal82, Zel87a, Zel80a, Zel84, Zel85, Zel86b, Zil80, Zin89, dC81, dL86a]. Unweighted [Nan80]. Updating [DZ82]. Upper [LNPS81, BE85, WF82]. ups [Suz85]. Use [DP84, Gas85, Gro86, JK87, Jus86, Nef85, Pro85, Sto83a, WMH85, Efr86a, Hab88, Sto82]. Used [KS80, Osb82, Tsi82]. useful [BE85, NS83, NS86, Wij85]. usefulness [Cla87b, Cla89]. User [Sch83b]. User-Oriented [Sch83b]. uses [Dic83b]. Using [BHF88, BC89, Ber81, BBL89, BS89, BM88, CL86, Dem80, DD83, Eri81a, GA82, HW88, HM85c, HM85d, Isa83, JJ88, Kay80, LO81, Law88, LNPS81, Pre86a, Rog80, RR84b, Rub80b, Rub80c, Sco80b, Sta86, Wei80c, AG83, Alb88b, BK89, BK85c, CSC86, FH82, GN82, GM86b, HW86, LLT89, LS86, Lou84, MRW87, Nai87b, Nef82, O’G82, SM83a, Smo82, Tam85, Zel86a, de 81, de 83a]. Utilities [KW88]. Utility [Sha84a].

V [Daw82b]. validated [NWGP84]. Validation [Gon86, MKR89, PC84, Efr83, Hur85b, ST87, Sil84]. Validity [Arn80, Dem80, Kay80, Rog80, Rub80b, Rub80c, Sco80b, Wei80c, Hoo89]. Value [CSS81, CSS86, WY89, Sch87g]. Values [BS87b, BMR88, CB87b, Dic87, DK80, Goo87b, GRZ82, Hin87b, Lit88, Mor87d, Pra87b, SO83, Sha81c, TW89, Var87, BS87a, Bor87, CS86, Lou84, Uus85]. Variability [Tsu88]. Variable [Ber88c, BMR88, CF80, CS80, fC81, CMM88, DM88b, FH86, GRZ82, Mal88b, MB88a, Bin82, But84, Krä86, MB88b, Nai87b, Pud82]. Variables [Ano88d, BL88a, BL88b, Bro82c, Bro83c, CF80, CMS76, CC89, GAF83, HW85, NS82, Sch84b, Agr83b, BT81, BF83, BK85c, CGG85, CMS87, Dyk82, FD85, Hwa86, KP81c, KP82, Mar84b, MZ89, Sch86a, Spl83, TH82, dAJ86]. Variance [Bet89, Biy80, Bro83b, BW85, DC87, DW87, Dun80a, Dun80b, Ful81, GN82, Isa83, JJ88, JV85a, Kap83, Kor80, LL89, RC81a, RC81b, Sär81, See80, Sha81a, Smi81, SM84b, Sne82, Tho81, Tib88, Tuk85c, Wan80b, Wre81, Zin80, BD88, BS85a, Flu85, Gol83b, HRR81, HV86, HS89c, IS83, JK89, Khu81, KR86, Loh86, Mar81b, MS88a, MT81, MT82, RC81c, Roy86, Ruk87, Sha89a, TV85, Val87b, Wes87b, Wes87a, Wil84, Wol84, Won82a, Zeh88]. Variances [BH83a, GW80, Kar80a, Pro85, RKC81, Kob86, Riv86, TKK89]. variant [KP81c, KP82]. Variate [GH80, Voo81, SS80b]. Variates [SL80, Wal80b]. Variation [BH83b, BJM81, GS80b]. Variational [LM85b]. Variety [KF81]. variograms [Ste87b]. Varying [BNT86]. Vector [JF80, Spl83]. Velleman [Tuk85c, TV85]. Version [Bar82a, Kar85b]. versions [FHO87]. Versus [Fra81a, RFPW80, Sär84, SM84c, Fli84, Spu88]. Via [fC81, Ger80, Har80b, Tib88, LV88, Sam85a, Won82a]. Viability [WABA89]. Victory [Ber84c]. view [Ber88b, JG83]. Vinod [Hil84, Kiv80]. Violence [BS88a]. Vision [MMP87]. Vol [Moo84b]. Volume [Ano80s, Ano81n, Ano82n, Ano83q, Ano84r, Ano85s, Ano86p, Ano88s, Ano89u]. Volumes [Ano80e, Har81b, Kau81, May81, Sch81c, Wio81a, Wio81b]. voting [BP86].

W [Blo84, Bro84, HHM88b, JV85b, Mye84, O’B84, Sim87b]. W. [God82c]. Wage [EW81]. Waksberg [Pot87]. Wald [Ber84a, HD80, MS84a, MS84b]. walk [NM83b]. Wallis [Wei81a]. Walsh [SSR^{+}88]. warning [Hab88]. Wasserman [HL81b]. Water [JK87, LWZ86a, LWZ86b, Mac86b, Pre86b, Rog86a, SR86, Whi86b]. Watson [WM82]. Way [BC89, EHK84, FK80, HC84, MO83, RS81, RKC81, Rin83a, Sie83, Sne82, WB85, PH86, Rin85a, Sin86, TA89]. Wearden [Tho84c]. Weather [EGRW86, GH83b]. Weibull [CSS81, CSS86, FG80, SY84]. weighing [GM86b]. weight [PS82]. Weighted [Ano87d, Ano88f, DR85, GL83, Mül87, Nan80, CWR88, Mac86a, Mac87, Tar87]. Weights [DD83, CWR88]. Welch [Rot83]. Well [Daw82a, Kad82, LWZ86a, LWZ86b, Lin82b, Mac86b, Pre86b, Rog86a, SR86, Whi86b, Daw82b]. Well-Calibrated [Daw82a, Kad82, Lin82b, Daw82b]. Welsch [Rup85]. Wetherill [How84]. Wheel [Eth82]. Where [And84b]. Which [Rub86]. White [Fab94, SB81]. Wichern [O’B84]. Width [Pie85, Kan88, Nai83]. Wilcoxon [Cha80a, FHO87, SW82]. Wilcoxon-Type [Cha80a]. Wiley [Kru80a]. Wilk [SS88a]. Will [MZ80]. William [Hol85]. win [Isr81]. Windows [Stu87]. Wins [BCF^{+}81]. without [AK86]. Witnesses [Mei86]. WLS [WM82]. Woburn [LWZ86a, LWZ86b, Mac86b, Pre86b, Rog86a, SR86, Whi86b]. Woollcott [Zah84b]. Wording [Ano81e, DS80]. Work [Ber84a, MS84a, MS84b]. Workweek [Orr89]. World [CV89, Coh89, Lun84, Sin89a, Tur89b, dGJ89].

xiv [Kru80a].

Yates [Biy80]. Year [Bai85, EK85a, EK85b, Fay85a, Fel85, Han85a, Hau85, Pas85, Pre85, Rol85, Sed85, Tuk85b, Tuk85a]. Years [And84a, LH89]. York [Kru80a].

Z [Muk84]. Zeev [Bra84b]. Zellner [Kar81]. Zidek [Tsu84, Tsu86].

References

Albers:1987:CRT

[AA87] Willem Albers and Michael G. Akritas. Combined rank tests for the two-sample problem with randomly censored data. Journal of the American Statistical Association, 82(398):648–655, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

            (electronic). URL http://www.jstor.org/stable/2289478.

Abbott:1983:BRB

[Abb83] Robert D. Abbott. Book review: [untitled]. Journal of the American Statistical Association, 78(381):207–208, March 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

            URL http://www.jstor.org/stable/2287142.

Abeles:1983:BRB

[Abe83] Francine Abeles. Book review: [untitled]. Journal of the American Statistical Association, 78(383):739–740, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

            URL http://www.jstor.org/stable/2288162.

Abelson:1989:BRB

[Abe89] Robert P. Abelson. Book review: [untitled]. Journal of the American Statistical Association, 84(405):345, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

            http://www.jstor.org/stable/2289903.

Abraham:1981:BRB

[Abr81] Bovas Abraham. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1018, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287622.

Abrahams:1984:BRB

[Abr84] Julia Abrahams. Book review: [untitled]. Journal of the American Statistical Association, 79(388):957, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288754.

Abramson:1988:RRH

[Abr88] Ian Abramson. A recursive regression for high-dimensional models, with application to growth curves and repeated measures. Journal of the American Statistical Association, 83(404):1073– 1077, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290138.

Ashikaga:1981:RFL

[AC81] T. Ashikaga and P. C. Chang. Robustness of Fisher’s linear discriminant function under two-component mixed normal models. Journal of the American Statistical Association, 76(375):676–680, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287529.

Agresti:1987:ORS

[ACK87] Alan Agresti, Christy Chuang, and Abbas Kezouh. Orderrestricted score parameters in association models for contingency tables. Journal of the American Statistical Association, 82(398): 619–623, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289473.

Adke:1984:BRB

[Adk84] S. R. Adke. Book review: [untitled]. Journal of the American Statistical Association, 79(388):952, December 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288743.

Adler:1984:BRB

[Adl84] Robert J. Adler. Book review: [untitled]. Journal of the American Statistical Association, 79(388):959, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288756.

Arnold:1980:SPA

[AG80] Barry C. Arnold and Richard A. Groeneveld. Some properties of the arcsine distribution. Journal of the American Statistical Association, 75(369):173–175, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287407.

Arnold:1981:MDB

[AG81] Barry C. Arnold and Richard A. Groeneveld. Maximal deviation between sample and population means in finite populations. Journal of the American Statistical Association, 76(374):443–445, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287847.

Ali:1982:IDH

[AG82a] Mukhtar M. Ali and Carmelo Giaccotto. The identical distribution hypothesis for stock market prices–location- and scale-shift alternatives. Journal of the American Statistical Association, 77 (377):19–28, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287765.

Ali:1982:CID

[AG82b] Mukhtar M. Ali and Carmelo Giacotto. Corrigenda: The identical distribution hypothesis for stock market prices– location- and scale-shift alternatives. Journal of the American Statistical Association, 77(380):955, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287368.

Albert:1983:BEM

[AG83] James H. Albert and Arjun K. Gupta. Bayesian estimation methods for 2 × 2 contingency tables using mixtures of Dirichlet distributions. Journal of the American Statistical Association, 78 (383):708–717, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288141.

Albert:1985:BMB

[AG85] James H. Albert and Arjun K. Gupta. Bayesian methods for binomial data with applications to a nonresponse problem. Journal of the American Statistical Association, 80(389):167–174, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288067.

Agnew:1985:MPA

[Agn85] Carson E. Agnew. Multiple probability assessments by dependent experts. Journal of the American Statistical Association, 80(390): 343–347, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287894.

Agresti:1981:MNO

[Agr81] Alan Agresti. Measures of nominal-ordinal association. Journal of the American Statistical Association, 76(375):524–529, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287505.

Agresti:1983:BRB

[Agr83a] Alan Agresti. Book review: [untitled]. Journal of the American Statistical Association, 78(383):738, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288159.

Agresti:1983:SSM

[Agr83b] Alan Agresti. A survey of strategies for modeling crossclassifications having ordinal variables. Journal of the American Statistical Association, 78(381):184–198, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287128.

Agresti:1988:BRB

[Agr88] Alan Agresti. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1221, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290172.

Anderson:1981:EDM

[AH81] T. W. Anderson and Cheng Hsiao. Estimation of dynamic models with error components. Journal of the American Statistical Association, 76(375):598–606, September 1981. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287517.

Andrews:1986:QSD

[AH86] Frank M. Andrews and A. Regula Herzog. The quality of survey data as related to age of respondent. Journal of the American Statistical Association, 81(394):403–410, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289229.

Aickin:1981:BRB

[Aic81] Mikel Aickin. Book review: [untitled]. Journal of the American Statistical Association, 76(375):742–743, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287544.

Andreatta:1986:EFP

[AK86] G. Andreatta and G. M. Kaufman. Estimation of finite population properties when sampling is without replacement and proportional to magnitude. Journal of the American Statistical Association, 81(395):657–666, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288993.

Akritas:1986:BKM

[Akr86] Michael G. Akritas. Bootstrapping the Kaplan–Meier estimator. Journal of the American Statistical Association, 81(396): 1032–1038, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289079.

Akritas:1988:PTG

[Akr88] Michael G. Akritas. Pearson-type goodness-of-fit tests: the univariate case. Journal of the American Statistical Association, 83 (401):222–230, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288944.

Antille:1982:TS

[AKZ82] A. Antille, G. Kersting, and W. Zucchini. Testing symmetry. Journal of the American Statistical Association, 77(379):639–646, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287727.

Alalouf:1980:MAM

[Ala80] I. S. Alalouf. A multivariate approach to a mixed linear model. Journal of the American Statistical Association, 75(369):194– 200, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287410.

Alam:1984:AGD

[Ala84] Khursheed Alam. Analysis of grouped data from multinomial populations. Journal of the American Statistical Association, 79 (385):132–135, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288347.

Albright:1985:BRB

[Alb85] S. Christian Albright. Book review: [untitled]. Journal of the American Statistical Association, 80(389):238–240, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288085.

Albers:1988:CRT

[Alb88a] Willem Albers. Combined rank tests for randomly censored paired data. Journal of the American Statistical Association, 83(404): 1159–1162, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290149.

Albert:1988:CMU

[Alb88b] James H. Albert. Computational methods using a Bayesian hierarchical generalized linear model. Journal of the American Statistical Association, 83(404):1037–1044, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290133.

Aldous:1987:BRB

[Ald87] David J. Aldous. Book review: [untitled]. Journal of the American Statistical Association, 82(397):343, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289174.

Aldous:1988:BRB

[Ald88] David Aldous. Book review: [untitled]. Journal of the American Statistical Association, 83(402):565–566, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288885.

Aleong:1980:BRB

[Ale80] John Aleong. Book review: [untitled]. Journal of the American Statistical Association, 75(370):474–475, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287490.

Aitkin:1983:RSHa

[ALF83a] Murray Aitkin, Nan Laird, and Brian Francis. A reanalysis of the Stanford heart transplant data. Journal of the American Statistical Association, 78(382):264–274, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288622.

Aitkin:1983:RSHb

[ALF83b] Murray Aitkin, Nan Laird, and Brian Francis. A reanalysis of the Stanford heart transplant data: Rejoinder. Journal of the American Statistical Association, 78(382):291–292, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288630.

Ali:1989:TAR

[Ali89] Mukhtar M. Ali. Tests for autocorrelation and randomness in multiple time series. Journal of the American Statistical Association, 84(406):533–540, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289939.

Allison:1985:SAB

[All85] Paul D. Allison. Survival analysis of backward recurrence times. Journal of the American Statistical Association, 80(390):315–322, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287888.

Alldredge:1987:BRB

[All87] J. Richard Alldredge. Book review: [untitled]. Journal of the American Statistical Association, 82(399):957–958, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288831.

Alwin:1988:BRB

[Alw88] Duane F. Alwin. Book review: [untitled]. Journal of the American Statistical Association, 83(401):276–277, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288966.

Aghevli:1981:OGI

[AM81a] B. B. Aghevli and F. Mehran. Optimal grouping of income distribution data. Journal of the American Statistical Association, 76(373):22–26, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287035.

Alam:1981:PTM

[AM81b] Khursheed Alam and Amitava Mitra. Polarization test for the multinomial distribution. Journal of the American Statistical Association, 76(373):107–109, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287053.

Atkinson:1984:GMA

[AM84] A. C. Atkinson and P. McCullagh. Graphical methods for assessing logistic regression models: Comment. Journal of the American Statistical Association, 79(385):72, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288335.

Amemiya:1983:BRB

[Ame83] Takeshi Amemiya. Book review: [untitled]. Journal of the American Statistical Association, 78(381):200–201, March 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287131.

Amemiya:1985:BRB

[Ame85] Yasuo Amemiya. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1068–1069, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288584.

Ammann:1981:BRB

[Amm81] Larry P. Ammann. Book review: [untitled]. Journal of the American Statistical Association, 76(373):208, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287073.

Ansley:1981:BEF

[AN81] Craig F. Ansley and Paul Newbold. On the bias in estimates of forecast mean squared error. Journal of the American Statistical Association, 76(375):569–578, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287514.

Andrews:1983:BRB

[And83] Margaret Andrews. Book review: [untitled]. Journal of the American Statistical Association, 78(381):202–203, March 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287134.

Andersen:1984:BRB

[And84a] Ronald M. Andersen. Book review: Health: United States, 1982, by National Center for Health Statistics; Hematologial and Nutritional Biochemistry Reference Data for Persons 6 Months–74 Years of Age: United States, 1976–80, by National Center for Health Statistics; Social Security Bulletin: Annual Statistical Supplement, 1981, by U.S. Department of Health and Human Services. Journal of the American Statistical Association, 79(385): 240–241, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288385.

Anderson:1984:GWW

[And84b] R. L. Anderson. Goals: Where are we and where should we be going? Journal of the American Statistical Association, 79(386): 253–258, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288256.

Angle:1983:BRB

[Ang83] John Angle. Book review: [untitled]. Journal of the American Statistical Association, 78(382):499–500, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288673.

Anonymous:1980:BMa

[Ano80a] Anonymous. Back matter. Journal of the American Statistical Association, 75(369):257–259, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287442.

Anonymous:1980:BMb

[Ano80b] Anonymous. Back matter. Journal of the American Statistical Association, 75(370):483–484, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287502.

Anonymous:1980:BMc

[Ano80c] Anonymous. Back matter. Journal of the American Statistical Association, 75(371):766, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287702.

Anonymous:1980:BMd

[Ano80d] Anonymous. Back matter. Journal of the American Statistical Association, 75(372):1047–1050, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287226.

Anonymous:1980:CAS

[Ano80e] Anonymous. Corrigenda: Author and subject index to volumes 51–73 (1956–1978). Journal of the American Statistical Association, 75(369):256, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287441.

Anonymous:1980:CCP

[Ano80f] Anonymous. Corrigenda: Computational probability and simulation. Journal of the American Statistical Association, 75(370):482, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287500.

Anonymous:1980:FMa

[Ano80g] Anonymous. Front matter. Journal of the American Statistical Association, 75(369):??, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287370.

Anonymous:1980:FMb

[Ano80h] Anonymous. Front matter. Journal of the American Statistical Association, 75(370):??, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287443.

Anonymous:1980:FMc

[Ano80i] Anonymous. Front matter. Journal of the American Statistical Association, 75(371):??, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287635.

Anonymous:1980:FMd

[Ano80j] Anonymous. Front matter. Journal of the American Statistical Association, 75(372):??, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287157.

Anonymous:1980:LBRa

[Ano80k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 75(369):238, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287417.

Anonymous:1980:LBRb

[Ano80l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 75(370):469, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287479.

Anonymous:1980:LBRc

[Ano80m] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 75(371):747, September 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287682.

Anonymous:1980:LBRd

[Ano80n] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 75(372):1031, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287200.

Anonymous:1980:PRa

[Ano80o] Anonymous. Publications received. Journal of the American Statistical Association, 75(369):254–255, March 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287440.

Anonymous:1980:PRb

[Ano80p] Anonymous. Publications received. Journal of the American Statistical Association, 75(370):480–481, June 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287499.

Anonymous:1980:PRc

[Ano80q] Anonymous. Publications received. Journal of the American Statistical Association, 75(371):762–764, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287699.

Anonymous:1980:PRd

[Ano80r] Anonymous. Publications received. Journal of the American Statistical Association, 75(372):1045–1046, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287225.

Anonymous:1980:VI

[Ano80s] Anonymous. Volume information. Journal of the American Statistical Association, 75(372):1051–1056, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287156.

Anonymous:1981:BMa

[Ano81a] Anonymous. Back matter. Journal of the American Statistical Association, 76(373):218, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287091.

Anonymous:1981:BMb

[Ano81b] Anonymous. Back matter. Journal of the American Statistical Association, 76(374):??, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287879.

Anonymous:1981:BMc

[Ano81c] Anonymous. Back matter. Journal of the American Statistical Association, 76(375):756, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287563.

Anonymous:1981:BMd

[Ano81d] Anonymous. Back matter. Journal of the American Statistical Association, 76(376):1024–1038, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287634.

Anonymous:1981:CEQ

[Ano81e] Anonymous. Corrigenda: Effects of question wording and context: An experiment with religious indicators. Journal of the American Statistical Association, 76(373):217, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287089.

Anonymous:1981:LBRa

[Ano81f] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 76(373):205, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287069.

Anonymous:1981:LBRb

[Ano81g] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 76(374):491, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287857.

Anonymous:1981:LBRc

[Ano81h] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 76(375):741, September 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287542.

Anonymous:1981:LBRd

[Ano81i] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 76(376):1010, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287604.

Anonymous:1981:PRa

[Ano81j] Anonymous. Publications received. Journal of the American Statistical Association, 76(373):215–216, March 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287088.

Anonymous:1981:PRb

[Ano81k] Anonymous. Publications received. Journal of the American Statistical Association, 76(374):502–503, June 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287878.

Anonymous:1981:PRc

[Ano81l] Anonymous. Publications received. Journal of the American Statistical Association, 76(375):754–755, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287562.

Anonymous:1981:PRd

[Ano81m] Anonymous. Publications received. Journal of the American Statistical Association, 76(376):1020–1021, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287625.

Anonymous:1981:VI

[Ano81n] Anonymous. Volume information. Journal of the American Statistical Association, 76(376):1029–1037, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287564.

Anonymous:1982:BMa

[Ano82a] Anonymous. Back matter. Journal of the American Statistical Association, 77(377):226–230, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287814.

Anonymous:1982:BMb

[Ano82b] Anonymous. Back matter. Journal of the American Statistical Association, 77(378):502, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287293.

Anonymous:1982:BMc

[Ano82c] Anonymous. Back matter. Journal of the American Statistical Association, 77(379):698, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287761.

Anonymous:1982:BMd

[Ano82d] Anonymous. Back matter. Journal of the American Statistical Association, 77(380):956–957, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287369.

Anonymous:1982:FM

[Ano82e] Anonymous. Front matter. Journal of the American Statistical Association, 77(377):??, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287762.

Anonymous:1982:LBRa

[Ano82f] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 77(377):211, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287792.

Anonymous:1982:LBRb

[Ano82g] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 77(378):488, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287272.

Anonymous:1982:LBRc

[Ano82h] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 77(379):683, September 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287737.

Anonymous:1982:LBRd

[Ano82i] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 77(380):936, December 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287332.

Anonymous:1982:PRa

[Ano82j] Anonymous. Publications received. Journal of the American Statistical Association, 77(377):223–225, March 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287813.

Anonymous:1982:PRb

[Ano82k] Anonymous. Publications received. Journal of the American Statistical Association, 77(378):500–501, June 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287292.

Anonymous:1982:PRc

[Ano82l] Anonymous. Publications received. Journal of the American Statistical Association, 77(379):696–697, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287760.

Anonymous:1982:PRd

[Ano82m] Anonymous. Publications received. Journal of the American Statistical Association, 77(380):952–953, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287357.

Anonymous:1982:VI

[Ano82n] Anonymous. Volume information. Journal of the American Statistical Association, 77(380):958–964, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287294.

Anonymous:1983:BMa

[Ano83a] Anonymous. Back matter. Journal of the American Statistical Association, 78(381):??, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287155.

Anonymous:1983:BMb

[Ano83b] Anonymous. Back matter. Journal of the American Statistical Association, 78(382):??, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288702.

Anonymous:1983:BMc

[Ano83c] Anonymous. Back matter. Journal of the American Statistical Association, 78(383):??, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288176.

Anonymous:1983:BMd

[Ano83d] Anonymous. Back matter. Journal of the American Statistical Association, 78(384):1009–1011, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288254.

Anonymous:1983:FMa

[Ano83e] Anonymous. Front matter. Journal of the American Statistical Association, 78(381):??, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287092.

Anonymous:1983:FMb

[Ano83f] Anonymous. Front matter. Journal of the American Statistical Association, 78(382):??, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288615.

Anonymous:1983:FMc

[Ano83g] Anonymous. Front matter. Journal of the American Statistical Association, 78(383):??, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288112.

Anonymous:1983:FMd

[Ano83h] Anonymous. Front matter. Journal of the American Statistical Association, 78(384):??, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288178.

Anonymous:1983:LBRa

[Ano83i] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 78(381):199, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287129.

Anonymous:1983:LBRb

[Ano83j] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 78(382):491, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288665.

Anonymous:1983:LBRc

[Ano83k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 78(383):730, September 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288144.

Anonymous:1983:LBRd

[Ano83l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 78(384):984, December 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288214.

Anonymous:1983:PRa

[Ano83m] Anonymous. Publications received. Journal of the American Statistical Association, 78(381):214–215, March 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287154.

Anonymous:1983:PRb

[Ano83n] Anonymous. Publications received. Journal of the American Statistical Association, 78(382):513–514, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288701.

Anonymous:1983:PRc

[Ano83o] Anonymous. Publications received. Journal of the American Statistical Association, 78(383):745–746, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288175.

Anonymous:1983:PRd

[Ano83p] Anonymous. Publications received. Journal of the American Statistical Association, 78(384):1006–1007, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288248.

Anonymous:1983:VI

[Ano83q] Anonymous. Volume information. Journal of the American Statistical Association, 78(384):1012–1018, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288177.

Anonymous:1984:BMa

[Ano84a] Anonymous. Back matter. Journal of the American Statistical Association, 79(385):??, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288393.

Anonymous:1984:BMb

[Ano84b] Anonymous. Back matter. Journal of the American Statistical Association, 79(386):??, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288323.

Anonymous:1984:BMc

[Ano84c] Anonymous. Back matter. Journal of the American Statistical Association, 79(387):??, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288461.

Anonymous:1984:BMd

[Ano84d] Anonymous. Back matter. Journal of the American Statistical Association, 79(388):966–969, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288772.

Anonymous:1984:CPR

[Ano84e] Anonymous. Corrigenda: Publications received. Journal of the American Statistical Association, 79(388):965, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288770.

Anonymous:1984:FMa

[Ano84f] Anonymous. Front matter. Journal of the American Statistical Association, 79(385):??, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288324.

Anonymous:1984:FMb

[Ano84g] Anonymous. Front matter. Journal of the American Statistical Association, 79(386):??, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288255.

Anonymous:1984:FMc

[Ano84h] Anonymous. Front matter. Journal of the American Statistical Association, 79(387):??, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288394.

Anonymous:1984:FMd

[Ano84i] Anonymous. Front matter. Journal of the American Statistical Association, 79(388):747, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288704.

Anonymous:1984:LBRa

[Ano84j] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 79(385):222, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288361.

Anonymous:1984:LBRb

[Ano84k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 79(386):466, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288292.

Anonymous:1984:LBRc

[Ano84l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 79(387):724, September 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288425.

Anonymous:1984:LBRd

[Ano84m] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 79(388):940, December 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288728.

Anonymous:1984:PRa

[Ano84n] Anonymous. Publications received. Journal of the American Statistical Association, 79(385):244–245, March 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288392.

Anonymous:1984:PRb

[Ano84o] Anonymous. Publications received. Journal of the American Statistical Association, 79(386):485–487, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288322.

Anonymous:1984:PRc

[Ano84p] Anonymous. Publications received. Journal of the American Statistical Association, 79(387):745–746, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288460.

Anonymous:1984:PRd

[Ano84q] Anonymous. Publications received. Journal of the American Statistical Association, 79(388):964–965, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288769.

Anonymous:1984:VI

[Ano84r] Anonymous. Volume information. Journal of the American Statistical Association, 79(388):970–976, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288703.

Anonymous:1985:BMa

[Ano85a] Anonymous. Back matter. Journal of the American Statistical Association, 80(389):??, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288111.

Anonymous:1985:BMb

[Ano85b] Anonymous. Back matter. Journal of the American Statistical Association, 80(390):497–498, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287955.

Anonymous:1985:BMc

[Ano85c] Anonymous. Back matter. Journal of the American Statistical Association, 80(391):??, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288531.

Anonymous:1985:BMd

[Ano85d] Anonymous. Back matter. Journal of the American Statistical Association, 80(392):??, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288614.

Anonymous:1985:CQD

[Ano85e] Anonymous. Correction: Quotes, damned quotes, and. Journal of the American Statistical Association, 80(390):495, June 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287953.

Anonymous:1985:EC

[Ano85f] Anonymous. Editorial collaborators. Journal of the American Statistical Association, 80(392):1085–1088, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288613.

Anonymous:1985:FMa

[Ano85g] Anonymous. Front matter. Journal of the American Statistical Association, 80(389):??, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288028.

Anonymous:1985:FMb

[Ano85h] Anonymous. Front matter. Journal of the American Statistical Association, 80(390):??, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287880.

Anonymous:1985:FMc

[Ano85i] Anonymous. Front matter. Journal of the American Statistical Association, 80(391):i–iv, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288462.

Anonymous:1985:FMd

[Ano85j] Anonymous. Front matter. Journal of the American Statistical Association, 80(392):i–iv, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288533.

Anonymous:1985:LBRa

[Ano85k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 80(389):231, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288078.

Anonymous:1985:LBRb

[Ano85l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 80(390):478, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287919.

Anonymous:1985:LBRc

[Ano85m] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 80(391):767, September 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288498.

Anonymous:1985:LBRd

[Ano85n] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 80(392):1060, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288574.

Anonymous:1985:PRa

[Ano85o] Anonymous. Publications received. Journal of the American Statistical Association, 80(389):257, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288110.

Anonymous:1985:PRb

[Ano85p] Anonymous. Publications received. Journal of the American Statistical Association, 80(390):495–496, June 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287954.

Anonymous:1985:PRc

[Ano85q] Anonymous. Publications received. Journal of the American Statistical Association, 80(391):790–791, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288530.

Anonymous:1985:PRd

[Ano85r] Anonymous. Publications received. Journal of the American Statistical Association, 80(392):1082–1083, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288609.

Anonymous:1985:VI

[Ano85s] Anonymous. Volume information. Journal of the American Statistical Association, 80(392):1089–1095, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288532.

Anonymous:1986:BMa

[Ano86a] Anonymous. Back matter. Journal of the American Statistical Association, 81(393):??, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288027.

Anonymous:1986:BMb

[Ano86b] Anonymous. Back matter. Journal of the American Statistical Association, 81(394):??, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289278.

Anonymous:1986:BMc

[Ano86c] Anonymous. Back matter. Journal of the American Statistical Association, 81(395):??, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289053.

Anonymous:1986:BMd

[Ano86d] Anonymous. Back matter. Journal of the American Statistical Association, 81(396):1133–1135, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289116.

Anonymous:1986:FMa

[Ano86e] Anonymous. Front matter. Journal of the American Statistical Association, 81(393):i–iv, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287956.

Anonymous:1986:FMb

[Ano86f] Anonymous. Front matter. Journal of the American Statistical Association, 81(394):??, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289213.

Anonymous:1986:FMc

[Ano86g] Anonymous. Front matter. Journal of the American Statistical Association, 81(395):i–iv, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288981.

Anonymous:1986:FMd

[Ano86h] Anonymous. Front matter. Journal of the American Statistical Association, 81(396):i–iv, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289055.

Anonymous:1986:LBRa

[Ano86i] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 81(393):249, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287998.

Anonymous:1986:LBRb

[Ano86j] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 81(394):559, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289250.

Anonymous:1986:LBRc

[Ano86k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 81(395):856, September 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289019.

Anonymous:1986:LBRd

[Ano86l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 81(396):1109, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289091.

Anonymous:1986:PRa

[Ano86m] Anonymous. Publications received. Journal of the American Statistical Association, 81(393):267–268, March 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288026.

Anonymous:1986:PRb

[Ano86n] Anonymous. Publications received. Journal of the American Statistical Association, 81(394):581, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289277.

Anonymous:1986:PRc

[Ano86o] Anonymous. Publications received. Journal of the American Statistical Association, 81(395):875–877, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289052.

Anonymous:1986:VI

[Ano86p] Anonymous. Volume information. Journal of the American Statistical Association, 81(396):1136–1141, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289054.

Anonymous:1987:BMa

[Ano87a] Anonymous. Back matter. Journal of the American Statistical Association, 82(397):??, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289212.

Anonymous:1987:BMb

[Ano87b] Anonymous. Back matter. Journal of the American Statistical Association, 82(398):??, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289517.

Anonymous:1987:BMc

[Ano87c] Anonymous. Back matter. Journal of the American Statistical Association, 82(399):??, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288840.

Anonymous:1987:CEW

[Ano87d] Anonymous. Corrections: Efficiencies of weighted averages in stationary autoregressive processes. Journal of the American Statistical Association, 82(398):704, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289516.

Anonymous:1987:FMa

[Ano87e] Anonymous. Front matter. Journal of the American Statistical Association, 82(397):??, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289117.

Anonymous:1987:FMb

[Ano87f] Anonymous. Front matter. Journal of the American Statistical Association, 82(398):i–iv, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289437.

Anonymous:1987:FMc

[Ano87g] Anonymous. Front matter. Journal of the American Statistical Association, 82(399):i–iv, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288773.

Anonymous:1987:LE

[Ano87h] Anonymous. Letters to the editor. Journal of the American Statistical Association, 82(398):703–704, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289514.

Anonymous:1987:LBRa

[Ano87i] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 82(397):340, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289171.

Anonymous:1987:LBRb

[Ano87j] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 82(398):676, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289481.

Anonymous:1987:LBRc

[Ano87k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 82(399):945, September 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288809.

Anonymous:1987:PRa

[Ano87l] Anonymous. Publications received. Journal of the American Statistical Association, 82(397):366–367, March 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289211.

Anonymous:1987:PRb

[Ano87m] Anonymous. Publications received. Journal of the American Statistical Association, 82(398):701–702, June 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289513.

Anonymous:1987:PRc

[Ano87n] Anonymous. Publications received. Journal of the American Statistical Association, 82(399):962–963, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288839.

Anonymous:1988:BMa

[Ano88a] Anonymous. Back matter. Journal of the American Statistical Association, 83(401):??, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288980.

Anonymous:1988:BMb

[Ano88b] Anonymous. Back matter. Journal of the American Statistical Association, 83(402):??, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288911.

Anonymous:1988:BMc

[Ano88c] Anonymous. Back matter. Journal of the American Statistical Association, 83(404):1233–1236, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290190.

Anonymous:1988:CMS

[Ano88d] Anonymous. Corrections: a method for simulating stable random variables. Journal of the American Statistical Association, 83 (402):581, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288909.

Anonymous:1988:CBR

[Ano88e] Anonymous. Corrections: On the bounds of the range of order statistic processes. Journal of the American Statistical Association, 83(402):581, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288908.

Anonymous:1988:CWL

[Ano88f] Anonymous. Corrections: Weighted local regression and kernel methods for nonparametric curve fitting. Journal of the American Statistical Association, 83(402):581, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288910.

Anonymous:1988:FMa

[Ano88g] Anonymous. Front matter. Journal of the American Statistical Association, 83(401):i–iv, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288912.

Anonymous:1988:FMb

[Ano88h] Anonymous. Front matter. Journal of the American Statistical Association, 83(402):i–iv, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288841.

Anonymous:1988:FMc

[Ano88i] Anonymous. Front matter. Journal of the American Statistical Association, 83(404):??, December 1988. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290116.

Anonymous:1988:LEa

[Ano88j] Anonymous. Letters to the editor. Journal of the American Statistical Association, 83(402):581, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288907.

Anonymous:1988:LEb

[Ano88k] Anonymous. Letters to the editor. Journal of the American Statistical Association, 83(404):1231–1232, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290186.

Anonymous:1988:LBRa

[Ano88l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 83(401):265, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288949.

Anonymous:1988:LBRb

[Ano88m] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 83(402):561, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288878.

Anonymous:1988:LBRc

[Ano88n] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 83(404):1211, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290159.

Anonymous:1988:PRa

[Ano88o] Anonymous. Publications received. Journal of the American Statistical Association, 83(401):284–285, March 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288979.

Anonymous:1988:PRb

[Ano88p] Anonymous. Publications received. Journal of the American Statistical Association, 83(402):579–580, June 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288906.

Anonymous:1988:PRc

[Ano88q] Anonymous. Publications received. Journal of the American Statistical Association, 83(404):1229–1230, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290185.

Anonymous:1988:U

[Ano88r] Anonymous. [untitled]. Journal of the American Statistical Association, 83(404):1229, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2290184.

Anonymous:1988:VI

[Ano88s] Anonymous. Volume information. Journal of the American Statistical Association, 83(404):1237–1243, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290115.

Anonymous:1989:BMa

[Ano89a] Anonymous. Back matter. Journal of the American Statistical Association, 84(405):??, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289915.

Anonymous:1989:BMb

[Ano89b] Anonymous. Back matter. Journal of the American Statistical Association, 84(406):??, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289990.

Anonymous:1989:BMc

[Ano89c] Anonymous. Back matter. Journal of the American Statistical Association, 84(407):??, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289704.

Anonymous:1989:BMd

[Ano89d] Anonymous. Back matter. Journal of the American Statistical Association, 84(408):1108–1110, December 1989. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290114.

Anonymous:1989:FMa

[Ano89e] Anonymous. Front matter. Journal of the American Statistical Association, 84(405):??, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289835.

Anonymous:1989:FMb

[Ano89f] Anonymous. Front matter. Journal of the American Statistical Association, 84(406):??, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289916.

Anonymous:1989:FMc

[Ano89g] Anonymous. Front matter. Journal of the American Statistical Association, 84(407):??, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289643.

Anonymous:1989:FMd

[Ano89h] Anonymous. Front matter. Journal of the American Statistical Association, 84(408):??, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290057.

Anonymous:1989:LEa

[Ano89i] Anonymous. Letters to the editor. Journal of the American Statistical Association, 84(405):352, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289912.

Anonymous:1989:LEb

[Ano89j] Anonymous. Letters to the editor. Journal of the American Statistical Association, 84(406):636, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289989.

Anonymous:1989:LBRa

[Ano89k] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 84(405):331, March 1989. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289882.

Anonymous:1989:LBRb

[Ano89l] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 84(406):611, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289951.

Anonymous:1989:LBRc

[Ano89m] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 84(407):829, September 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289674.

Anonymous:1989:LBRd

[Ano89n] Anonymous. [list of book reviews]. Journal of the American Statistical Association, 84(408):1092, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290089.

Anonymous:1989:NEa

[Ano89o] Anonymous. New editions. Journal of the American Statistical Association, 84(405):350, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289910.

Anonymous:1989:NEb

[Ano89p] Anonymous. New editions. Journal of the American Statistical Association, 84(406):634, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289987.

Anonymous:1989:PRa

[Ano89q] Anonymous. Publications received. Journal of the American Statistical Association, 84(405):350–351, March 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289911.

Anonymous:1989:PRb

[Ano89r] Anonymous. Publications received. Journal of the American Statistical Association, 84(406):634–635, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289988.

Anonymous:1989:PRc

[Ano89s] Anonymous. Publications received. Journal of the American Statistical Association, 84(407):851–852, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289703.

Anonymous:1989:PRd

[Ano89t] Anonymous. Publications received. Journal of the American Statistical Association, 84(408):1106–1107, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290113.

Anonymous:1989:VI

[Ano89u] Anonymous. Volume information. Journal of the American Statistical Association, 84(408):1111–1116, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290056.

Anscombe:1980:SRG

[Ans80a] F. J. Anscombe. The Statistical Research Group, 1942–1945: Comment. Journal of the American Statistical Association, 75 (370):331, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287452.

Ansley:1980:BRB

[Ans80b] Craig F. Ansley. Book review: [untitled]. Journal of the American Statistical Association, 75(371):760–761, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287696.

Anscombe:1985:BRB

[Ans85] F. J. Anscombe. Book review: [untitled]. Journal of the American Statistical Association, 80(389):250, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288100.

Al-osh:1986:BFB

[Ao86] Mohamed Al-osh. Birth forecasting based on birth order probabilities, with applications to U.S. data. Journal of the American Statistical Association, 81(395):645–656, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288992.

Abdelbasit:1983:EDB

[AP83] K. M. Abdelbasit and R. L. Plackett. Experimental design for binary data. Journal of the American Statistical Association, 78 (381):90–98, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287114.

Andrews:1987:BMU

[AP87] Donald W. K. Andrews and Peter C. B. Phillips. Best medianunbiased estimation in linear regression with bounded asymmetric loss functions. Journal of the American Statistical Association, 82 (399):886–893, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288801.

Arnold:1989:BEP

[AP89a] Barry C. Arnold and S. James Press. Bayesian estimation and prediction for Pareto data. Journal of the American Statistical Association, 84(408):1079–1084, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290086.

Arnold:1989:CCD

[AP89b] Barry C. Arnold and S. James Press. Compatible conditional distributions. Journal of the American Statistical Association, 84 (405):152–156, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289858.

Arrow:1981:PSB

[APS81] Kenneth J. Arrow, Leon Pesotchinsky, and Milton Sobel. On partitioning a sample with binary-type questions in lieu of collecting observations. Journal of the American Statistical Association, 76 (374):402–409, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287842.

Archer:1989:AMC

[AR89] Linda E. Archer and Louise M. Ryan. Accounting for misclassification in the cause-of-death test for carcinogenicity. Journal of the American Statistical Association, 84(407):787–791, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289667.

Arjas:1988:GMA

[Arj88] Elja Arjas. A graphical method for assessing goodness of fit in Cox’s proportional hazards model. Journal of the American Statistical Association, 83(401):204–212, March 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288942.

Armstrong:1982:BRB

[Arm82] Thomas E. Armstrong. Book review: [untitled]. Journal of the American Statistical Association, 77(380):938, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287335.

Arnold:1980:AVT

[Arn80] Steven F. Arnold. Asymptotic validity of F tests for the ordinary linear model and the multiple correlation model. Journal of the American Statistical Association, 75(372):890–894, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287177.

Arnold:1981:BRB

[Arn81] Barry C. Arnold. Book review: [untitled]. Journal of the American Statistical Association, 76(374):492, June 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287859.

Arnold:1986:BRB

[Arn86] Barry C. Arnold. Book review: [untitled]. Journal of the American Statistical Association, 81(395):860, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289028.

Arnold:1987:BRB

[Arn87] Barry C. Arnold. Book review: [untitled]. Journal of the American Statistical Association, 82(399):951, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288819.

Arnold:1989:BRBa

[Arn89a] Barry C. Arnold. Book review: [untitled]. Journal of the American Statistical Association, 84(406):625, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289972.

Arnold:1989:BRBb

[Arn89b] Steven Arnold. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1104–1105, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290111.

Arthur:1982:BRB

[Art82] Jeffrey L. Arthur. Book review: [untitled]. Journal of the American Statistical Association, 77(377):217, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287801.

Arvesen:1986:BRB

[Arv86] James Arvesen. Book review: [untitled]. Journal of the American Statistical Association, 81(394):564, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289252.

Alho:1985:UPF

[AS85a] Juha M. Alho and Bruce D. Spencer. Uncertain population forecasting. Journal of the American Statistical Association, 80(390): 306–314, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287887.

Ali:1985:TEB

[AS85b] Mukhtar M. Ali and J. Lew Silver. Tests for equality between sets of coefficients in two linear regressions under heteroscedasticity. Journal of the American Statistical Association, 80(391): 730–735, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288492.

Arnold:1988:BDE

[AS88] Barry C. Arnold and David Strauss. Bivariate distributions with exponential conditionals. Journal of the American Statistical Association, 83(402):522–527, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288872.

Aschauer:1984:BRB

[Asc84] David Alan Aschauer. Book review: [untitled]. Journal of the American Statistical Association, 79(388):962, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288764.

Ashikaga:1982:BRBa

[Ash82a] Takamaru Ashikaga. Book review: [untitled]. Journal of the American Statistical Association, 77(378):498, June 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287287.

Ashikaga:1982:BRBb

[Ash82b] Takamaru Ashikaga. Book review: [untitled]. Journal of the American Statistical Association, 77(378):498–499, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287289.

Ashenfelter:1983:DPI

[Ash83] Orley Ashenfelter. Determining participation in income-tested social programs. Journal of the American Statistical Association, 78(383):517–525, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288113.

Askey:1980:DEB

[Ask80] Richard Askey. Density estimation and bump-hunting by the penalized likelihood method exemplified by scattering and meteorite data: Comment. Journal of the American Statistical Association, 75(369):69, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287384.

Assad:1983:BRB

[Ass83] Arjang A. Assad. Book review: [untitled]. Journal of the American Statistical Association, 78(381):210, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287145.

Athreya:1983:BRB

[Ath83] K. B. Athreya. Book review: [untitled]. Journal of the American Statistical Association, 78(381):210–211, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287147.

Athreya:1987:BRB

[Ath87] K. B. Athreya. Book review: [untitled]. Journal of the American Statistical Association, 82(397):347–348, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289182.

Attfield:1980:TAP

[Att80] C. L. F. Attfield. Testing the assumptions of the permanentincome model. Journal of the American Statistical Association, 75(369):32–38, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287375.

Atwood:1981:BRB

[Atw81] Corwin L. Atwood. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1012, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287609.

Atwood:1984:ATI

[Atw84] Corwin L. Atwood. Approximate tolerance intervals, based on maximum likelihood estimates. Journal of the American Statistical Association, 79(386):459–465, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288291.

Auer:1986:BRB

[Aue86] Richard E. Auer. Book review: [untitled]. Journal of the American Statistical Association, 81(394):564–565, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289253.

Austin:1989:BRB

[Aus89] James T. Austin. Book review: [untitled]. Journal of the American Statistical Association, 84(405):346–347, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289904.

Atkinson:1982:ISC

[AY82] A. C. Atkinson and Lam Yeh. Inference for Sichel’s compound Poisson distribution. Journal of the American Statistical Association, 77(377):153–158, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287782.

Byrne:1983:IAM

[BA83] Philip J. Byrne and Steven F. Arnold. Inference about multivariate means for a nonstationary autoregressive model. Journal of the American Statistical Association, 78(384):850–855, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288195.

Babu:1988:BRB

[Bab88] Gutti Jogesh Babu. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1217–1218, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290167.

Bacchetti:1989:AIM

[Bac89] Peter Bacchetti. Additive isotonic models. Journal of the American Statistical Association, 84(405):289–294, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289876.

Bailey:1982:BRB

[Bai82] Kent Bailey. Book review: [untitled]. Journal of the American Statistical Association, 77(380):946–947, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287349.

Bailar:1985:EPC

[Bai85] Barbara A. Bailar. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):109–114, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288052. See [EK85a].

Bailey:1987:RRP

[Bai87] R. A. Bailey. Restricted randomization: a practical example. Journal of the American Statistical Association, 82(399):712–719, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288775.

Bailar:1988:SPR

[Bai88] Barbara A. Bailar. Statistical practice and research: The essential interactions. Journal of the American Statistical Association, 83(401):1–8, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288913.

Bain:1989:BRB

[Bai89] Lee Bain. Book review: [untitled]. Journal of the American Statistical Association, 84(405):336, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289890.

Baltagi:1986:BRB

[Bal86] Badi H. Baltagi. Book review: [untitled]. Journal of the American Statistical Association, 81(395):868–869, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289039.

Bancroft:1981:BRB

[Ban81] T. A. Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 76(375):744, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287548.

Bancroft:1982:BRB

[Ban82] T. A. Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 77(377):212, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287793.

Bancroft:1986:BRB

[Ban86a] Diccon Bancroft. Book review: [untitled]. Journal of the American Statistical Association, 81(393):252–253, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288002.

Bankier:1986:EBS

[Ban86b] Michael D. Bankier. Estimators based on several stratified samples with applications to multiple frame surveys. Journal of the American Statistical Association, 81(396):1074–1079, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289086.

Banks:1988:BRB

[Ban88] David L. Banks. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1214, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290162.

Banks:1989:POE

[Ban89] David L. Banks. Patterns of oppression: An exploratory analysis of human-rights data. Journal of the American Statistical Association, 84(407):674–681, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289648.

Bartels:1982:RVN

[Bar82a] Robert Bartels. The rank version of von Neumann’s ratio test for randomness. Journal of the American Statistical Association, 77(377):40–46, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287767.

Bartholdi:1982:BRB

[Bar82b] John J. Bartholdi, III. Book review: [untitled]. Journal of the American Statistical Association, 77(377):223, March 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287812.

Barbara:1986:BRB

[Bar86a] Vincent P. Barbara. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1114–1116, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289094.

Barlow:1986:BRB

[Bar86b] Richard E. Barlow. Book review: [untitled]. Journal of the American Statistical Association, 81(393):258, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288009.

Barlow:1989:BRB

[Bar89] Richard E. Barlow. Book review: [untitled]. Journal of the American Statistical Association, 84(407):843–844, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289689.

Basu:1980:RAEa

[Bas80a] D. Basu. Randomization analysis of experimental data: The Fisher randomization test. Journal of the American Statistical Association, 75(371):575–582, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287648.

Basu:1980:RAEb

[Bas80b] D. Basu. Randomization analysis of experimental data: The Fisher randomization test rejoinder. Journal of the American Statistical Association, 75(371):593–595, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287654.

Basu:1983:BRB

[Bas83] A. K. Basu. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1003, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288241.

Basu:1985:BRBa

[Bas85a] Asit P. Basu. Book review: [untitled]. Journal of the American Statistical Association, 80(389):249, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288099.

Basu:1985:BRBb

[Bas85b] Asit P. Basu. Book review: [untitled]. Journal of the American Statistical Association, 80(390):492, June 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287945.

Basu:1987:BRBa

[Bas87a] Asit P. Basu. Book review: [untitled]. Journal of the American Statistical Association, 82(397):363, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289207.

Basu:1987:BRBb

[Bas87b] Asit P. Basu. Book review: [untitled]. Journal of the American Statistical Association, 82(398):699, June 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289508.

Bryant:1980:MCP

[BB80] John L. Bryant and Norman T. Bruvold. Multiple comparison procedures in the analysis of covariance. Journal of the American Statistical Association, 75(372):874–880, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287175.

Brockwell:1981:HEE

[BB81] P. J. Brockwell and B. M. Brown. High-efficiency estimation for the positive stable laws. Journal of the American Statistical Association, 76(375):626–631, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287521.

Baker:1985:OCC

[BB85a] Fred D. Baker and Rolf E. Bargmann. Orthogonal central composite designs of the third order in the evaluation of sensitivity and plant growth simulation models. Journal of the American Statistical Association, 80(391):574–579, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288472.

Brindley:1985:SNR

[BB85b] Dennis A. Brindley and Ralph A. Bradley. Some new results on Grubbs’s estimators. Journal of the American Statistical Association, 80(391):711–714, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288489.

Bjornstad:1988:EBE

[BB88] Jan F. Bjørnstad and Ronald W. Butler. The equivalence of backward elimination and multiple comparisons. Journal of the American Statistical Association, 83(401):136–144, March 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288932.

Berger:1989:EPM

[BB89a] James O. Berger and Jose M. Bernardo. Estimating a product of means: Bayesian analysis with reference priors. Journal of the American Statistical Association, 84(405):200–207, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289864.

Budde:1989:MTP

[BB89b] Michael Budde and Peter Bauer. Multiple test procedures in clinical dose finding studies. Journal of the American Statistical Association, 84(407):792–796, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289668.

Berlin:1989:APB

[BBL89] Jesse A. Berlin, Colin B. Begg, and Thomas A. Louis. An assessment of publication bias using a sample of published clinical trials. Journal of the American Statistical Association, 84(406):381–392, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289921.

Box:1982:ATR

[BC82a] G. E. P. Box and D. R. Cox. An analysis of transformations revisited, rebutted. Journal of the American Statistical Association, 77 (377):209–210, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287791.

Box:1982:CAT

[BC82b] G. E. P. Box and D. R. Cox. Comment on: “An analysis of transformations revisited” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 296–311, MR 83b:62048] by P. J. Bickel and K. A. Doksum. Journal of the American Statistical Association, 77(377):209–210, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [BD81a].

Brookmeyer:1982:SMT

[BC82c] Ron Brookmeyer and John Crowley. A k-sample median test for censored data. Journal of the American Statistical Association, 77(378):433–440, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287264.

Browdy:1982:BPC

[BC82d] Ben L. Browdy and Potter C. Chang. Bayes procedures for the classification of multiple polynomial trends with dependent residuals. Journal of the American Statistical Association, 77(378): 483–487, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287271.

Becker:1989:AST

[BC89] Mark P. Becker and Clifford C. Clogg. Analysis of sets of two-way contingency tables using association models. Journal of the American Statistical Association, 84(405):142–151, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289857.

Bohrer:1981:MTD

[BCF^{+}81] Robert Bohrer, Winston Chow, Ray Faith, V. M. Joshi, and C.-F. Wu. Multiple three-decision rules for factorial simple effects: Bonferroni wins again! Journal of the American Statistical Association, 76(373):119–124, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287056.

Bickel:1981:ATR

[BD81a] Peter J. Bickel and Kjell A. Doksum. An analysis of transformations revisited. Journal of the American Statistical Association, 76(374):296–311, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2287831. See comment [BC82b].

Blumenthal:1981:EBP

[BD81b] Saul Blumenthal and Ram C. Dahiya. Estimating the binomial parameter n. Journal of the American Statistical Association, 76 (376):903–909, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287586.

Berger:1988:BAR

[BD88] James O. Berger and John Deely. A Bayesian approach to ranking and selection of related means with alternatives to analysis-of-variance methodology. Journal of the American Statistical Association, 83(402):364–373, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288851.

Bayarri:1989:ORP

[BD89] M. J. Bayarri and M. H. DeGroot. Optimal reporting of predictions. Journal of the American Statistical Association, 84(405): 214–222, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289866.

Barnett:1982:QTD

[BE82] Arnold Barnett and Ellen Eisen. A quartile test for differences in distribution. Journal of the American Statistical Association, 77(377):47–51, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287768.

Berman:1985:UUB

[BE85] M. Berman and G. K. Eagleson. A useful upper bound for the tail probabilities of the scan statistic when the sample size is large. Journal of the American Statistical Association, 80(392): 886–889, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288548.

Beaver:1980:BRB

[Bea80] Robert J. Beaver. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1038–1039, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287212.

Beaver:1983:BRBa

[Bea83a] Robert J. Beaver. Book review: [untitled]. Journal of the American Statistical Association, 78(382):504–505, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288685.

Beaver:1983:BRBb

[Bea83b] Robert J. Beaver. Book review: [untitled]. Journal of the American Statistical Association, 78(382):506, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288687.

Bechhofer:1980:BRB

[Bec80] Robert E. Bechhofer. Book review: [untitled]. Journal of the American Statistical Association, 75(371):751–756, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287687.

Becker:1981:SPF

[Bec81] Stan Becker. Seasonal patterns of fertility measures: Theory and data. Journal of the American Statistical Association, 76(374): 249–259, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287819.

Beck:1983:RSH

[Bec83] Gerald J. Beck. A reanalysis of the Stanford heart transplant data: Comment. Journal of the American Statistical Association, 78 (382):290, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288629.

Becker:1984:BRB

[Bec84] William E. Becker, Jr. Book review: [untitled]. Journal of the American Statistical Association, 79(387):729–731, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288430.

Becker:1989:MAA

[Bec89] Mark P. Becker. Models for the analysis of association in multivariate contingency tables. Journal of the American Statistical Association, 84(408):1014–1019, December 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2290077.

Bellhouse:1980:LMR

[Bel80] David R. Bellhouse. Linear models for randomized response designs. Journal of the American Statistical Association, 75(372): 1001–1004, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287195.

Bellhouse:1981:BRB

[Bel81] D. R. Bellhouse. Book review: [untitled]. Journal of the American Statistical Association, 76(375):750–751, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287557.

Bell:1989:BRB

[Bel89a] William Bell. Book review: [untitled]. Journal of the American Statistical Association, 84(405):340–341, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289896.

Bell:1989:SAS

[Bel89b] William R. Bell. Sensitivity analysis of seasonal adjustments: Empirical case studies: Comment: Sources of uncertainty in seasonal adjustment. Journal of the American Statistical Association, 84 (405):22–24, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289839.

Benson:1982:BRB

[Ben82] Charles S. Benson. Book review: [untitled]. Journal of the American Statistical Association, 77(379):685, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287739.

Benjamin:1983:TRC

[Ben83a] Yoav Benjamin. Is the t test really conservative when the parent distribution is long-tailed? Journal of the American Statistical Association, 78(383):645–654, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288133.

Benjamini:1983:TRC

[Ben83b] Yoav Benjamini. Is the t test really conservative when the parent distribution is long-tailed. Journal of the American Statistical Association, 78(383):645–654, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1981:NSC

[Ber81] Roger L. Berger. A necessary and sufficient condition for reaching a consensus using DeGroot’s method. Journal of the American Statistical Association, 76(374):415–418, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287844.

Berger:1982:BRS

[Ber82] James Berger. Bayesian robustness and the Stein effect. Journal of the American Statistical Association, 77(378):358–368, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287253.

Beran:1983:BRB

[Ber83a] Rudolf J. Beran. Book review: [untitled]. Journal of the American Statistical Association, 78(382):503–504, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288683.

Berger:1983:PEB

[Ber83b] James Berger. Parametric empirical Bayes inference: Theory and applications: Comment. Journal of the American Statistical Association, 78(381):55–57, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287099.

Bergstrom:1983:BRB

[Ber83c] Ted Bergstrom. Book review: [untitled]. Journal of the American Statistical Association, 78(383):741–742, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288164.

Berk:1983:BRB

[Ber83d] Kenneth N. Berk. Book review: [untitled]. Journal of the American Statistical Association, 78(381):213–214, March 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287153.

Berger:1984:AWW

[Ber84a] James O. Berger. Abraham Wald’s work on aircraft survivability: Comment. Journal of the American Statistical Association, 79 (386):267–269, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288258.

Berger:1984:BRB

[Ber84b] James O. Berger. Book review: [untitled]. Journal of the American Statistical Association, 79(388):941–942, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288729.

Bernardo:1984:MSS

[Ber84c] Jose M. Bernardo. Monitoring the 1982 Spanish Socialist victory: a Bayesian analysis. Journal of the American Statistical Association, 79(387):510–515, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288397.

Berger:1985:BRB

[Ber85a] James O. Berger. Book review: [untitled]. Journal of the American Statistical Association, 80(389):232–233, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288079.

Berman:1985:BRB

[Ber85b] Simeon M. Berman. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1070–1071, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288586.

Berndt:1985:BRB

[Ber85c] E. R. Berndt. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1063–1064, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288578.

Berry:1986:BRB

[Ber86] Donald A. Berry. Book review: [untitled]. Journal of the American Statistical Association, 81(394):573, June 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289262.

Beran:1987:BBC

[Ber87] Rudolf Beran. Better bootstrap confidence intervals: Comment. Journal of the American Statistical Association, 82(397):195– 196, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289151.

Beran:1988:BSC

[Ber88a] Rudolf Beran. Balanced simultaneous confidence sets. Journal of the American Statistical Association, 83(403):679–686, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Beran:1988:PTS

[Ber88b] Rudolf Beran. Prepivoting test statistics: a bootstrap view of asymptotic refinements. Journal of the American Statistical Association, 83(403):687–697, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Berger:1988:BVS

[Ber88c] James Berger. Bayesian variable selection in linear regression: Comment. Journal of the American Statistical Association, 83 (404):1033–1034, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290130.

Berger:1989:UMP

[Ber89] Roger L. Berger. Uniformly more powerful tests for hypotheses concerning linear inequalities and normal means. Journal of the American Statistical Association, 84(405):192–199, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289863.

Bethel:1989:MVE

[Bet89] James Bethel. Minimum variance estimation in stratified sampling. Journal of the American Statistical Association, 84(405): 260–265, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289872.

Bain:1985:TIT

[BEW85] Lee J. Bain, Max Engelhardt, and F. T. Wright. Tests for an increasing trend in the intensity of a Poisson process: a power study. Journal of the American Statistical Association, 80(390): 419–422, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287907.

Breiman:1983:HMV

[BF83] L. Breiman and D. Freedman. How many variables should be entered in a regression equation. Journal of the American Statistical Association, 78(381):131–136, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287119.

Breiman:1985:EOTb

[BF85a] Leo Breiman and Jerome H. Friedman. Estimating optimal transformations for multiple regression and correlation: Rejoinder. Journal of the American Statistical Association, 80(391):614–619, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288477.

Breiman:1985:EOTa

[BF85b] Leo Breiman and Jerome H. Friedman. Estimating optimal transformations for multiple regression and correlation. With discussion and with a reply by the authors. Journal of the American Statistical Association, 80(391):580–598, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288473.

Basu:1980:APS

[BG80] A. P. Basu and J. K. Ghosh. Asymptotic properties of a solution to the likelihood equation with life- testing applications. Journal of the American Statistical Association, 75(370):410–414, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287468.

Begun:1981:CNK

[BG81] Janet M. Begun and K. Ruben Gabriel. Closure of the Newman– Keuls multiple comparisons procedure. Journal of the American Statistical Association, 76(374):241–245, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287817.

Brookmeyer:1988:MOS

[BG88] Ron Brookmeyer and Mitchell H. Gail. A method for obtaining short-term projections and lower bounds on the size of the AIDS epidemic. Journal of the American Statistical Association, 83(402):301–308, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288844.

Bhalerao:1980:MIP

[BGT80] Narayan R. Bhalerao, John Gurland, and Ram C. Tripathi. A method of increasing power of a test for the negative binomial and Neyman Type A distributions. Journal of the American Statistical Association, 75(372):934–938, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287184.

Bancroft:1983:NPV

[BH83a] T. A. Bancroft and Chien-Pai Han. A note on pooling variances. Journal of the American Statistical Association, 78(384):981–983, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288213.

Bell:1983:MTS

[BH83b] W. R. Bell and S. C. Hillmer. Modeling time series with calendar variation. Journal of the American Statistical Association, 78 (383):526–534, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288114.

Bye:1985:BRB

[BH85] Barry V. Bye and John C. Hennessey. Book review: [untitled]. Journal of the American Statistical Association, 80(391):775–776, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288506.

Berliner:1988:BNS

[BH88] L. Mark Berliner and Bruce M. Hill. Bayesian nonparametric survival analysis. With comments by Christian Genest, Jack Kalbfleisch and Seymour Geisser and with a reply by the authors. Journal of the American Statistical Association, 83(403):772–784, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Bhansali:1981:EKO

[Bha81] R. J. Bhansali. Effects of not knowing the order of an autoregressive process on the mean squared error of prediction. I. Journal of the American Statistical Association, 76(375):588–597, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287516.

Bhattacharyya:1982:BRB

[Bha82] B. B. Bhattacharyya. Book review: [untitled]. Journal of the American Statistical Association, 77(379):691, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287751.

Bhattacharyya:1985:AML

[Bha85] Gouri K. Bhattacharyya. The asymptotics of maximum likelihood and related estimators based on Type II censored data. Journal of the American Statistical Association, 80(390):398–404, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287904.

Bhattacharya:1987:BRB

[Bha87] R. N. Bhattacharya. Book review: [untitled]. Journal of the American Statistical Association, 82(399):949, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288816.

Battese:1988:ECM

[BHF88] George E. Battese, Rachel M. Harter, and Wayne A. Fuller. An error-components model for prediction of county crop areas using survey and satellite data. Journal of the American Statistical Association, 83(401):28–36, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288915.

Brownie:1983:EPC

[BHR83] Cavell Brownie, Jean-Pierre Habicht, and Douglas S. Robson. An estimation procedure for the contaminated normal distributions arising in clinical chemistry. Journal of the American Statistical Association, 78(382):228–237, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288617.

Bickel:1981:RDA

[BHS81] P. J. Bickel, Agnes M. Herzberg, and M. F. Schilling. Robustness of design against autocorrelation in time. II. optimality, theoretical and numerical results for the first-order autoregressive process. Journal of the American Statistical Association, 76(376): 870–877, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287582.

Brewer:1988:HNC

[BHT88] K. R. W. Brewer, Muhammad Hanif, and S. M. Tam. How nearly can model-based prediction and design-based estimation be reconciled. Journal of the American Statistical Association, 83(401): 128–132, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288930.

Bickel:1983:MAB

[Bic83] Peter J. Bickel. Minimax aspects of bounded-influence regression: Comment. Journal of the American Statistical Association, 78 (381):75–77, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287109.

Bickel:1984:ATD

[Bic84] Peter J. Bickel. The analysis of transformed data: Comment. Journal of the American Statistical Association, 79(386):315–316, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288268.

Bickel:1987:BBC

[Bic87] P. J. Bickel. Better boostrap confidence intervals: Comment. Journal of the American Statistical Association, 82(397):191, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289149.

Biemer:1983:BRB

[Bie83a] Paul Biemer. Book review: [untitled]. Journal of the American Statistical Association, 78(384):996–997, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288230.

Bierens:1983:UCK

[Bie83b] Herman J. Bierens. Uniform consistency of kernel estimators of a regression function under generalized conditions. Journal of the American Statistical Association, 78(383):699–707, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288140.

Billard:1981:BRB

[Bil81] Lynne Billard. Book review: [untitled]. Journal of the American Statistical Association, 76(374):493, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287862.

Billard:1983:BRBb

[Bil83a] L. Billard. Book review: [untitled]. Journal of the American Statistical Association, 78(384):991, December 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288223.

Billard:1983:BRBc

[Bil83b] L. Billard. Book review: [untitled]. Journal of the American Statistical Association, 78(384):992–993, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288225.

Billard:1983:BRBa

[Bil83c] Lynne Billard. Book review: [untitled]. Journal of the American Statistical Association, 78(382):501, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288676.

Billard:1984:BRBa

[Bil84a] L. Billard. Book review: [untitled]. Journal of the American Statistical Association, 79(388):949–951, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288739.

Billard:1984:BRBb

[Bil84b] L. Billard. Book review: [untitled]. Journal of the American Statistical Association, 79(388):951, December 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288740.

Billard:1985:BRB

[Bil85] Lynne Billard. Book review: [untitled]. Journal of the American Statistical Association, 80(391):788–789, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288526.

Binkley:1982:EVC

[Bin82] James K. Binkley. The effect of variable correlation on the efficiency of seemingly unrelated regression in a two-equation model. Journal of the American Statistical Association, 77(380):890–895, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287323.

Bingham:1985:BRB

[Bin85] Christopher Bingham. Book review: [untitled]. Journal of the American Statistical Association, 80(389):243, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288090.

Bishop:1988:BRB

[Bis88] Yvonne M. Bishop. Book review: [untitled]. Journal of the American Statistical Association, 83(401):273, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288961.

Biyani:1980:IYG

[Biy80] Shriram H. Biyani. On inadmissibility of the Yates–Grundy variance estimator in unequal probability sampling. Journal of the American Statistical Association, 75(371):709–712, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287674.

Beckman:1981:RPP

[BJ81] Richard J. Beckman and Mark E. Johnson. A ranking procedure for partial discriminant analysis. Journal of the American Statistical Association, 76(375):671–675, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287528.

Brown:1981:VDT

[BJM81] Lawrence D. Brown, Iain M. Johnstone, and K. Brenda MacGibbon. Variation diminishing transformations: a direct approach to total positivity and its statistical applications. Journal of the American Statistical Association, 76(376):824–832, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287577.

Bassett:1982:EQF

[BK82] Gilbert Bassett, Jr. and Roger Koenker. An empirical quantile function for linear models with iid errors. Journal of the American Statistical Association, 77(378):407–415, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287261.

Bayless:1984:BRB

[BK84a] David Bayless and Richard Kulka. Book review: [untitled]. Journal of the American Statistical Association, 79(387):727–728, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288428.

Bonney:1984:IPC

[BK84b] G. E. Bonney and G. E. Kissling. On the inverse of a patterned covariance matrix. Journal of the American Statistical Association, 79(387):722–723, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288424.

Bendre:1985:MET

[BK85a] S. M. Bendre and B. K. Kale. Masking effect on tests for outliers in exponential models. Journal of the American Statistical Association, 80(392):1020–1025, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288569. See correction [BK86].

Buja:1985:EOT

[BK85b] Andreas Buja and Robert E. Kass. Estimating optimal transformations for multiple regression and correlation: Comment. Journal of the American Statistical Association, 80(391):602–607, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288475.

Butler:1985:DPP

[BK85c] William J. Butler and Richard A. Kronmal. Discrimination with polychotomous predictor variables using orthogonal functions.

Journal of the American Statistical Association, 80(390):443–448, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287912.

Bendre:1986:CME

[BK86] S. M. Bendre and B. K. Kale. Correction: “Masking effect on tests for outliers in exponential samples” [J. Amer. Statist. Assoc. 80 (1985), no. 392, 1020–1025, MR 87c:62036]. Journal of the American Statistical Association, 81(396):1132, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289115. See [BK85a].

Bickel:1989:CBD

[BK89] P. J. Bickel and A. M. Krieger. Confidence bands for a distribution function using the bootstrap. Journal of the American Statistical Association, 84(405):95–100, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289850.

Baker:1988:CRA

[BL88a] S. G. Baker and N. M. Laird. Correction: Regression analysis for categorical variables with outcome subject to nonignorable nonresponse. Journal of the American Statistical Association, 83 (404):1232, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290189.

Baker:1988:RAC

[BL88b] Stuart G. Baker and Nan M. Laird. Regression analysis for categorical variables with outcome subject to nonignorable nonresponse. Journal of the American Statistical Association, 83(401): 62–69, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288919.

Bar-Lev:1988:SPU

[BLE88] Shaul K. Bar-Lev and Peter Enis. Sign-preserving unbiased estimators in linear exponential families. Journal of the American Statistical Association, 83(404):1187–1189, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290154.

Breslow:1983:MMC

[BLML83] N. E. Breslow, J. H. Lubin, P. Marek, and B. Langholz. Multiplicative models and cohort analysis. Journal of the American Statistical Association, 78(381):1–12, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287093.

Blom:1980:ELE

[Blo80] Gunnar Blom. Extrapolation of linear estimates to larger sample sizes. Journal of the American Statistical Association, 75(372): 912–917, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287181.

Bloomfield:1984:BRD

[Blo84] Peter Bloomfield. Book review: Digital Filters (2nd ed.) by R. W. Hamming. Journal of the American Statistical Association, 79 (387):736–737, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288442.

Blyth:1986:ABC

[Bly86] Colin R. Blyth. Approximate binomial confidence limits. Journal of the American Statistical Association, 81(395):843–855, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289018.

Bhattacharyya:1981:TET

[BM81] Gouri K. Bhattacharyya and Kishan G. Mehrotra. On testing equality of two exponential distributions under combined Type II censoring. Journal of the American Statistical Association, 76 (376):886–894, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287584.

Bhattacharyya:1983:UTL

[BM83] Gouri K. Bhattacharyya and Kishan G. Mehrotra. A unified treatment of locally most powerful rank tests under Type II censoring. Journal of the American Statistical Association, 78(382):375–381, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288643.

Basford:1985:EAR

[BM85] K. E. Basford and G. J. McLachlan. Estimation of allocation rates in a cluster analysis context. Journal of the American Statistical Association, 80(390):286–293, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287884.

Bourke:1988:EPR

[BM88] Patrick D. Bourke and Michael A. Moran. Estimating proportions from randomized response data using the EM algorithm. Journal of the American Statistical Association, 83(404):964–968, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290120.

Bolk:1989:ANM

[BM89] Robert J. Bolk and Mervyn G. Marasinghe. Analysis of nonadditive multiway classifications. Journal of the American Statistical Association, 84(408):1059–1064, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290083.

Burbidge:1988:ATH

[BMR88] John B. Burbidge, Lonnie Magee, and A. Leslie Robb. Alternative transformations to handle extreme values of the dependent variable. Journal of the American Statistical Association, 83(401): 123–127, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288929.

Barndorff-Nielsen:1982:SAS

[BN82] O. Barndorff-Nielsen. Some ancillary statistics and their properties: Comment. Journal of the American Statistical Association, 77(379):590–591, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287715.

Barndorff-Nielsen:1984:BRB

[BN84] O. E. Barndorff-Nielsen. Book review: [untitled]. Journal of the American Statistical Association, 79(387):734–735, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288437.

Blackstone:1986:DTV

[BNT86] Eugene H. Blackstone, David C. Naftel, and Malcolm E. Turner, Jr. The decomposition of time-varying hazard into phases, each incorporating a separate stream of concomitant information. Journal of the American Statistical Association, 81(395):615–624, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288989.

Boardman:1983:RSH

[Boa83] Thomas J. Boardman. A reanalysis of the Stanford heart transplant data: Comment. Journal of the American Statistical Association, 78(382):282–285, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288625.

Bockenholt:1989:BRB

[Boc89] Ulf Bockenholt. Book review: [untitled]. Journal of the American Statistical Association, 84(406):619–620, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289961.

Bodian:1985:BRB

[Bod85] Carol A. Bodian. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1080–1081, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288605.

Boen:1983:BRB

[Boe83] James R. Boen. Book review: [untitled]. Journal of the American Statistical Association, 78(381):206–207, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287140.

Bofinger:1985:MCT

[Bof85] Eve Bofinger. Multiple comparisons and Type III errors. Journal of the American Statistical Association, 80(390):433–437, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287910.

Boik:1986:TRM

[Boi86] Robert J. Boik. Testing the rank of a matrix with applications to the analysis of interaction in ANOVA. Journal of the American Statistical Association, 81(393):243–248, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287997.

Bondar:1987:BRB

[Bon87] James V. Bondar. Book review: [untitled]. Journal of the American Statistical Association, 82(399):950–951, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288818.

Book:1980:BRB

[Boo80a] Stephen A. Book. Book review: [untitled]. Journal of the American Statistical Association, 75(369):247, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287432.

Boos:1980:NMC

[Boo80b] Dennis D. Boos. A new method for constructing approximate confidence intervals from M estimates. Journal of the American Statistical Association, 75(369):142–145, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287402.

Boos:1981:MDE

[Boo81a] Dennis D. Boos. Minimum distance estimators for location and goodness of fit. Journal of the American Statistical Association, 76 (375):663–670, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287527.

Booth:1981:BRB

[Boo81b] Gordon D. Booth. Book review: [untitled]. Journal of the American Statistical Association, 76(375):753–754, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287561.

Boos:1982:TAA

[Boo82] Dennis D. Boos. A test for asymmetry associated with the Hodges-Lehmann estimator. Journal of the American Statistical Association, 77(379):647–651, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287728.

Boos:1986:CPL

[Boo86] Dennis D. Boos. Comparing K populations with linear rank statistics. Journal of the American Statistical Association, 81(396): 1018–1025, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289077.

Baglivo:1988:MAC

[BOP88] Jenny Baglivo, Donald Olivier, and Marcello Pagano. Methods for the analysis of contingency tables with large and small cell counts. Journal of the American Statistical Association, 83(404): 1006–1013, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290127.

Bornholt:1987:CJB

[Bor87] Graham Bornholt. Comment on J. O. Berger and T. Sellke: “Testing a point null hypothesis: irreconcilability of P values and evidence” [J. Amer. Statist. Assoc. 82 (1987), no. 397, 112–139, MR 89a:62009a]. With a reply by Sellke. Journal of the American Statistical Association, 82(400):1201, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [BS87a].

Bowerman:1983:BRB

[Bow83] Bruce L. Bowerman. Book review: [untitled]. Journal of the American Statistical Association, 78(381):207, March 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287141.

Bowie:1986:BRB

[Bow86] Lee Bowie. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1116, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289095.

Bowden:1987:RSP

[Bow87] Roger J. Bowden. Repeated sampling in the presence of publication effects. With comments by Lawrence S. Cahoon and Peter C. Ordeshook. Journal of the American Statistical Association, 82(398):476–484, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289449.

Bozdogan:1985:BRB

[Boz85] Hamparsum Bozdogan. Book review: [untitled]. Journal of the American Statistical Association, 80(389):252, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288104.

Brown:1986:ADE

[BP86] Philip J. Brown and Clive D. Payne. Aggregate data, ecological regression, and voting transitions. Journal of the American Statistical Association, 81(394):452–460, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289235.

Box:1987:ETG

[BPN87] George E. P. Box, David A. Pierce, and Paul Newbold. Estimating trend and growth rates in seasonal time series. Journal of the American Statistical Association, 82(397):276–282, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289164.

Butler:1980:PIB

[BR80] R. Butler and E. D. Rothman. Predictive intervals based on reuse of the sample. Journal of the American Statistical Association, 75(372):881–889, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287176.

Begun:1983:ERR

[BR83a] J. M. Begun and N. Reid. Estimating the relative risk with censored data. Journal of the American Statistical Association, 78 (382):337–341, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288638.

Berk:1983:CNA

[BR83b] Richard A. Berk and David Rauma. Capitalizing on nonrandom assignment to treatments: a regression-discontinuity evaluation of a crime-control program. Journal of the American Statistical Association, 78(381):21–27, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287095.

Bramson:1980:BRB

[Bra80] Maury Bramson. Book review: [untitled]. Journal of the American Statistical Association, 75(370):470, June 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287480.

Bradburn:1981:BRB

[Bra81] Norman M. Bradburn. Book review: [untitled]. Journal of the American Statistical Association, 76(373):212, March 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287082.

Bradley:1982:FSD

[Bra82] Ralph A. Bradley. The future of statistics as a discipline. Journal of the American Statistical Association, 77(377):1–10, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287763.

Braden:1984:BRB

[Bra84a] Magdalen Braden. Book review: [untitled]. Journal of the American Statistical Association, 79(387):745, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288459.

Bramson:1984:BRB

[Bra84b] Maury Bramson. Book review: Theory and Applications of Stochastic Differential Equations, by Zeev Schuss. Journal of the American Statistical Association, 79(385):243–244, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288391.

Braden:1985:BRBa

[Bra85a] Magdalen Braden. Book review: [untitled]. Journal of the American Statistical Association, 80(389):242, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288089.

Braden:1985:BRBb

[Bra85b] Magdalen Braden. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1073–1074, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288590.

Brazee:1985:BRB

[Bra85c] Richard Brazee. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1063, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288577.

Braun:1986:BRB

[Bra86] Henry Braun. Book review: [untitled]. Journal of the American Statistical Association, 81(395):857, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289021.

Bradley:1987:BRB

[Bra87] Richard C. Bradley. Book review: [untitled]. Journal of the American Statistical Association, 82(399):946–947, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288811.

Brant:1988:BRB

[Bra88] Rollin Brant. Book review: [untitled]. Journal of the American Statistical Association, 83(402):563, June 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288882.

Breiger:1981:EFP

[Bre81] Ronald L. Breiger. An exponential family of probability distributions for directed graphs: Comment. Journal of the American Statistical Association, 76(373):51–53, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287038.

Breen:1986:BRB

[Bre86] T. J. Breen. Book review: [untitled]. Journal of the American Statistical Association, 81(394):577, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289268.

Brilmayer:1982:SED

[Bri82] Lea Brilmayer. Statistical evidence of discrimination: Comment. Journal of the American Statistical Association, 77(380):789–790, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287308.

Brillinger:1983:BRB

[Bri83] David R. Brillinger. Book review: [untitled]. Journal of the American Statistical Association, 78(384):986–987, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288217.

Brockett:1980:BRBa

[Bro80a] Patrick L. Brockett. Book review: [untitled]. Journal of the American Statistical Association, 75(369):241–242, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287422.

Brockett:1980:BRBb

[Bro80b] Patrick L. Brockett. Book review: [untitled]. Journal of the American Statistical Association, 75(370):470, June 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287481.

Brockett:1980:BRBc

[Bro80c] Patrick L. Brockett. Book review: [untitled]. Journal of the American Statistical Association, 75(370):473–474, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287487.

Brown:1980:BRB

[Bro80d] Kenneth G. Brown. Book review: [untitled]. Journal of the American Statistical Association, 75(370):476–477, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287493.

Brockett:1981:BRB

[Bro81a] Patrick L. Brockett. Book review: [untitled]. Journal of the American Statistical Association, 76(373):209, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287076.

Brostrom:1981:OCC

[Bro81b] Göran Broström. On optimal choice of components for parallel systems. Journal of the American Statistical Association, 76(374): 486–490, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287856.

Brockett:1982:BRBa

[Bro82a] Patrick L. Brockett. Book review: [untitled]. Journal of the American Statistical Association, 77(379):688, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287743.

Brockett:1982:BRBb

[Bro82b] Patrick L. Brockett. Book review: [untitled]. Journal of the American Statistical Association, 77(379):688, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287744.

Brown:1982:RLE

[Bro82c] Michael L. Brown. Robust line estimation with errors in both variables. Journal of the American Statistical Association, 77(377): 71–79, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287771.

Brockett:1983:BRB

[Bro83a] Patrick L. Brockett. Book review: [untitled]. Journal of the American Statistical Association, 78(382):503, June 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288682.

Brown:1983:SBD

[Bro83b] Kenneth G. Brown. Sub-balanced data and the mixed analysis of variance. Journal of the American Statistical Association, 78 (381):162–167, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287124.

Brown:1983:CRL

[Bro83c] Michael L. Brown. Corrigenda: Robust line estimation with errors in both variables. Journal of the American Statistical Association, 78(384):1008, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288250.

Brown:1983:BRB

[Bro83d] R. Grover Brown. Book review: [untitled]. Journal of the American Statistical Association, 78(383):744, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288171.

Brockett:1984:BRB

[Bro84] Patrick L. Brockett. Book review: Stochastic Methods in Economics and Finance, by A. G. Malliaris; W. A. Brock. Journal of the American Statistical Association, 79(385):236–237, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288379.

Brockwell:1987:BRB

[Bro87] Peter J. Brockwell. Book review: [untitled]. Journal of the American Statistical Association, 82(397):346, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289180.

Brown:1988:BRB

[Bro88] Lawrence D. Brown. Book review: [untitled]. Journal of the American Statistical Association, 83(402):564–565, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288884.

Bryant:1984:BRB

[Bry84] Edward C. Bryant. Book review: [untitled]. Journal of the American Statistical Association, 79(386):479, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288311.

Blyth:1983:BCI

[BS83a] Colin R. Blyth and Harold A. Still. Binomial confidence intervals. Journal of the American Statistical Association, 78(381): 108–116, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287116.

Boyles:1983:MLE

[BS83b] Russell A. Boyles and Francisco J. Samaniego. Maximum likelihood estimation for a discrete multivariate shock model. Journal of the American Statistical Association, 78(382):445–448, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288655.

Berger:1984:THC

[BS84a] Roger L. Berger and Dennis F. Sinclair. Testing hypotheses concerning unions of linear subspaces. Journal of the American Statistical Association, 79(385):158–163, March 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288351.

Boyles:1984:MIM

[BS84b] Russell A. Boyles and Francisco J. Samaniego. Modeling and inference for multivariate binary data with positive dependence. Journal of the American Statistical Association, 79(385):188– 193, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288355.

Biemer:1985:ODI

[BS85a] Paul P. Biemer and S. Lynne Stokes. Optimal design of interviewer variance experiments in complex surveys. Journal of the American Statistical Association, 80(389):158–166, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288066.

Boyle:1985:SMF

[BS85b] Kerrie Eileen Boyle and Thomas B. Starr. Survival models for fertility evaluation. Journal of the American Statistical Association, 80(392):823–827, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288539.

Boyd:1986:UIR

[BS86a] Michael N. Boyd and Pranab K. Sen. Union-intersection rank tests for ordered alternatives in ANOCOVA. Journal of the American Statistical Association, 81(394):526–532, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289245.

Boyles:1986:EDF

[BS86b] Russell A. Boyles and Francisco J. Samaniego. Estimating a distribution function based on nomination sampling. Journal of the American Statistical Association, 81(396):1039–1045, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289080.

Bye:1986:LMM

[BS86c] Barry V. Bye and Evan S. Schechter. A latent Markov model approach to the estimation of response errors in multiwave panel data. Journal of the American Statistical Association, 81(394): 375–380, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289226.

Berger:1987:TPNa

[BS87a] James O. Berger and Thomas Sellke. Testing a point null hypothesis: irreconcilability of P values and evidence. With comments and a rejoinder by the authors. Journal of the American Statistical Association, 82(397):112–122, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289131. See comment [Bor87].

Berger:1987:TPNb

[BS87b] James O. Berger and Thomas Sellke. Testing a point null hypothesis: The irreconcilability of P values and evidence: Rejoinder. Journal of the American Statistical Association, 82(397):135– 139, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289139.

Berk:1988:PRF

[BS88a] Richard A. Berk and Lawrence W. Sherman. Police responses to family violence incidents: An analysis of an experimental design with incomplete randomization. Journal of the American Statistical Association, 83(401):70–76, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288920.

Boyett:1988:CNO

[BS88b] J. M. Boyett and J. J. Shuster. Correction: Nonparametric onesided tests in multivariate analysis with medical application. Journal of the American Statistical Association, 83(404):1232, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290187.

Bloch:1989:ECA

[BS89] Daniel A. Bloch and Mark R. Segal. Empirical comparison of approaches to forming strata: Using classification trees to adjust for covariates. Journal of the American Statistical Association, 84 (408):897–905, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290064.

Beckett:1980:SDR

[BSB80] James Beckett, William R. Schucany, and James D. Broffitt. A special distributional result for bilinear forms. Journal of the American Statistical Association, 75(370):466–468, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287478.

Bowden:1981:CSI

[BT81] R. J. Bowden and D. A. Turkington. A comparative study of instrumental variables estimators for nonlinear simultaneous models. Journal of the American Statistical Association, 76(376):988– 995, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287600.

Bubolz:1980:BRB

[Bub80] Thomas A. Bubolz. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1041, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287218.

Bubolz:1983:BRB

[Bub83] Thomas A. Bubolz. Book review: [untitled]. Journal of the American Statistical Association, 78(382):511, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288698.

Bubolz:1984:BRB

[Bub84] Thomas A. Bubolz. Book review: [untitled]. Journal of the American Statistical Association, 79(386):477–478, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288308.

Buckland:1989:BRB

[Buc89] Stephen T. Buckland. Book review: [untitled]. Journal of the American Statistical Association, 84(406):628, June 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289978.

Buehler:1981:BRB

[Bue81] Robert J. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 76(373):207–208, March 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287072.

Buehler:1982:SASb

[Bue82a] Robert J. Buehler. Some ancillary statistics and their properties: Rejoinder. Journal of the American Statistical Association, 77 (379):593–594, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287718.

Buehler:1982:SASa

[Bue82b] Robert J. Buehler. Some ancillary statistics and their properties. With comments by O. Barndorff-Nielsen, D. R. Cox and David Hinkley and a rejoinder by the author. Journal of the American Statistical Association, 77(379):581–589, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287714.

Buehler:1983:BRBa

[Bue83a] Robert J. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 78(381):210, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287146.

Buehler:1983:BRBb

[Bue83b] Robert J. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 78(382):512–513, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288700.

Buehler:1984:BRB

[Bue84] Robert J. Buehler. Book review: [untitled]. Journal of the American Statistical Association, 79(388):942–943, December 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288731.

Buncher:1989:I

[Bun89] C. Ralph Buncher. [introduction]. Journal of the American Statistical Association, 84(407):775, September 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289664.

Buse:1984:BRB

[Bus84] Adolf Buse. Book review: [untitled]. Journal of the American Statistical Association, 79(386):481–482, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288315.

Busch:1987:BRB

[Bus87] Paul S. Busch. Book review: [untitled]. Journal of the American Statistical Association, 82(399):956–957, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288829.

Butler:1984:SAB

[But84] Ronald W. Butler. The significance attained by the best-fitting regressor variable. Journal of the American Statistical Association, 79(386):341–348, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288274.

Butz:1985:BRB

[But85] William P. Butz. Book review: [untitled]. Journal of the American Statistical Association, 80(390):484, June 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287928.

Burridge:1985:CVS

[BW85] Peter Burridge and Kenneth F. Wallis. Calculating the variance of seasonally adjusted series. Journal of the American Statistical Association, 80(391):541–552, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288467.

Bustos:1986:REA

[BY86] Oscar H. Bustos and Víctor J. Yohai. Robust estimates for ARMA models. Journal of the American Statistical Association, 81(393): 155–168, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287983.

Bickel:1988:REB

[BY88] Peter J. Bickel and Joseph A. Yahav. Richardson extrapolation and the bootstrap. Journal of the American Statistical Association, 83(402):387–393, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288854.

Cahoon:1987:RSP

[Cah87] Lawrence S. Cahoon. Repeated sampling in the presence of publication effects: Comment. Journal of the American Statistical Association, 82(398):484–486, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289450.

Calhoun:1985:BRB

[Cal85] Charles A. Calhoun. Book review: [untitled]. Journal of the American Statistical Association, 80(389):234–235, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288081.

Caldwell:1987:BRB

[Cal87] Steven B. Caldwell. Book review: [untitled]. Journal of the American Statistical Association, 82(398):682–683, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289489.

Camero:1984:CSP

[Cam84] Murray A. Camero. Choosing a symmetrizing power transformation. Journal of the American Statistical Association, 79(385): 107–108, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288342.

Campbell:1986:BRB

[Cam86] Richard Campbell. Book review: [untitled]. Journal of the American Statistical Association, 81(393):256–257, March 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288006.

Cannell:1982:BRB

[Can82] Charles F. Cannell. Book review: [untitled]. Journal of the American Statistical Association, 77(380):944, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287347.

Cargill:1980:BRB

[Car80] Thomas F. Cargill. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1037, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287209.

Carroll:1982:PPT

[Car82] Raymond J. Carroll. Prediction and power transformations when the choice of power is restricted to a finite set. Journal of the American Statistical Association, 77(380):908–915, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287326.

Carroll:1983:MAB

[Car83] Raymond J. Carroll. Minimax aspects of bounded-influence regression: Comment. Journal of the American Statistical Association, 78(381):78–79, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287111.

Carlson:1986:BRBa

[Car86a] James Carlson. Book review: [untitled]. Journal of the American Statistical Association, 81(395):863–864, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289033.

Carlson:1986:BRBb

[Car86b] Roger Carlson. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1117–1118, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289097.

Carroll:1987:BRB

[Car87] Raymond J. Carroll. Book review: [untitled]. Journal of the American Statistical Association, 82(399):952, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288821.

Cassel:1982:ESS

[Cas82] C. M. Cassel. Estimation in survey sampling: Robustness and optimality: Comment. Journal of the American Statistical Association, 77(378):404–405, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287259.

Casella:1985:CNM

[Cas85] George Casella. Condition numbers and minimax ridge regression estimators. Journal of the American Statistical Association, 80 (391):753–758, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288496.

Casella:1986:BRB

[Cas86a] George Casella. Book review: [untitled]. Journal of the American Statistical Association, 81(395):859–860, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289026.

Casella:1986:SBE

[Cas86b] George Casella. Stabilizing binomial n estimators. Journal of the American Statistical Association, 81(393):172–175, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287985.

Casella:1987:BRB

[Cas87] George Casella. Book review: [untitled]. Journal of the American Statistical Association, 82(397):348, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289183.

Casella:1988:BRB

[Cas88] George Casella. Book review: [untitled]. Journal of the American Statistical Association, 83(402):567, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288887.

Casella:1989:BRB

[Cas89] George Casella. Book review: [untitled]. Journal of the American Statistical Association, 84(407):844–845, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289690.

Causey:1980:FAC

[Cau80] Beverly D. Causey. A frequentist analysis of a class of ridge regression estimators. Journal of the American Statistical Association, 75(371):736–738, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287680.

Casella:1987:RBF

[CB87a] George Casella and Roger L. Berger. Reconciling Bayesian and frequentist evidence in the one-sided testing problem. With comments and a rejoinder by the authors. Journal of the American Statistical Association, 82(397):106–111, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289130.

Casella:1987:TPN

[CB87b] George Casella and Roger L. Berger. Testing a point null hypothesis: The irreconcilability of P values and evidence: Rejoinder. Journal of the American Statistical Association, 82(397):133– 135, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289138.

Cobb:1983:SSQ

[CC83] E. Benton Cobb and J. D. Church. Small-sample quantal response methods for estimating the location parameter for a location-scale family of dose-response curves. Journal of the American Statistical Association, 78(381):99–107, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287115.

Cressie:1989:SMR

[CC89] Noel Cressie and Ngai H. Chan. Spatial modeling of regional variables. Journal of the American Statistical Association, 84(406):

393–401, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289922.

Causey:1985:ATT

[CCE85] Beverley D. Causey, Lawrence H. Cox, and Lawrence R. Ernst. Applications of transportation theory to statistical problems. Journal of the American Statistical Association, 80(392):903–909, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288551.

Cleveland:1980:CEM

[CD80] William S. Cleveland and Susan J. Devlin. Calendar effects in monthly time series: Detection by spectrum analysis and graphical methods. Journal of the American Statistical Association, 75 (371):487–496, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287636.

Cleveland:1982:CEM

[CD82] William S. Cleveland and Susan J. Devlin. Calendar effects in monthly time series: Modeling and adjustment. Journal of the American Statistical Association, 77(379):520–528, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287705.

Campbell:1989:CEM

[CD89a] M. Karen Campbell and Allan Donner. Classification efficiency of multinomial logistic regression relative to ordinal logistic regression. Journal of the American Statistical Association, 84(406): 587–591, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289946.

Carlin:1989:SASa

[CD89b] J. B. Carlin and A. P. Dempster. Sensitivity analysis of seasonal adjustments: Empirical case studies. Journal of the American Statistical Association, 84(405):6–20, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289837.

Carlin:1989:SASb

[CD89c] J. B. Carlin and A. P. Dempster. Sensitivity analysis of seasonal adjustments: Empirical case studies: Rejoinder. Journal of the American Statistical Association, 84(405):30–32, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289843.

Chernick:1982:DOT

[CDP82] Michael R. Chernick, Darryl J. Downing, and David H. Pike. Detecting outliers in time series data. Journal of the American Statistical Association, 77(380):743–747, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287300.

Carter:1980:IVE

[CF80] R. L. Carter and Wayne A. Fuller. Instrumental variable estimation of the simple errors-in-variables model. Journal of the American Statistical Association, 75(371):687–692, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287670.

Chua:1987:MMR

[CF87a] Tin Chiu Chua and Wayne A. Fuller. A model for multinomial response error applied to labor flows. Journal of the American Statistical Association, 82(397):46–51, March 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289123.

Clayton:1987:NEP

[CF87b] Murray K. Clayton and Edward W. Frees. Nonparametric estimation of the probability of discovering a new species. Journal of the American Statistical Association, 82(397):305–311, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289168.

Crook:1982:PST

[CG82] James Flinn Crook and Irving John Good. The powers and strengths of tests for multinomials and contingency tables. Journal of the American Statistical Association, 77(380):793–802, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287310.

Clogg:1984:LSA

[CG84] Clifford C. Clogg and Leo A. Goodman. Latent structure analysis of a set of multidimensional contingency tables. Journal of the American Statistical Association, 79(388):762–771, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288706.

Cobb:1986:BRB

[CG86] George Cobb and Janice Gifford. Book review: Publications Received. Journal of the American Statistical Association, 81(396): 1129–1131, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289110.

Carroll:1985:CLS

[CGG85] Raymond J. Carroll, Paul Gallo, and Leon Jay Gleser. Comparison of least squares and errors-in-variables regression, with special reference to randomized analysis of covariance. Journal of the American Statistical Association, 80(392):929–932, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288556.

Cohen:1983:HTM

[CGM83] Arthur Cohen, Constantine Gatsonis, and John I. Marden. Hypothesis testing for marginal probabilities in a 2 × 2 × 2 contingency table with conditional independence. Journal of the American Statistical Association, 78(384):920–929, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288205.

Casella:1983:EBC

[CH83] George Casella and Jiunn Tzon Hwang. Empirical Bayes confidence sets for the mean of a multivariate normal distribution. Journal of the American Statistical Association, 78(383):688–698, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288139.

Carroll:1988:ORC

[CH88] Raymond J. Carroll and Peter Hall. Optimal rates of convergence for deconvolving a density. Journal of the American Statistical Association, 83(404):1184–1186, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290153.

Chanda:1980:APC

[Cha80a] Kamal C. Chanda. Asymptotic properties of classification rules based on Wilcoxon-type statistics. Journal of the American Statistical Association, 75(371):726–728, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287677.

Chatterjee:1980:BRB

[Cha80b] Samprit Chatterjee. Book review: [untitled]. Journal of the American Statistical Association, 75(369):247–248, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287433.

Chao:1981:AMS

[Cha81] Anne Chao. Approximate mean squared errors of estimators of reliability for k-out-of-m systems in the independent exponential case. Journal of the American Statistical Association, 76(375): 720–724, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287537.

Chaiken:1984:BRB

[Cha84a] Jan Chaiken. Book review: Indicators of Crime and Criminal Justice: Quantitative Studies, by Stephen E. Fienberg; Albert J. Reiss Jr. Journal of the American Statistical Association, 79 (385):238–239, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288382.

Chanda:1984:BRB

[Cha84b] Anup K. Chanda. Book review: [untitled]. Journal of the American Statistical Association, 79(388):961–962, December 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288763.

Chang:1984:BRB

[Cha84c] Theodore Chang. Book review: The Mathematical Experience, by Philip J. Davis; Reuben Hersh. Journal of the American Statistical Association, 79(385):225–226, March 1984. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288364.

Chang:1984:TML

[Cha84d] Y. C. Chang. A truncated maximum likelihood estimator of a constrained bivariate linear regression coefficient. Journal of the American Statistical Association, 79(386):454–458, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288290.

Chatterjee:1984:BRB

[Cha84e] Samprit Chatterjee. Book review: [untitled]. Journal of the American Statistical Association, 79(388):951–952, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288741.

Chatterjee:1985:BRBa

[Cha85a] Samprit Chatterjee. Book review: [untitled]. Journal of the American Statistical Association, 80(390):495, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287952.

Chatterjee:1985:BRBc

[Cha85b] Samprit Chatterjee. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1082, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288608.

Chatterjee:1985:BRBb

[Cha85c] Sangit Chatterjee. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1078–1079, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288600.

Chaloner:1986:BRB

[Cha86a] Kathryn Chaloner. Book review: [untitled]. Journal of the American Statistical Association, 81(394):572–573, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289261.

Chambers:1986:ORF

[Cha86b] Raymond L. Chambers. Outlier robust finite population estimation. Journal of the American Statistical Association, 81(396):

1063–1069, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289084.

Chatterjee:1986:BRB

[Cha86c] Samprit Chatterjee. Book review: [untitled]. Journal of the American Statistical Association, 81(395):857, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289020.

Chaloner:1987:BRB

[Cha87] Kathryn Chaloner. Book review: [untitled]. Journal of the American Statistical Association, 82(397):342–343, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289173.

Chan:1988:PIN

[Cha88a] Ngai Hang Chan. The parameter inference for nearly nonstationary time series. Journal of the American Statistical Association, 83(403):857–862, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Chang:1988:ERR

[Cha88b] Ted Chang. Estimating the relative rotation of two tectonic plates from boundary crossings. Journal of the American Statistical Association, 83(404):1178–1183, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290152.

Chan:1989:BRB

[Cha89] Wai Chan. Book review: [untitled]. Journal of the American Statistical Association, 84(407):846, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289694.

Chen:1983:STR

[Che83] Chan-Fu Chen. Score tests for regression models. Journal of the American Statistical Association, 78(381):158–161, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287123.

Cherlin:1989:BRB

[Che89] Andrew Cherlin. Book review: [untitled]. Journal of the American Statistical Association, 84(406):631–632, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289983.

Childs:1985:BRB

[Chi85] Lindsay N. Childs. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1072–1073, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288589.

Carr:1989:ARM

[CHK89] Gregory J. Carr, Kerry B. Hafner, and Gary G. Koch. Analysis of rank measures of association for ordinal data from longitudinal studies. Journal of the American Statistical Association, 84(407): 797–804, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289669.

Chen:1982:SSR

[CHL82] Y. Y. Chen, M. Hollander, and N. A. Langberg. Small-sample results for the Kaplan–Meier estimator. Journal of the American Statistical Association, 77(377):141–144, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287780.

Cleveland:1982:JCS

[CHM82] William S. Cleveland, Charles S. Harris, and Robert McGill. Judgments of circle sizes on statistical maps. Journal of the American Statistical Association, 77(379):541–547, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287708.

Christensen:1984:NOL

[Chr84] Ronald Christensen. A note on ordinary least squares methods for two-stage sampling. Journal of the American Statistical Association, 79(387):720–721, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288423.

Christ:1985:BRB

[Chr85] Carl F. Christ. Book review: [untitled]. Journal of the American Statistical Association, 80(391):781–783, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288515.

Christensen:1987:ATS

[Chr87] Ronald Christensen. The analysis of two-stage sampling data by ordinary least squares. Journal of the American Statistical Association, 82(398):492–498, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289452.

Christakos:1988:BRB

[Chr88a] George Christakos. Book review: [untitled]. Journal of the American Statistical Association, 83(401):279–280, March 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288972.

Chromy:1988:BRB

[Chr88b] James P. Chromy. Book review: [untitled]. Journal of the American Statistical Association, 83(401):283, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288977.

Cheng:1988:CTF

[CJ88] Ching-Shui Cheng and Mike Jacroux. The construction of trendfree run orders of two-level factorial designs. Journal of the American Statistical Association, 83(404):1152–1158, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290148.

Chen:1989:QRA

[CK89] James J. Chen and Ralph L. Kodell. Quantitative risk assessment for teratological effects. Journal of the American Statistical Association, 84(408):966–971, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290071.

Cobb:1983:EMR

[CKC83] Loren Cobb, Peter Koppstein, and Neng Hsin Chen. Estimation and moment recursion relations for multimodal distributions of the exponential family. Journal of the American Statistical Association, 78(381):124–130, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287118.

Carroll:1985:NEB

[CL85a] Raymond J. Carroll and F. Lombard. A note on N estimators for the binomial distribution. Journal of the American Statistical Association, 80(390):423–426, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287908.

Carson:1985:SMP

[CL85b] Richard E. Carson and Kenneth Lange. A statistical model for positron emission tomography: Comment. Journal of the American Statistical Association, 80(389):20–22, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288031.

Carter:1986:JFU

[CL86] Lawrence R. Carter and Ronald D. Lee. Joint forecasts of U.S. marital fertility, nuptiality, births, and marriages using time series models. Journal of the American Statistical Association, 81(396): 902–911, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289058.

Clarke:1987:ACL

[Cla87a] G. P. Y. Clarke. Approximate confidence limits for a parameter function in nonlinear regression. Journal of the American Statistical Association, 82(397):221–230, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289158.

Clarke:1987:MCT

[Cla87b] G. P. Y. Clarke. Marginal curvatures and their usefulness in the analysis of nonlinear regression models. Journal of the American Statistical Association, 82(399):844–850, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288795. See correction

[Cla89].

Clayton:1987:BRB

[Cla87c] Murray K. Clayton. Book review: [untitled]. Journal of the American Statistical Association, 82(397):356–357, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289195.

Clarke:1989:CMC

[Cla89] G. P. Y. Clarke. Correction to: “Marginal curvatures and their usefulness in the analysis of nonlinear regression models” [J. Amer. Statist. Assoc. 82 (1987), no. 399, 844–850, MR 88k:62104]. Journal of the American Statistical Association, 84 (405):352, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289913. See [Cla87b].

Cleveland:1982:EHH

[Cle82] William S. Cleveland. Estimating historical heights: Comment. Journal of the American Statistical Association, 77(378):294–295, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287232.

Cleveland:1987:RSG

[Cle87] William S. Cleveland. Research in statistical graphics. Journal of the American Statistical Association, 82(398):419–423, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289443.

Cleveland:1989:SAS

[Cle89] William S. Cleveland. Sensitivity analysis of seasonal adjustments: Empirical case studies: Comment. Journal of the American Statistical Association, 84(405):24–27, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289840.

Carr:1987:SMT

[CLNL87] D. B. Carr, R. J. Littlefield, W. L. Nicholson, and J. S. Littlefield. Scatterplot matrix techniques for large N. Journal of the American Statistical Association, 82(398):424–436, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289444.

Clogg:1980:BRBa

[Clo80a] Clifford C. Clogg. Book review: [untitled]. Journal of the American Statistical Association, 75(370):471–472, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287483.

Clogg:1980:BRBb

[Clo80b] Clifford C. Clogg. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1032, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287201.

Clogg:1982:SMA

[Clo82] Clifford C. Clogg. Some models for the analysis of association in multiway cross-classifications having ordered categories. Journal of the American Statistical Association, 77(380):803–815, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287311.

Clogg:1983:BRBa

[Clo83a] Clifford C. Clogg. Book review: [untitled]. Journal of the American Statistical Association, 78(382):505–506, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288686.

Clogg:1983:BRBb

[Clo83b] Clifford C. Clogg. Book review: [untitled]. Journal of the American Statistical Association, 78(383):737–738, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288158.

Chhikara:1984:LDA

[CM84a] Raj S. Chhikara and Jim McKeon. Linear discriminant analysis with misallocation in training samples. Journal of the American Statistical Association, 79(388):899–906, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288722.

Cleveland:1984:GPT

[CM84b] William S. Cleveland and Robert McGill. Graphical perception: Theory, experimentation, and application to the development of graphical methods. Journal of the American Statistical Association, 79(387):531–554, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288400.

Cleveland:1984:MFS

[CM84c] William S. Cleveland and Robert McGill. The many faces of a scatterplot. Journal of the American Statistical Association, 79 (388):807–822, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288711.

Collings:1985:TGF

[CM85] Bruce J. Collings and Barry H. Margolin. Testing goodness of fit for the Poisson assumption when observations are not identically distributed. Journal of the American Statistical Association, 80 (390):411–418, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287906.

Cowan:1986:CRM

[CM86] Charles D. Cowan and Donald Malec. Capture-recapture models when both sources have clustered observations. Journal of the American Statistical Association, 81(394):347–353, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289223.

Cleveland:1988:SPT

[CMM88] William S. Cleveland, Marylyn E. McGill, and Robert McGill. The shape parameter of a two-variable graph. Journal of the American Statistical Association, 83(402):289–300, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288843.

Chambers:1976:MSS

[CMS76] J. M. Chambers, C. L. Mallows, and B. W. Stuck. A method for simulating stable random variables. Journal of the American Statistical Association, 71(354):340–344, June 1976. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2285309. See correction [CMS87].

Chambers:1987:CMS

[CMS87] J. M. Chambers, C. L. Mallows, and B. W. Stuck. Correction to: “A method for simulating stable random variables” [J. Amer. Statist. Assoc. 71 (1976), no. 354, 340–344, MR 54 #4059]. Journal of the American Statistical Association, 82(398):704, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289515. See

[CMS76].

Cheng:1988:OST

[CMST88] Ching-Shui Cheng, Dibyen Majumdar, John Stufken, and Tahsin Erkan Türe. Optimal step-type designs for comparing test treatments with a control. Journal of the American Statistical Association, 83(402):477–482, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288865.

Cobb:1982:BRB

[Cob82] E. Benton Cobb. Book review: [untitled]. Journal of the American Statistical Association, 77(379):687–688, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287742.

Cobb:1984:BRBa

[Cob84a] George W. Cobb. Book review: [untitled]. Journal of the American Statistical Association, 79(386):469–470, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288295.

Cobb:1984:BRBb

[Cob84b] George W. Cobb. Book review: [untitled]. Journal of the American Statistical Association, 79(386):472, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288298.

Cobb:1987:BRBb

[Cob87a] George Cobb. Book review: [untitled]. Journal of the American Statistical Association, 82(399):954, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288826.

Cobb:1987:BRBa

[Cob87b] George W. Cobb. Book review: Introductory Textbooks: A Framework for Evaluation. Journal of the American Statistical Association, 82(397):321–339, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289170.

Cohen:1981:IMM

[Coh81] Arthur Cohen. Inference for marginal means in contingency tables with conditional independence. Journal of the American Statistical Association, 76(376):895–902, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287585.

Cohen:1982:BRB

[Coh82] A. Clifford Cohen. Book review: [untitled]. Journal of the American Statistical Association, 77(378):497–498, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287285.

Cohen:1989:WFS

[Coh89] Joel E. Cohen. The world fertility survey: An appraisal of methodology: Comment. Journal of the American Statistical Association, 84(407):772–774, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289663.

Collings:1987:CRN

[Col87] Bruce Jay Collings. Compound random number generators. Journal of the American Statistical Association, 82(398):525–527, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289456.

Conaway:1989:ARC

[Con89a] Mark R. Conaway. Analysis of repeated categorical measurements with conditional likelihood methods. Journal of the American Statistical Association, 84(405):53–62, March 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289846.

Conway:1989:BRB

[Con89b] Delores A. Conway. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1095–1096, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290094.

Cooil:1983:BRB

[Coo83] Bruce Cooil. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1001–1002, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288238.

Cordray:1988:BRB

[Cor88a] David S. Cordray. Book review: [untitled]. Journal of the American Statistical Association, 83(402):571–572, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288893.

Cornell:1988:BRBa

[Cor88b] Richard G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 83(402):575–576, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288900.

Cornell:1988:BRBb

[Cor88c] Richard G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 83(402):578, June 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288904.

Cornell:1989:BRBb

[Cor89a] Richard G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 84(407):838–839, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289680.

Cornell:1989:BRBc

[Cor89b] Richard G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1097, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290097.

Cornell:1989:BRBd

[Cor89c] Richard G. Cornell. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1097–1098, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290098.

Costanza:1982:BRBa

[Cos82a] Michael C. Costanza. Book review: [untitled]. Journal of the American Statistical Association, 77(379):694–695, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287756.

Costanza:1982:BRBb

[Cos82b] Michael C. Costanza. Book review: [untitled]. Journal of the American Statistical Association, 77(380):947–948, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287351.

Costanza:1983:BRB

[Cos83] Michael C. Costanza. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1000–1001, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288236.

Coveyou:1983:BRB

[Cov83] Michael R. Coveyou. Book review: [untitled]. Journal of the American Statistical Association, 78(382):498–499, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288672.

Coveyou:1985:BRB

[Cov85] Michael R. Coveyou. Book review: [untitled]. Journal of the American Statistical Association, 80(390):492–493, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287947.

Cover:1986:BRB

[Cov86] Thomas M. Cover. Book review: [untitled]. Journal of the American Statistical Association, 81(393):265–266, March 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288023.

Cowan:1982:BRB

[Cow82] Charles D. Cowan. Book review: [untitled]. Journal of the American Statistical Association, 77(379):686–687, September 1982.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287741.

Cox:1980:SMS

[Cox80a] Lawrence H. Cox. Suppression methodology and statistical disclosure control. Journal of the American Statistical Association, 75(370):377–385, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287463.

Cox:1980:BRB

[Cox80b] Philip Cox. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1040, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287215.

Cox:1982:SAS

[Cox82] D. R. Cox. Some ancillary statistics and their properties: Comment. Journal of the American Statistical Association, 77(379): 591, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287716.

Cox:1985:BRB

[Cox85] D. F. Cox. Book review: [untitled]. Journal of the American Statistical Association, 80(390):486, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287932.

Cox:1986:SCI

[Cox86a] D. R. Cox. Statistics and causal inference: Comment. Journal of the American Statistical Association, 81(396):963–964, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289066.

Cox:1986:DLD

[Cox86b] Lawrence H. Cox. Disclosure-limited data dissemination: Comment. Journal of the American Statistical Association, 81(393): 19–21, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287960.

Cox:1987:BBC

[Cox87a] D. R. Cox. Better bootstrap confidence intervals: Comment. Journal of the American Statistical Association, 82(397):190, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289148.

Cox:1987:CPU

[Cox87b] Lawrence H. Cox. A constructive procedure for unbiased controlled rounding. Journal of the American Statistical Association, 82(398):520–524, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289455.

Carroll:1982:CBM

[CR82] R. J. Carroll and David Ruppert. A comparison between maximum likelihood and generalized least squares in a heteroscedastic linear model. Journal of the American Statistical Association, 77 (380):878–882, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287321.

Carroll:1984:ATD

[CR84a] Raymond J. Carroll and David Ruppert. The analysis of transformed data: Comment. Journal of the American Statistical Association, 79(386):312–313, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288266.

Carroll:1984:PTW

[CR84b] Raymond J. Carroll and David Ruppert. Power transformations when fitting theoretical models to data. Journal of the American Statistical Association, 79(386):321–328, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288271.

Chi:1989:MLD

[CR89] Eric M. Chi and Gregory C. Reinsel. Models for longitudinal data with random effects and ar(1) errors. Journal of the American Statistical Association, 84(406):452–459, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289929.

Cressie:1982:PSM

[Cre82] Noel Cressie. Playing safe with misweighted means. Journal of the American Statistical Association, 77(380):754–759, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287302.

Cressie:1984:BRB

[Cre84] Noel A. C. Cressie. Book review: [untitled]. Journal of the American Statistical Association, 79(387):733, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288434.

Cressie:1986:KND

[Cre86] Noel Cressie. Kriging nonstationary data. Journal of the American Statistical Association, 81(395):625–634, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288990.

Cressie:1988:GPD

[Cre88] Noel Cressie. A graphical procedure for determining nonstationarity in time series. Journal of the American Statistical Association, 83(404):1108–1116, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290143.

Cressie:1989:EBE

[Cre89] Noel Cressie. Empirical Bayes estimation of undercount in the decennial census. Journal of the American Statistical Association, 84 (408):1033–1044, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290080.

Crockett:1985:ADH

[Cro85] Patrick W. Crockett. Asymptotic distribution of the Hildreth-Houck estimator. Journal of the American Statistical Association, 80(389):202–204, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288072.

Casella:1980:CBL

[CS80] George Casella and William E. Strawderman. Confidence bands for linear regression with restricted predictor variable. Journal of the American Statistical Association, 75(372):862–868, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287173.

Crowley:1983:RSH

[CS83] John Crowley and Barry E. Storer. A reanalysis of the Stanford heart transplant data: Comment. Journal of the American Statistical Association, 78(382):277–281, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288624.

Campbell:1985:NSM

[CS85] Gregory Campbell and John H. Skillings. Nonparametric stepwise multiple comparison procedures. Journal of the American Statistical Association, 80(392):998–1003, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288566.

Chiu:1986:BPI

[CS86] H. Y. Chiu and J. Sedransk. A Bayesian procedure for imputing missing values in sample surveys. Journal of the American Statistical Association, 81(395):667–676, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288994.

Cohen:1987:AIF

[CS87] Arthur Cohen and H. B. Sackrowitz. An approach to inference following model selection with applications to transformation-based and adaptive inference. Journal of the American Statistical Association, 82(400):1123–1130, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Cohen:1989:ETR

[CS89] Arthur Cohen and H. B. Sackrowitz. Exact tests that recover interblock information in balanced incomplete blocks designs. Journal of the American Statistical Association, 84(406):556–559, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289942.

Czaja:1986:RBS

[CSC86] Ronald F. Czaja, Cecelia B. Snowden, and Robert J. Casady. Reporting bias and sampling errors in a survey of a rare population using multiplicity counting rules. Journal of the American Statistical Association, 81(394):411–419, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289230.

Chandra:1981:KST

[CSS81] M. Chandra, N. D. Singpurwalla, and M. A. Stephens. Kolmogorov statistics for tests of fit for the extreme value and Weibull distributions. Journal of the American Statistical Association, 76 (375):729–731, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287539.

Chandra:1986:CKS

[CSS86] M. Chandra, N. D. Singpurwalla, and M. A. Stephens. Corrections: Kolmogorov statistics for tests of fit for the extreme-value and Weibull distributions. Journal of the American Statistical Association, 81(396):1132, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289111.

Chinchilli:1989:IBR

[CSS89] Vernon M. Chinchilli, Barry H. Schwab, and Pranab K. Sen. Inference based on ranks for the multiple-design multivariate linear model. Journal of the American Statistical Association, 84(406): 517–524, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289937.

Cook:1980:MRO

[CT80] R. Dennis Cook and L. A. Thibodeau. Marginally restricted Doptimal designs. Journal of the American Statistical Association, 75(370):366–371, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287460.

Cleveland:1982:GMS

[CT82] William S. Cleveland and Irma J. Terpenning. Graphical methods for seasonal adjustment. Journal of the American Statistical Association, 77(377):52–62, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287769.

Cleland:1989:WFS

[CV89] J. Cleland and V. Verma. The world fertility survey: An appraisal of methodology. Journal of the American Statistical Association, 84(407):756–767, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289659.

Cohen:1980:ETP

[CW80] A. Clifford Cohen and Betty Jones Whitten. Estimation in the three-parameter lognormal distribution. Journal of the American Statistical Association, 75(370):399–404, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287466.

Cook:1983:MAB

[CW83] R. Dennis Cook and Sanford Weisberg. Minimax aspects of bounded-influence regression: Comment. Journal of the American Statistical Association, 78(381):74–75, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287108.

Cook:1985:NPE

[CW85] R. Dennis Cook and Jeffrey A. Witmer. A note on parametereffects curvature. Journal of the American Statistical Association, 80(392):872–878, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288546.

Cornell:1989:BRBa

[CW89] Richard G. Cornell and Fredric M. Wolf. Book review: [untitled]. Journal of the American Statistical Association, 84(406):629–631, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289981.

Carroll:1988:EEW

[CWR88] Raymond J. Carroll, C.-F. Jeff Wu, and David Ruppert. The effect of estimating weights in weighted least squares. Journal of the American Statistical Association, 83(404):1045–1054, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290134.

Cobb:1985:ACT

[CZ85] Loren Cobb and Shelemyahu Zacks. Applications of catastrophe theory for statistical modeling in the biosciences. Journal of the American Statistical Association, 80(392):793–802, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288534.

Dagum:1982:EAF

[Dag82] Estela Bee Dagum. The effects of asymmetric filters on seasonal factor revisions. Journal of the American Statistical Association, 77(380):732–738, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287298.

deAmorim:1986:EDE

[dAJ86] Sebastiao de Amorim and Richard A. Johnson. Experimental designs for estimating the correlation between two destructively tested variables. Journal of the American Statistical Association, 81(395):807–812, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289013.

Dallal:1982:SIM

[Dal82] Gerard E. Dallal. A simple interaction model for two-dimensional contingency tables. Journal of the American Statistical Association, 77(378):425–432, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287263.

Dalenius:1983:BRB

[Dal83a] Tore Dalenius. Book review: [untitled]. Journal of the American Statistical Association, 78(384):995, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288228.

Dalenius:1983:EMD

[Dal83b] Tore Dalenius. An evaluation of model-dependent and probabilitysampling inferences in sample surveys: Comment. Journal of the American Statistical Association, 78(384):799–800, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288185.

Dalenius:1984:BRB

[Dal84] Tore Dalenius. Book review: [untitled]. Journal of the American Statistical Association, 79(387):725, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288426.

Danaher:1989:MMM

[Dan89] Peter J. Danaher. A Markov mixture model for magazine exposure. Journal of the American Statistical Association, 84(408): 922–926, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290067.

Darmstadter:1980:BRB

[Dar80] Joel Darmstadter. Book review: [untitled]. Journal of the American Statistical Association, 75(371):761, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287697.

Davis:1981:ESP

[Dav81] Kathryn Bullock Davis. Estimation of the scaling parameter for a kernel type density estimate. Journal of the American Statistical Association, 76(375):632–636, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287522.

Davis:1982:EMM

[Dav82] A. W. Davis. On the effects of moderate multivariate nonnormality on Roy’s largest root test. Journal of the American Statistical Association, 77(380):896–900, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287324.

David:1984:BRBb

[Dav84a] H. A. David. Book review: [untitled]. Journal of the American Statistical Association, 79(387):728–729, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288429.

David:1984:BRBa

[Dav84b] H. T. David. Book review: [untitled]. Journal of the American Statistical Association, 79(386):470–471, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288296.

Davis:1984:BRBb

[Dav84c] Richard A. Davis. Book review: [untitled]. Journal of the American Statistical Association, 79(386):484, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288320.

Davis:1984:BRBa

[Dav84d] William W. Davis. Book review: [untitled]. Journal of the American Statistical Association, 79(386):483–484, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288319.

Davis:1984:BRBc

[Dav84e] William W. Davis. Book review: [untitled]. Journal of the American Statistical Association, 79(387):739, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288447.

Davis:1985:BRB

[Dav85a] Richard A. Davis. Book review: [untitled]. Journal of the American Statistical Association, 80(389):251, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288102.

Davis:1985:DGL

[Dav85b] William W. Davis. Dynamic generalized linear models and Bayesian forecasting: Comment. Journal of the American Statistical Association, 80(389):84–85, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288043.

Dawid:1982:WCBb

[Daw82a] A. P. Dawid. The well-calibrated Bayesian: Rejoinder. Journal of the American Statistical Association, 77(379):612–613, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287723.

Dawid:1982:WCBa

[Daw82b] A. P. Dawid. The well-calibrated Bayesian. With comments by Joseph B. Kadane and D. V. Lindley and a rejoinder by the author. Journal of the American Statistical Association, 77(379): 605–610, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287720.

Dawid:1984:NDS

[Daw84] A. P. Dawid. On the nature and discovery of structure: Comment. Journal of the American Statistical Association, 79(385): 22–24, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288327.

Dawid:1985:SCP

[Daw85] A. P. Dawid. Self-calibrating priors do not exist: Comment. Journal of the American Statistical Association, 80(390):340–341, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287892.

Dey:1983:TSE

[DB83] Dipak K. Dey and James O. Berger. On truncation of shrinkage estimators in simultaneous estimation of normal means. Journal of the American Statistical Association, 78(384):865–869, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288197.

deBeer:1985:TSM

[dB85] Joop de Beer. A time series model for cohort data. Journal of the American Statistical Association, 80(391):525–530, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288465.

deCarvalho:1981:BRB

[dC81] Jose Ferreira de Carvalho. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1012, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287608.

Dempster:1985:DGL

[DC85] A. P. Dempster and J. B. Carlin. Dynamic generalized linear models and Bayesian forecasting: Comment. Journal of the American Statistical Association, 80(389):85–88, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288044.

Davidian:1987:VFE

[DC87] M. Davidian and R. J. Carroll. Variance function estimation. Journal of the American Statistical Association, 82(400):1079– 1091, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dhariyal:1981:OSF

[DD81] Ishwari D. Dhariyal and Edward J. Dudewicz. Optimal selection from a finite sequence with sampling cost. Journal of the American Statistical Association, 76(376):952–959, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287595. See corrigendum

[DD82].

Dhariyal:1982:COS

[DD82] Ishwari D. Dhariyal and Edward J. Dudewicz. Corrigendum: “Optimal selection from a finite sequence with sampling cost” [J. Amer. Statist. Assoc. 76 (1981), no. 376, 954–959, MR 83e:62111]. Journal of the American Statistical Association, 77 (380):955, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287367. See [DD81].

DuMouchel:1983:USS

[DD83] William H. DuMouchel and Greg J. Duncan. Using sample survey weights in multiple regression analyses of stratified samples. Journal of the American Statistical Association, 78(383):535–543, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288115.

Dabrowska:1988:ETT

[DD88] Dorota M. Da̧browska and Kjell A. Doksum. Estimation and testing in a two-sample generalized odds-rate model. Journal of the American Statistical Association, 83(403):744–749, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deHaan:1981:EMF

[de 81] Laurens de Haan. Estimation of the minimum of a function using order statistics. Journal of the American Statistical Association, 76(374):467–469, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2287851. See corrigenda [de 83a].

deHaan:1983:CEM

[de 83a] Laurens de Haan. Corrigenda: “Estimation of the minimum of a function using order statistics” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 467–469, MR 83a:62103]. Journal of the American Statistical Association, 78(384):1008, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288249. See [de 81].

DeRiggi:1983:ULF

[De 83b] Dennis F. De Riggi. Unimodality of likelihood functions for the binomial distribution. Journal of the American Statistical Association, 78(381):181–183, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287127.

Dean:1989:BRBa

[Dea89a] Angela Dean. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1102–1103, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290107.

Dean:1989:BRBb

[Dea89b] Angela Dean. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1103, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290108.

Deaton:1989:BRB

[Dea89c] Michael L. Deaton. Book review: [untitled]. Journal of the American Statistical Association, 84(405):332–333, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289884.

DeGroot:1982:LPC

[DeG82] Morris H. DeGroot. Lindley’s paradox: Comment. Journal of the American Statistical Association, 77(378):336–339, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287246.

DeGroot:1983:BRB

[DeG83] Morris H. DeGroot. Book review: [untitled]. Journal of the American Statistical Association, 78(381):200, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287130.

Delampady:1989:LBB

[Del89] Mohan Delampady. Lower bounds on Bayes factors for interval null hypotheses. Journal of the American Statistical Association, 84(405):120–124, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289854.

Dempster:1980:UEB

[Dem80] A. P. Dempster. Using empirical Bayes techniques in the law school validity studies: Comment. Journal of the American Statistical Association, 75(372):817, December 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287163.

Dempster:1982:LPC

[Dem82] A. P. Dempster. Lindley’s paradox: Comment. Journal of the American Statistical Association, 77(378):339–341, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287247.

Dempster:1983:PEB

[Dem83] A. P. Dempster. Parametric empirical Bayes inference: Theory and applications: Comment. Journal of the American Statistical Association, 78(381):57–58, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287100.

Dempster:1987:CPD

[Dem87] A. P. Dempster. The calculation of posterior distributions by data augmentation: Comment. Journal of the American Statistical Association, 82(398):541, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289458.

Dent:1984:BRB

[Den84] Richard J. Dent. Book review: [untitled]. Journal of the American Statistical Association, 79(387):742, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288453.

DerSimonian:1987:EBC

[Der87] Rebecca DerSimonian. Empirical Bayes confidence intervals based on bootstrap samples: Comment. Journal of the American Statistical Association, 82(399):755–756, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288782.

Devroye:1980:BRB

[Dev80] Luc Devroye. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1040–1041, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287216.

Devault:1984:BRB

[Dev84] Arthur R. Devault. Book review: [untitled]. Journal of the American Statistical Association, 79(386):482–483, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288317.

Dickie:1987:MTH

[DFG87] Mark Dickie, Ann Fisher, and Shelby Gerking. Market transactions and hypothetical demand data: a comparative study. Journal of the American Statistical Association, 82(397):69–75, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289126.

Dalal:1989:RAS

[DFH89] Siddhartha R. Dalal, Edward B. Fowlkes, and Bruce Hoadley. Risk analysis of the Space Shuttle: Pre-Challenger prediction of failure. Journal of the American Statistical Association, 84(408): 945–957, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290069.

Diggle:1985:RSS

[DG85] Peter J. Diggle and David J. Gates. Regularity and structure of the spatial pattern of blue cones of macaque retina: Comment. Journal of the American Statistical Association, 80(392): 813, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288536.

Draper:1986:RSD

[DG86] N. R. Draper and I. Guttman. Response surface designs in flexible regions. Journal of the American Statistical Association, 81(396): 1089–1094, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289088.

deGraft-Johnson:1989:WFS

[dGJ89] Kweku T. de Graft-Johnson. The world fertility survey: An appraisal of methodology: Comment. Journal of the American Statistical Association, 84(407):769–770, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289661.

Devlin:1981:RED

[DGK81] S. J. Devlin, R. Gnandesikan, and J. R. Kettenring. Robust estimation of dispersion matrices and principal components. Journal of the American Statistical Association, 76(374):354–362, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287836.

Dallal:1980:NTM

[DH80] G. E. Dallal and J. A. Hartigan. Note on a test of monotone association insensitive to outliers. Journal of the American Statistical Association, 75(371):722–725, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287676.

DuMouchel:1983:BMCb

[DH83a] William H. DuMouchel and Jeffrey E. Harris. Bayes methods for combining the results of cancer studies in humans and other species: Rejoinder. Journal of the American Statistical Association, 78(382):313–315, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288635.

DuMouchel:1983:BMCa

[DH83b] William H. DuMouchel and Jeffrey E. Harris. Bayes methods for combining the results of cancer studies in humans and other species. With comments by Daniel Krewski, Adrian F. M. Smith and Robert E. Kass, and with a reply by the authors. Journal of the American Statistical Association, 78(382):293–308, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288631.

Diggle:1986:STO

[DH86] Peter J. Diggle and Peter Hall. The selection of terms in an orthogonal series density estimator. Journal of the American Statistical Association, 81(393):230–233, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287995.

Dickey:1984:TUR

[DHF84] D. A. Dickey, D. P. Hasza, and W. A. Fuller. Testing for unit roots in seasonal time series. Journal of the American Statistical Association, 79(386):355–367, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288276.

Dickey:1981:BRB

[Dic81] David A. Dickey. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1016–1017, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287620.

Dickey:1982:BRBa

[Dic82a] David A. Dickey. Book review: [untitled]. Journal of the American Statistical Association, 77(377):214, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287796.

Dickey:1982:BRBb

[Dic82b] David A. Dickey. Book review: [untitled]. Journal of the American Statistical Association, 77(380):941, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287342.

Dickey:1983:BRB

[Dic83a] David A. Dickey. Book review: [untitled]. Journal of the American Statistical Association, 78(382):510, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288696.

Dickey:1983:MHF

[Dic83b] James M. Dickey. Multiple hypergeometric functions: probabilistic interpretations and statistical uses. Journal of the American Statistical Association, 78(383):628–637, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288131.

Dickey:1984:BRB

[Dic84] David A. Dickey. Book review: Applied Time Series Analysis, by David F. Findley. Journal of the American Statistical Association, 79(385):234, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288375.

Dickey:1987:TPN

[Dic87] James M. Dickey. Testing a point null hypothesis: The irreconcilability of P values and evidence: Comment. Journal of the American Statistical Association, 82(397):129–130, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289135.

Dietz:1982:BNT

[Die82] E. Jacquelin Dietz. Bivariate nonparametric tests for the onesample location problem. Journal of the American Statistical Association, 77(377):163–169, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287784.

Diebold:1989:BRB

[Die89] Francis X. Diebold. Book review: [untitled]. Journal of the American Statistical Association, 84(405):339–340, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289895.

Dinse:1986:NPM

[Din86] Gregg E. Dinse. Nonparametric prevalence and mortality estimators for animal experiments with incomplete cause-of-death data. Journal of the American Statistical Association, 81(394):328–336, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289220.

deJong:1989:SIS

[dJ89] Piet de Jong. Smoothing and interpolation with the state-space model. Journal of the American Statistical Association, 84(408): 1085–1088, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290087.

deJongh:1988:MTB

[dJdWW88] P. J. de Jongh, T. de Wet, and A. H. Welsh. Mallows-type bounded-influence-regression trimmed means. Journal of the American Statistical Association, 83(403):805–810, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

deJong:1983:TRP

[dJGM83] Piet de Jong, Malcolm Greig, and Dilip Madan. Testing for random pairing. Journal of the American Statistical Association, 78 (382):332–336, ???? 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dickey:1987:BMC

[DJK87] James M. Dickey, Jhy-Ming Jiang, and Joseph B. Kadane. Bayesian methods for censored categorical data. Journal of the American Statistical Association, 82(399):773–781, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288786.

Dyer:1980:DCV

[DK80] Danny D. Dyer and Jerome P. Keating. On the determination of critical values for Bartlett’s test. Journal of the American Statistical Association, 75(370):313–319, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287450.

Dietz:1981:NMT

[DK81] E. Jacquelin Dietz and Timothy J. Killeen. A nonparametric multivariate test for monotone trend with pharmaceutical applications. Journal of the American Statistical Association, 76(373): 169–174, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287063.

Drost:1989:PAM

[DKMO89] F. C. Drost, W. C. M. Kallenberg, D. S. Moore, and J. Oosterhoff. Power approximations to multinomial tests of fit. Journal of the American Statistical Association, 84(405):130–141, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289856.

Dunn:1980:ACE

[DL80] Douglas M. Dunn and James M. Landwehr. Analyzing clustering effects across time. Journal of the American Statistical Association, 75(369):8–15, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287372.

Deely:1981:BEB

[DL81] J. J. Deely and D. V. Lindley. Bayes empirical Bayes. Journal of the American Statistical Association, 76(376):833–841, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287578.

deLeeuw:1986:BRB

[dL86a] Jan de Leeuw. Book review: [untitled]. Journal of the American Statistical Association, 81(394):580, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289276.

Duncan:1986:DLDa

[DL86b] George T. Duncan and Diane Lambert. Disclosure-limited data dissemination. Journal of the American Statistical Association, 81(393):10–18, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287959.

Duncan:1986:DLDb

[DL86c] George T. Duncan and Diane Lambert. Disclosure-limited data dissemination: Rejoinder. Journal of the American Statistical Association, 81(393):27–28, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287964.

Dykstra:1988:DPM

[DL88] Richard L. Dykstra and Jon H. Lemke. Duality of I projections and maximum likelihood estimation for log-linear models under cone constraints. Journal of the American Statistical Association, 83(402):546–554, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288876.

Dean:1989:TDO

[DL89] C. Dean and J. F. Lawless. Tests for detecting overdispersion in Poisson regression models. Journal of the American Statistical Association, 84(406):467–472, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289931.

David:1986:AMC

[DLST86] Martin David, Roderick J. A. Little, Michael E. Samuhel, and Robert K. Triest. Alternative methods for CPS income imputation. Journal of the American Statistical Association, 81(393): 29–41, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287965.

Dalal:1988:WSO

[DM88a] S. R. Dalal and C. L. Mallows. When should one stop testing software. Journal of the American Statistical Association, 83(403): 872–879, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dayton:1988:CVL

[DM88b] C. Mitchell Dayton and George B. Macready. Concomitantvariable latent-class models. Journal of the American Statistical Association, 83(401):173–178, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288938.

Diggle:1988:ESP

[DM88c] Peter Diggle and J. S. Marron. Equivalence of smoothing parameter selectors in density and intensity estimation. Journal of the American Statistical Association, 83(403):793–800, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Diaconis:1989:MSC

[DM89] Persi Diaconis and Frederick Mosteller. Methods for studying coincidences. Journal of the American Statistical Association, 84 (408):853–861, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290058.

Doksum:1984:ATD

[Dok84] Kjell A. Doksum. The analysis of transformed data: Comment. Journal of the American Statistical Association, 79(386):316–319, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288269.

Domowitz:1988:BRB

[Dom88] Ian Domowitz. Book review: [untitled]. Journal of the American Statistical Association, 83(401):280–281, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288974.

Donnelly:1987:BRB

[Don87] Peter Donnelly. Book review: [untitled]. Journal of the American Statistical Association, 82(399):948–949, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288815.

Davies:1984:UGP

[DP84] Neville Davies and Joseph D. Petruccelli. On the use of the general partial autocorrelation function for order determination in processes on the use of the general partial autocorrelation function for order determination in processes. Journal of the American Statistical Association, 79(386):374–377, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288278.

Dielman:1989:EOL

[DP89] Terry E. Dielman and Roger C. Pfaffenberger. Efficiency of ordinary least squares for linear models with autocorrelation. Journal of the American Statistical Association, 84(405):248, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289870.

Dunsmuir:1981:ETS

[DR81] W. Dunsmuir and P. M. Robinson. Estimation of time series models in the presence of missing data. Journal of the American Statistical Association, 76(375):560–568, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287513.

Dykstra:1983:TMT

[DR83] Richard L. Dykstra and Tim Robertson. On testing monotone tendencies. Journal of the American Statistical Association, 78 (382):342–350, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288639.

Dempster:1985:WNP

[DR85] A. P. Dempster and Louise M. Ryan. Weighted normal plots. Journal of the American Statistical Association, 80(392):845–850, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288543.

Draper:1987:BRB

[Dra87] David Draper. Book review: [untitled]. Journal of the American Statistical Association, 82(397):349–350, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289185.

Dempster:1981:ECC

[DRT81] A. P. Dempster, D. B. Rubin, and R. K. Tsutakawa. Estimation in covariance components models. Journal of the American Statistical Association, 76(374):341–353, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287835.

Duncan:1980:EQW

[DS80] Otis Dudley Duncan and Howard Schuman. Effects of question wording and context: An experiment with religious indicators. Journal of the American Statistical Association, 75(370):269–275, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287445.

Damsleth:1982:ETC

[DS82] E. Damsleth and E. Spjotvoll. Estimation of trigonometric components in time series. Journal of the American Statistical Association, 77(378):381–387, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287255.

Dawson-Saunders:1987:BRB

[DS87] Beth Dawson-Saunders. Book review: [untitled]. Journal of the American Statistical Association, 82(397):359–360, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289199.

Dempster:1983:CHR

[DSW83] Arthur P. Dempster, Murray R. Selwyn, and Barbara J. Weeks. Combining historical and randomized controls for assessing trends in proportions. Journal of the American Statistical Association, 78(382):221–227, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288616.

Dwyer:1980:EEP

[DT80] Paul S. Dwyer and Derrick S. Tracy. Expectation and estimation of product moments in sampling from a finite population. Journal of the American Statistical Association, 75(370):431–437, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287472.

DiCiccio:1987:BBC

[DT87a] Thomas DiCiccio and Robert Tibshirani. Better bootstrap confidence intervals: Comment. Journal of the American Statistical Association, 82(397):187–188, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289146.

DiCiccio:1987:BCI

[DT87b] Thomas DiCiccio and Robert Tibshirani. Bootstrap confidence intervals and bootstrap approximations. Journal of the American Statistical Association, 82(397):163–170, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289143.

Duan:1983:SEN

[Dua83] Naihua Duan. Smearing estimate: a nonparametric retransformation method. Journal of the American Statistical Association, 78 (383):605–610, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288126.

Dubin:1984:BRB

[Dub84] Neil Dubin. Book review: [untitled]. Journal of the American Statistical Association, 79(388):959–960, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288758.

Duffua:1982:BRB

[Duf82] Salih O. Duffua. Book review: [untitled]. Journal of the American Statistical Association, 77(379):688, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287745.

Dunnett:1980:PMCa

[Dun80a] Charles W. Dunnett. Pairwise multiple comparisons in the homogeneous variance, unequal sample size case. Journal of the American Statistical Association, 75(372):789–795, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287160.

Dunnett:1980:PMCb

[Dun80b] Charles W. Dunnett. Pairwise multiple comparisons in the unequal variance case. Journal of the American Statistical Association, 75(372):796–800, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287161.

Duncan:1981:BRB

[Dun81] George T. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 76(373):214, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287085.

Duncan:1985:BRBa

[Dun85a] George Duncan. Book review: [untitled]. Journal of the American Statistical Association, 80(390):483, June 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287926.

Duncan:1985:BRBb

[Dun85b] Otis Dudley Duncan. Book review: [untitled]. Journal of the American Statistical Association, 80(391):776–777, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288507.

Duncan:1987:BRB

[Dun87] Greg J. Duncan. Book review: [untitled]. Journal of the American Statistical Association, 82(399):959–960, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288834.

Dudewicz:1981:EBT

[DvdM81] Edward J. Dudewicz and Edward C. van der Meulen. Entropybased tests of uniformity. Journal of the American Statistical Association, 76(376):967–974, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287597.

Doksum:1983:STB

[DW83] Kjell A. Doksum and Chi-Wing Wong. Statistical tests based on transformed data. Journal of the American Statistical Association, 78(382):411–417, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288649.

Deng:1987:EVR

[DW87] Lih-Yuan Deng and C. F. J. Wu. Estimation of variance of the regression estimator. Journal of the American Statistical Association, 82(398):568–576, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289466.

DeGruttola:1987:IAG

[DWL87] Victor De Gruttola, James H. Ware, and Thomas A. Louis. Influence analysis of generalized least squares estimators. Journal of the American Statistical Association, 82(399):911–917, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Dyer:1982:CGD

[Dye82] Danny Dyer. The convolution of generalized F distributions. Journal of the American Statistical Association, 77(377):184– 189, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287787.

Dykstra:1980:PII

[Dyk80] R. L. Dykstra. Product inequalities involving the multivariate normal distribution. Journal of the American Statistical Association, 75(371):646–650, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287662.

Dykstra:1982:MLE

[Dyk82] Richard L. Dykstra. Maximum likelihood estimation of the survival functions of stochastically ordered random variables. Journal of the American Statistical Association, 77(379):621–628, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287725.

Dykstra:1983:ARL

[Dyk83] Richard L. Dykstra. An algorithm for restricted least squares regression. Journal of the American Statistical Association, 78 (384):837–842, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288193.

Dykstra:1988:BRB

[Dyk88] Richard Dykstra. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1219, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290169.

Dykstra:1989:BRBa

[Dyk89a] Richard Dykstra. Book review: [untitled]. Journal of the American Statistical Association, 84(405):338–339, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289894.

Dykstra:1989:BRBb

[Dyk89b] Richard Dykstra. Book review: [untitled]. Journal of the American Statistical Association, 84(406):627, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289975.

Diaconis:1982:USP

[DZ82] Persi Diaconis and Sandy L. Zabell. Updating subjective probability. Journal of the American Statistical Association, 77(380):

822–830, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287313.

Diggle:1989:NGM

[DZ89] Peter J. Diggle and Scott L. Zeger. A non-Gaussian model for time series with pulses. Journal of the American Statistical Association, 84(406):354–359, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289918.

Easton:1980:BRB

[Eas80] Stephen T. Easton. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1035–1036, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287206.

Eaton:1982:BRB

[Eat82] Morris L. Eaton. Book review: [untitled]. Journal of the American Statistical Association, 77(379):690, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287749.

Eaton:1983:BRB

[Eat83] Morris L. Eaton. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1004, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288243.

Eccleston:1985:BRB

[Ecc85] John A. Eccleston. Book review: [untitled]. Journal of the American Statistical Association, 80(391):785, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288520.

Eddy:1980:DEB

[Edd80] William F. Eddy. Density estimation and bump-hunting by the penalized likelihood method exemplified by scattering and meteorite data: Comment. Journal of the American Statistical Association, 75(369):59–61, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287379.

Eddy:1981:GMT

[Edd81] William F. Eddy. Graphics for the multivariate two-sample problem: Comment. Journal of the American Statistical Association, 76(374):287–289, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287826.

Eddy:1982:BRB

[Edd82] William F. Eddy. Book review: [untitled]. Journal of the American Statistical Association, 77(379):692, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287753.

Eddy:1983:BRB

[Edd83] William F. Eddy. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1002, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288239.

Eddy:1984:BRB

[Edd84] William F. Eddy. Book review: [untitled]. Journal of the American Statistical Association, 79(386):483, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288318.

Eddy:1988:I

[Edd88] William F. Eddy. [introduction]. Journal of the American Statistical Association, 83(404):975, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290122.

Efron:1981:CDB

[Efr81] Bradley Efron. Censored data and the bootstrap. Journal of the American Statistical Association, 76(374):312–319, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287832.

Efron:1983:EER

[Efr83] Bradley Efron. Estimating the error rate of a prediction rule: improvement on cross-validation. Journal of the American Statistical Association, 78(382):316–331, June 1983. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288636.

Efron:1984:CNN

[Efr84] Bradley Efron. Comparing non-nested linear models. Journal of the American Statistical Association, 79(388):791–803, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288709.

Efron:1986:DEF

[Efr86a] Bradley Efron. Double exponential families and their use in generalized linear regression. Journal of the American Statistical Association, 81(395):709–721, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289002.

Efron:1986:HBA

[Efr86b] Bradley Efron. How biased is the apparent error rate of a prediction rule. Journal of the American Statistical Association, 81 (394):461–470, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289236.

Efron:1987:EBC

[Efr87a] B. Efron. Empirical Bayes confidence intervals based on bootstrap samples: Comment. Journal of the American Statistical Association, 82(399):754, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288781.

Efron:1987:BBCb

[Efr87b] Bradley Efron. Better bootstrap confidence intervals: Rejoinder. Journal of the American Statistical Association, 82(397): 198–200, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289153.

Efron:1987:BBCa

[Efr87c] Bradley Efron. Better bootstrap confidence intervals. With comments and a rejoinder by the author. Journal of the American Statistical Association, 82(397):171–185, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289144.

Efron:1988:LRS

[Efr88] Bradley Efron. Logistic regression, survival analysis, and the Kaplan–Meier curve. Journal of the American Statistical Association, 83(402):414–425, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288857.

Egeth:1981:RPM

[Ege81] Howard Egeth. Representing points in many dimensions by trees and castles: Comment. Journal of the American Statistical Association, 76(374):269–270, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287821.

Engle:1986:SER

[EGRW86] Robert F. Engle, C. W. J. Granger, John Rice, and Andrew Weiss. Semiparametric estimates of the relation between weather and electricity sales. Journal of the American Statistical Association, 81(394):310–320, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289218.

Edwards:1983:MCB

[EH83] Donald G. Edwards and Jason C. Hsu. Multiple comparisons with the best treatment. Journal of the American Statistical Association, 78(384):965–971, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288211. See correction [EH84].

Edwards:1984:CMC

[EH84] Donald G. Edwards and Jason C. Hsu. Corrigenda: “Multiple comparisons with the best treatment”. Journal of the American Statistical Association, 79(388):965, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288771. See [EH83].

Edwards:1987:FMS

[EH87] David Edwards and Tomáš Havránek. A fast model selection procedure for large families of models. Journal of the American Statistical Association, 82(397):205–213, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289155.

Emerson:1984:LLS

[EHK84] John D. Emerson, David C. Hoaglin, and Peter J. Kempthorne. Leverage in least squares additive-plus-multiplicative fits for twoway tables. Journal of the American Statistical Association, 79 (386):329–335, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288272.

Eddy:1982:CDO

[EK82] William F. Eddy and Joseph B. Kadane. The cost of drilling for oil and gas: An application of constrained robust regression. Journal of the American Statistical Association, 77(378):262–269, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287229.

Ericksen:1985:EPCa

[EK85a] Eugene P. Ericksen and Joseph B. Kadane. Estimating the population in a census year: 1980 and beyond. Journal of the American Statistical Association, 80(389):98–109, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288051. See comment [Bai85].

Ericksen:1985:EPCb

[EK85b] Eugene P. Ericksen and Joseph B. Kadane. Estimating the population in a census year: 1980 and beyond: Rejoinder. Journal of the American Statistical Association, 80(389):129–131, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288062.

Ericksen:1989:ACH

[EKT89a] E. Ericksen, J. B. Kadane, and John W. Tukey. Adjusting the 1980 Census of Housing and Population. Journal of the American Statistical Association, 84(408):927–944, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ericksen:1989:ACP

[EKT89b] Eugene P. Ericksen, Joseph B. Kadane, and John W. Tukey. Adjusting the 1980 census of population and housing. Journal of the American Statistical Association, 84(408):927–944, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290068.

Elder:1981:GPB

[Eld81] Robert S. Elder. The grouping problem and beef quality standards. Journal of the American Statistical Association, 76(374): 246–248, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287818.

Eliashberg:1983:BRB

[Eli83] Jehoshua Eliashberg. Book review: [untitled]. Journal of the American Statistical Association, 78(383):742–743, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288167.

Eubank:1987:TSD

[ELR87] R. L. Eubank, V. N. LaRiccia, and R. B. Rosenstein. Test statistics derived as components of Pearson’s phi-squared distance measure. Journal of the American Statistical Association, 82(399): 816–825, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288791.

Emerson:1981:BRB

[Eme81] John Emerson. Book review: [untitled]. Journal of the American Statistical Association, 76(374):498–499, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287874.

Emigh:1983:BRB

[Emi83] Ted H. Emigh. Book review: [untitled]. Journal of the American Statistical Association, 78(384):993–994, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288226.

Espeland:1985:LLM

[EO85] Mark A. Espeland and Charles L. Odoroff. Log-linear models for doubly sampled categorical data fitted by the EM algorithm. Journal of the American Statistical Association, 80(391):663–670, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288482.

Espeland:1989:JEI

[EPG89] Mark A. Espeland, Orah S. Platt, and Dianne Gallagher. Joint estimation of incidence and diagnostic error rates from irregular longitudinal data. Journal of the American Statistical Association, 84(408):972–979, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290072.

Easton:1986:GSA

[ER86] George S. Easton and Elvezio Ronchetti. General saddlepoint approximations with applications to L statistics. Journal of the American Statistical Association, 81(394):420–430, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289231.

Erickson:1981:URR

[Eri81a] Gary M. Erickson. Using ridge regression to estimate directly lagged effects in marketing. Journal of the American Statistical Association, 76(376):766–773, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287566.

Ericson:1981:BRB

[Eri81b] W. A. Ericson. Book review: [untitled]. Journal of the American Statistical Association, 76(373):213–214, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287084.

Ericson:1982:ESS

[Eri82] W. A. Ericson. Estimation in survey sampling: Robustness and optimality: Comment. Journal of the American Statistical Association, 77(378):403–404, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287258.

El-Sabbagh:1981:BRB

[ES81] Mohammad F. El-Sabbagh. Book review: [untitled]. Journal of the American Statistical Association, 76(375):749, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287554.

El-Sabbagh:1982:BRB

[ES82a] Mohamed F. El-Sabbagh. Book review: [untitled]. Journal of the American Statistical Association, 77(378):499, June 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287290.

Emerson:1982:EMC

[ES82b] John D. Emerson and Michael A. Stoto. Exploratory methods for choosing power transformations. Journal of the American Statistical Association, 77(377):103–108, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287775.

Emerson:1984:CSP

[ES84] John D. Emerson and Michael A. Stoto. Choosing a symmetrizing power transformation: Rejoinder. Journal of the American Statistical Association, 79(385):108–109, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288343.

Escobar:1986:ERS

[Esc86] Luis A. Escobar. The equivalence of regression-simple and bestlinear-unbiased estimators with Type II censored data from a location scale distribution. Journal of the American Statistical Association, 81(393):210–214, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287992.

Espenshade:1984:BRB

[Esp84] Thomas J. Espenshade. Book review: [untitled]. Journal of the American Statistical Association, 79(387):742–743, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288454.

Espenshade:1987:BRB

[Esp87] Thomas J. Espenshade. Book review: [untitled]. Journal of the American Statistical Association, 82(398):687–688, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289496.

Ethier:1982:TFN

[Eth82] S. N. Ethier. Testing for favorable numbers on a roulette wheel. Journal of the American Statistical Association, 77(379):660–665, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287733.

Eubank:1986:BRB

[Eub86a] R. Eubank. Book review: [untitled]. Journal of the American Statistical Association, 81(395):858, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289023.

Eubank:1986:NSP

[Eub86b] R. L. Eubank. A note on smoothness priors and nonlinear regression. Journal of the American Statistical Association, 81(394): 514–517, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289243.

Engle:1981:OFM

[EW81] Robert Engle and Mark Watson. A one-factor multivariate time series model of metropolitan wage rates. Journal of the American Statistical Association, 76(376):774–781, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287567.

Ewbank:1985:BRB

[Ewb85] Douglas C. Ewbank. Book review: [untitled]. Journal of the American Statistical Association, 80(391):781, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288514.

Fabian:1994:CSA

[Fab94] Václav Fabian. Comment on: “Some asymptotic results for learning in single hidden-layer feedforward network models” [J. Amer. Statist. Assoc. 84 (1989), no. 408, 1003–1013, MR 92e:62119] by H. White. Journal of the American Statistical Association, 89 (428):1571, December 1994. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2291041. See [Whi89, Whi92].

Fahrenholtz:1981:BRB

[Fah81] Steven Fahrenholtz. Book review: [untitled]. Journal of the American Statistical Association, 76(373):209–210, March 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287077.

Fairley:1985:BRB

[Fai85] William B. Fairley. Book review: [untitled]. Journal of the American Statistical Association, 80(390):492, June 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287946.

Fairley:1986:BRB

[Fai86] William Fairley. Book review: [untitled]. Journal of the American Statistical Association, 81(395):873, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289046.

Farewell:1983:BRB

[Far83a] Vern Farewell. Book review: [untitled]. Journal of the American Statistical Association, 78(384):991, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288222.

Farley:1983:EHC

[Far83b] Dean E. Farley. Evaluating a hospital cost-containment program in a paired experiment: Comment. Journal of the American Statistical Association, 78(382):256–257, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288620.

Farrell:1987:BRB

[Far87] Roger Farrell. Book review: [untitled]. Journal of the American Statistical Association, 82(397):341–342, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289172.

Faris:1988:BRB

[Far88] William G. Faris. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1219–1220, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290170.

Fatti:1983:REM

[Fat83] L. P. Fatti. The random-effects model in discriminant analysis. Journal of the American Statistical Association, 78(383):679–687, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288138.

Fay:1980:BRB

[Fay80] Robert E. Fay, III. Book review: [untitled]. Journal of the American Statistical Association, 75(369):239, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287419.

Fay:1985:EPC

[Fay85a] Robert E. Fay. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):114–116, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288053.

Fay:1985:JCS

[Fay85b] Robert E. Fay. A jackknifed chi-squared test for complex samples. Journal of the American Statistical Association, 80(389): 148–157, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288065.

Fay:1986:CMP

[Fay86] Robert E. Fay. Causal models for patterns of nonresponse. Journal of the American Statistical Association, 81(394):354–365, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289224.

Fries:1983:ATF

[FB83] Arthur Fries and Gouri K. Bhattacharyya. Analysis of twofactor experiments under an inverse Gaussian model. Journal of the American Statistical Association, 78(384):820–826, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288191.

Chen:1981:EAM

[fC81] Chan fu Chen. The EM approach to the multiple indicators and multiple causes model via the estimation of the latent variable. Journal of the American Statistical Association, 76(375):704–708, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287534.

Feltz:1985:MLE

[FD85] Carol J. Feltz and Richard L. Dykstra. Maximum likelihood estimation of the survival functions of N stochastically ordered random variables. Journal of the American Statistical Association, 80 (392):1012–1019, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288568.

Ferenstein:1988:OSS

[FE88] E. Z. Ferenstein and E. G. Enns. Optimal sequential selection from a known distribution with holding costs. Journal of the American Statistical Association, 83(402):382–386, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288853.

Feigin:1982:BRB

[Fei82] Paul Feigin. Book review: [untitled]. Journal of the American Statistical Association, 77(377):214–215, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287797.

Feingold:1988:BRB

[Fei88] Marcia Feingold. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1225–1226, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290179.

Fellegi:1980:ATI

[Fel80] Ivan P. Fellegi. Approximate tests of independence and goodness of fit based on stratified multistage samples. Journal of the American Statistical Association, 75(370):261–268, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287444.

Fellegi:1985:EPC

[Fel85] I. P. Fellegi. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):116–119, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288054.

Feldman:1986:BRB

[Fel86] Kenneth A. Feldman. Book review: [untitled]. Journal of the American Statistical Association, 81(393):253–255, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288004.

Feldman:1988:BRB

[Fel88] Mark Feldman. Book review: [untitled]. Journal of the American Statistical Association, 83(401):267, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288951.

Fenech:1989:BRB

[Fen89] Alan Fenech. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1100–1101, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290103.

Ferguson:1982:IML

[Fer82] Thomas S. Ferguson. An inconsistent maximum likelihood estimate. Journal of the American Statistical Association, 77(380): 831–834, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287314.

Friedman:1980:MLE

[FG80] Lea Friedman and Ilya B. Gertsbakh. Maximum likelihood estimation in a minimum-type model with exponential and Weibull failure modes. Journal of the American Statistical Association, 75(370):460–465, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287477.

Fienberg:1984:GMA

[FG84] Stephen E. Fienberg and Gail D. Gong. Graphical methods for assessing logistic regression models: Comment. Journal of the American Statistical Association, 79(385):72–77, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288336.

Fuller:1978:REA

[FH78] Wayne A. Fuller and Michael A. Hidiroglou. Regression estimation after correcting for attenuation. Journal of the American Statistical Association, 73(361):99–104, March 1978. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2286529. See correction [FH81c].

Fuller:1981:CPP

[FH81a] Wayne A. Fuller and David P. Hasza. Correction: “Properties of predictors for autoregressive time series” [J. Amer. Statist. Assoc. 76 (1981), no. 373, 155–161, MR 82c:62132]. Journal of the American Statistical Association, 76(376):1023, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287633. See [FH81b].

Fuller:1981:PPA

[FH81b] Wayne A. Fuller and David P. Hasza. Properties of predictors for autoregressive time series. Journal of the American Statistical Association, 76(373):155–161, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287061. See correction [FH81a].

Fuller:1981:CRE

[FH81c] Wayne A. Fuller and Michael A. Hidiroglou. Correction: “Regression estimation after correcting for attenuation” [J. Amer. Statist. Assoc. 73 (1978), no. 361, 99–104, MR 80k:62087]. Journal of the American Statistical Association, 76(376):1022, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287626. See

[FH78, JTB81].

Frank:1982:CIU

[FH82] Ove Frank and Frank Harary. Cluster inference by using transitivity indices in empirical graphs. Journal of the American Statistical Association, 77(380):835–840, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287315.

Fatti:1986:VSH

[FH86] L. Paul Fatti and Douglas M. Hawkins. Variable selection in heteroscedastic discriminant analysis. Journal of the American Statistical Association, 81(394):494–500, June 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289240.

Fisher:1989:BCR

[FH89] Nicholas I. Fisher and Peter Hall. Bootstrap confidence regions for directional data. Journal of the American Statistical Association, 84(408):996–1002, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2290075. See correction [FH90].

Fisher:1990:CBC

[FH90] Nicholas I. Fisher and Peter Hall. Correction: “Bootstrap confidence regions for directional data”. Journal of the American Statistical Association, 85(410):608, June 1990. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289832. See [FH89].

Fleming:1987:SVL

[FHO87] Thomas R. Fleming, David P. Harrington, and Margaret O’Sullivan. Supremum versions of the log-rank and generalized Wilcoxon statistics. Journal of the American Statistical Association, 82(397):312–320, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289169.

Fienberg:1982:SED

[Fie82] Stephen E. Fienberg. Statistical evidence of discrimination: Comment. Journal of the American Statistical Association, 77(380): 784–787, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287306.

Fienberg:1987:BRB

[Fie87] Stephen E. Fienberg. Book review: [untitled]. Journal of the American Statistical Association, 82(397):353, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289191.

Finster:1983:FAS

[Fin83] Mark Finster. A frequentistic approach to sequential estimation in the general linear model. Journal of the American Statistical Association, 78(382):403–407, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288647.

Finney:1984:IPM

[Fin84] D. J. Finney. Improvement by planned multistage selection. Journal of the American Statistical Association, 79(387):501–509, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288396.

Findley:1986:BRB

[Fin86] Sally E. Findley. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1121–1122, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289102.

Finster:1987:BRB

[Fin87] Mark P. Finster. Book review: [untitled]. Journal of the American Statistical Association, 82(397):363–364, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289208.

Finster:1988:BRB

[Fin88] Mark P. Finster. Book review: [untitled]. Journal of the American Statistical Association, 83(401):280, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288973.

Firebaugh:1988:BRB

[Fir88] Glenn Firebaugh. Book review: [untitled]. Journal of the American Statistical Association, 83(402):570–571, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288892.

Fishman:1980:BRB

[Fis80] George S. Fishman. Book review: [untitled]. Journal of the American Statistical Association, 75(369):243, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287425.

Fisher:1986:SEA

[Fis86] Franklin M. Fisher. Statisticians, econometricians, and adversary proceedings. Journal of the American Statistical Association, 81 (394):277–286, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289215.

Field:1989:IAI

[FJV89] J. B. F. Field, A. T. James, and W. N. Venables. The influence of assumptions implicit in a model on parametric inference. Journal of the American Statistical Association, 84(405):101–106, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289851.

Fuchs:1980:TDO

[FK80] Camil Fuchs and Ron Kenett. A test for detecting outlying cells in the multinomial distribution and two-way contingency tables. Journal of the American Statistical Association, 75(370):395–398, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287465.

Fowlkes:1985:EOT

[FK85] E. B. Fowlkes and J. R. Kettenring. Estimating optimal transformations for multiple regression and correlation: Comment. Journal of the American Statistical Association, 80(391):607–613, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288476.

Follmann:1989:GLR

[FL89] Dean A. Follmann and Diane Lambert. Generalizing logistic regression by nonparametric mixing. Journal of the American Statistical Association, 84(405):295–300, March 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289877.

Flack:1986:BRB

[Fla86] Virginia Flack. Book review: [untitled]. Journal of the American Statistical Association, 81(394):573–574, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289263.

Fleiss:1984:BRB

[Fle84] Joseph L. Fleiss. Book review: [untitled]. Journal of the American Statistical Association, 79(386):474–475, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288302.

Fligner:1981:CCT

[Fli81] Michael Fligner. Corrigenda: a class of two-sample distributionfree tests for scale. Journal of the American Statistical Association, 76(373):217, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287090.

Fligner:1984:NTS

[Fli84] Michael A. Fligner. A note on two-sided distribution-free treatment versus control multiple comparisons. Journal of the American Statistical Association, 79(385):208–211, March 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288358.

Flinn:1986:BRB

[Fli86] Christopher J. Flinn. Book review: [untitled]. Journal of the American Statistical Association, 81(393):250, March 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287999.

Flury:1984:CPC

[Flu84] Bernhard N. Flury. Common principal components in K groups. Journal of the American Statistical Association, 79(388):892–898, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288721.

Flury:1985:ALC

[Flu85] Bernhard N. Flury. Analysis of linear combinations with extreme ratios of variance. Journal of the American Statistical Association, 80(392):915–922, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288553.

Feuerverger:1981:SFM

[FM81] Andrey Feuerverger and Philip McDunnough. On some Fourier methods for inference. Journal of the American Statistical Association, 76(374):379–387, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287839.

Fishman:1982:SEM

[FM82] George S. Fishman and Louis R. Moore. A statistical evaluation of multiplicative congruential random number generators with modulus 2^{31} − 1. Journal of the American Statistical Association, 77(377):129–136, March 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic).  URL http://links.jstor.org/sici?sici=0162-
1459%28198203%2977%3A377%3C129%3AASEOMC%3E2.0.CO%3B2-
Q; http://www.jstor.org/stable/2287778.

Faltin:1983:SSP

[FM83a] Frederick W. Faltin and Charles E. McCulloch. On the smallsample properties of the Olkin-Sobel-Tong estimator of the probability of correct selection. Journal of the American Statistical Association, 78(382):464–467, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288659.

Fowlkes:1983:MCTa

[FM83b] E. B. Fowlkes and C. L. Mallows. A method for comparing two hierarchical clusterings. Journal of the American Statistical Association, 78(383):553–569, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288117.

Fowlkes:1983:MCTb

[FM83c] E. B. Fowlkes and C. L. Mallows. A method for comparing two hierarchical clusterings: Rejoinder. Journal of the American Statistical Association, 78(383):584, September 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288123.

Fairbanks:1982:CIE

[FMD82] Kenneth Fairbanks, Richard Madsen, and Richard Dykstra. A confidence interval for an exponential parameter from a hybrid life test. Journal of the American Statistical Association, 77(377): 137–140, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287779.

Finch:1989:PMA

[FMT89] Stephen J. Finch, Nancy R. Mendell, and Henry C. Thode, Jr. Probabilistic measures of adequacy of a numerical search for a global maximum. Journal of the American Statistical Association, 84(408):1020–1023, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290078.

Fienberg:1985:SAM

[FMW85] Stephen E. Fienberg, Michael M. Meyer, and Stanley S. Wasserman. Statistical analysis of multiple sociometric relations. Journal of the American Statistical Association, 80(389):51–67, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288040.

Folks:1985:BRB

[Fol85] J. Leroy Folks. Book review: [untitled]. Journal of the American Statistical Association, 80(391):784–785, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288519.

Folks:1988:BRB

[Fol88] J. Leroy Folks. Book review: [untitled]. Journal of the American Statistical Association, 83(401):277, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288967.

Fomby:1986:BRB

[Fom86] Thomas B. Fomby. Book review: [untitled]. Journal of the American Statistical Association, 81(394):567–569, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289256.

Fong:1985:BRB

[Fon85] Geoffrey T. Fong. Book review: [untitled]. Journal of the American Statistical Association, 80(390):485–486, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287930.

Forrester:1980:ISMa

[For80a] Jay W. Forrester. Information sources for modeling the national economy. Journal of the American Statistical Association, 75 (371):555–566, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287644.

Forrester:1980:ISMb

[For80b] Jay W. Forrester. Information sources for modeling the national economy: Rejoinder. Journal of the American Statistical Association, 75(371):572–574, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287647.

Foutz:1980:ECG

[Fou80] Robert V. Foutz. Estimation of a common group delay between two multiple time series. Journal of the American Statistical Association, 75(372):779–788, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287159.

Fowler:1988:BRB

[Fow88] Floyd J. Fowler, Jr. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1223–1224, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290176.

Fligner:1981:RRP

[FP81] Michael A. Fligner and George E. Policello, II. Robust rank procedures for the Behrens–Fisher problem. Journal of the American Statistical Association, 76(373):162–168, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287062.

Freedman:1984:BRE

[FP84] David A. Freedman and Stephen C. Peters. Bootstrapping a regression equation: some empirical results. Journal of the American Statistical Association, 79(385):97–106, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288341.

Flury:1981:GRM

[FR81a] Bernhard Flury and Hans Riedwyl. Graphical representation of multivariate data by means of asymmetrical faces. Journal of the American Statistical Association, 76(376):757–765, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287565.

Friedman:1981:GMTa

[FR81b] Jerome H. Friedman and Lawrence C. Rafsky. Graphics for the multivariate two-sample problem. Journal of the American Statistical Association, 76(374):277–287, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287825.

Friedman:1981:GMTb

[FR81c] Jerome H. Friedman and Lawrence C. Rafsky. Graphics for the multivariate two-sample problem: Rejoinder. Journal of the American Statistical Association, 76(374):294–295, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287830.

Fielitz:1983:SDM

[FR83] Bruce D. Fielitz and James P. Rozelle. Stable distributions and the mixtures of distributions hypotheses for common stock returns. Journal of the American Statistical Association, 78(381): 28–36, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287096.

Franck:1981:MPI

[Fra81a] Wallace E. Franck. The most powerful invariant test of normal versus Cauchy with applications to stable alternatives. Journal of the American Statistical Association, 76(376):1002–1005, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287602.

Frank:1981:EFP

[Fra81b] Ove Frank. An exponential family of probability distributions for directed graphs: Comment. Journal of the American Statistical Association, 76(373):58–59, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287040.

Frankel:1982:BRB

[Fra82] Marvin Frankel. Book review: [untitled]. Journal of the American Statistical Association, 77(380):950, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287355.

Franck:1984:BRB

[Fra84a] Wallace E. Franck. Book review: [untitled]. Journal of the American Statistical Association, 79(388):956–957, December 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288752.

Franck:1984:LRT

[Fra84b] Wallace E. Franck. A likelihood ratio test for stochastic ordering. Journal of the American Statistical Association, 79(387): 686–691, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288417.

Franck:1986:BRB

[Fra86a] Wallace Franck. Book review: [untitled]. Journal of the American Statistical Association, 81(393):257, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288008.

Frank:1986:DLD

[Fra86b] Ove Frank. Disclosure-limited data dissemination: Comment. Journal of the American Statistical Association, 81(393):21– 22, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287961.

Franck:1988:BRB

[Fra88a] Wallace Franck. Book review: [untitled]. Journal of the American Statistical Association, 83(401):268–269, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288954.

Frankel:1988:BRB

[Fra88b] Lester R. Frankel. Book review: [untitled]. Journal of the American Statistical Association, 83(402):570, June 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288891.

Freebairn:1983:BRB

[Fre83] John Freebairn. Book review: [untitled]. Journal of the American Statistical Association, 78(381):201–202, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287132.

Fristedt:1986:BRB

[Fri86] Bert Fristedt. Book review: [untitled]. Journal of the American Statistical Association, 81(393):264–265, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288022.

Friedman:1987:EPP

[Fri87] Jerome H. Friedman. Exploratory projection pursuit. Journal of the American Statistical Association, 82(397):249–266, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289161.

Friedman:1989:RDA

[Fri89] Jerome H. Friedman. Regularized discriminant analysis. Journal of the American Statistical Association, 84(405):165–175, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289860.

Frydman:1984:MLE

[Fry84] Halina Frydman. Maximum likelihood estimation in the moverstayer model. Journal of the American Statistical Association, 79 (387):632–638, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288410.

Friedman:1981:PPR

[FS81] Jerome H. Friedman and Werner Stuetzle. Projection pursuit regression. Journal of the American Statistical Association, 76 (376):817–823, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287576.

Frank:1986:MG

[FS86] Ove Frank and David Strauss. Markov graphs. Journal of the American Statistical Association, 81(395):832–842, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289017.

Fitzpatrick:1987:QSC

[FS87] Simon Fitzpatrick and Alastair Scott. Quick simultaneous confidence intervals for multinomial proportions. Journal of the American Statistical Association, 82(399):875–878, September 1987.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288799.

Research:1984:BRB

[fSMR84] The Center for Survey Methods Research. Book review: [untitled]. Journal of the American Statistical Association, 79(386):478–479, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288310.

Friedman:1984:PPD

[FSS84] Jerome H. Friedman, Werner Stuetzle, and Anne Schroeder. Projection pursuit density estimation. Journal of the American Statistical Association, 79(387):599–608, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288406.

Fienberg:1985:DGL

[FT85] Stephen E. Fienberg and Ruey S. Tsay. Dynamic generalized linear models and Bayesian forecasting: Comment. Journal of the American Statistical Association, 80(389):89–90, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288045.

Fu:1987:BRB

[Fu87] James C. Fu. Book review: [untitled]. Journal of the American Statistical Association, 82(397):343–344, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289175.

Fuchs:1982:MLE

[Fuc82] Camil Fuchs. Maximum likelihood estimation and model selection in contingency tables with missing data. Journal of the American Statistical Association, 77(378):270–278, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287230.

Fuller:1981:ESR

[Ful81] Wayne A. Fuller. An empirical study of the ratio estimator and estimators of its variance: Comment. Journal of the American Statistical Association, 76(373):78–80, March 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287044.

Fligner:1988:MRM

[FV88] Michael A. Fligner and Joseph S. Verducci. Multistage ranking models. Journal of the American Statistical Association, 83(403): 892–901, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Fienberg:1981:EFP

[FW81] Stephen E. Fienberg and Stanley Wasserman. An exponential family of probability distributions for directed graphs: Comment. Journal of the American Statistical Association, 76(373): 54–57, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287039.

Fenech:1988:PFC

[FW88] Alan P. Fenech and Peter H. Westfall. The power function of conditional log-linear model tests. Journal of the American Statistical Association, 83(401):198–203, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288941.

Griffiths:1982:UTS

[GA82] William E. Griffiths and Jock R. Anderson. Using time-series and cross-section data to estimate a production function with positive and negative marginal risks. Journal of the American Statistical Association, 77(379):529–536, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287706.

Ganse:1983:PWB

[GAF83] Robert A. Ganse, Yasuo Amemiya, and Wayne A. Fuller. Prediction when both variables are subject to error, with application to earthquake magnitudes. Journal of the American Statistical Association, 78(384):761–765, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288180.

Gail:1983:RSH

[Gai83] Mitchell H. Gail. A reanalysis of the Stanford heart transplant data: Comment. Journal of the American Statistical Association, 78(382):275–277, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288623.

Gallant:1980:EEP

[Gal80] A. Ronald Gallant. Explicit estimators of parametric functions in nonlinear regression. Journal of the American Statistical Association, 75(369):182–193, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287409.

Gardner:1983:BRBb

[Gar83a] Roger Gardner. Book review: [untitled]. Journal of the American Statistical Association, 78(382):502–503, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288681.

Gardner:1983:BRBa

[Gar83b] Roy Gardner. Book review: [untitled]. Journal of the American Statistical Association, 78(382):502, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288680.

Gardiner:1984:BRB

[Gar84] Joseph C. Gardiner. Book review: [untitled]. Journal of the American Statistical Association, 79(387):734, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288436.

Gass:1983:BRB

[Gas83] Saul I. Gass. Book review: [untitled]. Journal of the American Statistical Association, 78(383):744, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288170.

Gastwirth:1985:UME

[Gas85] Joseph L. Gastwirth. The use of maximin efficiency robust tests in combining contingency tables and survival analysis. Journal of the American Statistical Association, 80(390):380–384, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287901.

Gastwirth:1986:DLD

[Gas86] Joseph L. Gastwirth. Disclosure-limited data dissemination: Comment. Journal of the American Statistical Association, 81 (393):23–25, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287962.

Gatsonis:1989:BRB

[Gat89] Constantine A. Gatsonis. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1096–1097, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290096.

Gaver:1989:BRBa

[Gav89] Donald P. Gaver, Jr. Book review: [untitled]. Journal of the American Statistical Association, 84(405):342–343, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289899.

Gaynor:1983:RSH

[Gay83] Jeffrey J. Gaynor. A reanalysis of the Stanford heart transplant data: Comment. Journal of the American Statistical Association, 78(382):288–290, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288628.

Grice:1980:ICM

[GB80] John V. Grice and Lee J. Bain. Inferences concerning the mean of the gamma distribution. Journal of the American Statistical Association, 75(372):929–933, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287183.

Goldstein:1982:SSS

[GC82] Michael Goldstein and David Collett. On sample-size selection and the evaluation of discriminability in the model choice problem. Journal of the American Statistical Association, 77(380): 922–928, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287328.

Gennings:1989:RSA

[GCC89] Chris Gennings, Vernon M. Chinchilli, and Walter H. Carter, Jr. Response surface analysis with correlated data: a nonlinear model approach. Journal of the American Statistical Association, 84 (407):805–809, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289670.

Goel:1980:OND

[GD80] Prem K. Goel and Morris H. DeGroot. Only normal distributions have linear posterior expectations in linear regression. Journal of the American Statistical Association, 75(372):895–900, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287178.

Goel:1981:IAH

[GD81] Prem K. Goel and Morris H. Degroot. Information about hyperparamters in hierarchical models. Journal of the American Statistical Association, 76(373):140–147, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287059.

Geisser:1980:CPA

[GE80] Seymour Geisser and William F. Eddy. Corrigenda: a predictive approach to model selection. Journal of the American Statistical Association, 75(371):765, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287700.

Gelfand:1986:BRBa

[Gel86a] Alan E. Gelfand. Book review: [untitled]. Journal of the American Statistical Association, 81(393):250–251, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288000.

Gelfand:1986:BRBb

[Gel86b] Alan E. Gelfand. Book review: [untitled]. Journal of the American Statistical Association, 81(395):872–873, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289045.

Gentleman:1981:GMT

[Gen81] Jane F. Gentleman. Graphics for the multivariate two-sample problem: Comment. Journal of the American Statistical Association, 76(374):289–291, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287827.

George:1984:BRB

[Geo84] Edward I. George. Book review: [untitled]. Journal of the American Statistical Association, 79(387):735–736, September 1984.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288439.

George:1986:CMS

[Geo86] Edward I. George. Combining minimax shrinkage estimators. Journal of the American Statistical Association, 81(394):437–445, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289233.

George:1987:BRB

[Geo87] Stephen L. George. Book review: [untitled]. Journal of the American Statistical Association, 82(397):358, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289197.

Gerber:1980:CCF

[Ger80] Hans U. Gerber. A characterization of certain families of distributions via essche transforms and independence. Journal of the American Statistical Association, 75(372):1015–1018, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287198.

Gertsbakh:1982:CLH

[Ger82] I. Gertsbakh. Confidence limits for highly reliable coherent systems with exponentially distributed component life. Journal of the American Statistical Association, 77(379):673–678, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287735.

Geweke:1981:CTI

[Gew81] John Geweke. A comparison of tests of the independence of two covariance-stationary time series. Journal of the American Statistical Association, 76(374):363–373, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287837.

Geweke:1982:MLDb

[Gew82a] John Geweke. Measurement of linear dependence and feedback between multiple time series: Rejoinder. Journal of the American Statistical Association, 77(378):323–324, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287243.

Geweke:1982:MLDa

[Gew82b] John Geweke. Measurement of linear dependence and feedback between multiple time series. With discussion and with a reply by the author. Journal of the American Statistical Association, 77(378):304–313, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287238.

Geweke:1984:NDS

[Gew84a] John Geweke. On the nature and discovery of structure: Comment. Journal of the American Statistical Association, 79(385): 25–26, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288328.

Geweke:1984:MCL

[Gew84b] John F. Geweke. Measures of conditional linear dependence and feedback between time series. Journal of the American Statistical Association, 79(388):907–915, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288723.

Geweke:1989:SAS

[Gew89] John Geweke. Sensitivity analysis of seasonal adjustments: Empirical case studies: Comment. Journal of the American Statistical Association, 84(405):28–30, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289842.

Good:1980:DEBa

[GG80a] I. J. Good and R. A. Gaskins. Density estimation and bumphunting by the penalized likelihood method exemplified by scattering and meteorite data. Journal of the American Statistical Association, 75(369):42–56, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287377.

Good:1980:DEBb

[GG80b] I. J. Good and R. A. Gaskins. Density estimation and bumphunting by the penalized likelihood method exemplified by scattering and meteorite data: Rejoinder. Journal of the American Statistical Association, 75(369):70–73, March 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287385.

Gheysens:1983:BRB

[GG83] Filip Gheysens and Bruce L. Golden. Book review: [untitled]. Journal of the American Statistical Association, 78(383):744–745, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288172.

Gastwirth:1987:ECR

[GG87] Joseph L. Gastwirth and Samuel W. Greenhouse. Estimating a common relative risk: application in equal employment. Journal of the American Statistical Association, 82(397):38–45, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289122.

Gilliland:1980:IOB

[GH80] Dennis C. Gilliland and James Hannan. Identification of the ordered bivariate normal distribution by minimum variate. Journal of the American Statistical Association, 75(371):651–654, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287663.

Gabriel:1983:RIR

[GH83a] K. R. Gabriel and W. J. Hall. Rerandomization inference on regression and shift effects: Computationally feasible methods. Journal of the American Statistical Association, 78(384):827–836, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288192.

Gabriel:1983:EPR

[GH83b] K. Ruben Gabriel and Chin-Fei Hsu. Evaluation of the power of rerandomization tests, with application to weather modification experiments. Journal of the American Statistical Association, 78 (384):766–775, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288181.

Gilula:1986:CAC

[GH86] Zvi Gilula and Shelby J. Haberman. Canonical analysis of contingency tables by maximum likelihood. Journal of the American Statistical Association, 81(395):780–788, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289010.

Gideon:1987:RCC

[GH87a] Rudy A. Gideon and Robert A. Hollister. A rank correlation coefficient resistant to outliers. Journal of the American Statistical Association, 82(398):656–666, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289479.

Greenacre:1987:GIC

[GH87b] Michael Greenacre and Trevor Hastie. The geometric interpretation of correspondence analysis. Journal of the American Statistical Association, 82(398):437–447, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289445.

Gilula:1988:AMC

[GH88] Zvi Gilula and Shelby J. Haberman. The analysis of multivariate contingency tables by restricted canonical and restricted association models. Journal of the American Statistical Association, 83 (403):760–771, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Ghosh:1983:BRB

[Gho83] J. K. Ghosh. Book review: [untitled]. Journal of the American Statistical Association, 78(383):731–732, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288146.

Ghosh:1988:BRB

[Gho88] Malay Ghosh. Book review: [untitled]. Journal of the American Statistical Association, 83(401):266–267, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288950.

Griffiths:1987:SSP

[GHP87] William E. Griffiths, R. Carter Hill, and Peter J. Pope. Small sample properties of probit model estimators. Journal of the American Statistical Association, 82(399):929–937, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288807.

Ghysels:1988:STD

[Ghy88] Eric Ghysels. A study toward a dynamic theory of seasonality for economic time series. Journal of the American Statistical Association, 83(401):168–172, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288937.

Gibbons:1981:SSS

[Gib81] Diane Galarneau Gibbons. A simulation study of some ridge estimators. Journal of the American Statistical Association, 76(373): 131–139, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287058.

Gibbons:1988:BRB

[Gib88] Robert D. Gibbons. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1221–1222, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290173.

Gifford:1987:BRB

[Gif87] Janice A. Gifford. Book review: [untitled]. Journal of the American Statistical Association, 82(398):677, June 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289482.

Gill:1984:UCR

[Gil84] Richard D. Gill. Understanding Cox’s regression model: a martingale approach. Journal of the American Statistical Association, 79(386):441–447, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288288.

Gilula:1986:GAC

[Gil86] Zvi Gilula. Grouping and association in contingency tables: an exploratory canonical correlation approach. Journal of the American Statistical Association, 81(395):773–779, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289009.

Glaz:1984:PIM

[GJ84] Joseph Glaz and Bruce McK. Johnson. Probability inequalities for multivariate distributions with dependence structures. Journal of the American Statistical Association, 79(386):436–440, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288287.

Gaver:1989:BRBb

[GJ89] Donald P. Gaver, Jr. and Patricia A. Jacobs. Book review: [untitled]. Journal of the American Statistical Association, 84(407): 843, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289687.

Gilula:1983:DMP

[GK83] Zvi Gilula and Abba M. Krieger. The decomposability and monotonicity of Pearson’s chi-square for collapsed contingency tables with applications. Journal of the American Statistical Association, 78(381):176–180, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287126.

Gilula:1988:OAC

[GKR88] Zvi Gilula, Abba M. Krieger, and Yaakov Ritov. Ordinal association in contingency tables: some interpretive aspects. Journal of the American Statistical Association, 83(402):540–545, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288875.

Gilstein:1983:SWR

[GL83] C. Zachary Gilstein and Edward E. Leamer. The set of weighted regression estimates. Journal of the American Statistical Association, 78(384):942–948, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288208.

Ghosh:1987:REB

[GL87] Malay Ghosh and Parthasarathi Lahiri. Robust empirical Bayes estimation of means from stratified samples. Journal of the American Statistical Association, 82(400):1153–1162, ???? 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Guttorp:1988:FLS

[GL88] Peter Guttorp and Richard A. Lockhart. Finding the location of a signal: a Bayesian analysis. Journal of the American Statistical Association, 83(402):322–330, June 1988. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288846.

Glaser:1980:BRF

[Gla80] Ronald E. Glaser. Bathtub and related failure rate characterizations. Journal of the American Statistical Association, 75(371): 667–672, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287666.

Glazebrook:1988:BRB

[Gla88] Kevin D. Glazebrook. Book review: [untitled]. Journal of the American Statistical Association, 83(401):274–275, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288963.

Glaz:1989:ABD

[Gla89] Joseph Glaz. Approximations and bounds for the distribution of the scan statistic. Journal of the American Statistical Association, 84(406):560–566, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289943.

Gleser:1985:BRB

[Gle85a] Leon Jay Gleser. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1069–1070, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288585.

Gleser:1985:EPG

[Gle85b] Leon Jay Gleser. Exact power of goodness-of-fit tests of Kolmogorov type for discontinuous distributions. Journal of the American Statistical Association, 80(392):954–958, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288560.

Glenn:1986:BRB

[Gle86] Norval D. Glenn. Book review: [untitled]. Journal of the American Statistical Association, 81(395):870, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289041.

Glymour:1986:SCI

[Gly86] Clark Glymour. Statistics and causal inference: Comment: Statistics and metaphysics. Journal of the American Statistical Association, 81(396):964–966, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289067.

Gebski:1984:RMR

[GM84] Val Gebski and Don McNeil. A refined method of robust smoothing. Journal of the American Statistical Association, 79(387):616– 623, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288408.

Ghosh:1986:EBE

[GM86a] Malay Ghosh and Glen Meeden. Empirical Bayes estimation in finite population sampling. Journal of the American Statistical Association, 81(396):1058–1062, December 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289083.

Guttman:1986:CWA

[GM86b] Irwin Guttman and Ulrich Menzefricke. Counting by weighing: an approach using renewal theory. Journal of the American Statistical Association, 81(393):129–131, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287979.

Golub:1982:NAV

[GN82] Gene H. Golub and Stephen G. Nash. Nonorthogonal analysis of variance using a generalized conjugate-gradient algorithm. Journal of the American Statistical Association, 77(377):109– 116, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287776.

Godambe:1982:APS

[God82a] V. P. Godambe. Ancillarity principle and a statistical paradox. Journal of the American Statistical Association, 77(380):931–933, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287330.

Godambe:1982:ESSb

[God82b] V. P. Godambe. Estimation in survey sampling: Robustness and optimality: Rejoinder. Journal of the American Statistical Association, 77(378):405–406, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287260.

Godambe:1982:ESSa

[God82c] V. P. Godambe. Estimation in survey sampling: robustness and optimality. With comments by W. A. Ericson and C.-M. Cassel and with a reply by the author. Journal of the American Statistical Association, 77(378):393–403, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287257.

Godbold:1985:BRB

[God85] James Godbold. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1080, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288603.

Goel:1981:BRB

[Goe81] Prem K. Goel. Book review: [untitled]. Journal of the American Statistical Association, 76(373):214, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287086.

Goel:1983:IMB

[Goe83] Prem K. Goel. Information measures and Bayesian hierarchical models. Journal of the American Statistical Association, 78(382): 408–410, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288648.

Golden:1983:BRBc

[Gol83a] Bruce Golden. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1005–1006, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288247.

Goldstein:1983:GVM

[Gol83b] Michael Goldstein. General variance modifications for linear Bayes estimators. Journal of the American Statistical Association, 78 (383):616–618, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288128.

Goldstein:1983:PP

[Gol83c] Michael Goldstein. The prevision of a prevision. Journal of the American Statistical Association, 78(384):817–819, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288190.

Goldie:1985:BRB

[Gol85] Charles M. Goldie. Book review: [untitled]. Journal of the American Statistical Association, 80(391):785–786, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288521.

Goldman:1986:BRB

[Gol86a] Noreen Goldman. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1120, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289100.

Goldstein:1986:EBS

[Gol86b] Michael Goldstein. Exchangeable belief structures. Journal of the American Statistical Association, 81(396):971–976, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289070.

Golomski:1988:BRB

[Gol88] William A. Golomski. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1227–1228, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290181.

Gong:1986:CVJ

[Gon86] Gail Gong. Cross-validation, the jackknife, and the bootstrap: Excess error estimation in forward logistic regression. Journal of the American Statistical Association, 81(393):108–113, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287975.

Good:1980:BRBa

[Goo80a] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 75(369):239, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287418.

Good:1980:BRBb

[Goo80b] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 75(370):480, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287498.

Goodman:1981:AMC

[Goo81] Leo A. Goodman. Association models and canonical correlation in the analysis of cross- classifications having ordered categories. Journal of the American Statistical Association, 76(374):320–334, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287833.

Good:1982:DCM

[Goo82a] I. J. Good. Diversity as a concept and its measurement: Comment. Journal of the American Statistical Association, 77(379): 561–563, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287710.

Good:1982:LPC

[Goo82b] I. J. Good. Lindley’s paradox: Comment. Journal of the American Statistical Association, 77(378):342–344, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287248.

Good:1983:BRB

[Goo83] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 78(384):987–989, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288219.

Good:1985:BRB

[Goo85] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1067–1068, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288583.

Good:1987:BRB

[Goo87a] I. J. Good. Book review: [untitled]. Journal of the American Statistical Association, 82(398):697–698, June 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289507.

Good:1987:TPN

[Goo87b] I. J. Good. Testing a point null hypothesis: The irreconcilability of P values and evidence: Comment. Journal of the American Statistical Association, 82(397):125–128, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289133.

Guttman:1985:DGL

[GP85] I. Guttman and D. Pena. Dynamic generalized linear models and Bayesian forecasting: Comment. Journal of the American Statistical Association, 80(389):91–92, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288046.

Gans:1981:DGK

[GR81] Lydia P. Gans and C. A. Robertson. Distributions of Goodman and Kruskal’s gamma and Spearman’s rho in 2×2 tables for small and moderate sample sizes. Journal of the American Statistical Association, 76(376):942–946, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287592.

Gray:1982:BRB

[Gra82] Lawrence F. Gray. Book review: [untitled]. Journal of the American Statistical Association, 77(380):937, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287334.

Granger:1984:BRB

[Gra84] Clive W. J. Granger. Book review: [untitled]. Journal of the American Statistical Association, 79(388):946, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288734.

Granger:1986:SCI

[Gra86] Clive Granger. Statistics and causal inference: Comment. Journal of the American Statistical Association, 81(396):967–968, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289068.

Green:1981:BRBa

[Gre81a] Bert F. Green. Book review: Randomization Tests, by Eugene S. Edgington. Journal of the American Statistical Association, 76 (374):495, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287866.

Green:1981:BRBb

[Gre81b] Richard F. Green. Book review: [untitled]. Journal of the American Statistical Association, 76(374):498, June 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287873.

Green:1981:BRBc

[Gre81c] Stephanie Green. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1015, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287617.

Green:1983:BRB

[Gre83] Stephanie Green. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1003–1004, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288242.

Greenacre:1984:BRB

[Gre84] Michael J. Greenacre. Book review: [untitled]. Journal of the American Statistical Association, 79(388):953–954, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288745.

Greenberg:1985:OSG

[Gre85a] Irwin Greenberg. A one-sided goodness-of-fit test for a multinomial population. Journal of the American Statistical Association, 80(391):558–562, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288469.

Greenhouse:1985:BRB

[Gre85b] Samuel W. Greenhouse. Book review: [untitled]. Journal of the American Statistical Association, 80(391):780–781, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288513.

Green:1986:BRB

[Gre86a] R. Jeffery Green. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1124–1125, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289106.

Greenberg:1986:BRB

[Gre86b] David F. Greenberg. Book review: [untitled]. Journal of the American Statistical Association, 81(395):870–871, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289042.

Greenhouse:1988:BRB

[Gre88] Joel B. Greenhouse. Book review: [untitled]. Journal of the American Statistical Association, 83(401):277–278, March 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288969.

Greenwood:1989:BRB

[Gre89] Michael J. Greenwood. Book review: [untitled]. Journal of the American Statistical Association, 84(407):849–850, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289700.

Griffin:1980:BRB

[Gri80] Diane Griffin. Book review: [untitled]. Journal of the American Statistical Association, 75(370):474, June 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287489.

Griffith:1989:BRB

[Gri89] William S. Griffith. Book review: [untitled]. Journal of the American Statistical Association, 84(406):625, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289971.

Grove:1980:TIA

[Gro80] D. M. Grove. A test of independence against a class of ordered alternatives in a 2 × C contingency table. Journal of the American Statistical Association, 75(370):454–459, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287476.

Groeneveld:1981:BRB

[Gro81a] Richard A. Groeneveld. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1011–1012, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287607.

Gross:1981:APE

[Gro81b] Shulamith T. Gross. On asymptotic power and efficiency of tests of independence in contingency tables with ordered classifications. Journal of the American Statistical Association, 76(376):935–941, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287591.

Grootaert:1986:UMD

[Gro86] Christiaan Grootaert. The use of multiple diaries in a household expenditure survey in Hong Kong. Journal of the American Statistical Association, 81(396):938–944, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289063.

Groshen:1987:BRB

[Gro87] Susan Groshen. Book review: [untitled]. Journal of the American Statistical Association, 82(397):360–361, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289200.

Greenlees:1982:IMV

[GRZ82] John S. Greenlees, William S. Reece, and Kimberly D. Zieschang. Imputation of missing values when the probability of response depends on the variable being imputed. Journal of the American Statistical Association, 77(378):251–261, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287228.

Gastwirth:1980:RPT

[GS80a] Joseph L. Gastwirth and Murray R. Selwyn. The robustness properties of two tests for serial correlation. Journal of the American Statistical Association, 75(369):138–141, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287401.

Gerig:1980:MTN

[GS80b] Thomas M. Gerig and A. R. Sen. MLE in two normal samples with equal but unknown population coefficients o variation. Journal of the American Statistical Association, 75(371):704–708, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287673.

Geweke:1980:ILR

[GS80c] John F. Geweke and Kenneth J. Singleton. Interpreting the likelihood ratio statistic in factor models when sample size is small. Journal of the American Statistical Association, 75(369): 133–137, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287400.

Grimes:1980:RTE

[GS80d] J. E. Grimes and B. V. Sukhatme. A regression-type estimator based on preliminary test of significance. Journal of the American Statistical Association, 75(372):957–962, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287188.

Ghosh:1989:MUP

[GS89] Malay Ghosh and Pranab Kumar Sen. Median unbiasedness and Pitman closeness. Journal of the American Statistical Association, 84(408):1089–1091, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290088.

Gupta:1989:IEC

[GSS89] Arjun K. Gupta, A. K. Md. Ehsanes Saleh, and Pranab K. Sen. Improved estimation in a contingency table: independence structure. Journal of the American Statistical Association, 84(406): 525–532, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289938.

Guilbaud:1983:NPI

[Gui83] Olivier Guilbaud. Nonparametric prediction intervals for sample medians in the general case. Journal of the American Statistical Association, 78(384):937–941, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288207.

Guilkey:1985:BRB

[Gui85] David K. Guilkey. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1065–1066, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288580.

Guilbaud:1988:EKT

[Gui88] Olivier Guilbaud. Exact Kolmogorov-type tests for left-truncated and/or right-censored data. Journal of the American Statistical Association, 83(401):213–221, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288943.

Gunst:1980:CSR

[Gun80] Richard F. Gunst. A critique of some ridge regression methods: Comment. Journal of the American Statistical Association, 75 (369):98–100, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287394.

Gupta:1981:BRB

[Gup81] Somesh Das Gupta. Book review: [untitled]. Journal of the American Statistical Association, 76(373):207, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287071.

Gupta:1984:BRB

[Gup84] Prithwis Das Gupta. Book review: [untitled]. Journal of the American Statistical Association, 79(388):961, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288761.

Gutmann:1982:SPI

[Gut82] Sam Gutmann. Stein’s paradox is impossible in the nonanticipative context. Journal of the American Statistical Association, 77 (380):934–935, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287331.

Gupta:1986:LOT

[GV86] Ashis Sen Gupta and Lea Vermeire. Locally optimal tests for multiparameter hypotheses. Journal of the American Statistical Association, 81(395):819–825, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289015.

Graybill:1980:CIN

[GW80] Franklin A. Graybill and Chih-Ming Wang. Confidence intervals on nonnegative linear combinations of variances. Journal of the American Statistical Association, 75(372):869–873, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287174.

Golden:1983:BRBa

[GW83a] Bruce L. Golden and Edward A. Wasil, Jr. Book review: [untitled]. Journal of the American Statistical Association, 78(381): 203–204, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287136.

Golden:1983:BRBb

[GW83b] Bruce L. Golden and Edward A. Wasil, Jr. Book review: [untitled]. Journal of the American Statistical Association, 78(382):496–497, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288668.

Gates:1984:DSS

[GW84] David J. Gates and Mark Westcott. On the distributions of scan statistics. Journal of the American Statistical Association, 79 (386):423–429, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288285.

Gray:1986:NAS

[GW86] H. L. Gray and Wayne A. Woodward. A new ARMA spectral estimator. Journal of the American Statistical Association, 81 (396):1100–1108, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289090.

Gay:1988:MLQ

[GW88] David M. Gay and Roy E. Welsch. Maximum likelihood and quasi-likelihood for nonlinear exponential family regression models. Journal of the American Statistical Association, 83(404):990– 998, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290125.

Gruttola:1987:IAG

[GWL87] Victor De Gruttola, James H. Ware, and Thomas A. Louis. Influence analysis of generalized least squares estimators. Journal of the American Statistical Association, 82(399):911–917, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288804.

Haack:1988:BRB

[Haa88] Dennis G. Haack. Book review: [untitled]. Journal of the American Statistical Association, 83(402):563, June 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288881.

Haber:1980:CSC

[Hab80] Michael Haber. A comparison of some continuity corrections for the chi-squared test on 2 × 2 tables. Journal of the American Statistical Association, 75(371):510–515, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287639.

Haberman:1981:EFP

[Hab81] Shelby J. Haberman. An exponential family of probability distributions for directed graphs: Comment. Journal of the American Statistical Association, 76(373):60–61, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287041.

Haberman:1982:ADM

[Hab82] Shelby J. Haberman. Analysis of dispersion of multinomial responses. Journal of the American Statistical Association, 77(379): 568–580, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287713.

Haberman:1987:CPD

[Hab87] Shelby J. Haberman. The calculation of posterior distributions by data augmentation: Comment. Journal of the American Statistical Association, 82(398):546, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289461.

Haberman:1988:WUC

[Hab88] Shelby J. Haberman. A warning on the use of chi-squared statistics with frequency tables with small expected cell counts. Journal of the American Statistical Association, 83(402):555–560, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288877.

Haines:1981:BRB

[Hai81] Michael R. Haines. Book review: [untitled]. Journal of the American Statistical Association, 76(373):211, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287081.

Hauck:1983:CFS

[HAJ83] Walter W. Hauck, Sharon Anderson, and Francis J. Leahy, III. Corrigenda: Finite-sample properties of some old and some new estimators of a common odds ratio from multiple 2×2 tables. Journal of the American Statistical Association, 78(384):1008, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288252. See

[HAL82].

Hauck:1982:FSP

[HAL82] Walter W. Hauck, Sharon Anderson, and Francis J. Leahy, III. Finite-sample properties of some old and some new estimators of a common odds ratio from multiple 2 × 2 tables. Journal of the American Statistical Association, 77(377):145–152, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287781. See

corrigenda [HAJ83].

Han:1981:BRB

[Han81] Chien-Pai Han. Book review: [untitled]. Journal of the American Statistical Association, 76(375):751–752, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287558.

Han:1983:BRB

[Han83] Chien-Pai Han. Book review: [untitled]. Journal of the American Statistical Association, 78(381):212–213, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287152.

Hanumara:1984:BRB

[Han84] R. Choudary Hanumara. Book review: [untitled]. Journal of the American Statistical Association, 79(388):952–953, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288744.

Hansen:1985:EPC

[Han85a] Morris H. Hansen. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):119–122, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288055.

Hanssens:1985:BRB

[Han85b] D. M. Hanssens. Book review: [untitled]. Journal of the American Statistical Association, 80(390):494, June 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287949.

Hartley:1980:SSP

[Har80a] H. O. Hartley. Statistics as a science and as a profession. Journal of the American Statistical Association, 75(369):1–7, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287371.

Harville:1980:PNF

[Har80b] David Harville. Predictions for national football league games via linear-model methodology. Journal of the American Statistical Association, 75(371):516–524, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287640.

Hartigan:1981:CSL

[Har81a] J. A. Hartigan. Consistency of single linkage for high-density clusters. Journal of the American Statistical Association, 76(374): 388–394, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287840.

Hartigan:1981:EVR

[Har81b] J. A. Hartigan. Estimating volumes of remaining fossil fuel resources: a critical review: Comment. Journal of the American Statistical Association, 76(375):548, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287508.

Hardy:1982:BRB

[Har82] Robert J. Hardy. Book review: [untitled]. Journal of the American Statistical Association, 77(380):947, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287350.

Hart:1984:EKD

[Har84] Jeffrey D. Hart. Efficiency of a kernel density estimator under an autoregressive dependence model. Journal of the American Statistical Association, 79(385):110–117, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288344.

Harrington:1985:BRB

[Har85a] David Harrington. Book review: [untitled]. Journal of the American Statistical Association, 80(389):244–245, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288093.

Harville:1985:DPE

[Har85b] David A. Harville. Decomposition of prediction error. Journal of the American Statistical Association, 80(389):132–138, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288063.

Harshman:1986:BRB

[Har86] Richard A. Harshman. Book review: [untitled]. Journal of the American Statistical Association, 81(395):864–865, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289034.

Hartigan:1987:ECD

[Har87] J. A. Hartigan. Estimation of a convex density contour in two dimensions. Journal of the American Statistical Association, 82 (397):267–270, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289162.

Hart:1989:BRB

[Har89] Marilyn Hart. Book review: [untitled]. Journal of the American Statistical Association, 84(405):343–344, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289900.

Hasza:1980:ADS

[Has80] David P. Hasza. The asymptotic distribution of the sample autocorrelations for an integrated ARMA process. Journal of the American Statistical Association, 75(370):349–352, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287457.

Hastie:1984:GMA

[Has84] Trevor J. Hastie. Graphical methods for assessing logistic regression models: Comment. Journal of the American Statistical Association, 79(385):77–78, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288337.

Haugh:1982:BRB

[Hau82] Larry D. Haugh. Book review: [untitled]. Journal of the American Statistical Association, 77(378):492–493, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287278.

Hauser:1985:EPC

[Hau85] Philip M. Hauser. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):122–123, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288056.

Hawkins:1989:BRB

[Haw89] Douglas M. Hawkins. Book review: [untitled]. Journal of the American Statistical Association, 84(405):334–335, March 1989.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289888.

Hayter:1986:MFE

[Hay86] Anthony J. Hayter. The maximum familywise error rate of Fisher’s least significant difference test. Journal of the American Statistical Association, 81(396):1000–1004, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289074.

Hayter:1989:PCG

[Hay89] A. J. Hayter. Pairwise comparisons of generally correlated means. Journal of the American Statistical Association, 84(405):208– 213, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289865.

Hui:1983:EBE

[HB83] Siu L. Hui and James O. Berger. Empirical Bayes estimation of rates in longitudinal studies. Journal of the American Statistical Association, 78(384):753–760, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288179.

Haber:1986:MLM

[HB86] Michael Haber and Morton B. Brown. Maximum likelihood methods for log-linear models when expected frequencies are subjected to linear constraints. Journal of the American Statistical Association, 81(394):477–482, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289238.

Hardle:1988:BNR

[HB88] Wolfgang Härdle and Adrian W. Bowman. Bootstrapping in nonparametric regression: local adaptive smoothing and confidence bands. Journal of the American Statistical Association, 83(401): 102–110, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288926.

Havenner:1981:EAC

[HC81] Arthur Havenner and Roger Craine. Estimation analogies in control. Journal of the American Statistical Association, 76(376):

850–859, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287580.

Hora:1984:STW

[HC84] Stephen C. Hora and W. J. Conover. The F statistic in the two-way layout with rank-score transformed data. Journal of the American Statistical Association, 79(387):668–673, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288415.

Herman:1985:SMP

[HCGL85] Gabor T. Herman, Yair Censor, Dan Gordon, and Robert M. Lewitt. A statistical model for positron emission tomography: Comment. Journal of the American Statistical Association, 80 (389):22–25, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288032.

Hauck:1980:CWT

[HD80] Walter W. Hauck and Allan Donner. Corrigenda: Wald’s test as applied to hypotheses in logit analysis. Journal of the American Statistical Association, 75(370):482, June 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287501.

Hong:1988:GE

[HD88] Chong Sun Hong and H. T. David. Granularity and efficiency. Journal of the American Statistical Association, 83(403):880–885, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hsu:1983:SMC

[HE83] Jason C. Hsu and Donald G. Edwards. Sequential multiple comparisons with the best. Journal of the American Statistical Association, 78(384):958–964, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288210.

Heckman:1988:MES

[Hec88] Nancy E. Heckman. Minimax estimates in a semiparametric model. Journal of the American Statistical Association, 83(404): 1090–1096, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290141.

Heltshe:1983:BRB

[Hel83] James F. Heltshe. Book review: [untitled]. Journal of the American Statistical Association, 78(383):735, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288154.

Henze:1979:END

[Hen79] Fritz H.-H. Henze. The exact noncentral distributions of Spearman’s r and other related correlation coefficients. Journal of the American Statistical Association, 74(366):459–464, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2286356. See cor-

rigenda [Hen80].

Henze:1980:CEN

[Hen80] Fritz H.-H. Henze. Corrigenda: The exact noncentral distribution of Spearman’s r and other related correlation coefficients. Journal of the American Statistical Association, 75(371):765, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287701. See

[Hen79].

Hendrickson:1983:BRB

[Hen83] John A. Hendrickson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 78(384):991–992, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288224.

Hermann:1980:BRB

[Her80] Johannes Hermann. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1042–1043, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287220.

Hernandez:1985:BRB

[Her85] Donald J. Hernandez. Book review: [untitled]. Journal of the American Statistical Association, 80(389):254–255, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288107.

Hettmansperger:1987:BRB

[Het87] Thomas P. Hettmansperger. Book review: [untitled]. Journal of the American Statistical Association, 82(398):691–693, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289500.

Heyde:1982:BRBa

[Hey82a] C. C. Heyde. Book review: [untitled]. Journal of the American Statistical Association, 77(379):690, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287748.

Heyde:1982:BRBb

[Hey82b] C. C. Heyde. Book review: [untitled]. Journal of the American Statistical Association, 77(380):937, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287333.

Hayre:1981:SSL

[HG81] Lakhbir S. Hayre and John C. Gittins. Sequential selection of the larger of two normal means. Journal of the American Statistical Association, 76(375):696–700, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287532.

Huang:1982:NSU

[HG82] J. S. Huang and Malay Ghosh. A note on strong unimodality of order statistics. Journal of the American Statistical Association, 77(380):929–930, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2287329. See corrigenda [HG83].

Huang:1983:CNS

[HG83] J. S. Huang and M. Ghosh. Corrigenda: “A note on strong unimodality of order statistics” [J. Amer. Statist. Assoc. 77 (1982), no. 380, 929–930, MR 84d:62098]. Journal of the American Statistical Association, 78(384):1008, December 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288251. See [HG82].

Heckman:1989:CAAa

[HH89a] James J. Heckman and V. Joseph Hotz. Choosing among alternative nonexperimental methods for estimating the impact of social programs: The case of manpower training. Journal of the American Statistical Association, 84(408):862–874, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290059.

Heckman:1989:CAAb

[HH89b] James J. Heckman and V. Joseph Hotz. Choosing among alternative nonexperimental methods for estimating the impact of social programs: The case of manpower training: Rejoinder. Journal of the American Statistical Association, 84(408):878–880, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290062.

Hardle:1988:HFAb

[HHM88a] Wolfgang Hardle, Peter Hall, and J. S. Marron. How far are automatically chosen regression smoothing parameters from their optimum?: Rejoinder. Journal of the American Statistical Association, 83(401):100–101, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288925.

Hardle:1988:HFAa

[HHM88b] Wolfgang Härdle, Peter Hall, and J. S. Marron. How far are automatically chosen regression smoothing parameters from their optimum? with comments by David W. Scott and Iain Johnstone and a reply by the authors. Journal of the American Statistical Association, 83(401):86–95, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288922.

Hansen:1988:EBI

[HHO88] Lars Peter Hansen, John C. Heaton, and Masao Ogaki. Efficiency bounds implied by multiperiod conditional moment restrictions. Journal of the American Statistical Association, 83(403):863–871, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hora:1988:ARE

[HI88] Stephen C. Hora and Ronald L. Iman. Asymptotic relative efficiencies of the rank-transformation procedure in randomized complete block designs. Journal of the American Statistical Association, 83(402):462–470, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288863.

Hill:1980:BRB

[Hil80] Bruce Hill. Book review: [untitled]. Journal of the American Statistical Association, 75(369):252–253, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287437.

Hill:1982:LPC

[Hil82] Bruce M. Hill. Lindley’s paradox: Comment. Journal of the American Statistical Association, 77(378):344–347, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287249.

Hill:1983:BRB

[Hil83] Bruce M. Hill. Book review: [untitled]. Journal of the American Statistical Association, 78(382):498, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288671.

Hill:1984:BRB

[Hil84] R. Carter Hill. Book review: Recent Advances in Regression Methods, by Hrishikesh D. Vinod; Aman Ullah. Journal of the American Statistical Association, 79(385):236, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288378.

Hill:1985:BRBa

[Hil85] Bruce M. Hill. Book review: [untitled]. Journal of the American Statistical Association, 80(389):253–254, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288105.

Hill:1986:BRBa

[Hil86a] Bruce M. Hill. Book review: [untitled]. Journal of the American Statistical Association, 81(394):569–571, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289257.

Hill:1986:BRBb

[Hil86b] Walter W. Hill, Jr. Book review: [untitled]. Journal of the American Statistical Association, 81(394):578, June 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289272.

Hill:1986:BRBc

[Hil86c] Walter W. Hill, Jr. Book review: [untitled]. Journal of the American Statistical Association, 81(394):578–579, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289273.

Hill:1987:EBC

[Hil87a] Joe R. Hill. Empirical Bayes confidence intervals based on bootstrap samples: Comment. Journal of the American Statistical Association, 82(399):752–754, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288780.

Hill:1987:BRB

[Hil87b] M. Anne Hill. Book review: [untitled]. Journal of the American Statistical Association, 82(398):684–685, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289492.

Hildebrand:1989:BRB

[Hil89] David K. Hildebrand. Book review: [untitled]. Journal of the American Statistical Association, 84(405):347–348, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289906.

Hinkley:1980:RAE

[Hin80] David V. Hinkley. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):582–584, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287649.

Hinich:1982:BRB

[Hin82a] Melvin J. Hinich. Book review: [untitled]. Journal of the American Statistical Association, 77(380):941–942, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287343.

Hinkley:1982:SAS

[Hin82b] David Hinkley. Some ancillary statistics and their properties: Comment. Journal of the American Statistical Association, 77 (379):592–593, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287717.

Hinkley:1983:PEB

[Hin83] David Hinkley. Parametric empirical Bayes inference: Theory and applications: Comment. Journal of the American Statistical Association, 78(381):58–59, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287101.

Hinkley:1984:ATDb

[Hin84] David Hinkley. The analysis of transformed data: Rejoinder. Journal of the American Statistical Association, 79(386):319–320, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288270.

Hinich:1987:BRB

[Hin87a] Melvin J. Hinich. Book review: [untitled]. Journal of the American Statistical Association, 82(398):694, June 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289502.

Hinkley:1987:TPN

[Hin87b] David V. Hinkley. Testing a point null hypothesis: The irreconcilability of P values and evidence: Comment. Journal of the American Statistical Association, 82(397):128–129, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289134.

Hirsch:1981:BRB

[Hir81] Werner Z. Hirsch. Book review: [untitled]. Journal of the American Statistical Association, 76(374):500–501, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287877.

Hirsch:1982:BRB

[Hir82] Werner Z. Hirsch. Book review: [untitled]. Journal of the American Statistical Association, 77(379):684–685, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287738.

Hirsch:1983:BRB

[Hir83] Werner Z. Hirsch. Book review: [untitled]. Journal of the American Statistical Association, 78(384):985–986, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288216.

Hoaglin:1986:PSR

[HIT86] David C. Hoaglin, Boris Iglewicz, and John W. Tukey. Performance of some resistant rules for outlier labeling. Journal of the American Statistical Association, 81(396):991–999, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289073. See

comment [Sim87b].

Hoaglin:1987:HIT

[HIT87] D. C. Hoaglin, B. Iglewicz, and J. W. Tukey. Hoaglin, Iglewicz, and Tukey — reply. Journal of the American Statistical Association, 82(398):704, ???? 1987. CODEN JSTNAL. ISSN 0162-1459.

Hernandez:1980:LSB

[HJ80] Fabian Hernandez and Richard A. Johnson. The large-sample behavior of transformations to normality. Journal of the American Statistical Association, 75(372):855–861, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287172.

Holland:1981:EFPb

[HL81a] Paul W. Holland and Samuel Leinhardt. An exponential family of probability distributions for directed graphs: Rejoinder. Journal of the American Statistical Association, 76(373):62–65, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287042.

Holland:1981:EFPa

[HL81b] Paul W. Holland and Samuel Leinhardt. An exponential family of probability distributions for directed graphs. With comments by Ronald L. Breiger, Stephen E. Fienberg, Stanley Wasserman, Ove Frank and Shelby J. Haberman and a reply by the authors. Journal of the American Statistical Association, 76(373):33–50, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287037.

Hettmansperger:1983:GII

[HM83] Thomas P. Hettmansperger and Joseph W. McKean. A geometric interpretation of inferences based on ranks in the linear model. Journal of the American Statistical Association, 78(384):885–893, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288200.

Hedayat:1985:CEU

[HM85a] A. S. Hedayat and Dibyen Majumdar. Combining experiments under Gauss–Markov models. Journal of the American Statistical Association, 80(391):698–703, September 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288487.

Hill:1985:BRBb

[HM85b] Martha S. Hill and James N. Morgan. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1061– 1062, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288575.

Hsiao:1985:CES

[HM85c] Cheng Hsiao and Dean Mountain. Correction: Estimating the short-run income elasticity of demand for elasticity by using crosssectional categorized data. Journal of the American Statistical Association, 80(392):1084, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288611.

Hsiao:1985:ESR

[HM85d] Cheng Hsiao and Dean Mountain. Estimating the short-run income elasticity of demand for electricity by using cross-sectional categorized data. Journal of the American Statistical Association, 80(390):259–265, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287881.

Hallin:1988:RBT

[HM88a] Marc Hallin and Guy Mélard. Rank-based tests for randomness against first-order serial dependence. Journal of the American Statistical Association, 83(404):1117–1128, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290144.

Hannan:1988:RPA

[HM88b] E. J. Hannan and A. J. McDougall. Regression procedures for ARMA estimation. Journal of the American Statistical Association, 83(402):490–498, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2288867. See correction [HM89].

Hannan:1989:CRP

[HM89] E. J. Hannan and A. J. McDougall. Correction: “Regression procedures for ARMA estimation”. Journal of the American Statistical Association, 84(406):636, ???? 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [HM88b].

Hsieh:1985:ERP

[HMM85] David A. Hsieh, Charles F. Manski, and Daniel McFadden. Estimation of response probabilities from augmented retrospective observations. Journal of the American Statistical Association, 80 (391):651–662, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288481.

Hirji:1987:CDE

[HMP87] Karim F. Hirji, Cyrus R. Mehta, and Nitin R. Patel. Computing distributions for exact logistic regression. Journal of the American Statistical Association, 82(400):1110–1117, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Hale:1980:FSA

[HMR80] C. Hale, R. S. Mariano, and J. G. Ramage. Finite sample analysis of misspecification in simultaneous equation models. Journal of the American Statistical Association, 75(370):418–427, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287470.

Hansen:1983:EMDa

[HMT83a] Morris H. Hansen, William G. Madow, and Benjamin J. Tepping. An evaluation of model-dependent and probability-sampling inferences in sample surveys. Journal of the American Statistical Association, 78(384):776–793, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288182.

Hansen:1983:EMDb

[HMT83b] Morris H. Hansen, William G. Madow, and Benjamin J. Tepping. An evaluation of model-dependent and probability-sampling inferences in sample surveys: Rejoinder. Journal of the American Statistical Association, 78(384):805–807, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288188.

Hettmansperger:1987:TPA

[HN87] Thomas P. Hettmansperger and Robert M. Norton. Tests for patterned alternatives in k-sample problems. Journal of the American Statistical Association, 82(397):292–299, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289166.

Ho:1984:BRB

[Ho84] Irwin Ho. Book review: [untitled]. Journal of the American Statistical Association, 79(386):476, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288305.

Hodges:1987:AAN

[Hod87a] James S. Hodges. Assessing the accuracy of normal approximations. Journal of the American Statistical Association, 82(397): 149–154, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289141.

Hodges:1987:BRB

[Hod87b] James S. Hodges. Book review: [untitled]. Journal of the American Statistical Association, 82(398):689–691, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289498.

Hodoshima:1988:ETI

[Hod88] Jiro Hodoshima. The effect of truncation on the identifiability of regression coefficients. Journal of the American Statistical Association, 83(404):1055–1056, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290135.

Hoeffding:1984:RPU

[Hoe84] Wassily Hoeffding. Range preserving unbiased estimators in the multinomial case. Journal of the American Statistical Association, 79(387):712–714, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288421.

Hogg:1985:BRB

[Hog85] Robert V. Hogg. Book review: [untitled]. Journal of the American Statistical Association, 80(391):784, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288518.

Hogg:1989:HHS

[Hog89] Robert V. Hogg. How to hope with statistics. Journal of the American Statistical Association, 84(405):1–5, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289836.

Holmes:1983:BRB

[Hol83] Paul T. Holmes. Book review: [untitled]. Journal of the American Statistical Association, 78(382):500–501, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288675.

Holt:1985:BRB

[Hol85] D. Holt. Book review: Planning and Analysis of Observational Studies, by William G. Cochran. Journal of the American Statistical Association, 80(391):772–773, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288502.

Holland:1986:SCIb

[Hol86a] Paul W. Holland. Statistics and causal inference: Rejoinder. Journal of the American Statistical Association, 81(396):968–970, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289069.

Holland:1986:SCIa

[Hol86b] Paul W. Holland. Statistics and causal inference. With discussion and a reply by the author. Journal of the American Statistical Association, 81(396):945–960, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289064.

Holden:1987:TSA

[Hol87] Robert T. Holden. Time series analysis of a contagious process. Journal of the American Statistical Association, 82(400):1019– 1026, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Holland:1989:CAA

[Hol89] Paul W. Holland. Choosing among alternative nonexperimental methods for estimating the impact of social programs: The case of manpower training: Comment. Journal of the American Statistical Association, 84(408):875–877, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290060.

Hooke:1987:BRB

[Hoo87] Robert Hooke. Book review: [untitled]. Journal of the American Statistical Association, 82(399):953, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288824.

Hooper:1989:ERV

[Hoo89] Peter M. Hooper. Experimental randomization and the validity of normal-theory inference. Journal of the American Statistical Association, 84(406):576–586, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289945.

Horn:1983:SES

[Hor83] Paul S. Horn. Some easy t statistics. Journal of the American Statistical Association, 78(384):930–936, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288206.

Hora:1984:BRB

[Hor84] Stephen C. Hora. Book review: [untitled]. Journal of the American Statistical Association, 79(386):473–474, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288300.

Horn:1988:BPI

[Hor88] Paul S. Horn. A biweight prediction interval for random samples. Journal of the American Statistical Association, 83(401): 249–256, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288947.

Horiuchi:1989:BRB

[Hor89] Shiro Horiuchi. Book review: [untitled]. Journal of the American Statistical Association, 84(405):344, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289901.

Hosking:1980:MPS

[Hos80] J. R. M. Hosking. The multivariate portmanteau statistic. Journal of the American Statistical Association, 75(371):602–608, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287656.

Hosmer:1986:BRB

[Hos86] David Hosmer. Book review: [untitled]. Journal of the American Statistical Association, 81(393):259–260, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288013.

Howard:1984:BRB

[How84] Diantha Howard. Book review: Elementary Statistical Methods, by G. Barrie Wetherill. Journal of the American Statistical Association, 79(385):242–243, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288388.

Howrey:1986:BRB

[How86] E. Philip Howrey. Book review: [untitled]. Journal of the American Statistical Association, 81(393):261–262, March 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288018.

Howrey:1989:BRB

[How89] E. P. Howrey. Book review: [untitled]. Journal of the American Statistical Association, 84(405):341–342, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289897.

Harvey:1984:EMO

[HP84] A. C. Harvey and R. G. Pierse. Estimating missing observations in economic time series. Journal of the American Statistical Association, 79(385):125–131, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288346.

Hollander:1988:NTU

[HP88] Myles Hollander and Edsel Peña. Nonparametric tests under restricted treatment-assignment rules. Journal of the American Statistical Association, 83(404):1144–1151, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290147.

Hwang:1981:OHP

[HPE81] Frank K. Hwang, Charles G. Pfeifer, and Peter Enis. An optimal hierarchical procedure for a modified binomial group-testing problem. Journal of the American Statistical Association, 76(376): 947–949, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287593.

Hollander:1986:CLD

[HPP86] Myles Hollander, Dong Ho Park, and Frank Proschan. A class of life distributions for aging. Journal of the American Statistical Association, 81(393):91–95, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287972.

Hinkley:1984:ATDa

[HR84] D. V. Hinkley and G. Runger. The analysis of transformed data. With discussion. Journal of the American Statistical Association, 79(386):302–309, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288264.

Hampel:1981:CVC

[HRR81] Frank R. Hampel, Peter J. Rousseeuw, and Elvezio Ronchetti. The change-of-variance curve and optimal redescending M-estimators. Journal of the American Statistical Association, 76(375):643–648, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287524.

Hidiroglou:1981:SEP

[HS81] Michael A. Hidiroglou and Kadaba P. Srinath. Some estimators of a population total from simple random samples containing large units. Journal of the American Statistical Association, 76(375): 690–695, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287531.

Hardle:1989:ISM

[HS89a] Wolfgang Härdle and Thomas M. Stoker. Investigating smooth multiple regression by the method of average derivatives. Journal of the American Statistical Association, 84(408):986–995, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290074.

Hastie:1989:PC

[HS89b] Trevor Hastie and Werner Stuetzle. Principal curves. Journal of the American Statistical Association, 84(406):502–516, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289936.

Hedayat:1989:RBP

[HS89c] A. Hedayat and J. Stufken. A relation between pairwise balanced and variance balanced block designs. Journal of the American Statistical Association, 84(407):753–755, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289658.

Hasselblad:1980:ACG

[HSG80] Victor Hasselblad, Andrew G. Stead, and Warren Galke. Analysis of coarsely grouped data from the lognormal distribution. Journal of the American Statistical Association, 75(372):771–778, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287158.

Hasselblad:1982:CAC

[HSG82] Victor Hasselblad, Andrew G. Stead, and Warren Galke. Corrigenda: Analysis of coarsely grouped data from the lognormal distribution. Journal of the American Statistical Association, 77 (380):954, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287358.

Hsiao:1980:BRB

[Hsi80] Cheng Hsiao. Book review: [untitled]. Journal of the American Statistical Association, 75(371):757–758, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287691.

Hsu:1980:BRB

[Hsu80] D. A. Hsu. Book review: [untitled]. Journal of the American Statistical Association, 75(369):254, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287439.

Hsu:1982:BRD

[Hsu82a] D. A. Hsu. A Bayesian robust detection of shift in the risk structure of stock market returns. Journal of the American Statistical Association, 77(377):29–39, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287766.

Hsu:1982:SIR

[Hsu82b] Jason C. Hsu. Simultaneous inference with respect to the best treatment in block designs. Journal of the American Statistical Association, 77(378):461–467, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287268.

Hsu:1985:BRB

[Hsu85] Jason C. Hsu. Book review: [untitled]. Journal of the American Statistical Association, 80(390):489, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287935.

Hsu:1989:BRB

[Hsu89] Jason C. Hsu. Book review: [untitled]. Journal of the American Statistical Association, 84(406):626–627, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289974.

Hudson:1981:SPE

[HT81] H. Malcolm Hudson and Kam Wah Tsui. Simultaneous Poisson estimators for a priori hypotheses about means. Journal of the American Statistical Association, 76(373):182–187, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287065.

Hillmer:1982:AMB

[HT82] S. C. Hillmer and G. C. Tiao. An ARIMA-model-based approach to seasonal adjustment. Journal of the American Statistical Association, 77(377):63–70, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287770.

Hastie:1987:GAM

[HT87a] Trevor Hastie and Robert Tibshirani. Generalized additive models: Some applications. Journal of the American Statistical Association, 82(398):371–386, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289439.

Hillmer:1987:BET

[HT87b] Steven C. Hillmer and Abdelwahed Trabelsi. Benchmarking of economic time series. Journal of the American Statistical Association, 82(400):1064–1071, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Huber:1983:MABb

[Hub83a] Peter J. Huber. Minimax aspects of bounded-influence regression: Rejoinder. Journal of the American Statistical Association, 78 (381):80, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287112.

Huber:1983:MABa

[Hub83b] Peter J. Huber. Minimax aspects of bounded-influence regression. With discussion. Journal of the American Statistical Association, 78(381):66–72, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287106.

Huber:1987:ETD

[Hub87] Peter J. Huber. Experiences with three-dimensional scatterplots. Journal of the American Statistical Association, 82(398):448–453, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289446.

Hurd:1985:NEM

[Hur85a] Cuthbert C. Hurd. A note on early Monte Carlo computations and scientific meetings. Annals of the History of Computing, 7(2):141–155, April/June 1985. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1985/pdf/a2141.pdf;   http://www.computer.org/
annals/an1985/a2141abs.htm.

Hurvich:1985:DDC

[Hur85b] Clifford M. Hurvich. Data-driven choice of a spectrum estimate: extending the applicability of cross-validation methods. Journal of the American Statistical Association, 80(392):933–940, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288557.

Hartung:1986:BIU

[HV86] Joachim Hartung and Bernard Voet. Best invariant unbiased estimators for the mean squared error of variance component estimators. Journal of the American Statistical Association, 81(395): 689–691, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288997.

Hochberg:1984:SPC

[HVS84] Yosef Hochberg and Yetty Varon-Salomon. On simultaneous pairwise comparisons in analysis of covariance. Journal of the American Statistical Association, 79(388):863–866, December 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288716.

Heiberger:1983:GTD

[HVY83] Richard M. Heiberger, Paul F. Velleman, and M. Agelia Ypelaar. Generating test data with independently controllable features for multivariate general linear forms. Journal of the American Statistical Association, 78(383):585–595, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288124.

Hinich:1984:MED

[HW84] Melvin J. Hinich and Warren E. Weber. A method for estimating distributed lags when observations are randomly missing. Journal of the American Statistical Association, 79(386):368–373, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288277.

Hausman:1985:EVS

[HW85] Jerry A. Hausman and Mark W. Watson. Errors in variables and seasonal adjustment procedures. Journal of the American Statistical Association, 80(391):531–540, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288466.

Hart:1986:KRE

[HW86] Jeffrey D. Hart and Thomas E. Wehrly. Kernel regression estimation using repeated measurements data. Journal of the American Statistical Association, 81(396):1080–1088, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289087.

Hinich:1988:TAU

[HW88] Melvin J. Hinich and Murray A. Wolinsky. A test for aliasing using bispectral analysis. Journal of the American Statistical Association, 83(402):499–502, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288868.

Heckman:1989:FAP

[HW89] James J. Heckman and James R. Walker. Forecasting aggregate period-specific birth rates: The time series properties of a microdynamic neoclassical model of fertility. Journal of the American Statistical Association, 84(408):958–965, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290070.

Hwang:1986:MEV

[Hwa86] Jiunn T. Hwang. Multiplicative errors-in-variables models with applications to recent data released by the U.S. Department of Energy. Journal of the American Statistical Association, 81(395): 680–688, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288996.

Hochberg:1982:GPM

[HWH82] Yosef Hochberg, Gideon Weiss, and Sergiu Hart. On graphical procedures for multiple comparisons. Journal of the American Statistical Association, 77(380):767–772, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287304.

Halperin:1980:CDF

[HWW80] Max Halperin, James H. Ware, and Margaret Wu. Conditional distribution-free tests for the two-sample problem in the presence of right censoring. Journal of the American Statistical Association, 75(371):638–645, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287661.

Hoel:1986:IHC

[HY86] D. G. Hoel and T. Yanagawa. Incorporating historical controls in testing for a trend in proportions. Journal of the American Statistical Association, 81(396):1095–1099, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289089.

Iachan:1982:BRBa

[Iac82a] Ronaldo Iachan. Book review: [untitled]. Journal of the American Statistical Association, 77(377):213, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287795.

Iachan:1982:BRBb

[Iac82b] Ronaldo Iachan. Book review: [untitled]. Journal of the American Statistical Association, 77(380):943–944, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287346.

Lee:1987:CST

[iCL87] Chu in Charles Lee. Chi-squared tests for and against an order restriction on multinomial parameters. Journal of the American Statistical Association, 82(398):611–618, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289472.

Isaki:1982:SDU

[IF82] Cary T. Isaki and Wayne A. Fuller. Survey design under the regression superpopulation model. Journal of the American Statistical Association, 77(377):89–96, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287773.

Iman:1984:CAD

[IHC84] Ronald L. Iman, Stephen C. Hora, and W. J. Conover. Comparison of asymptotically distribution-free procedures for the analysis of complete blocks. Journal of the American Statistical Association, 79(387):674–685, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288416.

Isserman:1985:FIM

[IPRB85] Andrew M. Isserman, David A. Plane, Peter A. Rogerson, and Paul M. Beaumont. Forecasting interstate migration with limited data: a demographic-economic approach. Journal of the American Statistical Association, 80(390):277–285, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287883.

Iwase:1983:UMV

[IS83] Kōsei Iwase and Noriaki Setô. Uniformly minimum variance unbiased estimation for the inverse Gaussian distribution. Journal of the American Statistical Association, 78(383):660–663, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288135.

Izenman:1987:SCR

[IS87] Alan J. Izenman and Sanat K. Sarkar. Simultaneous confidence regions for the frequency analysis of multiple time series. Journal of the American Statistical Association, 82(397):271–275, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289163.

Izenman:1988:PMM

[IS88] Alan J. Izenman and Charles J. Sommer. Philatelic mixtures and multimodal densities. Journal of the American Statistical Association, 83(404):941–953, December 1988. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290118.

Isaacson:1981:BRB

[Isa81] Dean Isaacson. Book review: [untitled]. Journal of the American Statistical Association, 76(374):493, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287861.

Isaki:1983:VEU

[Isa83] Cary T. Isaki. Variance estimation using auxiliary information. Journal of the American Statistical Association, 78(381):117– 123, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287117.

Israel:1981:SPN

[Isr81] Robert B. Israel. Stronger players need not win more knockout tournaments. Journal of the American Statistical Association, 76 (376):950–951, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287594.

Iverson:1985:BRB

[Ive85] Geoffrey J. Iverson. Book review: [untitled]. Journal of the American Statistical Association, 80(390):481–482, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287923.

Iversen:1986:BRBa

[Ive86a] Gudmund R. Iversen. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1116–1117, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289096.

Iversen:1986:BRBb

[Ive86b] Gudmund R. Iversen. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1123, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289104.

Iversen:1987:BRB

[Ive87] Gudmund R. Iversen. Book review: [untitled]. Journal of the American Statistical Association, 82(398):699, June 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289509.

Iversen:1988:BRBa

[Ive88a] Gudmund R. Iversen. Book review: [untitled]. Journal of the American Statistical Association, 83(401):277, March 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288968.

Iversen:1988:BRBb

[Ive88b] Gudmund R. Iversen. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1228–1229, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290183.

Iversen:1989:BRBa

[Ive89a] Gudmund R. Iversen. Book review: [untitled]. Journal of the American Statistical Association, 84(407):838, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289679.

Iversen:1989:BRBb

[Ive89b] Gudmund R. Iversen. Book review: [untitled]. Journal of the American Statistical Association, 84(407):839–840, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289682.

Iyer:1989:BRB

[Iye89] Hari K. Iyer. Book review: [untitled]. Journal of the American Statistical Association, 84(405):333, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289885.

Izenman:1988:BRB

[Ize88] Alan Julian Izenman. Book review: [untitled]. Journal of the American Statistical Association, 83(401):269–270, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288956.

Jacob:1981:RPM

[Jac81] Robert J. K. Jacob. Representing points in many dimensions by trees and castles: Comment. Journal of the American Statistical Association, 76(374):270–272, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287822.

Jacobsen:1984:BRB

[Jac84] Martin Jacobsen. Book review: [untitled]. Journal of the American Statistical Association, 79(388):957–959, December 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288755.

Jacobs:1985:BRB

[Jac85] Eva E. Jacobs. Book review: [untitled]. Journal of the American Statistical Association, 80(389):254, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288106.

Jacka:1987:BRB

[Jac87] S. D. Jacka. Book review: [untitled]. Journal of the American Statistical Association, 82(399):948, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288814.

Jacquez:1989:BRB

[Jac89a] John A. Jacquez. Book review: [untitled]. Journal of the American Statistical Association, 84(406):628–629, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289979.

Jacroux:1989:OBD

[Jac89b] Mike Jacroux. The A-optimality of block designs for comparing test treatments with a control. Journal of the American Statistical Association, 84(405):310–317, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289879.

Jarrett:1986:BRBa

[Jar86a] Jeffrey Jarrett. Book review: [untitled]. Journal of the American Statistical Association, 81(395):874–875, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289049.

Jarrett:1986:BRBb

[Jar86b] Jeffrey Jarrett. Book review: [untitled]. Journal of the American Statistical Association, 81(395):875, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289050.

Jaro:1989:ARL

[Jar89] Matthew A. Jaro. Advances in record-linkage methodology as applied to matching the 1985 census of Tampa, Florida. Journal of the American Statistical Association, 84(406):414–420, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289924.

Jaynes:1981:BRB

[Jay81] E. T. Jaynes. Book review: [untitled]. Journal of the American Statistical Association, 76(375):742, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287543.

Jayachandran:1982:BRB

[Jay82] Toke Jayachandran. Book review: [untitled]. Journal of the American Statistical Association, 77(378):499–500, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287291.

Jebe:1982:BRB

[Jeb82] Emil H. Jebe. Book review: [untitled]. Journal of the American Statistical Association, 77(379):688–689, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287746.

Jennings:1986:JIA

[Jen86a] Dennis E. Jennings. Judging inference adequacy in logistic regression. Journal of the American Statistical Association, 81(394): 471–476, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289237.

Jennings:1986:ORD

[Jen86b] Dennis E. Jennings. Outliers and residual distributions in logistic regression. Journal of the American Statistical Association, 81 (396):987–990, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289072.

Jobson:1980:LSE

[JF80] J. D. Jobson and W. A. Fuller. Least squares estimation when the covariance matrix and parameter vector are functionally related. Journal of the American Statistical Association, 75(369): 176–181, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287408.

Johnson:1983:PVD

[JG83] Wesley Johnson and Seymour Geisser. A predictive view of the detection and characterization of influential observations in regression analysis. Journal of the American Statistical Association, 78 (381):137–144, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287120.

Jong:1983:TRP

[JGM83] Piet De Jong, Malcolm Greig, and Dilip Madan. Testing for random pairing. Journal of the American Statistical Association, 78 (382):332–336, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288637.

Jensen:1982:EFT

[JH82] D. R. Jensen and Y. V. Hui. Efficiency of Friedman’s χ_{r}^{2} test under dependence. Journal of the American Statistical Association, 77 (378):468–474, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287269.

Johnson:1984:END

[JH84] Robert E. Johnson and David G. Herr. The effect of nonorthogonality on the dependence of F ratios sharing a common denominator. Journal of the American Statistical Association, 79(387): 702–708, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288419.

Jamshidian:1988:NAV

[JJ88] Mortaza Jamshidian and Robert I. Jennrich. Nonorthogonal analysis of variance using gradient methods. Journal of the American Statistical Association, 83(402):483–489, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288866.

Juritz:1983:ASP

[JJS83] J. M. Juritz, J. W. F. Juritz, and M. A. Stephens. On the accuracy of simulated percentage points. Journal of the American Statistical Association, 78(382):441–444, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288654.

James:1984:EE

[JJW84a] Barry R. James, Kang Ling James, and Henrique Westenberger. An efficient R-estimator for the ed50. Journal of the American Statistical Association, 79(385):164–173, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

James:1984:EEE

[JJW84b] Barry R. James, Kang Ling James, and Henrique Westenberger. An efficient R-estimator for the ED50. Journal of the American Statistical Association, 79(385):164–173, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288352.

Jobson:1980:EME

[JK80] J. D. Jobson and B. Korkie. Estimation for Markowitz efficient portfolios. Journal of the American Statistical Association, 75 (371):544–554, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287643.

James:1987:ABM

[JK87] I. R. James and M. W. Knuiman. An application of Bayes methodology to the analysis of diary records from a water use study. Journal of the American Statistical Association, 82(399):705–711, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288774.

Jorn:1989:CAU

[JK89] Hongsuk Jorn and Jerome Klotz. The contrast algorithm for unbalanced multifactor analysis of variance. Journal of the American Statistical Association, 84(405):318–324, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289880.

Jolliffe:1983:MCT

[JM83] Ian T. Jolliffe and Bryon J. T. Morgan. A method for comparing two hierarchical clusterings: Comment. Journal of the American Statistical Association, 78(383):580–581, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288120.

Joe:1989:REM

[Joe89] Harry Joe. Relative entropy measures of multivariate dependence. Journal of the American Statistical Association, 84(405): 157–164, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289859.

Johnson:1980:BRB

[Joh80] Mark E. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 75(369):242–243, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287424.

John:1981:BRB

[Joh81] Peter W. M. John. Book review: [untitled]. Journal of the American Statistical Association, 76(373):208–209, March 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287074.

Johnson:1984:ATD

[Joh84a] Richard A. Johnson. The analysis of transformed data: Comment. Journal of the American Statistical Association, 79(386):314–315, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288267.

Johnson:1984:BRB

[Joh84b] Robert A. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 79(388):954–955, December 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288747.

Johnson:1985:BRB

[Joh85] Robert A. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 80(389):236–237, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288083.

Johnson:1987:BRBb

[Joh87a] Douglas H. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 82(398):697, June 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289506.

Johnson:1987:BRBc

[Joh87b] Mark E. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 82(399):946, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288810.

Johnson:1987:BRBa

[Joh87c] Richard A. Johnson. Book review: [untitled]. Journal of the American Statistical Association, 82(397):361, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289202.

Johnstone:1988:HFA

[Joh88] Iain Johnstone. How far are automatically chosen regression smoothing parameters from their optimum?: Comment. Journal of the American Statistical Association, 83(401):99, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288924.

Joiner:1982:BRB

[Joi82] Brian L. Joiner. Book review: [untitled]. Journal of the American Statistical Association, 77(379):692, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287754.

Jones:1982:SED

[Jon82] Douglas H. Jones. Statistical evidence of discrimination: Comment. Journal of the American Statistical Association, 77(380):

787–788, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287307.

Jones:1989:DIK

[Jon89] M. C. Jones. Discretized and interpolated kernel density estimates. Journal of the American Statistical Association, 84(407): 733–741, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289655.

Joe:1985:ENF

[JR85] H. Joe and N. Reid. Estimating the number of faults in a system. Journal of the American Statistical Association, 80(389): 222–226, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288076.

Jennen-Steinmetz:1988:UAN

[JSG88] Christine Jennen-Steinmetz and Theo Gasser. A unifying approach to nonparametric regression estimation. Journal of the American Statistical Association, 83(404):1084–1089, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290140.

Johnson:1981:BDF

[JT81] Mark E. Johnson and Aaron Tenenbein. A bivariate distribution family with specified marginals. Journal of the American Statistical Association, 76(373):198–201, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287067.

Johnson:1980:NFP

[JTB80] Mark E. Johnson, Gary L. Tietjen, and Richard J. Beckman. A new family of probability distributions with applications to Monte Carlo studies. Journal of the American Statistical Association, 75 (370):276–279, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2287446. See correction [JTB81].

Johnson:1981:CNF

[JTB81] Mark E. Johnson, Gary L. Tietjen, and Richard J. Beckman. Correction: “A new family of probability distributions with applications to Monte Carlo studies” [J. Amer. Statist. Assoc. 75 (1980), no. 370, 276–279, MR 81f:62032]. Journal of the American Statistical Association, 76(376):1022, December 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287627. See [FH81c, JTB80].

Juster:1986:REM

[Jus86] F. Thomas Juster. Response errors in the measurement of time use. Journal of the American Statistical Association, 81(394): 390–402, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289228.

Johnstone:1985:ESV

[JV85a] Iain M. Johnstone and Paul F. Velleman. Efficient scores, variance decompositions, and Monte Carlo swindles. Journal of the American Statistical Association, 80(392):851–862, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288544.

Johnstone:1985:RLR

[JV85b] Iain M. Johnstone and Paul F. Velleman. The resistant line and related regression methods. With a comment by John W. Tukey. Journal of the American Statistical Association, 80(392): 1041–1054, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288572.

Johnson:1988:BSM

[JW88] Richard A. Johnson and Samaradasa Weerahandi. A Bayesian solution to the multivariate Behrens–Fisher problem. Journal of the American Statistical Association, 83(401):145–149, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288933.

Kohn:1986:EPI

[KA86] Robert Kohn and Craig F. Ansley. Estimation, prediction, and interpolation for ARIMA models with missing data. Journal of the American Statistical Association, 81(395):751–761, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289007.

Kadane:1981:MDR

[Kad81] Joseph B. Kadane. Modeling demographic relationships: An analysis of forecast functions for Australian births: Comment. Journal of the American Statistical Association, 76(376):792–793, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287569.

Kadane:1982:WCB

[Kad82] Joseph B. Kadane. The well-calibrated Bayesian: Comment. Journal of the American Statistical Association, 77(379):610–611, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287721.

Kadane:1983:JHD

[Kad83] Joseph B. Kadane. Juries hearing death penalty cases: Statistical analysis of a legal procedure. Journal of the American Statistical Association, 78(383):544–552, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288116.

Kadane:1984:BRB

[Kad84] Joseph B. Kadane. Book review: [untitled]. Journal of the American Statistical Association, 79(386):467–469, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288293.

Kafadar:1982:BAO

[Kaf82] Karen Kafadar. A biweight approach to the one-sample problem. Journal of the American Statistical Association, 77(378):416–424, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287262.

Kafadar:1988:BRB

[Kaf88] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 83(401):270–271, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288957.

Kafadar:1989:BRB

[Kaf89] Karen Kafadar. Book review: [untitled]. Journal of the American Statistical Association, 84(406):617–618, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289958.

Kanoh:1988:RWC

[Kan88] Satoru Kanoh. The reduction of the width of confidence bands in linear regression. Journal of the American Statistical Association, 83(401):116–122, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288928.

Kaplan:1983:MCM

[Kap83] Jack S. Kaplan. A method for calculating MINQUE estimators of variance components. Journal of the American Statistical Association, 78(382):476–477, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288662.

Kariya:1980:NCE

[Kar80a] Takeaki Kariya. Note on a condition for equality of sample variances in a linear model. Journal of the American Statistical Association, 75(371):701–703, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287672.

Karr:1980:BRB

[Kar80b] Alan F. Karr. Book review: [untitled]. Journal of the American Statistical Association, 75(371):750–751, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287686.

Kariya:1981:BCM

[Kar81] Takeaki Kariya. Bounds for the covariance matrices of Zellner’s estimator in the SUR model and the 2SAE in a heteroscedastic model. Journal of the American Statistical Association, 76(376): 975–979, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287598.

Karatzas:1985:BRB

[Kar85a] Ioannis Karatzas. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1072, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288588.

Kariya:1985:NVG

[Kar85b] Takeaki Kariya. A nonlinear version of the Gauss–Markov theorem. Journal of the American Statistical Association, 80(390): 476–477, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287918.

Karr:1987:BRB

[Kar87a] Alan F. Karr. Book review: [untitled]. Journal of the American Statistical Association, 82(397):346–347, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289181.

Karrison:1987:RML

[Kar87b] Theodore Karrison. Restricted mean life with adjustment for covariates. Journal of the American Statistical Association, 82(400): 1169–1176, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kass:1982:LPC

[Kas82] Robert E. Kass. Lindley’s paradox: Comment. Journal of the American Statistical Association, 77(378):347–349, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287250.

Kass:1983:BMC

[Kas83] Robert E. Kass. Bayes methods for combining the results of cancer studies in humans and other species: Comment. Journal of the American Statistical Association, 78(382):312–313, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288634.

Kass:1985:BRB

[Kas85] Robert E. Kass. Book review: Numerical Methods for Unconstrained Optimization and Nonlinear Equations, by John E. Dennis, Jr.; Robert B. Schnabel. Journal of the American Statistical Association, 80(389):247–248, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288097.

Katti:1981:BRB

[Kat81] S. K. Katti. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1014, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287615.

Katti:1986:BRB

[Kat86] S. K. Katti. Book review: [untitled]. Journal of the American Statistical Association, 81(393):263–264, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288020.

Kaufman:1981:EVR

[Kau81] Gordon M. Kaufman. Estimating volumes of remaining fossil fuel resources: a critical review: Comment. Journal of the American Statistical Association, 76(375):549–550, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287509.

Kaye:1980:UEB

[Kay80] David Kaye. Using empirical Bayes techniques in the law school validity studies: Comment. Journal of the American Statistical Association, 75(372):818–819, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287164.

Kaye:1981:BRB

[Kay81] David Kaye. Book review: [untitled]. Journal of the American Statistical Association, 76(373):210–211, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287080.

Kaye:1982:SEDa

[Kay82a] David Kaye. Statistical evidence of discrimination. Journal of the American Statistical Association, 77(380):773–783, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287305.

Kaye:1982:SEDb

[Kay82b] David Kaye. Statistical evidence of discrimination: Rejoinder. Journal of the American Statistical Association, 77(380):790–792, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287309.

Kaye:1989:BRB

[Kay89] David H. Kaye. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1094–1095, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290093.

Kshirsagar:1989:BRB

[KBD89] Anant M. Kshirsagar, G. E. P. Box, and N. R. Draper. Book review: Correction: Book Reviews. Journal of the American Statistical Association, 84(405):352, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289914.

Kramer:1987:SAA

[KD87] Walter Krämer and Christian Donninger. Spatial autocorrelation among errors and the relative efficiency of OLS in the linear regression model. Journal of the American Statistical Association, 82(398):577–579, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289467.

Kao:1989:ANB

[KD89] Chihwa Kao and Donald H. Dutkowsky. An application of nonlinear bounded influence estimation to aggregate bank borrowing from the Federal Reserve. Journal of the American Statistical Association, 84(407):700–709, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289651.

Kadane:1980:IEO

[KDW^{+}80] Joseph B. Kadane, James M. Dickey, Robert L. Winkler, Wayne S. Smith, and Stephen C. Peters. Interactive elicitation of opinion for a normal linear model. Journal of the American Statistical Association, 75(372):845–854, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287171.

King:1986:TBE

[KE86] Maxwell L. King and Merran A. Evans. Testing for block effects in regression models based on survey data. Journal of the American Statistical Association, 81(395):677–679, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288995.

Kearfott:1985:SMP

[Kea85] K. J. Kearfott. A statistical model for positron emission tomography: Comment. Journal of the American Statistical Association, 80(389):26–28, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288033.

Kedem:1980:EPS

[Ked80] Benjamin Kedem. Estimation of the parameters in stationary autoregressive processes after hard limiting. Journal of the American Statistical Association, 75(369):146–153, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287403.

Keenan:1982:TSA

[Kee82] Daniel MacRae Keenan. A time series analysis of binary data. Journal of the American Statistical Association, 77(380):816–821, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287312.

Kelsey:1986:BRB

[Kel86] Jennifer L. Kelsey. Book review: [untitled]. Journal of the American Statistical Association, 81(394):576, June 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289266.

Kempthorne:1980:RAE

[Kem80] Oscar Kempthorne. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):584–587, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287650.

Kempthorne:1982:BRB

[Kem82] Oscar Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 77(378):492, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287277.

Kempthorne:1983:BRB

[Kem83] Oscar Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 78(382):482–490, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288664.

Kempthorne:1988:BRB

[Kem88] Peter J. Kempthorne. Book review: [untitled]. Journal of the American Statistical Association, 83(402):567–568, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288888.

Kennedy:1986:BRB

[Ken86] D. P. Kennedy. Book review: [untitled]. Journal of the American Statistical Association, 81(393):266, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288024.

Kestenbaum:1985:MER

[Kes85] Bert Kestenbaum. The measurement of early retirement. Journal of the American Statistical Association, 80(389):38–45, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288038.

Kershner:1981:TTC

[KF81] Ronald P. Kershner and Walter T. Federer. Two-treatment crossover designs for estimating a variety of effects. Journal of the American Statistical Association, 76(375):612–619, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287519.

Kitagawa:1984:SPS

[KG84] Genshiro Kitagawa and Will Gersch. A smoothness priors-state space modeling of time series with trend and seasonality. Journal of the American Statistical Association, 79(386):378–389, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288279.

Kleiner:1981:RPMa

[KH81a] B. Kleiner and J. A. Hartigan. Representing points in many dimensions by trees and castles. Journal of the American Statistical Association, 76(374):260–269, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287820.

Kleiner:1981:RPMb

[KH81b] B. Kleiner and J. A. Hartigan. Representing points in many dimensions by trees and castles: Rejoinder. Journal of the American Statistical Association, 76(374):276, June 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287824.

Kackar:1984:ASE

[KH84] Raghu N. Kackar and David A. Harville. Approximations for standard errors of estimators of fixed and random effects in mixed linear models. Journal of the American Statistical Association, 79 (388):853–862, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288715.

Khuri:1981:SCI

[Khu81] A. I. Khuri. Simultaneous confidence intervals for functions of variance components in random models. Journal of the American Statistical Association, 76(376):878–885, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287583.

Khuri:1989:BRB

[Khu89] Andre I. Khuri. Book review: [untitled]. Journal of the American Statistical Association, 84(406):621–622, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289965.

Kiefer:1986:BRB

[Kie86] Nicholas Kiefer. Book review: [untitled]. Journal of the American Statistical Association, 81(394):571, June 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289258.

Kieffer:1988:BRB

[Kie88] John C. Kieffer. Book review: [untitled]. Journal of the American Statistical Association, 83(402):563–564, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288883.

Kim:1986:LCB

[Kim86] Woo Chul Kim. A lower confidence bound on the probability of a correct selection. Journal of the American Statistical Association, 81(396):1012–1017, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289076.

Kimber:1989:BRBa

[Kim89a] Alan Kimber. Book review: [untitled]. Journal of the American Statistical Association, 84(406):627–628, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289977.

Kimber:1989:BRBb

[Kim89b] Alan Kimber. Book review: [untitled]. Journal of the American Statistical Association, 84(407):845, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289692.

King:1981:BRB

[Kin81] Benjamin King. Book review: [untitled]. Journal of the American Statistical Association, 76(373):212–213, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287083.

King:1983:BRB

[Kin83] Donald A. King. Book review: [untitled]. Journal of the American Statistical Association, 78(381):205–206, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287139.

King:1985:SCP

[Kin85] Benjamin F. King. Surveys combining probability and quota methods of sampling. Journal of the American Statistical Association, 80(392):890–896, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288549.

King:1987:BRB

[Kin87] Benjamin F. King. Book review: [untitled]. Journal of the American Statistical Association, 82(398):688–689, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289497.

King:1989:BRB

[Kin89] Benjamin King. Book review: [untitled]. Journal of the American Statistical Association, 84(406):612–613, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289952.

Kirmani:1984:BRB

[Kir84] Syed N. U. A. Kirmani. Book review: [untitled]. Journal of the American Statistical Association, 79(387):736, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288440.

Kitagawa:1987:NGS

[Kit87] Genshiro Kitagawa. Non-Gaussian state-space modeling of nonstationary time series. With comments and a reply by the author. Journal of the American Statistical Association, 82(400):1032– 1063, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kiviet:1980:EAE

[Kiv80] Jan F. Kiviet. Effects of ARMA errors on tests for regression coefficients: Comments on Vinod’s article; improved and additional results. Journal of the American Statistical Association, 75(370): 353–358, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287458.

Karmel:1987:CPR

[KJ87] T. S. Karmel and Malti Jain. Comparison of purposive and random sampling schemes for estimating capital expenditure. Journal of the American Statistical Association, 82(397):52–57, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289124.

Kanoh:1984:OSS

[KK84a] Satoru Kanoh and Yoshiatsu Kusunoki. One-sided simultaneous bounds in linear regression. Journal of the American Statistical Association, 79(387):715–719, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288422.

Kelton:1984:HTM

[KK84b] W. David Kelton and Christina M. L. Kelton. Hypothesis tests for Markov process models estimated from aggregate frequency data. Journal of the American Statistical Association, 79(388): 922–928, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288725.

Koehler:1980:EIG

[KL80] Kenneth J. Koehler and Kinley Larntz. An empirical investigation of goodness-of-fit statistics for sparse multinomials. Journal of the American Statistical Association, 75(370):336–344, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287455.

Kalbfleisch:1985:APD

[KL85] J. D. Kalbfleisch and J. F. Lawless. The analysis of panel data under a Markov assumption. Journal of the American Statistical Association, 80(392):863–871, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288545.

Kalbfleisch:1989:IBR

[KL89] J. D. Kalbfleisch and J. F. Lawless. Inference based on retrospective ascertainment: An analysis of the data on transfusion-related AIDS. Journal of the American Statistical Association, 84(406): 360–372, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289919.

Kleinbaum:1984:BRB

[Kle84] Robert Kleinbaum. Book review: [untitled]. Journal of the American Statistical Association, 79(388):962–963, December 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288765.

Klugman:1987:BRB

[Klu87] Stuart Klugman. Book review: [untitled]. Journal of the American Statistical Association, 82(399):951–952, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288820.

Klumpner:1988:BRB

[Klu88] James H. Klumpner. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1224–1225, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290177.

Krantz:1983:PLR

[KM83] David H. Krantz and John Miyamoto. Priors and likelihood ratios as evidence. Journal of the American Statistical Association, 78 (382):418–423, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288650.

Kadane:1986:ROE

[KM86] Joseph B. Kadane and Carl N. Morris. Report by the outgoing editors. Journal of the American Statistical Association, 81 (393):9, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287958.

Kmenta:1980:BRB

[Kme80] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1039, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287213.

Kmenta:1982:BRB

[Kme82] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 77(380):939, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287338.

Kmenta:1983:BRBa

[Kme83a] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 78(381):202, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287133.

Kmenta:1983:BRBb

[Kme83b] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 78(381):212, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287150.

Kmenta:1984:BRB

[Kme84] J. Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 79(387):738–739, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288446.

Kmenta:1985:BRBa

[Kme85a] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 80(391):790, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288529.

Kmenta:1985:BRBb

[Kme85b] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1064–1065, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288579.

Kmenta:1985:BRBc

[Kme85c] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1076, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288596.

Kmenta:1988:BRBa

[Kme88a] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 83(401):272–273, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288960.

Kmenta:1988:BRBb

[Kme88b] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 83(401):281–282, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288975.

Kmenta:1989:BRB

[Kme89] Jan Kmenta. Book review: [untitled]. Journal of the American Statistical Association, 84(406):618–619, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289959.

Knapp:1982:EHH

[Kna82] Thomas R. Knapp. Estimating historical heights: Comment. Journal of the American Statistical Association, 77(378):295–296, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287233.

Knoke:1983:BRB

[Kno83] James D. Knoke. Book review: [untitled]. Journal of the American Statistical Association, 78(384):990–991, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288221.

Kobayashi:1985:CES

[Kob85] Masahito Kobayashi. Comparison of efficiencies of several estimators for linear regressions with autocorrelated errors. Journal of the American Statistical Association, 80(392):951–953, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288559.

Kobayashi:1986:BTE

[Kob86] Masahito Kobayashi. A bounds test of equality between sets of coefficients in two linear regressions when disturbance variances are unequal. Journal of the American Statistical Association, 81 (394):510–513, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289242.

Kochar:1989:BRB

[Koc89] Subhash C. Kochar. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1098–1099, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290100.

Koehler:1981:BRB

[Koe81] Anne B. Koehler. Book review: [untitled]. Journal of the American Statistical Association, 76(373):214–215, March 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287087.

Koehler:1982:BRBb

[Koe82a] Anne B. Koehler. Book review: [untitled]. Journal of the American Statistical Association, 77(379):695–696, September 1982.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287758.

Koehler:1982:BRBa

[Koe82b] Kenneth J. Koehler. Book review: [untitled]. Journal of the American Statistical Association, 77(378):494–496, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287281.

Koehler:1986:GFT

[Koe86] Kenneth J. Koehler. Goodness-of-fit tests for log-linear models in sparse contingency tables. Journal of the American Statistical Association, 81(394):483–493, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289239.

Kohler:1988:BRB

[Koh88] Heinz Kohler. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1228, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290182.

Koop:1988:BRB

[Koo88a] J. C. Koop. Book review: [untitled]. Journal of the American Statistical Association, 83(402):578–579, June 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288905.

Koopmans:1988:BRB

[Koo88b] L. H. Koopmans. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1212–1214, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290161.

Korwar:1980:UMV

[Kor80] Ramesh M. Korwar. On the uniformly minimum variance unbiased estimators of the variance and its reciprocal of an inverse Gaussian distribution. Journal of the American Statistical Association, 75(371):734–735, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287679.

Korn:1981:HLL

[Kor81] Edward L. Korn. Hierarchical log-linear models not preserved by classification error. Journal of the American Statistical Association, 76(373):110–113, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287054.

Kallenberg:1985:NCC

[KOS85a] W. C. M. Kallenberg, J. Oosterhoff, and B. F. Schriever. The number of classes in chi-squared goodness-of-fit tests. Journal of the American Statistical Association, 80(392):959–968, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288561.

Kosslyn:1985:BRB

[Kos85b] Stephen M. Kosslyn. Book review: Graphics and Human Information Processing: A Review of Five Books. Journal of the American Statistical Association, 80(391):499–512, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288463.

Kott:1989:CCR

[Kot89] Phillip S. Kott. Comment to: “Conditioning ratio estimates under simple random sampling” [J. Amer. Statist. Assoc. 82 (1987), no. 399, 826–831, MR 88h:62016] by J. Robinson. Journal of the American Statistical Association, 84(405):352, ???? 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Rob87].

Koutrouvelis:1980:RTE

[Kou80] Ioannis A. Koutrouvelis. Regression-type estimation of the parameters of stable laws. Journal of the American Statistical Association, 75(372):918–928, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287182.

Koul:1982:BRB

[Kou82] H. L. Koul. Book review: [untitled]. Journal of the American Statistical Association, 77(380):938, December 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287336.

Koziol:1980:CMT

[Koz80] James A. Koziol. On a Cramer–von Mises-type statistic for testing symmetry. Journal of the American Statistical Association, 75 (369):161–167, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287405.

Kourouklis:1981:CAG

[KP81a] S. Kourouklis and C. C. Paige. A constrained approach to the general Gauss–Markov linear model. Journal of the American Statistical Association, 76(375):620–625, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Kourouklis:1981:CLS

[KP81b] S. Kourouklis and C. C. Paige. A constrained least squares approach to the general Gauss–Markov linear model. Journal of the American Statistical Association, 76(375):620–625, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287520.

Kronmal:1981:VAR

[KP81c] Richard A. Kronmal and Arthur V. Peterson, Jr. A variant of the acceptance-rejection method for computer generation of random variables. Journal of the American Statistical Association, 76(374):446–451, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2287848. See corrigendum [KP82].

Kronmal:1982:CVA

[KP82] Richard A. Kronmal and Arthur V. Peterson, Jr. Corrigendum: “A variant of the acceptance-rejection method for computer generation of random variables” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 446–451, MR 82h:62037]. Journal of the American Statistical Association, 77(380):954, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287364. See [KP81c].

Kenett:1983:SDS

[KP83] Ron Kenett and Moshe Pollak. On sequential detection of a shift in the probability of a rare event. Journal of the American Statistical Association, 78(382):389–395, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288645.

Koenker:1987:ELM

[KP87] Roger Koenker and Stephen Portnoy. L-estimation for linear models. Journal of the American Statistical Association, 82(399):851– 857, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288796.

Kepner:1982:DUM

[KR82] James L. Kepner and Ronald H. Randles. Detecting unequal marginal scales in a bivariate population. Journal of the American Statistical Association, 77(378):475–482, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287270.

Kepner:1984:DFR

[KR84] James L. Kepner and David H. Robinson. A distribution-free rank test for ordered alternatives in randomized complete block designs. Journal of the American Statistical Association, 79(385): 212–217, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288359.

Kleffe:1986:EAU

[KR86] J. Kleffe and J. N. K. Rao. The existence of asymptotically unbiased nonnegative quadratic estimates of variance components in ANOVA models. Journal of the American Statistical Association, 81(395):692–698, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288998.

Kepner:1988:NMD

[KR88] James L. Kepner and David H. Robinson. Nonparametric methods for detecting treatment effects in repeated-measures designs. Journal of the American Statistical Association, 83(402):456–461, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288862.

Kramer:1980:FSE

[Krä80] Walter Krämer. Finite sample efficiency of ordinary least squares in the linear regression model with autocorrelated errors. Journal of the American Statistical Association, 75(372):1005–1009, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287196.

Kramer:1986:LSR

[Krä86] Walter Krämer. Least squares regression when the independent variable follows an ARIMA process. Journal of the American Statistical Association, 81(393):150–154, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287982.

Kraemer:1989:BRB

[Kra89] Helena Chmura Kraemer. Book review: [untitled]. Journal of the American Statistical Association, 84(407):840–841, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289683.

Krewski:1983:BMC

[Kre83] Daniel Krewski. Bayes methods for combining the results of cancer studies in humans and other species: Comment. Journal of the American Statistical Association, 78(382):308–310, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288632.

Kremers:1986:CUE

[Kre86] Walter K. Kremers. Completeness and unbiased estimation for sum-quota sampling. Journal of the American Statistical Association, 81(396):1070–1073, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289085.

Krippendorff:1987:BRB

[Kri87] Klaus Krippendorff. Book review: [untitled]. Journal of the American Statistical Association, 82(397):354–355, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289192.

Kruskal:1980:SFR

[Kru80a] William Kruskal. The significance of Fisher: A review of R. A. Fisher: The Life of a Scientist, and Joan Fisher Box, New York: John Wiley & Sons, 1978. xiv + 512 pp. Journal of the American Statistical Association, 75(372):1019–1030, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287199.

Kruskal:1980:SRG

[Kru80b] William H. Kruskal. The Statistical Research Group, 1942–1945: Comment. Journal of the American Statistical Association, 75 (370):331–333, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287453.

Kruskal:1981:SSP

[Kru81] William Kruskal. Statistics in society: Problems unsolved and unformulated. Journal of the American Statistical Association, 76 (375):505–515, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287503.

Kruskal:1985:BRB

[Kru85] William Kruskal. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1074–1075, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288592.

Kruskal:1988:MSC

[Kru88] William Kruskal. Miracles and statistics: the casual assumption of independence. Journal of the American Statistical Association, 83(404):929–940, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290117.

Krzanowski:1979:BGC

[Krz79] W. J. Krzanowski. Between-groups comparison of principal components. Journal of the American Statistical Association, 74(367): 703–707, September 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2286995. See correction [Krz81].

Krzanowski:1981:CBG

[Krz81] W. J. Krzanowski. Correction: “Between-groups comparison of principal components” [J. Amer. Statist. Assoc. 74 (1979), no. 367, 703–707, MR 81a:62049]. Journal of the American Statistical Association, 76(376):1022, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287629. See [Krz79].

Krzysztofowicz:1987:MFP

[Krz87] Roman Krzysztofowicz. Markovian forecast processes. Journal of the American Statistical Association, 82(397):31–37, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289121.

Koul:1980:TNB

[KS80] H. L. Koul and V. Susarla. Testing for new better than used in expectation with incomplete data. Journal of the American Statistical Association, 75(372):952–956, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287187.

Kass:1989:ABI

[KS89] Robert E. Kass and Duane Steffey. Approximate Bayesian inference in conditionally independent hierarchical models (parametric empirical Bayes models). Journal of the American Statistical Association, 84(407):717–726, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289653.

Kunsch:1989:CUB

[KSC89] Hans R. Künsch, Leonard A. Stefanski, and Raymond J. Carroll. Conditionally unbiased bounded-influence estimation in general regression models, with applications to generalized linear models. Journal of the American Statistical Association, 84(406):460–466, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289930.

Kshirsagar:1988:BRB

[Ksh88] Anant M. Kshirsagar. Book review: [untitled]. Journal of the American Statistical Association, 83(402):569–570, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288890.

Knafl:1985:CBR

[KSY85] George Knafl, Jerome Sacks, and Donald Ylvisaker. Confidence bands for regression functions. Journal of the American Statistical Association, 80(391):683–691, September 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288485.

Kafadar:1988:BT

[KT88] Karen Kafadar and John W. Tukey. A bidec t table. Journal of the American Statistical Association, 83(402):532–539, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288874.

Kuechler:1987:BRB

[Kue87] Manfred Kuechler. Book review: [untitled]. Journal of the American Statistical Association, 82(398):677–678, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289483.

Kuklinski:1987:BRB

[Kuk87] James H. Kuklinski. Book review: [untitled]. Journal of the American Statistical Association, 82(398):687, June 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289495.

Kulkarni:1984:BRB

[Kul84] V. G. Kulkarni. Book review: [untitled]. Journal of the American Statistical Association, 79(388):955–956, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288750.

Kulley:1986:BRB

[Kul86] Andrew M. Kulley. Book review: [untitled]. Journal of the American Statistical Association, 81(393):261, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288016.

Kunitomo:1980:AED

[Kun80] Naoto Kunitomo. Asymptotic expansions of the distributions of estimators in a linear functional relationship and simultaneous equations. Journal of the American Statistical Association, 75 (371):693–700, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287671.

Kuo:1989:CET

[Kuo89] Lynn Kuo. Composite estimation of totals for livestock surveys. Journal of the American Statistical Association, 84(406):421–429, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289925.

Krasker:1982:EBI

[KW82] William S. Krasker and Roy E. Welsch. Efficient boundedinfluence regression estimation. Journal of the American Statistical Association, 77(379):595–604, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287719.

Krasker:1983:MAB

[KW83] William S. Krasker and Roy E. Welsch. Minimax aspects of bounded-influence regression: Comment. Journal of the American Statistical Association, 78(381):72–73, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287107.

Keener:1985:MLR

[KW85] Robert W. Keener and Donald M. Waldman. Maximum likelihood regression of rank-censored data. Journal of the American Statistical Association, 80(390):385–392, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287902.

Kadane:1988:SPE

[KW88] Joseph B. Kadane and Robert L. Winkler. Separating probability elicitation from utilities. Journal of the American Statistical Association, 83(402):357–363, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288850.

Kunitomo:1985:PPM

[KY85] Naoto Kunitomo and Taku Yamamoto. Properties of predictors in misspecified autoregressive time series models. Journal of the American Statistical Association, 80(392):941–950, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288558.

Koch:1986:MTI

[KY86] Paul D. Koch and Shie-Shien Yang. A method for testing the independence of two time series that accounts for a potential pattern in the cross-correlation function. Journal of the American Statistical Association, 81(394):533–544, June 1986. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289246.

Lachenbruch:1989:BRB

[Lac89] Peter A. Lachenbruch. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1096, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290095.

Lad:1984:BRB

[Lad84] Frank Lad. Book review: [untitled]. Journal of the American Statistical Association, 79(387):731–732, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288431.

Lahiff:1989:BRB

[Lah89] Maureen Lahiff. Book review: [untitled]. Journal of the American Statistical Association, 84(405):332, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289883.

Lambert:1981:IFT

[Lam81] Diane Lambert. Influence functions for testing. Journal of the American Statistical Association, 76(375):649–657, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287525.

Lambert:1982:BRB

[Lam82a] Diane Lambert. Book review: [untitled]. Journal of the American Statistical Association, 77(380):938–939, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287337.

Lambert:1982:QRT

[Lam82b] Diane Lambert. Qualitative robustness of tests. Journal of the American Statistical Association, 77(378):352–357, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287252.

Lam:1986:BRB

[Lam86a] David Lam. Book review: [untitled]. Journal of the American Statistical Association, 81(395):869–870, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289040.

Lambert:1986:BRB

[Lam86b] A. Lambert. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1125–1126, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289107.

Lambert:1988:BRB

[Lam88] Diane Lambert. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1226–1227, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290180.

Lamborn:1989:BRB

[Lam89] Kathleen R. Lamborn. Book review: [untitled]. Journal of the American Statistical Association, 84(406):629, June 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289980.

Lane:1980:RAE

[Lan80] David A. Lane. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):587–589, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287651.

Lane:1984:BRB

[Lan84] David A. Lane. Book review: [untitled]. Journal of the American Statistical Association, 79(388):959, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288757.

Landwehr:1985:BRB

[Lan85] James M. Landwehr. Book review: [untitled]. Journal of the American Statistical Association, 80(390):482, June 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287924.

Land:1986:MNP

[Lan86] Kenneth C. Land. Methods for national population forecasts: a review. Journal of the American Statistical Association, 81(396):

888–901, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289057.

Land:1987:BRB

[Lan87] Kenneth C. Land. Book review: [untitled]. Journal of the American Statistical Association, 82(397):352–353, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289190.

Larson:1980:BRB

[Lar80] Harold J. Larson. Book review: [untitled]. Journal of the American Statistical Association, 75(370):475, June 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287491.

Larntz:1984:BRBa

[Lar84a] Kinley Larntz. Book review: Experiments with Mixtures: Designs, Models, and the Analysis of Mixture Data, by John A. Cornell. Journal of the American Statistical Association, 79(385): 232, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288370.

Larntz:1984:BRBb

[Lar84b] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 79(388):954, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288746.

Larntz:1985:BRB

[Lar85] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 80(390):486, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287931.

LaRiccia:1986:ACS

[LaR86a] Vincent N. LaRiccia. Asymptotically chi-squared distributed tests of normality for Type II censored samples. Journal of the American Statistical Association, 81(396):1026–1031, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289078.

Larntz:1986:BRBa

[Lar86b] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 81(393):251–252, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288001.

Larntz:1986:BRBb

[Lar86c] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 81(394):577, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289270.

Larntz:1987:BRB

[Lar87] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 82(399):960, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288835.

Larntz:1989:BRBa

[Lar89a] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 84(406):624, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289969.

Larntz:1989:BRBb

[Lar89b] Kinley Larntz. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1101–1102, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290105.

Laskey:1984:BRB

[Las84] Kathryn Blackmond Laskey. Book review: [untitled]. Journal of the American Statistical Association, 79(387):741–742, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288452.

Latta:1981:MCS

[Lat81] Robert B. Latta. A Monte Carlo study of some two-sample rank tests with censored data. Journal of the American Statistical Association, 76(375):713–719, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287536.

Lawless:1981:MSE

[Law81] J. F. Lawless. Mean squared error properties of generalized ridge estimators. Journal of the American Statistical Association, 76 (374):462–466, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2287850. See corrigendum [Law82].

Lawless:1982:CMS

[Law82] J. F. Lawless. Corrigendum: “Mean squared error properties of generalized ridge estimators” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 462–466, MR 82h:62113]. Journal of the American Statistical Association, 77(380):954, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287365. See [Law81].

Lawless:1987:RMP

[Law87] J. F. Lawless. Regression methods for Poisson process data. Journal of the American Statistical Association, 82(399):808–815, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288790.

Lawrance:1988:RTD

[Law88] Anthony J. Lawrance. Regression transformation diagnostics using local influence. Journal of the American Statistical Association, 83(404):1067–1072, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290137.

Lax:1985:RES

[Lax85] David A. Lax. Robust estimators of scale: Finite-sample performance in long-tailed symmetric distributions. Journal of the American Statistical Association, 80(391):736–741, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288493.

Lindstrom:1988:NRE

[LB88] Mary J. Lindstrom and Douglas M. Bates. Newton–Raphson and EM algorithms for linear mixed-effects models for repeatedmeasures data. Journal of the American Statistical Association, 83 (404):1014–1022, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290128.

Li:1985:CPP

[LC85a] Guoying Li and Zhonglian Chen. Correction: Projection-pursuit approach to robust dispersion matrices and principal components: Primary theory and Monte Carlo. Journal of the American Statistical Association, 80(392):1084, December 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288612.

Li:1985:PPA

[LC85b] Guoying Li and Zhonglian Chen. Projection-pursuit approach to robust dispersion matrices and principal components: Primary theory and Monte Carlo. Journal of the American Statistical Association, 80(391):759–766, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288497.

Leamer:1981:CFR

[Lea81] Edward E. Leamer. Coordinate-free ridge regression bounds. Journal of the American Statistical Association, 76(376):842–849, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287579.

Leamer:1983:BRB

[Lea83] Edward E. Leamer. Book review: [untitled]. Journal of the American Statistical Association, 78(384):997–998, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288231.

Leamer:1984:GSR

[Lea84] Edward E. Leamer. Global sensitivity results for generalized least squares estimates. Journal of the American Statistical Association, 79(388):867–870, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288717.

Ledolter:1981:BRBa

[Led81a] Johannes Ledolter. Book review: [untitled]. Journal of the American Statistical Association, 76(375):749–750, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287555.

Ledolter:1981:BRBb

[Led81b] Johannes Ledolter. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1017–1018, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287621.

Lee:1980:TTC

[Lee80] Young Jack Lee. Test of trend in count data: Multinomial distribution case. Journal of the American Statistical Association, 75 (372):1010–1014, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287197.

Lee:1981:MDR

[Lee81] Ronald Lee. Modeling demographic relationships: An analysis of forecast functions for Australian births: Comment. Journal of the American Statistical Association, 76(376):793–795, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287570.

Lee:1988:PEG

[Lee88] Jack C. Lee. Prediction and estimation of growth curves with special covariance structures. Journal of the American Statistical Association, 83(402):432–440, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288859.

Lehmann:1981:ICB

[Leh81a] E. L. Lehmann. An interpretation of completeness and Basu’s theorem. Journal of the American Statistical Association, 76(374): 335–340, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287834.

Lehoczky:1981:BRB

[Leh81b] John P. Lehoczky. Book review: [untitled]. Journal of the American Statistical Association, 76(375):743, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287546.

Lehoczky:1982:BRB

[Leh82] J. P. Lehoczky. Book review: [untitled]. Journal of the American Statistical Association, 77(377):220–221, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287806.

Lehmann:1983:EII

[Leh83] E. L. Lehmann. Estimation with inadequate information. Journal of the American Statistical Association, 78(383):624–627, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288130.

Lehoczky:1984:BRB

[Leh84] John P. Lehoczky. Book review: [untitled]. Journal of the American Statistical Association, 79(388):956, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288751.

Lehoczky:1989:BRB

[Leh89] John P. Lehoczky. Book review: [untitled]. Journal of the American Statistical Association, 84(407):848–849, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289698.

Lenk:1986:BRB

[Len86] Peter J. Lenk. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1123–1124, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289105.

Lenk:1988:LND

[Len88] Peter J. Lenk. The logistic normal distribution for Bayesian, nonparametric, predictive densities. Journal of the American Statistical Association, 83(402):509–516, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288870.

Leonard:1982:SPD

[Leo82] Tom Leonard. A simple predictive density function: Comment. Journal of the American Statistical Association, 77(379):657–658, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287731.

Leonard:1983:PEB

[Leo83] Tom Leonard. Parametric empirical Bayes inference: Theory and applications: Comment. Journal of the American Statistical Association, 78(381):59–60, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287102.

Leonard:1989:BRB

[Leo89] Tom Leonard. Book review: [untitled]. Journal of the American Statistical Association, 84(406):625–626, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289973.

Lepage:1985:BRB

[Lep85] Yves Lepage. Book review: [untitled]. Journal of the American Statistical Association, 80(391):780, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288512.

Leurgans:1985:BRB

[Leu85] Sue Leurgans. Book review: [untitled]. Journal of the American Statistical Association, 80(389):245, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288094.

Levin:1987:BRB

[Lev87] Bruce Levin. Book review: [untitled]. Journal of the American Statistical Association, 82(399):958, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288832.

Lejeune:1982:SPDb

[LF82a] Michel Lejeune and G. David Faulkenberry. A simple predictive density function: Rejoinder. Journal of the American Statistical Association, 77(379):659, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287732.

Lejeune:1982:SPDa

[LF82b] Michel Lejeune and G. David Faulkenberry. A simple predictive density function. With a comment by Tom Leonard and a rejoinder by the authors. Journal of the American Statistical Association, 77(379):654–657, September 1982. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287730.

Lepkowski:1986:MSE

[LG86] James M. Lepkowski and Robert M. Groves. A mean squared error model for dual frame, mixed mode survey design. Journal of the American Statistical Association, 81(396):930–937, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289062.

Losh-Hesselbart:1985:BRB

[LH85] Susan Losh-Hesselbart. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1077–1078, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288598.

Land:1986:IAI

[LH86] Kenneth C. Land and George C. Hough, Jr. Improving the accuracy of intercensal estimates and postcensal projections of the civilian noninstitutional population: a parameterization of institutional prevalence rates. Journal of the American Statistical Association, 81(393):62–74, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287968.

Land:1989:NMT

[LH89] Kenneth C. Land and George C. Hough, Jr. New methods for tables of school life, with applications to U.S. data from recent school years. Journal of the American Statistical Association, 84 (405):63–75, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289847.

Leonard:1989:BMI

[LHT89] Tom Leonard, John S. J. Hsu, and Kam-Wah Tsui. Bayesian marginal inference. Journal of the American Statistical Association, 84(408):1051–1058, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290082.

Lii:1980:BRB

[Lii80] Keh-Shin Lii. Book review: [untitled]. Journal of the American Statistical Association, 75(369):241, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287421.

Lindley:1980:CSR

[Lin80a] D. V. Lindley. A critique of some ridge regression methods: Comment. Journal of the American Statistical Association, 75(369): 94–95, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287390.

Lindley:1980:RAE

[Lin80b] D. V. Lindley. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):589–590, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287652.

Ling:1980:BRB

[Lin80c] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 75(369):253–254, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287438.

Ling:1981:BRB

[Lin81a] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 76(374):496–497, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287870.

Ling:1981:CSA

[Lin81b] Robert F. Ling. Corrigenda: a study of the accuracy of some approximations for t, χ^{2}, F tail probabilities. Journal of the American Statistical Association, 76(376):1022, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287628.

Lindley:1982:LPC

[Lin82a] D. V. Lindley. Lindley’s paradox: Comment. Journal of the American Statistical Association, 77(378):334–336, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287245.

Lindley:1982:WCB

[Lin82b] D. V. Lindley. The well-calibrated Bayesian: Comment. Journal of the American Statistical Association, 77(379):611–612, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287722.

Ling:1982:BRB

[Lin82c] Robert F. Ling. Book review: Correlation and Causation, by David A. Kenny. Journal of the American Statistical Association, 77(378):489–491, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287275.

Lindley:1983:PEB

[Lin83a] Dennis V. Lindley. Parametric empirical Bayes inference: Theory and applications: Comment. Journal of the American Statistical Association, 78(381):61–62, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287103.

Ling:1983:BRBa

[Lin83b] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 78(381):209–210, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287144.

Ling:1983:BRBb

[Lin83c] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 78(382):511–512, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288699.

Ling:1984:BRBa

[Lin84a] Robert F. Ling. Book review: Applied Statistics: Principles and Examples, by D. R. Cox; E. J. Snell. Journal of the American Statistical Association, 79(385):229–231, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288367.

Ling:1984:BRBb

[Lin84b] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 79(388):948–949, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288737.

Lindsay:1985:EII

[Lin85] Bruce G. Lindsay. Errors in inspection: integer parameter maximum likelihood in a finite population. Journal of the American Statistical Association, 80(392):879–885, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288547.

Ling:1986:BRB

[Lin86] Robert F. Ling. Book review: [untitled]. Journal of the American Statistical Association, 81(394):574–575, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289264.

Lindsay:1987:BRB

[Lin87] Bruce G. Lindsay. Book review: [untitled]. Journal of the American Statistical Association, 82(398):694–695, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289504.

Lindsay:1989:BRB

[Lin89a] Bruce Lindsay. Book review: [untitled]. Journal of the American Statistical Association, 84(405):337–338, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289892.

Link:1989:MIC

[Lin89b] William A. Link. A model for informative censoring. Journal of the American Statistical Association, 84(407):749–752, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289657.

Little:1981:MDR

[Lit81] Roderick J. A. Little. Modeling demographic relationships: An analysis of forecast functions for Australian births: Comment. Journal of the American Statistical Association, 76(376):795–796, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287571.

Little:1982:MNS

[Lit82] Roderick J. A. Little. Models for nonresponse in sample surveys. Journal of the American Statistical Association, 77(378):237–250, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287227.

Little:1983:EFP

[Lit83a] Roderick J. A. Little. Estimating a finite population mean from unequal probability samples. Journal of the American Statistical Association, 78(383):596–604, September 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288125.

Little:1983:EMD

[Lit83b] Roderick J. A. Little. An evaluation of model-dependent and probability-sampling inferences in sample surveys: Comment. Journal of the American Statistical Association, 78(384):797–799, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288184.

Litterman:1986:BRB

[Lit86] Robert B. Litterman. Book review: [untitled]. Journal of the American Statistical Association, 81(393):262–263, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288019.

Little:1988:TMC

[Lit88] Roderick J. A. Little. A test of missing completely at random for multivariate data with missing values. Journal of the American Statistical Association, 83(404):1198–1202, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290157.

Liu:1983:EHC

[Liu83] Ben-Chieh Liu. Evaluating a hospital cost-containment program in a paired experiment. Journal of the American Statistical Association, 78(382):249–255, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288619.

Littell:1981:CRB

[LL81] Ramon C. Littell and William C. Louv. Confidence regions based on methods of combining test statistics. Journal of the American Statistical Association, 76(373):125–130, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287057.

Louv:1986:COS

[LL86] William C. Louv and Ramon C. Littell. Combining one-sided binomial tests. Journal of the American Statistical Association, 81(394):550–554, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289248.

Laird:1987:EBCb

[LL87a] Nan M. Laird and Thomas A. Louis. Empirical Bayes confidence intervals based on bootstrap samples: Rejoinder. Journal of the American Statistical Association, 82(399):756–757, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288783.

Laird:1987:EBCa

[LL87b] Nan M. Laird and Thomas A. Louis. Empirical Bayes confidence intervals based on bootstrap samples. With discussion and with a reply by the authors. Journal of the American Statistical Association, 82(399):739–750, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288778.

Lange:1989:ECS

[LL89] Nicholas Lange and Nan M. Laird. The effect of covariance structure on variance estimation in balanced growth-curve models with random parameters. Journal of the American Statistical Association, 84(405):241–247, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289869.

Laird:1987:MLC

[LLS87] Nan Laird, Nicholas Lange, and Daniel Stram. Maximum likelihood computations with repeated measures: application of the EM algorithm. Journal of the American Statistical Association, 82(397):97–105, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289129.

Lange:1989:RSM

[LLT89] Kenneth L. Lange, Roderick J. A. Little, and Jeremy M. G. Taylor. Robust statistical modeling using the t distribution. Journal of the American Statistical Association, 84(408):881–896, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290063.

Leeuw:1983:TTS

[LM83] Frank De Leeuw and Michael J. McKelvey. A “true” time series and its indicators. Journal of the American Statistical Association, 78(381):37–46, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287097.

Laird:1985:SMP

[LM85a] Nan M. Laird and Edward F. Maher. A statistical model for positron emission tomography: Comment. Journal of the American Statistical Association, 80(389):29–30, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288034.

Laska:1985:VAO

[LM85b] Eugene M. Laska and Morris Meisner. A variational approach to optimal two-treatment crossover designs: Application to carryover-effect models. Journal of the American Statistical Association, 80(391):704–710, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288488.

Lustbader:1985:DSS

[LM85c] Edward D. Lustbader and Suresh H. Moolgavkar. A diagnostic statistic for the score test. Journal of the American Statistical Association, 80(390):375–379, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287900.

Lefkopoulou:1989:AMC

[LMR89] Myrto Lefkopoulou, Dirk Moore, and Louise Ryan. The analysis of multiple correlated binary outcomes: Application to rodent teratology experiments. Journal of the American Statistical Association, 84(407):810–815, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289671.

Leitch:1981:IUM

[LNPS81] Robert A. Leitch, John Neter, Robert Plante, and Prabhakant Sinha. Implementation of upper multinomial bound using clustering. Journal of the American Statistical Association, 76(375): 530–533, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287506.

Laird:1981:CAC

[LO81] Nan Laird and Donald Olivier. Covariance analysis of censored survival data using log-linear analysis techniques. Journal of the American Statistical Association, 76(374):231–240, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287816.

Loh:1985:NMT

[Loh85] Wei-Yin Loh. A new method for testing separate families of hypotheses. Journal of the American Statistical Association, 80 (390):362–368, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287898.

Loh:1986:IER

[Loh86] Wei-Yin Loh. Improved estimators for ratios of variance components. Journal of the American Statistical Association, 81(395): 699–702, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288999.

Loh:1987:CCC

[Loh87] Wei-Yin Loh. Calibrating confidence coefficients. Journal of the American Statistical Association, 82(397):155–162, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289142.

Lohr:1989:BRB

[Loh89] Sharon Lohr. Book review: [untitled]. Journal of the American Statistical Association, 84(406):613–614, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289953.

Long:1981:MDR

[Lon81] John F. Long. Modeling demographic relationships: An analysis of forecast functions for Australian births: Comment. Journal of the American Statistical Association, 76(376):796–798, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287572.

Lorenz:1986:BRB

[Lor86] Frederick Lorenz. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1126–1127, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289108.

Los:1985:DGL

[Los85] Cornelis A. Los. Dynamic generalized linear models and Bayesian forecasting: Comment. Journal of the American Statistical Association, 80(389):92–93, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288047.

Louis:1984:EPP

[Lou84] Thomas A. Louis. Estimating a population of parameter values using Bayes and empirical Bayes methods. Journal of the American Statistical Association, 79(386):393–398, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288281.

Low:1982:BRB

[Low82] Leone Y. Low. Book review: [untitled]. Journal of the American Statistical Association, 77(378):498, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287286.

Low:1983:BRBa

[Low83a] Leone Y. Low. Book review: [untitled]. Journal of the American Statistical Association, 78(381):212, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287149.

Low:1983:BRBb

[Low83b] Leone Y. Low. Book review: [untitled]. Journal of the American Statistical Association, 78(384):987, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288218.

Ling:1984:APA

[LP84] Robert F. Ling and John W. Pratt. The accuracy of Peizer approximations to the hypergeometric distribution, with comparisons to some other approximations. Journal of the American Statistical Association, 79(385):49–60, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288333.

Levy:1986:OPF

[LP86] Martin S. Levy and S. K. Perng. An optimal prediction function for the normal linear model. Journal of the American Statistical Association, 81(393):196–198, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287989.

Landwehr:1984:GMAa

[LPS84a] James M. Landwehr, Daryl Pregibon, and Anne C. Shoemaker. Graphical methods for assessing logistic regression models. Journal of the American Statistical Association, 79(385):61–71, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288334.

Landwehr:1984:GMAb

[LPS84b] James M. Landwehr, Daryl Pregibon, and Anne C. Shoemaker. Graphical methods for assessing logistic regression models: Rejoinder. Journal of the American Statistical Association, 79(385): 81–83, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288339.

Lancaster:1985:NTL

[LQ85] J. F. Lancaster and Dana Quade. A nonparametric test for linear regression based on combining Kendall’s tau with the sign test. Journal of the American Statistical Association, 80(390):393–397, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287903.

Lindsay:1987:UTI

[LR87] Bruce G. Lindsay and Kathryn Roeder. A unified treatment of integer parameter models. Journal of the American Statistical Association, 82(399):758–764, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288784.

Liao:1983:SSS

[LS83] H. Liao and J. Sedransk. Selection of strata sample sizes for the comparison of domain means. Journal of the American Statistical Association, 78(384):870–878, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288198.

Lindley:1986:RFT

[LS86] Dennis V. Lindley and Nozer D. Singpurwalla. Reliability (and fault tree) analysis using expert opinions. Journal of the American Statistical Association, 81(393):87–90, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287971.

Little:1987:EIQ

[LS87] Roderick J. A. Little and Philip J. Smith. Editing and imputation for quantitative survey data. Journal of the American Statistical Association, 82(397):58–68, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289125.

Lewandowsky:1989:DSS

[LS89] Stephan Lewandowsky and Ian Spence. Discriminating strata in scatterplots. Journal of the American Statistical Association, 84 (407):682–688, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289649.

Limam:1988:STI

[LT88] Mohamed M. T. Limam and David R. Thomas. Simultaneous tolerance intervals for the linear regression model. Journal of the American Statistical Association, 83(403):801–804, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Lemeshow:1988:POI

[LTAP88] Stanley Lemeshow, Daniel Teres, Jill Spitz Avrunin, and Harris Pastides. Predicting the outcome of intensive care unit patients. Journal of the American Statistical Association, 83(402):348–356, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288849.

Lucas:1982:BRB

[Luc82] William F. Lucas. Book review: [untitled]. Journal of the American Statistical Association, 77(377):215–216, March 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287799.

Lundborg:1984:BRB

[Lun84] Per Lundborg. Book review: An Integrated System of World Models, by Ram Dayal. Journal of the American Statistical Association, 79(385):237–238, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288380.

Lurie:1982:BRB

[Lur82] Dan Lurie. Book review: [untitled]. Journal of the American Statistical Association, 77(378):497, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287284.

Lustbader:1985:BRB

[Lus85] Edward Lustbader. Book review: [untitled]. Journal of the American Statistical Association, 80(390):494–495, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287951.

Lutkepohl:1980:AAD

[Lut80] Helmut Lutkepohl. Approximation of arbitrary distributed lag structures by a modified polynomial lag: An extension. Journal of the American Statistical Association, 75(370):428–430, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287471.

Loh:1988:TSC

[LV88] Wei-Yin Loh and Nunta Vanichsetakul. Tree-structured classification via generalized discriminant analysis. With a comment by Leo Breiman and Jerome H. Friedman and with a reply by the authors. Journal of the American Statistical Association, 83(403): 715–728, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

LaRiccia:1985:APF

[LW85] Vincent N. LaRiccia and Thomas E. Wehrly. Asymptotic properties of a family of minimum quantile distance estimators. Journal of the American Statistical Association, 80(391):742–747, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288494.

Loh:1987:BBC

[LW87] Wei-Yin Loh and C. F. J. Wu. Better bootstrap confidence intervals: Comment. Journal of the American Statistical Association, 82(397):188–190, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289147.

Lin:1989:RIC

[LW89] D. Y. Lin and L. J. Wei. The robust inference for the Cox proportional hazards model. Journal of the American Statistical Association, 84(408):1074–1078, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290085.

Lagakos:1986:ACWa

[LWZ86a] S. W. Lagakos, B. J. Wessen, and M. Zelen. An analysis of contaminated well water and health effects in Woburn, Massachusetts. Journal of the American Statistical Association, 81(395):583–596, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288982.

Lagakos:1986:ACWb

[LWZ86b] S. W. Lagakos, B. J. Wessen, and M. Zelen. An analysis of contaminated well water and health effects in Woburn, Massachusetts: Rejoinder. Journal of the American Statistical Association, 81 (395):611–614, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288988.

Liang:1989:CLR

[LZ89] Kung-Yee Liang and Scott L. Zeger. A class of logistic regression models for multivariate binary time series. Journal of the American Statistical Association, 84(406):447–451, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289928.

Mack:1986:EWA

[Mac86a] Michael E. Mack. Efficiencies of weighted averages in stationary autoregressive processes. Journal of the American Statistical Association, 81(395):730–735, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289004. See correction [Mac87].

MacMahon:1986:ACW

[Mac86b] Brian MacMahon. An analysis of contaminated well water and health effects in Woburn, Massachusetts: Comment. Journal of the American Statistical Association, 81(395):597–599, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288983.

Mack:1987:CEW

[Mac87] Michael E. Mack. Correction to: “Efficiencies of weighted averages in stationary autoregressive processes” [J. Amer. Statist. Assoc. 81 (1986), no. 395, 730–735, MR 87j:62174]. Journal of the American Statistical Association, 82(398):704, ???? 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Mac86a].

MacCallum:1989:BRB

[Mac89] Robert C. MacCallum. Book review: [untitled]. Journal of the American Statistical Association, 84(406):622–623, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289967.

Madansky:1985:BRB

[Mad85] Albert Madansky. Book review: [untitled]. Journal of the American Statistical Association, 80(391):774–775, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288504.

Madsen:1989:BRB

[Mad89] Richard Madsen. Book review: [untitled]. Journal of the American Statistical Association, 84(405):335, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289889.

Majumdar:1988:BRB

[Maj88] Dibyen Majumdar. Book review: [untitled]. Journal of the American Statistical Association, 83(401):275, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288964.

Mallows:1981:GMT

[Mal81] C. L. Mallows. Graphics for the multivariate two-sample problem: Comment. Journal of the American Statistical Association, 76 (374):294, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287829.

Malley:1982:SCI

[Mal82] James D. Malley. Simultaneous confidence intervals for ratios of normal means. Journal of the American Statistical Association, 77 (377):170–176, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287785.

Mallows:1983:MAB

[Mal83] C. L. Mallows. Minimax aspects of bounded-influence regression: Comment. Journal of the American Statistical Association, 78 (381):77, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287110.

Mallows:1985:BRB

[Mal85] C. L. Mallows. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1067, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288582.

Malik:1988:BRB

[Mal88a] Henrick J. Malik. Book review: [untitled]. Journal of the American Statistical Association, 83(401):269, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288955.

Mallows:1988:BVS

[Mal88b] C. L. Mallows. Bayesian variable selection in linear regression: Comment. Journal of the American Statistical Association, 83 (404):1034–1035, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290131.

Mandrekar:1984:BRB

[Man84] V. Mandrekar. Book review: [untitled]. Journal of the American Statistical Association, 79(387):735, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288438.

Manski:1988:IBR

[Man88] Charles F. Manski. Identification of binary response models. Journal of the American Statistical Association, 83(403):729–738, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mare:1980:SBS

[Mar80a] Robert D. Mare. Social background and school continuation decisions. Journal of the American Statistical Association, 75(370): 295–305, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287448.

Marquardt:1980:CSC

[Mar80b] D. W. Marquardt. Comment on Smith and Campbell (1980). Journal of the American Statistical Association, 75(369):87–91, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Marquardt:1980:CSR

[Mar80c] Donald W. Marquardt. A critique of some ridge regression methods: Comment. Journal of the American Statistical Association, 75(369):87–91, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287388.

Martin:1981:SPB

[Mar81a] Margaret E. Martin. Statistical practice in bureaucracies. Journal of the American Statistical Association, 76(373):1–8, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287032.

Martinsek:1981:NVH

[Mar81b] Adam T. Martinsek. A note on the variance and higher central moments of the stopping time of an SPRT. Journal of the American Statistical Association, 76(375):701–703, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287533.

Maravall:1984:BRB

[Mar84a] Agustin Maravall. Book review: [untitled]. Journal of the American Statistical Association, 79(387):737–738, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288444.

Marsaglia:1984:EAM

[Mar84b] George Marsaglia. The exact-approximation method for generating random variables in a computer. Journal of the American Statistical Association, 79(385):218–221, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288360.

Maravall:1986:RAS

[Mar86] Agustín Maravall. Revisions in ARIMA signal extraction. Journal of the American Statistical Association, 81(395):736–740, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289005.

Marquardt:1987:IS

[Mar87a] Donald W. Marquardt. The importance of statisticians. Journal of the American Statistical Association, 82(397):1–7, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289118.

Martin:1987:BRB

[Mar87b] Margaret E. Martin. Book review: [untitled]. Journal of the American Statistical Association, 82(398):681–682, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289488.

Martinsek:1988:NRO

[Mar88] Adam T. Martinsek. Negative regret, optional stopping, and the elimination of outliers. Journal of the American Statistical Association, 83(401):160–163, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288935.

Maskus:1985:BRB

[Mas85] Keith E. Maskus. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1062–1063, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288576.

Mason:1988:BRB

[Mas88] Robert L. Mason. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1215–1216, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290164.

Mathew:1983:LEI

[Mat83] Thomas Mathew. Linear estimation with an incorrect dispersion matrix in linear models with a common linear part. Journal of the American Statistical Association, 78(382):468–471, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288660.

Mathew:1989:BRB

[Mat89] Thomas Mathew. Book review: [untitled]. Journal of the American Statistical Association, 84(406):622, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289966.

Mauskopf:1983:RRC

[Mau83] Josephine Mauskopf. Reproductive response to child mortality: a maximum likelihood estimation model. Journal of the American Statistical Association, 78(382):238–248, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288618.

Maxwell:1988:BRB

[Max88] Scott E. Maxwell. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1214–1215, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290163.

Mayer:1981:EVR

[May81] Lawrence S. Mayer. Estimating volumes of remaining fossil fuel resources: a critical review: Comment. Journal of the American Statistical Association, 76(375):551–554, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287510.

Mayer:1983:BRBc

[May83a] Lawrence S. Mayer. Book review: Corrigenda: [Book Review]. Journal of the American Statistical Association, 78(384):1008, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288253.

Mayer:1983:BRBa

[May83b] Lawrence S. Mayer. Book review: [untitled]. Journal of the American Statistical Association, 78(382):509–510, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288694.

Mayer:1983:BRBb

[May83c] Lawrence S. Mayer. Book review: [untitled]. Journal of the American Statistical Association, 78(383):734, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288151.

Mayer:1984:BRB

[May84] Lawrence S. Mayer. Book review: [untitled]. Journal of the American Statistical Association, 79(387):725–727, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288427.

Mehrotra:1982:CIJ

[MB82] Kishan G. Mehrotra and Gouri K. Bhattacharyya. Confidence intervals with jointly type-II censored samples from two exponential distributions. Journal of the American Statistical Association, 77 (378):441–446, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287265.

Mitchell:1988:BVSb

[MB88a] T. J. Mitchell and J. J. Beauchamp. Bayesian variable selection in linear regression: Rejoinder. Journal of the American Statistical Association, 83(404):1035–1036, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290132.

Mitchell:1988:BVSa

[MB88b] T. J. Mitchell and J. J. Beauchamp. Bayesian variable selection in linear regression. With comments by James Berger and C. L. Mallows and with a reply by the authors. Journal of the American Statistical Association, 83(404):1023–1032, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290129.

Monlezun:1988:GNS

[MB88c] Charles J. Monlezun and David C. Blouin. A general nested splitplot analysis of covariance. Journal of the American Statistical Association, 83(403):818–823, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Muller:1989:APR

[MB89] Keith E. Muller and Curtis N. Barton. Approximate power for repeated-measures ANOVA lacking sphericity. Journal of the American Statistical Association, 84(406):549–555, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289941.

McKnight:1984:TDT

[MC84] Barbara McKnight and John Crowley. Tests for differences in tumor incidence based on animal carcinogenesis experiments. Journal of the American Statistical Association, 79(387):639–648, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288411.

McClure:1981:BRB

[McC81a] Richard McClure. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1016, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287619.

Mcconway:1981:MLO

[Mcc81b] K. J. Mcconway. Marginalization and linear opinion pools. Journal of the American Statistical Association, 76(374):410–414, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287843.

McCulloch:1982:SMD

[McC82] Charles E. McCulloch. Symmetric matrix derivatives with applications. Journal of the American Statistical Association, 77(379): 679–682, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287736.

McCleary:1983:BRB

[McC83] Richard McCleary. Book review: [untitled]. Journal of the American Statistical Association, 78(383):738–739, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288160.

McCormick:1984:BRB

[McC84] William P. McCormick. Book review: [untitled]. Journal of the American Statistical Association, 79(386):484–485, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288321.

McCullagh:1986:CDG

[McC86] Peter McCullagh. The conditional distribution of goodness-of-fit statistics for discrete data. Journal of the American Statistical Association, 81(393):104–107, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287974.

McConnell:1987:BRB

[McC87] Terry R. McConnell. Book review: [untitled]. Journal of the American Statistical Association, 82(398):700, June 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289511.

McCulloch:1988:BRB

[McC88] Charles E. McCulloch. Book review: [untitled]. Journal of the American Statistical Association, 83(401):267–268, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288952.

McCullagh:1989:SSP

[McC89a] Peter McCullagh. Some statistical properties of a family of continuous univariate distributions. Journal of the American Statistical Association, 84(405):125–129, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289855.

McCulloch:1989:LMI

[McC89b] Robert E. McCulloch. Local model influence. Journal of the American Statistical Association, 84(406):473–478, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289932.

McDonald:1980:BTS

[McD80] John McDonald. Births time series models and structural interpretations. Journal of the American Statistical Association, 75 (369):39–41, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287376.

McDonald:1981:MDRa

[McD81a] John McDonald. Modeling demographic relationships: An analysis of forecast functions for Australian births. Journal of the American Statistical Association, 76(376):782–792, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287568.

McDonald:1981:MDRb

[McD81b] John McDonald. Modeling demographic relationships: An analysis of forecast functions for Australian births: Rejoinder. Journal of the American Statistical Association, 76(376):799–801, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287574.

McDonald:1989:BRB

[McD89] Gary C. McDonald. Book review: [untitled]. Journal of the American Statistical Association, 84(407):845–846, September 1989.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289693.

McKenzie:1984:BRB

[McK84] John McKenzie. Book review: The Professional Forecaster: The Forecasting Process Through Data Analysis, by James P. Cleary; Hans Levenbach. Journal of the American Statistical Association, 79(385):234–235, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288376.

McKean:1988:BRB

[McK88] Joseph W. McKean. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1217, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290166.

McLean:1980:BRB

[McL80] James E. McLean. Book review: [untitled]. Journal of the American Statistical Association, 75(371):759, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287693.

McLeish:1985:BRB

[McL85] D. L. McLeish. Book review: [untitled]. Journal of the American Statistical Association, 80(389):233–234, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288080.

Morgan:1986:ECC

[ME86] Timothy M. Morgan and Robert M. Elashoff. Effect of categorizing a continuous covariate on the comparison of survival time. Journal of the American Statistical Association, 81(396):917–921, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289060.

Meeden:1981:BAB

[Mee81a] Glen Meeden. Betting against a Bayesian bookie. Journal of the American Statistical Association, 76(373):202–204, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287068.

Meeden:1981:BRBa

[Mee81b] Glen Meeden. Book review: [untitled]. Journal of the American Statistical Association, 76(375):743, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287545.

Meeden:1981:BRBb

[Mee81c] Glen Meeden. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1011, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287605.

Meeden:1984:BRB

[Mee84] Glen Meeden. Book review: [untitled]. Journal of the American Statistical Association, 79(387):734, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288435.

Mehay:1982:BRB

[Meh82] Stephen L. Mehay. Book review: [untitled]. Journal of the American Statistical Association, 77(380):949–950, December 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287354.

Meier:1982:EHH

[Mei82] Paul Meier. Estimating historical heights: Comment. Journal of the American Statistical Association, 77(378):296–297, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287234.

Meier:1986:DLE

[Mei86] Paul Meier. Damned liars and expert witnesses. Journal of the American Statistical Association, 81(394):269–276, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289214.

Mensing:1980:BRB

[Men80] Richard W. Mensing. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1041–1042, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287219.

Mengle:1983:BRB

[Men83] David L. Mengle. Book review: [untitled]. Journal of the American Statistical Association, 78(383):740–741, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288163.

Metter:1984:BRB

[Met84] Gerald E. Metter. Book review: [untitled]. Journal of the American Statistical Association, 79(386):476, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288304.

Metter:1985:BRB

[Met85] Gerald E. Metter. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1080, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288604.

Meyer:1985:DGL

[Mey85] Michael M. Meyer. Dynamic generalized linear models and Bayesian forecasting: Comment. Journal of the American Statistical Association, 80(389):94–95, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288048.

Meyer:1986:BRB

[Mey86] Michael M. Meyer. Book review: [untitled]. Journal of the American Statistical Association, 81(395):862, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289031.

Meyer:1987:OIC

[Mey87a] John S. Meyer. Outer and inner confidence intervals for finite population quantile intervals. Journal of the American Statistical Association, 82(397):201–204, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289154.

Meyers:1987:BRB

[Mey87b] Glenn Meyers. Book review: [untitled]. Journal of the American Statistical Association, 82(397):362, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289204.

Mezrich:1984:DRM

[MFS84] J. J. Mezrich, S. P. Frysinger, and R. Slivjanovski. Dynamic representation of multivariate time series data. Journal of the American Statistical Association, 79(385):34–40, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288331.

Morton:1984:SE

[MG84] M. J. Morton and H. L. Gray. The G-spectral estimator. Journal of the American Statistical Association, 79(387):692–701, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288418.

Markowski:1982:IBS

[MH82] Edward P. Markowski and Thomas P. Hettmansperger. Inference based on simple rank step score statistics for the location model. Journal of the American Statistical Association, 77(380): 901–907, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287325.

Mountain:1989:CSF

[MH89] Dean Mountain and Cheng Hsiao. A combined structural and flexible functional approach for modeling energy substitution. Journal of the American Statistical Association, 84(405):76–87, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289848.

Miller:1982:GKG

[MHMEH82] Rupert G. Miller, Jr., Meredith Halks-Miller, Marlene J. Egger, and Jerry W. Halpern. Growth kinetics of glioma cells. Journal of the American Statistical Association, 77(379):505–514, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287703.

Milch:1981:BRB

[Mil81a] Paul R. Milch. Book review: [untitled]. Journal of the American Statistical Association, 76(374):494, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287863.

Miller:1981:BRB

[Mil81b] Robert B. Miller. Book review: [untitled]. Journal of the American Statistical Association, 76(375):750, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287556.

Miller:1985:BRB

[Mil85] Peter V. Miller. Book review: [untitled]. Journal of the American Statistical Association, 80(389):235–236, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288082.

Miller:1986:BRBa

[Mil86a] Robert B. Miller. Book review: [untitled]. Journal of the American Statistical Association, 81(394):577–578, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289271.

Miller:1986:BRBb

[Mil86b] Robert B. Miller. Book review: [untitled]. Journal of the American Statistical Association, 81(395):865–866, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289035.

Miller:1987:BRB

[Mil87] Robert B. Miller. Book review: [untitled]. Journal of the American Statistical Association, 82(397):361, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289203.

Minkin:1987:ODB

[Min87] Salomon Minkin. Optimal designs for binary data. Journal of the American Statistical Association, 82(400):1098–1103, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Mislevy:1985:ELG

[Mis85] Robert J. Mislevy. Estimation of latent group effects. Journal of the American Statistical Association, 80(392):993–997, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288565.

Mislevy:1987:BRB

[Mis87] Robert J. Mislevy. Book review: [untitled]. Journal of the American Statistical Association, 82(398):680–681, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289487.

Mitra:1984:BRB

[Mit84] S. Mitra. Book review: [untitled]. Journal of the American Statistical Association, 79(386):480–481, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288314.

Mittal:1986:BRB

[Mit86] Yashaswini Mittal. Book review: [untitled]. Journal of the American Statistical Association, 81(394):579–580, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289275.

Mitra:1988:BRB

[Mit88] S. Mitra. Book review: [untitled]. Journal of the American Statistical Association, 83(401):278–279, March 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288971.

Marasinghe:1982:TIA

[MJ82] Mervyn G. Marasinghe and Dallas E. Johnson. A test of incomplete additivity in the multiplicative interaction model. Journal of the American Statistical Association, 77(380):869–877, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287320.

Morimune:1980:IML

[MK80] Kimio Morimune and Naoto Kunitomo. Improving the maximum likelihood estimate in linear functional relationships for alternative parameter sequences. Journal of the American Statistical Association, 75(369):230–237, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287416.

Margolin:1989:ASM

[MKR89] Barry H. Margolin, Byung Soo Kim, and Kenneth J. Risko. The Ames Salmonella/Microsome Mutagenicity Assay: Issues of inference and validation. Journal of the American Statistical Association, 84(407):651–661, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289645.

Marquis:1986:RBR

[MMP86] Kent H. Marquis, M. Susan Marquis, and J. Michael Polich. Response bias and reliability in sensitive topic surveys. Journal of the American Statistical Association, 81(394):381–389, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289227.

Marroquin:1987:PSI

[MMP87] J. Marroquin, S. Mitter, and T. Poggio. Probabilistic solution of ill-posed problems in computational vision. Journal of the American Statistical Association, 82(397):76–89, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289127.

Mee:1983:IFO

[MO83] Robert W. Mee and D. B. Owen. Improved factors for one-sided tolerance limits for balanced one-way ANOVA random model. Journal of the American Statistical Association, 78(384):901–905, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288202.

Marshall:1985:FBD

[MO85] Albert W. Marshall and Ingram Olkin. A family of bivariate distributions generated by the bivariate Bernoulli distribution. Journal of the American Statistical Association, 80(390):332–338, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287890.

Marshall:1988:FMD

[MO88] Albert W. Marshall and Ingram Olkin. Families of multivariate distributions. Journal of the American Statistical Association, 83 (403):834–841, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Moffitt:1989:CAA

[Mof89] Robert Moffitt. Choosing among alternative nonexperimental methods for estimating the impact of social programs: The case of manpower training: Comment. Journal of the American Statistical Association, 84(408):877–878, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290061.

Monahan:1985:BRB

[Mon85] John F. Monahan. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1081–1082, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288607.

Monrad:1988:BRB

[Mon88] Ditlev Monrad. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1220, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290171.

Moore:1980:BRB

[Moo80] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 75(371):750, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287685.

Moore:1981:BRB

[Moo81a] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 76(374):497–498, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287872.

Moors:1981:ILI

[Moo81b] J. J. A. Moors. Inadmissibility of linearly invariant estimators in truncated parameter spaces. Journal of the American Statistical Association, 76(376):910–915, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287587.

Moore:1982:BRBa

[Moo82a] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 77(377):222, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287809.

Moore:1982:BRBb

[Moo82b] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 77(380):944–946, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287348.

Moore:1983:BRBa

[Moo83a] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 78(382):506–507, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288688.

Moore:1983:BRBb

[Moo83b] Roger H. Moore. Book review: [untitled]. Journal of the American Statistical Association, 78(382):507, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288690.

Moore:1984:BRBb

[Moo84a] Albert H. Moore. Book review: Life Testing and Reliability Estimation, by S. K. Sinha; B. K. Kale. Journal of the American Statistical Association, 79(385):233, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288372.

Moore:1984:BRBa

[Moo84b] Roger H. Moore. Book review: Encyclopedia of Statistical Sciences. Vol 1: A–Circular Probable Error, by Samuel Kotz; Norman L. Johnson; Encyclopedia of Statistical Sciences. Vol 2: Classification–Eye Estimate, by Samuel Kotz; Norman L. Johnson. Journal of the American Statistical Association, 79(385):226– 227, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288365.

Moore:1987:BRB

[Moo87] David S. Moore. Book review: [untitled]. Journal of the American Statistical Association, 82(398):699–700, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289510.

Moore:1989:BRB

[Moo89] Thomas L. Moore. Book review: [untitled]. Journal of the American Statistical Association, 84(407):834–836, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289676.

Morgan:1980:BRB

[Mor80] Frederick W. Morgan. Book review: [untitled]. Journal of the American Statistical Association, 75(371):760, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287695.

Morimune:1981:AED

[Mor81] Kimio Morimune. Asymptotic expansions of the distribution of an improved limited information maximum likelihood estimator. Journal of the American Statistical Association, 76(374):476–478, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287854.

Morris:1983:PEBb

[Mor83a] Carl N. Morris. Parametric empirical Bayes inference: Theory and applications: Rejoinder. Journal of the American Statistical Association, 78(381):63–65, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287105.

Morris:1983:PEBa

[Mor83b] Carl N. Morris. Parametric empirical Bayes inference: theory and applications. With discussion. Journal of the American Statistical Association, 78(381):47–55, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287098.

Morrison:1984:BRB

[Mor84] Donald F. Morrison. Book review: [untitled]. Journal of the American Statistical Association, 79(387):732, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288432.

Morgenthaler:1986:RCI

[Mor86] Stephan Morgenthaler. Robust confidence intervals for a location parameter: the configural approach. Journal of the American Statistical Association, 81(394):518–525, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289244.

Morgan:1987:BRB

[Mor87a] James N. Morgan. Book review: [untitled]. Journal of the American Statistical Association, 82(399):958–959, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288833.

Morris:1987:CPD

[Mor87b] C. N. Morris. The calculation of posterior distributions by data augmentation: Comment: Simulation in hierarchical models. Journal of the American Statistical Association, 82(398):542–543, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289459.

Morris:1987:EBC

[Mor87c] C. N. Morris. Empirical Bayes confidence intervals based on bootstrap samples: Comment. Journal of the American Statistical Association, 82(399):751–752, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288779.

Morris:1987:TPN

[Mor87d] C. N. Morris. Testing a point null hypothesis: The irreconcilability of P values and evidence: Comment. Journal of the American Statistical Association, 82(397):131–133, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289137.

Moskowitz:1980:BRB

[Mos80] Herbert Moskowitz. Book review: [untitled]. Journal of the American Statistical Association, 75(369):249–251, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287435.

Moses:1985:BRB

[Mos85] Lincoln E. Moses. Book review: [untitled]. Journal of the American Statistical Association, 80(390):479–480, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287920.

Mosteller:1986:BRB

[Mos86] Frederick Mosteller. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1118–1119, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289098.

Mehta:1983:NAP

[MP83] Cyrus R. Mehta and Nitin R. Patel. A network algorithm for performing Fisher’s exact test in r × c contingency tables. Journal of the American Statistical Association, 78(382):427–434, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288652.

Michelini:1985:EHI

[MP85] Claudio Michelini and Michael Pickford. Estimating the Herfindahl index from concentration ratio data. Journal of the American Statistical Association, 80(390):301–305, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287886.

Mehta:1985:CEC

[MPG85] Cyrus R. Mehta, Nitin R. Patel, and Robert Gray. Computing an exact confidence interval for the common odds ratio in several 2 × 2 contingency tables. Journal of the American Statistical Association, 80(392):969–973, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288562. See correction [MPG86].

Mehta:1986:CCE

[MPG86] C. R. Mehta, N. R. Patel, and R. Gray. Correction: “Computing an exact confidence interval for the common odds ratio in several 2 × 2 contingency tables” [J. Amer. Statist. Assoc. 80 (1985), no. 392, 969–973]. Journal of the American Statistical Association, 81 (396):1132, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289114. See [MPG85].

Mehta:1988:ISE

[MPS88] Cyrus R. Mehta, Nitin R. Patel, and Pralay Senchaudhuri. Importance sampling for estimating exact probabilities in permutational inference. Journal of the American Statistical Association, 83(404):999–1005, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290126.

Mendelsohn:1982:DMH

[MR82] John Mendelsohn and John Rice. Deconvolution of microfluorometric histograms with B splines. Journal of the American Statistical Association, 77(380):748–753, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287301.

Montgomery:1986:CHB

[MRB86] Mark R. Montgomery, Toni Richards, and Henry I. Braun. Child health, breast-feeding, and survival in Malaysia: a random-effects logit approach. Journal of the American Statistical Association, 81(394):297–309, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289217.

Mukerjee:1987:CST

[MRW87] Hari Mukerjee, Tim Robertson, and F. T. Wright. Comparison of several treatments with a control using multiple contrasts. Journal of the American Statistical Association, 82(399):902–910, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288803.

Mack:1980:FTR

[MS80] Gregory A. Mack and John H. Skillings. A Friedman-type rank test for main effects in a two-factor ANOVA. Journal of the American Statistical Association, 75(372):947–951, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287186.

Mangel:1984:AWWa

[MS84a] Marc Mangel and Francisco J. Samaniego. Abraham Wald’s work on aircraft survivability. Journal of the American Statistical Association, 79(386):259–267, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288257.

Mangel:1984:AWWb

[MS84b] Marc Mangel and Francisco J. Samaniego. Abraham Wald’s work on aircraft survivability: Rejoinder. Journal of the American Statistical Association, 79(386):270–271, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288259.

Malec:1985:BIF

[MS85a] Donald Malec and J. Sedransk. Bayesian inference for finite population parameters in multistage cluster sampling. Journal of the American Statistical Association, 80(392):897–902, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288550.

Michael:1985:ICG

[MS85b] John R. Michael and William R. Schucany. The influence curve and goodness of fit. Journal of the American Statistical Association, 80(391):678–682, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288484.

Montagne:1985:RSE

[MS85c] Ernest R. Montagne and Nozer D. Singpurwalla. Robustness of sequential exponential life-testing procedures. Journal of the American Statistical Association, 80(391):715–719, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288490.

Mathew:1988:OTF

[MS88a] Thomas Mathew and Bimal Kumar Sinha. Optimum tests for fixed effects and variance components in balanced models. Journal of the American Statistical Association, 83(401):133–135, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288931.

Muller:1988:KPE

[MS88b] Hans-Georg Müller and Thomas Schmitt. Kernel and probit estimates in quantal bioassay. Journal of the American Statistical Association, 83(403):750–759, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Meinhold:1989:RKF

[MS89] Richard J. Meinhold and Nozer D. Singpurwalla. Robustification of Kalman filter models. Journal of the American Statistical Association, 84(406):479–486, June 1989. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289933.

Manton:1986:AMH

[MSV86] Kenneth G. Manton, Eric Stallard, and James W. Vaupel. Alternative models for the heterogeneity of mortality risks among the aged. Journal of the American Statistical Association, 81(395): 635–644, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288991.

McDonald:1983:TPE

[MSW^{+}83] J. F. McDonald, M. A. Selby, C. S. Wong, L. D. Favro, and P. K. Kuo. Tests, point estimations, and confidence sets for a capturerecapture model. Journal of the American Statistical Association, 78(384):913–919, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288204.

Michalek:1980:EED

[MT80] Joel E. Michalek and Ram C. Tripathi. The effect of errors in diagnosis and measurement on the estimation of the probability of an event. Journal of the American Statistical Association, 75 (371):713–721, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287675.

Mudholkar:1981:GAD

[MT81] Govind S. Mudholkar and Madhusudan C. Trivedi. A Gaussian approximation to the distribution of the sample variance for nonnormal populations. Journal of the American Statistical Association, 76(374):479–485, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2287855. See corrigendum [MT82].

Mudholkar:1982:CGA

[MT82] Govind S. Mudholkar and Madhusudan C. Trivedi. Corrigendum: “A Gaussian approximation to the distribution of the sample variance for nonnormal populations” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 479–485, MR 82h:62034]. Journal of the American Statistical Association, 77(380):954–955, December 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287366. See [MT81].

Mueller:1980:BRB

[Mue80] Marnie W. Mueller. Book review: [untitled]. Journal of the American Statistical Association, 75(369):243–244, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287426.

Muirhead:1985:EPF

[Mui85] Robb J. Muirhead. Estimating a particular function of the multiple correlation coefficient. Journal of the American Statistical Association, 80(392):923–925, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288554.

Mukhopadhyay:1984:BRB

[Muk84] N. Mukhopadhyay. Book review: The Sequential Analysis of Hypothesis Testing, Point and Interval Estimation, and Decision Theory, by Z. Govindarajulu. Journal of the American Statistical Association, 79(385):233, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288373.

Muller:1983:BRB

[Mul83] Keith E. Muller. Book review: [untitled]. Journal of the American Statistical Association, 78(384):998, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288232.

Muller:1987:WLR

[Mül87] Hans-Georg Müller. Weighted local regression and kernel methods for nonparametric curve fitting. Journal of the American Statistical Association, 82(397):231–238, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289159.

Meeden:1985:BAS

[MV85] Glen Meeden and Stephen Vardeman. Bayes and admissible set estimation. Journal of the American Statistical Association, 80 (390):465–471, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287916.

Malani:1988:CTT

[MV88] Hina Mehta Malani and John Van Ryzin. Comparison of two treatments in animal carcinogenicity experiments. Journal of the American Statistical Association, 83(404):1171–1177, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290151.

Mack:1981:SRT

[MW81] Gregory A. Mack and Douglas A. Wolfe. K-sample rank tests for umbrella alternatives. Journal of the American Statistical Association, 76(373):175–181, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287064. See corrigendum [MW82].

Mack:1982:CSR

[MW82] Gregory A. Mack and Douglas A. Wolfe. Corrigendum: “k-sample rank tests for umbrella alternatives” [J. Amer. Statist. Assoc. 76 (1981), no. 373, 175–181, MR 82d:62084]. Journal of the American Statistical Association, 77(380):954, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287362. See [MW81].

Murphy:1984:PFM

[MW84] Allan H. Murphy and Robert L. Winkler. Probability forecasting in meteorology. Journal of the American Statistical Association, 79(387):489–500, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288395.

Manton:1989:EBP

[MWS^{+}89] Kenneth G. Manton, Max A. Woodbury, Eric Stallard, Wilson B. Riggan, John P. Creason, and Alvin C. Pellom. Empirical Bayes procedures for stabilizing maps of U.S. cancer mortality rates. Journal of the American Statistical Association, 84(407):637–650, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289644.

Myers:1984:BRB

[Mye84] Raymond H. Myers. Book review: Factorial Designs, by B. L. Raktoe; A. Hedayat; W. T. Federer. Journal of the American Statistical Association, 79(385):232, March 1984. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288371.

Meier:1980:BPH

[MZ80] Paul Meier and Sandy Zabell. Benjamin Peirce and the Howland Will. Journal of the American Statistical Association, 75(371): 497–506, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287637.

Martin:1989:AMM

[MZ89] R. Douglas Martin and Ruben H. Zamar. Asymptotically minmax bias robust M-estimates of scale for positive random variables. Journal of the American Statistical Association, 84(406): 494–501, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289935.

Naiman:1983:CST

[Nai83] Daniel Q. Naiman. Comparing Scheffé-type to constant-width confidence bounds in regression. Journal of the American Statistical Association, 78(384):906–912, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288203.

Nair:1984:BSE

[Nai84] Vijayan N. Nair. On the behavior of some estimators from probability plots. Journal of the American Statistical Association, 79 (388):823–831, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288712.

Naiman:1987:MRS

[Nai87a] Daniel Q. Naiman. Minimax regret simultaneous confidence bands for multiple regression functions. Journal of the American Statistical Association, 82(399):894–901, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288802.

Naiman:1987:SCB

[Nai87b] Daniel Q. Naiman. Simultaneous confidence bounds in multiple regression using predictor variable constraints. Journal of the American Statistical Association, 82(397):214–219, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289156.

Nair:1987:CST

[Nai87c] Vijayan N. Nair. Chi-squared-type tests for ordered alternatives in contingency tables. Journal of the American Statistical Association, 82(397):283–291, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289165.

Nam:1983:BRB

[Nam83] Charles B. Nam. Book review: [untitled]. Journal of the American Statistical Association, 78(382):496, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288667.

Nam:1984:BRB

[Nam84] Charles B. Nam. Book review: [untitled]. Journal of the American Statistical Association, 79(388):961, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288762.

Nam:1985:BRBa

[Nam85a] Charles B. Nam. Book review: [untitled]. Journal of the American Statistical Association, 80(391):779, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288511.

Nam:1985:BRBb

[Nam85b] Charles B. Nam. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1076, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288595.

Nam:1986:BRB

[Nam86] Charles B. Nam. Book review: [untitled]. Journal of the American Statistical Association, 81(395):873, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289047.

Nam:1987:BRBa

[Nam87a] Charles B. Nam. Book review: [untitled]. Journal of the American Statistical Association, 82(397):352, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289188.

Nam:1987:BRBb

[Nam87b] Charles B. Nam. Book review: [untitled]. Journal of the American Statistical Association, 82(397):352, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289189.

Nam:1987:BRBc

[Nam87c] Charles B. Nam. Book review: [untitled]. Journal of the American Statistical Association, 82(399):954, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288825.

Nam:1988:BRB

[Nam88] Charles B. Nam. Book review: [untitled]. Journal of the American Statistical Association, 83(401):278, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288970.

Nandi:1980:EDN

[Nan80] S. B. Nandi. On the exact distribution of a normalized ratio of the weighted geometric mean to the unweighted arithmetic mean in samples from gamma distributions. Journal of the American Statistical Association, 75(369):217–220, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287414.

Nathan:1988:BRB

[Nat88] Gad Nathan. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1223, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290175.

Naus:1982:ADS

[Nau82] Joseph I. Naus. Approximations for distributions of scan statistics. Journal of the American Statistical Association, 77(377): 177–183, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287786.

Norwood:1984:CMP

[NE84] Janet L. Norwood and John F. Early. A century of methodological progress at the U.S. Bureau of Labor Statistics. Journal of the American Statistical Association, 79(388):748–761, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288705.

Neftci:1982:SET

[Nef82] Salih N. Neftci. Specification of economic time series models using akaike’s criterion. Journal of the American Statistical Association, 77(379):537–540, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287707.

Neftci:1985:NUL

[Nef85] Salih N. Neftci. A note on the use of local maxima to predict turning points in related series. Journal of the American Statistical Association, 80(391):553–557, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288468.

Nelson:1985:MGP

[Nel85] James F. Nelson. Multivariate gamma-Poisson models. Journal of the American Statistical Association, 80(392):828–834, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288540.

Nelson:1989:BRB

[Nel89] Barry L. Nelson. Book review: [untitled]. Journal of the American Statistical Association, 84(405):334, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289887.

Nerman:1986:BRB

[Ner86] Olle Nerman. Book review: [untitled]. Journal of the American Statistical Association, 81(395):858–859, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289024.

Neter:1986:BSS

[Net86] John Neter. Boundaries of statistics-sharp or fuzzy? Journal of the American Statistical Association, 81(393):1–8, March 1986.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287957.

Neuts:1981:BRBa

[Neu81a] Marcel F. Neuts. Book review: [untitled]. Journal of the American Statistical Association, 76(374):492–493, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287860.

Neuts:1981:BRBb

[Neu81b] Marcel F. Neuts. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1018–1019, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287623.

Neuts:1983:BRB

[Neu83] Marcel F. Neuts. Book review: [untitled]. Journal of the American Statistical Association, 78(381):211–212, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287148.

Newhouse:1981:BRBa

[New81a] Joseph P. Newhouse. Book review: [untitled]. Journal of the American Statistical Association, 76(374):495–496, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287868.

Newhouse:1981:BRBb

[New81b] Joseph P. Newhouse. Book review: [untitled]. Journal of the American Statistical Association, 76(375):746–747, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287551.

Newbold:1984:BRB

[New84] Paul Newbold. Book review: [untitled]. Journal of the American Statistical Association, 79(388):943–945, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288732.

Nachtsheim:1988:NFM

[NJ88] Christopher J. Nachtsheim and Mark E. Johnson. A new family of multivariate distributions with applications to Monte Carlo studies. Journal of the American Statistical Association, 83(404): 984–989, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290124.

Noorbaloochi:1983:UDB

[NM83a] Siamak Noorbaloochi and Glen Meeden. Unbiasedness as the dual of being Bayes. Journal of the American Statistical Association, 78 (383):619–623, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288129.

Nyblom:1983:CTP

[NM83b] Jukka Nyblom and Timo Mäkeläinen. Comparisons of tests for the presence of random walk coefficients in a simple linear model. Journal of the American Statistical Association, 78(384):856–864, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288196.

Nakamura:1986:BRB

[NN86] Alice Nakamura and Masao Nakamura. Book review: [untitled]. Journal of the American Statistical Association, 81(394):566–567, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289255.

Noam:1981:BRB

[Noa81] Eli Noam. Book review: [untitled]. Journal of the American Statistical Association, 76(375):746, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287550.

Nobrega:1981:BRB

[Nob81] Fred Nobrega. Book review: [untitled]. Journal of the American Statistical Association, 76(375):747–748, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287552.

Noether:1980:BRB

[Noe80] Gottfried E. Noether. Book review: [untitled]. Journal of the American Statistical Association, 75(371):757, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287690.

Noether:1987:SSD

[Noe87] Gottfried E. Noether. Sample size determination for some common nonparametric tests. Journal of the American Statistical Association, 82(398):645–647, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289477.

Nordheim:1984:INM

[Nor84] Erik V. Nordheim. Inference from nonrandomly missing categorical data: An example from a genetic study on Turner’s syndrome. Journal of the American Statistical Association, 79(388): 772–780, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288707.

Norton:1985:BRB

[Nor85] Julia A. Norton. Book review: [untitled]. Journal of the American Statistical Association, 80(389):242, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288088.

Nordheim:1986:BRB

[Nor86] Erik V. Nordheim. Book review: [untitled]. Journal of the American Statistical Association, 81(395):862–863, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289032.

Notz:1989:BRB

[Not89] William I. Notz. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1102, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290106.

Newton:1983:MDP

[NP83] H. Joseph Newton and Marcello Pagano. A method for determining periods in time series. Journal of the American Statistical Association, 78(381):152–157, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287122.

Newhouse:1980:EDD

[NRMM80] Joseph P. Newhouse, John E. Rolph, Bryant Mori, and Maureen Murphy. The effect of deductibles on the demand for medical care services. Journal of the American Statistical Association, 75(371): 525–533, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287641.

Nelson:1982:TPR

[NS82] Charles R. Nelson and G. William Schwert. Tests for predictive relationships between time series variables: a Monte Carlo investigation. Journal of the American Statistical Association, 77(377): 11–18, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287764.

Nagarsenker:1983:DCS

[NS83] B. N. Nagarsenker and J. Suniaga. Distributions of a class of statistics useful in multivariate analysis. Journal of the American Statistical Association, 78(382):472–475, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288661. See correction [NS86].

Nagarsenker:1986:CDC

[NS86] B. N. Nagarsenker and J. Suniaga. Correction: “Distributions of a class of statistics useful in multivariate analysis” [J. Amer. Statist. Assoc. 78 (1983), no. 382, 472–475, MR 84i:62070]. Journal of the American Statistical Association, 81(396):1132, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289112. See

[NS83].

Nychka:1984:CVS

[NWGP84] Douglas Nychka, Grace Wahba, Stanley Goldfarb, and Thomas Pugh. Cross-validated spline methods for the estimation of threedimensional tumor size distributions from observations on twodimensional cross sections. Journal of the American Statistical Association, 79(388):832–846, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288713.

Nyblom:1986:TDL

[Nyb86] Jukka Nyblom. Testing for deterministic linear trend in time series. Journal of the American Statistical Association, 81(394): 545–549, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289247.

Nyblom:1989:TCP

[Nyb89] Jukka Nyblom. Testing for the constancy of parameters over time. Journal of the American Statistical Association, 84(405): 223–230, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289867.

Nychka:1988:BCI

[Nyc88] Douglas Nychka. Bayesian confidence intervals for smoothing splines. Journal of the American Statistical Association, 83(404): 1134–1143, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290146.

Nychka:1989:BRB

[Nyc89] Douglas Nychka. Book review: [untitled]. Journal of the American Statistical Association, 84(406):623–624, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289968.

Oakes:1983:RSH

[Oak83] D. Oakes. A reanalysis of the Stanford heart transplant data: Comment. Journal of the American Statistical Association, 78 (382):288, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288627.

Oakes:1985:SCP

[Oak85] David Oakes. Self-calibrating priors do not exist. With comments by A. P. Dawid and Mark J. Schervish. Journal of the American Statistical Association, 80(390):339, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287891.

Oakes:1989:BSM

[Oak89] David Oakes. Bivariate survival models induced by frailties. Journal of the American Statistical Association, 84(406):487–493, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289934.

OBrien:1980:BRBa

[O’B80a] Peter C. O’Brien. Book review: [untitled]. Journal of the American Statistical Association, 75(370):479–480, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287497.

OBrien:1980:BRBb

[O’B80b] Peter C. O’Brien. Book review: [untitled]. Journal of the American Statistical Association, 75(371):756–757, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287689.

OBrien:1981:BRBa

[O’B81a] Peter C. O’Brien. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1012–1013, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287610.

OBrien:1981:BRBb

[O’B81b] Peter C. O’Brien. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1014–1015, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287616.

OBrien:1982:BRB

[O’B82] Peter C. O’Brien. Book review: [untitled]. Journal of the American Statistical Association, 77(379):695, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287757.

OBrien:1983:BRB

[O’B83] Peter C. O’Brien. Book review: [untitled]. Journal of the American Statistical Association, 78(382):507, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288689.

OBrien:1984:BRB

[O’B84] Peter C. O’Brien. Book review: Applied Multivariate Statistical Analysis, by Richard A. Johnson; Dean W. Wichern. Journal of the American Statistical Association, 79(385):231, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288368.

OBrien:1988:CTS

[O’B88a] Peter C. O’Brien. Comparing two samples: Extensions of the t, rank-sum, and log-rank tests. Journal of the American Statistical Association, 83(401):52–61, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288918.

OHagan:1988:RPP

[OB88b] Anthony O’Hagan and James O. Berger. Ranges of posterior probabilities for quasiunimodal priors with specified quantiles. Journal of the American Statistical Association, 83(402):503–508, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288869.

Obenchain:1980:CSR

[Obe80] Robert L. Obenchain. A critique of some ridge regression methods: Comment. Journal of the American Statistical Association, 75 (369):95–96, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287391.

Oberschall:1986:BRB

[Obe86] Anthony Oberschall. Book review: [untitled]. Journal of the American Statistical Association, 81(393):255–256, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288005.

Odell:1981:BRB

[Ode81] Patrick L. Odell. Book review: [untitled]. Journal of the American Statistical Association, 76(373):209, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287075.

OFallon:1980:BRB

[O’F80] W. M. O’Fallon. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1039–1040, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287214.

OFallon:1984:BRB

[O’F84] William M. O’Fallon. Book review: [untitled]. Journal of the American Statistical Association, 79(386):482, June 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288316.

OGorman:1982:DSA

[O’G82] Thomas W. O’Gorman. On the design of seasonal adjustment methods using linear programming techniques. Journal of the American Statistical Association, 77(380):739–742, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287299.

Ogata:1988:SME

[Oga88] Yosihiko Ogata. Statistical models for earthquake occurrences and residual analysis for point processes. Journal of the American Statistical Association, 83(401):9–27, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288914.

OHagan:1987:BLE

[O’H87a] A. O’Hagan. Bayes linear estimators for randomized response models. Journal of the American Statistical Association, 82(398): 580–585, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289468.

OHagan:1987:CPD

[O’H87b] A. O’Hagan. The calculation of posterior distributions by data augmentation: Comment. Journal of the American Statistical Association, 82(398):547, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289462.

OLeary:1985:BRB

[O’L85] Dianne P. O’Leary. Book review: [untitled]. Journal of the American Statistical Association, 80(389):248–249, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288098.

Oler:1985:NPC

[Ole85] Jacqueline Oler. Noncentrality parameters in chi-squared goodness-of-fit analyses with an application to log-linear procedures. Journal of the American Statistical Association, 80(389): 181–189, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288069.

Oman:1981:CBA

[Oma81] Samuel D. Oman. A confidence bound approach to choosing the biasing parameter in ridge regression. Journal of the American Statistical Association, 76(374):452–461, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287849.

Oman:1985:SPD

[Oma85] Samuel D. Oman. Specifying a prior distribution in structured regression problems. Journal of the American Statistical Association, 80(389):190–195, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288070.

ONeill:1980:GDE

[O’N80] Terence J. O’Neill. The general distribution of the error rate of a classification procedure with application to logistic regression discrimination. Journal of the American Statistical Association, 75(369):154–160, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287404.

ONeill:1984:GFT

[O’N84] Terence J. O’Neill. A goodness-of-fit test for one-sample life table data. Journal of the American Statistical Association, 79(385): 194–199, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288356.

Oja:1989:BST

[ON89] Hannu Oja and Jukka Nyblom. Bivariate sign tests. Journal of the American Statistical Association, 84(405):249–259, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289871.

Olkin:1981:CEB

[OPZ81] Ingram Olkin, A. John Petkau, and James V. Zidek. A comparison of n estimators for the binomial distribution. Journal of the American Statistical Association, 76(375):637–642, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287523.

ORourke:1988:BRB

[O’R88] Diane O’Rourke. Book review: [untitled]. Journal of the American Statistical Association, 83(401):282–283, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288976.

Ord:1985:BRB

[Ord85] J. K. Ord. Book review: [untitled]. Journal of the American Statistical Association, 80(390):491, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287940.

Ordeshook:1987:RSP

[Ord87] Peter C. Ordeshook. Repeated sampling in the presence of publication effects: Comment: Public opinion polls and democratic processes. Journal of the American Statistical Association, 82 (398):486–491, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289451.

Orey:1983:BRBa

[Ore83a] Steven Orey. Book review: [untitled]. Journal of the American Statistical Association, 78(382):501, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288677.

Orey:1983:BRBb

[Ore83b] Steven Orey. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1002–1003, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288240.

Orr:1989:AWC

[Orr89] James A. Orr. The average workweek of capital in manufacturing, 1952–1984. Journal of the American Statistical Association, 84 (405):88–94, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289849.

Olkin:1987:SAD

[OS87a] Ingram Olkin and Clifford H. Spiegelman. A semiparametric approach to density estimation. Journal of the American Statistical Association, 82(399):858–865, September 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288797.

OSullivan:1987:BRB

[O’S87b] Patrick O’Sullivan. Book review: [untitled]. Journal of the American Statistical Association, 82(399):960–961, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288836.

Osborn:1982:CFU

[Osb82] Denise R. Osborn. On the criteria functions used for the estimation of moving average processes. Journal of the American Statistical Association, 77(378):388–392, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287256.

Orban:1982:CDF

[OW82] John Orban and Douglas A. Wolfe. A class of distribution-free two-sample tests based on placements. Journal of the American Statistical Association, 77(379):666–672, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287734.

OSullivan:1986:ASR

[OYR86] Finbarr O’Sullivan, Brian S. Yandell, and William J. Raynor, Jr. Automatic smoothing of regression functions in generalized linear models. Journal of the American Statistical Association, 81(393): 96–103, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287973.

Pack:1984:BRB

[Pac84] David J. Pack. Book review: [untitled]. Journal of the American Statistical Association, 79(388):945–946, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288733.

Padgett:1986:KTE

[Pad86] W. J. Padgett. A kernel-type estimator of a quantile function from right-censored data. Journal of the American Statistical Association, 81(393):215–222, March 1986. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287993.

Palloni:1986:BRB

[Pal86] Alberto Palloni. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1120–1121, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289101.

Panjer:1987:BRB

[Pan87] Harry H. Panjer. Book review: [untitled]. Journal of the American Statistical Association, 82(397):362, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289205.

Park:1980:BRB

[Par80a] C. J. Park. Book review: [untitled]. Journal of the American Statistical Association, 75(371):756, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287688.

Parzen:1980:BRB

[Par80b] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 75(370):472–473, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287485.

Parzen:1980:DEB

[Par80c] Emanuel Parzen. Density estimation and bump-hunting by the penalized likelihood method exemplified by scattering and meteorite data: Comment. Journal of the American Statistical Association, 75(369):56–59, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287378.

Parzen:1982:MLD

[Par82] Emanuel Parzen. Measurement of linear dependence and feedback between multiple time series: Comment. Journal of the American Statistical Association, 77(378):320–322, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287242.

Parzen:1983:BRB

[Par83] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 78(383):731, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288145.

Parzen:1984:BRB

[Par84] Emanuel Parzen. Book review: Spectral Analysis and Time Series, by M. B. Priestley. Journal of the American Statistical Association, 79(385):233–234, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288374.

Park:1985:BRB

[Par85a] Molly Park. Book review: [untitled]. Journal of the American Statistical Association, 80(389):243–244, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288092.

Parke:1985:BRBa

[Par85b] Robert Parke. Book review: [untitled]. Journal of the American Statistical Association, 80(390):480–481, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287921.

Parke:1985:BRBb

[Par85c] Robert Parke. Book review: [untitled]. Journal of the American Statistical Association, 80(391):771, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288500.

Parzen:1985:BRB

[Par85d] Emanuel Parzen. Book review: [untitled]. Journal of the American Statistical Association, 80(389):251–252, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288103.

Parker:1988:BRB

[Par88] Lloyd R. Parker, Jr. Book review: [untitled]. Journal of the American Statistical Association, 83(401):276, March 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288965.

Parr:1989:BRB

[Par89] William Parr. Book review: [untitled]. Journal of the American Statistical Association, 84(406):620, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289962.

Pascal:1980:BRB

[Pas80] Anthony Pascal. Book review: [untitled]. Journal of the American Statistical Association, 75(370):478–479, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287495.

Passel:1985:EPC

[Pas85] Jeffrey S. Passel. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):123–124, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288057.

Patel:1987:BRB

[Pat87] Kartik R. Patel. Book review: [untitled]. Journal of the American Statistical Association, 82(397):344–345, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289177.

Pena:1987:ISS

[PB87] Daniel Peña and George E. P. Box. Identifying a simplifying structure in time series. Journal of the American Statistical Association, 82(399):836–843, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288794.

Picard:1984:CVR

[PC84] Richard R. Picard and R. Dennis Cook. Cross-validation of regression models. Journal of the American Statistical Association, 79 (387):575–583, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288403.

Pearson:1987:BRB

[Pea87] Robert W. Pearson. Book review: [untitled]. Journal of the American Statistical Association, 82(398):678–679, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289484.

Pebley:1981:BRB

[Peb81] Anne R. Pebley. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1019–1020, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287624.

Peck:1980:BRB

[Pec80] Roxy Peck. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1038, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287211.

Pedler:1980:EDO

[Ped80] P. J. Pedler. Effect of dependence on the occupation time in a two-state stationar Markov chain. Journal of the American Statistical Association, 75(371):739–746, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287681.

Pederson:1988:BRB

[Ped88] Shane P. Pederson. Book review: [untitled]. Journal of the American Statistical Association, 83(401):271–272, March 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288959.

Pederson:1989:BRB

[Ped89] Shane P. Pederson. Book review: [untitled]. Journal of the American Statistical Association, 84(407):839, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289681.

Pektau:1980:BRB

[Pek80] A. John Pektau. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1041, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287217.

Peyrot:1989:BRB

[Pey89] Alain H. Peyrot. Book review: [untitled]. Journal of the American Statistical Association, 84(406):624–625, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289970.

Peters:1987:BBC

[PF87] S. C. Peters and D. A. Freedman. Better bootstrap confidence intervals: Comment. Journal of the American Statistical Association, 82(397):186–187, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289145.

Pfeifer:1987:BRB

[Pfe87] Charles G. Pfeifer. Book review: [untitled]. Journal of the American Statistical Association, 82(397):362–363, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289206.

Peck:1989:ACI

[PFV89] Roger Peck, Lloyd Fisher, and John Van Ness. Approximate confidence intervals for the number of clusters. Journal of the American Statistical Association, 84(405):184–191, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289862.

Petrondas:1983:MCR

[PG83] Demetrios A. Petrondas and K. Ruben Gabriel. Multiple comparisons by rerandomization tests. Journal of the American Statistical Association, 78(384):949–957, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288209.

Pagano:1981:AFE

[PH81] Marcello Pagano and Katherine Taylor Halvorsen. An algorithm for finding the exact significance levels of r × c contingency tables. Journal of the American Statistical Association, 76(376): 931–934, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287590.

Preston:1984:NEC

[PH84] Samuel H. Preston and Michael R. Haines. New estimates of child mortality in the United States at the turn of the century. Journal of the American Statistical Association, 79(386):272–281, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288260.

Peixoto:1986:CAP

[PH86] Julio L. Peixoto and David A. Harville. Comparisons of alternative predictors under the balanced one-way random model. Journal of the American Statistical Association, 81(394):431–436, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289232.

Pickard:1987:IDM

[Pic87] David K. Pickard. Inference for discrete Markov fields: the simplest nontrivial case. Journal of the American Statistical Association, 82(397):90–96, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289128.

Pickands:1989:BRB

[Pic89] James Pickands III. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1099, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290101.

Pielou:1981:BRBa

[Pie81a] E. C. Pielou. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1014, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287613.

Pielou:1981:BRBb

[Pie81b] E. C. Pielou. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1014, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287614.

Pierce:1982:MLD

[Pie82] David A. Pierce. Measurement of linear dependence and feedback between multiple time series: Comment. Journal of the American Statistical Association, 77(378):315–316, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287240.

Piegorsch:1985:AWO

[Pie85] Walter W. Piegorsch. Average-width optimality for confidence bands in simple linear regression. Journal of the American Statistical Association, 80(391):692–697, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288486.

Pieters:1986:BRB

[Pie86] Richard S. Pieters. Book review: [untitled]. Journal of the American Statistical Association, 81(395):873–874, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289048.

Piegorsch:1987:MRS

[Pie87] Walter W. Piegorsch. Model robustness for simultaneous confidence bands. Journal of the American Statistical Association, 82 (399):879–885, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288800.

Pierce:1989:SAS

[Pie89] David A. Pierce. Sensitivity analysis of seasonal adjustments: Empirical case studies: Comment. Journal of the American Statistical Association, 84(405):21–22, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289838.

Pijacki:1986:BRB

[Pij86] Roger J. Pijacki. Book review: [untitled]. Journal of the American Statistical Association, 81(393):260, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288015.

Plant:1980:BRB

[Pla80] Mark Plant. Book review: [untitled]. Journal of the American Statistical Association, 75(371):758–759, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287692.

Plakans:1983:BRB

[Pla83] Andrejs Plakans. Book review: [untitled]. Journal of the American Statistical Association, 78(381):205, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287138.

Pliskin:1987:BRB

[Pli87] Jeffrey Pliskin. Book review: [untitled]. Journal of the American Statistical Association, 82(399):955–956, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288828.

Pliskin:1988:BRB

[Pli88] Jeffrey Pliskin. Book review: [untitled]. Journal of the American Statistical Association, 83(402):574, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288897.

Ploch:1987:BRB

[Plo87] Donald R. Ploch. Book review: [untitled]. Journal of the American Statistical Association, 82(399):954–955, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288827.

Pierce:1987:CSA

[PM87] David A. Pierce and Sandra K. McKenzie. On concurrent seasonal adjustment. Journal of the American Statistical Association, 82 (399):720–732, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288776.

Patel:1988:SAT

[PMF88] Kartik R. Patel, Govind S. Mudholkar, and J. L. Indrasiri Fernando. Student’s t approximations for three simple robust estimators. Journal of the American Statistical Association, 83(404): 1203–1210, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290158.

Pfeffermann:1981:RAD

[PN81] Dan Pfeffermann and Gad Nathan. Regression analysis of data from a cluster sample. Journal of the American Statistical Association, 76(375):681–689, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287530.

Pointer:1980:DEB

[Poi80] Patrick C. Pointer. Density estimation and bump-hunting by the penalized likelihood method exemplified by scattering and meteorite data: Comment. Journal of the American Statistical Association, 75(369):68–69, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287383.

Polasek:1984:RDG

[Pol84a] Wolfgang Polasek. Regression diagnostics for general linear regression models. Journal of the American Statistical Association, 79(386):336–340, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288273.

Pollak:1984:BRB

[Pol84b] Edward Pollak. Book review: [untitled]. Journal of the American Statistical Association, 79(387):743, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288455.

Pollard:1985:BRB

[Pol85] David Pollard. Book review: [untitled]. Journal of the American Statistical Association, 80(390):490, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287938.

Pollak:1987:BRB

[Pol87] Edward Pollak. Book review: [untitled]. Journal of the American Statistical Association, 82(397):365, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289210.

Pool:1980:BRB

[Poo80] Karl Pool. Book review: [untitled]. Journal of the American Statistical Association, 75(371):761, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287698.

Portnoy:1986:BRB

[Por86] Stephen Portnoy. Book review: [untitled]. Journal of the American Statistical Association, 81(395):859, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289025.

Potthoff:1983:BRB

[Pot83] Richard F. Potthoff. Book review: [untitled]. Journal of the American Statistical Association, 78(381):204–205, March 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287137.

Potthoff:1987:SGM

[Pot87] Richard F. Potthoff. Some generalizations of the Mitofsky– Waksberg technique for random digit dialing. Journal of the American Statistical Association, 82(398):409–418, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289442.

Phillips:1988:AEO

[PP88] P. C. B. Phillips and Joon Y. Park. Asymptotic equivalence of ordinary least squares and generalized least squares in regressions with integrated regressors. Journal of the American Statistical Association, 83(401):111–115, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288927.

Peele:1980:CSR

[PR80] Lawrence C. Peele and Thomas P. Ryan. A critique of some ridge regression methods: Comment. Journal of the American Statistical Association, 75(369):96–97, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287392.

Pratt:1981:CLL

[Pra81] John W. Pratt. Concavity of the log likelihood. Journal of the American Statistical Association, 76(373):103–106, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287052. See cor-

rigendum [Pra82].

Pratt:1982:CCL

[Pra82] John W. Pratt. Corrigendum: “Concavity of the log likelihood” [J. Amer. Statist. Assoc. 76 (1981), no. 373, 103–106, MR 83h:62098]. Journal of the American Statistical Association, 77 (380):954, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287361. See [Pra81].

Pratt:1986:BRB

[Pra86] John W. Pratt. Book review: [untitled]. Journal of the American Statistical Association, 81(394):565–566, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289254.

Prabhu:1987:BRB

[Pra87a] N. U. Prabhu. Book review: [untitled]. Journal of the American Statistical Association, 82(399):947, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288813.

Pratt:1987:TPN

[Pra87b] John W. Pratt. Testing a point null hypothesis: The irreconcilability of P values and evidence: Comment. Journal of the American Statistical Association, 82(397):123–125, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289132.

Pregibon:1981:GMT

[Pre81a] Daryl Pregibon. Graphics for the multivariate two-sample problem: Comment. Journal of the American Statistical Association, 76(374):291–294, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287828.

Press:1981:BRB

[Pre81b] S. James Press. Book review: [untitled]. Journal of the American Statistical Association, 76(373):206–207, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287070.

Preston:1985:EPC

[Pre85] Samuel H. Preston. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):124–125, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288058.

Prentice:1986:BRU

[Pre86a] R. L. Prentice. Binary regression using an extended beta-binomial distribution, with discussion of correlation induced by covariate measurement errors. Journal of the American Statistical Association, 81(394):321–327, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289219.

Prentice:1986:ACW

[Pre86b] Ross L. Prentice. An analysis of contaminated well water and health effects in Woburn, Massachusetts: Comment. Journal of the American Statistical Association, 81(395):600–601, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288984.

Press:1989:BRB

[Pre89a] S. James Press. Book review: [untitled]. Journal of the American Statistical Association, 84(406):632, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289984.

Presser:1989:BRB

[Pre89b] Stanley Presser. Book review: [untitled]. Journal of the American Statistical Association, 84(406):633–634, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289986.

Proctor:1982:BRB

[Pro82] Charles H. Proctor. Book review: [untitled]. Journal of the American Statistical Association, 77(379):685–686, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287740.

Proctor:1983:BRBa

[Pro83a] Charles Proctor. Book review: [untitled]. Journal of the American Statistical Association, 78(384):996, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288229.

Proctor:1983:BRBb

[Pro83b] Charles Proctor. Book review: [untitled]. Journal of the American Statistical Association, 78(384):998–999, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288233.

Proctor:1984:BRB

[Pro84] Charles H. Proctor. Book review: Linear Panel Analysis: Models of Quantitative Change, by Ronald C. Kessler; David F. Greenberg. Journal of the American Statistical Association, 79(385): 239–240, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288383.

Proctor:1985:FHF

[Pro85] Charles H. Proctor. Fitting H. F. Smith’s empirical law to cluster variances for use in designing multi-stage sample surveys. Journal of the American Statistical Association, 80(390):294–300, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287885.

Pruzek:1984:BRB

[Pru84] Robert M. Pruzek. Book review: [untitled]. Journal of the American Statistical Association, 79(386):480, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288313.

Pruzek:1985:BRB

[Pru85] Robert M. Pruzek. Book review: [untitled]. Journal of the American Statistical Association, 80(391):777–778, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288508.

Pruitt:1988:BRB

[Pru88] Ronald C. Pruitt. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1216–1217, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290165.

Parr:1980:MDR

[PS80] William C. Parr and William R. Schucany. Minimum distance and robust estimation. Journal of the American Statistical Association, 75(371):616–624, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287658.

Parr:1982:JSS

[PS82] William C. Parr and William R. Schucany. Jackknifing L-statistics with smooth weight functions. Journal of the American Statistical Association, 77(379):629–638, September 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287726.

Pratt:1984:NDSa

[PS84a] J. W. Pratt and Robert Schlaifer. On the nature and discovery of structure. Journal of the American Statistical Association, 79(385):9–21, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288326.

Pratt:1984:NDSb

[PS84b] J. W. Pratt and Robert Schlaifer. On the nature and discovery of structure: Rejoinder. Journal of the American Statistical Association, 79(385):29–33, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288330.

Pierce:1986:RGL

[PS86] Donald A. Pierce and Daniel W. Schafer. Residuals in generalized linear models. Journal of the American Statistical Association, 81 (396):977–986, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289071.

Psacharopoulos:1986:BRB

[Psa86] George Psacharopoulos. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1120, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289099.

Patil:1982:DCMb

[PT82a] G. P. Patil and C. Taillie. Diversity as a concept and its measurement: Rejoinder. Journal of the American Statistical Association, 77(379):565–567, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287712.

Patil:1982:DCMa

[PT82b] G. P. Patil and C. Taillie. Diversity as a concept and its measurement. With comments by I. J. Good and George Sugihara and a rejoinder by the authors. Journal of the American Statistical Association, 77(379):548–561, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287709.

Pagano:1983:OPD

[PT83] Marcello Pagano and David Tritchler. On obtaining permutation distributions in polynomial time. Journal of the American Statistical Association, 78(382):435–440, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288653.

Pudney:1982:ELV

[Pud82] S. E. Pudney. Estimating latent variable systems when specification is uncertain: generalized component analysis and the eliminant method. Journal of the American Statistical Association, 77 (380):883–889, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287322.

Pullum:1985:BRB

[Pul85] Thomas W. Pullum. Book review: [untitled]. Journal of the American Statistical Association, 80(391):778–779, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288510.

Purdue:1988:BRB

[Pur88] Peter Purdue. Book review: [untitled]. Journal of the American Statistical Association, 83(402):562–563, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288880.

Purdue:1989:BRB

[Pur89] Peter Purdue. Book review: [untitled]. Journal of the American Statistical Association, 84(407):845, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289691.

Puterman:1980:BRB

[Put80] Martin L. Puterman. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1033–1035, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287204.

Puterman:1982:BRB

[Put82] Martin L. Puterman. Book review: [untitled]. Journal of the American Statistical Association, 77(378):489, June 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287274.

Puterman:1985:BRB

[Put85] Martin L. Puterman. Book review: [untitled]. Journal of the American Statistical Association, 80(390):491, June 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287942.

Pregibon:1985:EOT

[PV85] Daryl Pregibon and Yehuda Vardi. Estimating optimal transformations for multiple regression and correlation: Comment. Journal of the American Statistical Association, 80(391):598–601, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288474.

Quade:1980:BRB

[Qua80] Dana Quade. Book review: [untitled]. Journal of the American Statistical Association, 75(369):246, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287430.

Quigley:1982:BRB

[Qui82] John M. Quigley. Book review: [untitled]. Journal of the American Statistical Association, 77(380):950–951, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287356.

Ramsay:1989:BRM

[RA89a] J. O. Ramsay and M. Abrahamowicz. Binomial regression with monotone splines: a psychometric application. Journal of the American Statistical Association, 84(408):906–915, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290065.

Ruppert:1989:TSH

[RA89b] David Ruppert and Brian Aldershof. Transformations to symmetry and homoscedasticity. Journal of the American Statistical Association, 84(406):437–446, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289927.

Raftery:1987:IPG

[Raf87] Adrian E. Raftery. Inference and prediction for a general order statistic model with unknown population size. Journal of the American Statistical Association, 82(400):1163–1168, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raghavarao:1989:BRB

[Rag89] Damaraju Raghavarao. Book review: [untitled]. Journal of the American Statistical Association, 84(406):619, June 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289960.

Raj:1980:MCS

[Raj80] Baldev Raj. A Monte Carlo study of small-sample properties of simultaneous equation estimators with normal and nonnormal disturbances. Journal of the American Statistical Association, 75 (369):221–229, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287415.

Ramamoorthi:1985:BRB

[Ram85a] R. V. Ramamoorthi. Book review: [untitled]. Journal of the American Statistical Association, 80(390):489, June 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287936.

Ramaswami:1985:BRB

[Ram85b] V. Ramaswami. Book review: [untitled]. Journal of the American Statistical Association, 80(391):786–787, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288522.

Randles:1984:TAR

[Ran84] Ronald H. Randles. On tests applied to residuals. Journal of the American Statistical Association, 79(386):349–354, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288275.

Randles:1986:BRB

[Ran86] Ronald H. Randles. Book review: [untitled]. Journal of the American Statistical Association, 81(393):258, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288010.

Randles:1989:DFM

[Ran89] Ronald H. Randles. A distribution-free multivariate sign test based on interdirections. Journal of the American Statistical Association, 84(408):1045–1050, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2290081.

Rao:1980:BRB

[Rao80a] B. L. S. Prakasa Rao. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1044, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287224.

Rao:1980:ECM

[Rao80b] J. N. K. Rao. Estimating the common mean of possibly different normal populations: a simulation study. Journal of the American Statistical Association, 75(370):447–453, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287475.

Rao:1981:ENT

[Rao81] Poduri S. R. S. Rao. Efficiencies of nine two-phase ratio estimators for the mean. Journal of the American Statistical Association, 76 (374):434–442, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287846.

Rao:1984:BRB

[Rao84] C. R. Rao. Book review: [untitled]. Journal of the American Statistical Association, 79(388):942, December 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288730.

Raveh:1989:NAL

[Rav89] Adi Raveh. A nonmetric approach to linear discriminant analysis. Journal of the American Statistical Association, 84(405):176– 183, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289861.

Ruppert:1980:TLS

[RC80] David Ruppert and Raymond J. Carroll. Trimmed least squares estimation in the linear model. Journal of the American Statistical Association, 75(372):828–838, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287169. See corrigendum [RC82].

Royall:1981:ESRa

[RC81a] Richard M. Royall and William G. Cumberland. An empirical study of the ratio estimator and estimators of its variance. Journal of the American Statistical Association, 76(373):66–77, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287043.

Royall:1981:ESRb

[RC81b] Richard M. Royall and William G. Cumberland. An empirical study of the ratio estimator and estimators of its variance: Rejoinder. Journal of the American Statistical Association, 76(373): 87–88, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287049.

Royall:1981:FPL

[RC81c] Richard M. Royall and William G. Cumberland. The finitepopulation linear regression estimator and estimators of its variance—an empirical study. Journal of the American Statistical Association, 76(376):924–930, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287589.

Ruppert:1982:CTL

[RC82] David Ruppert and Raymond J. Carroll. Corrigendum: “Trimmed least squares estimation in the linear model” [J. Amer. Statist.

Assoc. 75 (1980), no. 372, 828–838, MR 82d:62073]. Journal of the American Statistical Association, 77(380):954, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287359. See

[RC80].

Royall:1985:CCP

[RC85] Richard M. Royall and William G. Cumberland. Conditional coverage properties of finite population confidence intervals. Journal of the American Statistical Association, 80(390):355–359, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287896.

Read:1981:RCC

[Rea81] R. R. Read. Representation of certain covariance matrices with application to asymptotic efficiency. Journal of the American Statistical Association, 76(373):148–154, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287060.

Read:1984:SSC

[Rea84] Timothy R. C. Read. Small-sample comparisons for the power divergence goodness-of-fit statistics. Journal of the American Statistical Association, 79(388):929–935, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288726.

Redfern:1986:BRB

[Red86] Philip Redfern. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1110–1112, December 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289092.

Regal:1980:BRB

[Reg80] Ronald Regal. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1032–1033, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287202.

Reinsel:1980:BRB

[Rei80] Gregory Reinsel. Book review: [untitled]. Journal of the American Statistical Association, 75(370):474, June 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287488.

Reinsel:1981:BRB

[Rei81] Gregory Reinsel. Book review: [untitled]. Journal of the American Statistical Association, 76(374):495, June 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287867.

Reinsel:1982:MRM

[Rei82] Greg Reinsel. Multivariate repeated-measurement or growth curve models with multivariate random-effects covariance structure. Journal of the American Statistical Association, 77(377): 190–195, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287788.

Reinsel:1984:EPM

[Rei84] Gregory Reinsel. Estimation and prediction in a multivariate random effects generalized linear model. Journal of the American Statistical Association, 79(386):406–414, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288283.

Reinsel:1985:MSE

[Rei85] Gregory C. Reinsel. Mean squared error properties of empirical Bayes estimators in a multivariate random effects general linear model. Journal of the American Statistical Association, 80(391): 642–650, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288480.

Renshaw:1985:BRB

[Ren85] Edward Renshaw. Book review: [untitled]. Journal of the American Statistical Association, 80(389):255–256, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288108.

Resnick:1985:BRB

[Res85] Sidney Resnick. Book review: [untitled]. Journal of the American Statistical Association, 80(390):489–490, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287937.

Randles:1980:ADF

[RFPW80] Ronald H. Randles, Michael A. Fligner, George E. Policello, II, and Douglas A. Wolfe. An asymptotically distribution-free test for symmetry versus asymmetry. Journal of the American Statistical Association, 75(369):168–172, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287406.

Richmond:1982:GMC

[Ric82] J. Richmond. A general method for constructing simultaneous confidence intervals. Journal of the American Statistical Association, 77(378):455–460, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287267.

Richards:1984:BRB

[Ric84] Dale O. Richards. Book review: [untitled]. Journal of the American Statistical Association, 79(388):949, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288738.

Richter:1985:BRB

[Ric85] Donald Richter. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1079, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288601.

Richard:1988:BRB

[Ric88a] J. F. Richard. Book review: [untitled]. Journal of the American Statistical Association, 83(402):573, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288895.

Riche:1988:BRB

[Ric88b] Martha Farnsworth Riche. Book review: [untitled]. Journal of the American Statistical Association, 83(402):572, June 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288894.

Riddell:1982:BRB

[Rid82] W. Craig Riddell. Book review: [untitled]. Journal of the American Statistical Association, 77(377):217–219, March 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287803.

Rinaman:1983:DFR

[Rin83a] William C. Rinaman, Jr. On distribution-free rank tests for twoway layouts. Journal of the American Statistical Association, 78 (383):655–659, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2288134. See corrections [Rin85a].

Ringland:1983:RMC

[Rin83b] James T. Ringland. Robust multiple comparisons. Journal of the American Statistical Association, 78(381):145–151, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287121.

Rinaman:1985:CDF

[Rin85a] William C. Rinaman, Jr. Corrections: “On distribution-free rank tests for two-way layouts” [J. Amer. Statist. Assoc. 78 (1983), no. 383, 655–659, MR 85a:62120]. Journal of the American Statistical Association, 80(392):1084, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288610. See [Rin83a].

Rindskopf:1985:BRB

[Rin85b] David Rindskopf. Book review: [untitled]. Journal of the American Statistical Association, 80(390):482, June 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287925.

Ripley:1984:BRB

[Rip84] Brian D. Ripley. Book review: Spatial Processes: Models and Applications, by A. D. Cliff; J. K. Ord. Journal of the American Statistical Association, 79(385):238, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288381.

Rivest:1986:BCH

[Riv86] Louis-Paul Rivest. Bartlett’s, Cochran’s, and Hartley’s tests on variances are liberal when the underlying distribution is longtailed. Journal of the American Statistical Association, 81(393): 124–128, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287978.

Rao:1981:EOW

[RKC81] Poduri S. R. S. Rao, Jack Kaplan, and William G. Cochran. Estimators for the one-way random effects model with unequal error variances. Journal of the American Statistical Association, 76 (373):89–97, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287050.

Ramsay:1980:PRB

[RN80] J. O. Ramsay and M. R. Novick. PLU robust Bayesian decision theory: Point estimation. Journal of the American Statistical Association, 75(372):901–907, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287179.

Roberts:1986:DLD

[Rob86] Harry V. Roberts. Disclosure-limited data dissemination: Comment. Journal of the American Statistical Association, 81(393): 25–27, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287963.

Robinson:1987:CRE

[Rob87] J. Robinson. Conditioning ratio estimates under simple random sampling. Journal of the American Statistical Association, 82 (399):826–831, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2288792. See comment [Kot89].

Robert:1989:BRB

[Rob89a] Christian Robert. Book review: [untitled]. Journal of the American Statistical Association, 84(407):843, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289688.

Roberts:1989:BRB

[Rob89b] Rosemary A. Roberts. Book review: [untitled]. Journal of the American Statistical Association, 84(407):842–843, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289686.

Rock:1985:BRB

[Roc85] Donald Rock. Book review: [untitled]. Journal of the American Statistical Association, 80(391):775, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288505.

Rocke:1989:BBA

[Roc89] David M. Rocke. Bootstrap Bartlett adjustment in seemingly unrelated regression. Journal of the American Statistical Association, 84(406):598–601, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289948.

Rodgers:1982:BRBa

[Rod82a] Jack Rodgers. Book review: [untitled]. Journal of the American Statistical Association, 77(377):222, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287810.

Rodgers:1982:BRBb

[Rod82b] Jack Rodgers. Book review: [untitled]. Journal of the American Statistical Association, 77(380):940–941, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287341.

Rodriguez:1985:BRB

[Rod85] Carlos C. Rodriguez. Book review: [untitled]. Journal of the American Statistical Association, 80(391):771–772, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288501.

Rodriguez:1986:BRB

[Rod86] Carlos C. Rodriguez. Book review: [untitled]. Journal of the American Statistical Association, 81(395):866–867, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289037.

Rogosa:1980:UEB

[Rog80] David Rogosa. Using empirical Bayes techniques in the law school validity studies: Comment. Journal of the American Statistical Association, 75(372):819–821, December 1980. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287165.

Rogers:1983:BRB

[Rog83] William H. Rogers. Book review: [untitled]. Journal of the American Statistical Association, 78(381):203, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287135.

Rogan:1986:ACW

[Rog86a] Walter J. Rogan. An analysis of contaminated well water and health effects in Woburn, Massachusetts: Comment. Journal of the American Statistical Association, 81(395):602–603, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288985.

Rogers:1986:PMP

[Rog86b] Andrei Rogers. Parameterized multistate population dynamics and projections. Journal of the American Statistical Association, 81(393):48–61, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287967.

Rogers:1988:BET

[Rog88] Alan J. Rogers. The Bahadur efficiency of tests of some joint hypotheses. Journal of the American Statistical Association, 83 (401):257–264, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288948.

Rohlf:1983:BRBa

[Roh83a] F. James Rohlf. Book review: [untitled]. Journal of the American Statistical Association, 78(382):508–509, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288693.

Rohlf:1983:BRBb

[Roh83b] F. James Rohlf. Book review: [untitled]. Journal of the American Statistical Association, 78(383):733, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288149.

Rohlf:1983:BRBc

[Roh83c] F. James Rohlf. Book review: [untitled]. Journal of the American Statistical Association, 78(383):735, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288152.

Rohlf:1983:BRBd

[Roh83d] F. James Rohlf. Book review: [untitled]. Journal of the American Statistical Association, 78(383):735, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288153.

Rolph:1985:EPC

[Rol85] John E. Rolph. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):125–126, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288059.

Romano:1988:BRS

[Rom88] Joseph P. Romano. A bootstrap revival of some nonparametric distance tests. Journal of the American Statistical Association, 83(403):698–708, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Rosenberger:1982:BRB

[Ros82] James L. Rosenberger. Book review: [untitled]. Journal of the American Statistical Association, 77(378):498, June 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287288.

Rose:1984:OSS

[Ros84a] John S. Rose. Optimal sequential selection based on relative ranks with renewable call options. Journal of the American Statistical Association, 79(386):430–435, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288286.

Rosenbaum:1984:CPT

[Ros84b] Paul R. Rosenbaum. Conditional permutation tests and the propensity score in observational studies. Journal of the American Statistical Association, 79(387):565–574, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288402.

Rosenbaum:1984:ACO

[Ros84c] Paul R. Rosenbaum. From association to causation in observational studies: The role of tests of strongly ignorable treatment assignment. Journal of the American Statistical Association, 79 (385):41–48, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288332.

Rosenberg:1984:BRB

[Ros84d] William J. Rosenberg. Book review: [untitled]. Journal of the American Statistical Association, 79(387):744, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288457.

Rosenfeld:1984:BRBa

[Ros84e] Bob Rosenfeld. Book review: [untitled]. Journal of the American Statistical Association, 79(388):960, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288759.

Rosenfeld:1984:BRBb

[Ros84f] Bob Rosenfeld. Book review: [untitled]. Journal of the American Statistical Association, 79(388):960, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288760.

Rosalsky:1987:BRB

[Ros87a] Andrew Rosalsky. Book review: [untitled]. Journal of the American Statistical Association, 82(398):700–701, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289512.

Rosenbaum:1987:BRB

[Ros87b] Paul R. Rosenbaum. Book review: [untitled]. Journal of the American Statistical Association, 82(397):350–351, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289186.

Rosenbaum:1987:MBD

[Ros87c] Paul R. Rosenbaum. Model-based direct adjustment. Journal of the American Statistical Association, 82(398):387–394, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289440.

Rosen:1988:BRB

[Ros88a] David Rosen. Book review: [untitled]. Journal of the American Statistical Association, 83(401):268, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288953.

Rosenbaum:1988:BRB

[Ros88b] Paul R. Rosenbaum. Book review: [untitled]. Journal of the American Statistical Association, 83(402):568–569, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288889.

Rosenbaum:1989:OMO

[Ros89a] Paul R. Rosenbaum. Optimal matching for observational studies. Journal of the American Statistical Association, 84(408): 1024–1032, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290079.

Rosner:1989:MMC

[Ros89b] Bernard Rosner. Multivariate methods for clustered binary data with more than one level of nesting. Journal of the American Statistical Association, 84(406):373–380, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289920.

Rossi:1989:BRB

[Ros89c] Peter H. Rossi. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1093–1094, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290092.

Rossman:1989:BRB

[Ros89d] Allan J. Rossman. Book review: [untitled]. Journal of the American Statistical Association, 84(406):621, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289964.

Rothman:1982:EHH

[Rot82] David Rothman. Estimating historical heights: Comment. Journal of the American Statistical Association, 77(378):298, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287235.

Roth:1983:RTT

[Rot83] Arthur J. Roth. Robust trend tests derived and simulated: analogs of the Welch and Brown-Forsythe tests. Journal of the American Statistical Association, 78(384):972–980, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288212.

Rousseeuw:1984:LMS

[Rou84] Peter J. Rousseeuw. Least median of squares regression. Journal of the American Statistical Association, 79(388):871–880, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288718.

Rowe:1987:BRB

[Row87] Judith S. Rowe. Book review: [untitled]. Journal of the American Statistical Association, 82(398):680, June 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289486.

Royall:1983:EMD

[Roy83] Richard M. Royall. An evaluation of model-dependent and probability-sampling inferences in sample surveys: Comment. Journal of the American Statistical Association, 78(384):794–796, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288183.

Royall:1986:PAR

[Roy86] Richard M. Royall. The prediction approach to robust variance estimation in two-stage cluster sampling. Journal of the American Statistical Association, 81(393):119–123, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287977.

Rosenbaum:1984:NDS

[RR84a] Paul R. Rosenbaum and Donald B. Rubin. On the nature and discovery of structure: Comment. Journal of the American Statistical Association, 79(385):26–28, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288329.

Rosenbaum:1984:RBO

[RR84b] Paul R. Rosenbaum and Donald B. Rubin. Reducing bias in observational studies using subclassification on the propensity score. Journal of the American Statistical Association, 79(387):516–524, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288398.

Rao:1981:ACD

[RS81] J. N. K. Rao and A. J. Scott. The analysis of categorical data from complex sample surveys: Chi-squared tests for goodness of fit and independence in two-way tables. Journal of the American Statistical Association, 76(374):221–230, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287815.

Rukhin:1982:EQE

[RS82] Andrew L. Rukhin and William E. Strawderman. Estimating a quantile of an exponential distribution. Journal of the American Statistical Association, 77(377):159–162, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287783.

Rubin:1986:MII

[RS86] Donald B. Rubin and Nathaniel Schenker. Multiple imputation for interval estimation from simple random samples with ignorable nonresponse. Journal of the American Statistical Association, 81 (394):366–374, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289225.

Ross:1980:DPP

[RSW80] Sheldon M. Ross, Mehrdad Shahshahani, and Gideon Weiss. On the duration of the problem of the points. Journal of the American Statistical Association, 75(371):663–666, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287665.

Reinsel:1987:ICS

[RT87] Gregory C. Reinsel and George C. Tiao. Impact of chlorofluoromethanes on stratospheric ozone: a statistical analysis of ozone data for trends. Journal of the American Statistical Association, 82(397):20–30, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289120.

Rubin:1980:RAE

[Rub80a] Donald B. Rubin. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):591–593, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287653.

Rubin:1980:UEBa

[Rub80b] Donald B. Rubin. Using empirical Bayes techniques in the law school validity studies. Journal of the American Statistical Association, 75(372):801–816, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287162.

Rubin:1980:UEBb

[Rub80c] Donald B. Rubin. Using empirical Bayes techniques in the law school validity studies: Rejoinder. Journal of the American Statistical Association, 75(372):825–827, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287168.

Rubin:1983:EMD

[Rub83] Donald B. Rubin. An evaluation of model-dependent and probability-sampling inferences in sample surveys: Comment. Journal of the American Statistical Association, 78(384):803–805, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288187.

Rubin:1984:ATD

[Rub84a] Donald B. Rubin. The analysis of transformed data: Comment. Journal of the American Statistical Association, 79(386):309–312, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288265.

Rubin:1984:GMA

[Rub84b] Donald B. Rubin. Graphical methods for assessing logistic regression models: Comment. Journal of the American Statistical Association, 79(385):79–80, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288338.

Rubin:1985:SMP

[Rub85] Donald B. Rubin. A statistical model for positron emission tomography: Comment. Journal of the American Statistical Association, 80(389):31–32, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288035.

Rubin:1986:SCI

[Rub86] Donald B. Rubin. Statistics and causal inference: Comment: Which ifs have causal answers. Journal of the American Statistical Association, 81(396):961–962, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289065.

Rubin:1987:CPD

[Rub87] Donald B. Rubin. The calculation of posterior distributions by data augmentation: Comment: a noniterative sampling/importance resampling alternative to the data augmentation algorithm for creating a few imputations when fractions of missing information are modest: The SIR algorithm. Journal of the American Statistical Association, 82(398):543–546, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289460.

Ruberg:1989:CIM

[Rub89] Stephen J. Ruberg. Contrasts for identifying the minimum effective dose. Journal of the American Statistical Association, 84 (407):816–822, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289672.

Rukhin:1984:ACP

[Ruk84] Andrew L. Rukhin. Adaptive classification procedures. Journal of the American Statistical Association, 79(386):415–422, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288284.

Rukhin:1986:IEL

[Ruk86] Andrew L. Rukhin. Improved estimation in lognormal models. Journal of the American Statistical Association, 81(396):1046– 1049, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289081.

Rukhin:1987:HMB

[Ruk87] Andrew L. Rukhin. How much better are better estimators of a normal variance. Journal of the American Statistical Association, 82(399):925–928, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288806.

Ruppert:1985:BIR

[Rup85] David Ruppert. On the bounded-influence regression estimator of Krasker and Welsch. Journal of the American Statistical Association, 80(389):205–208, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288073.

Russek:1982:BRB

[Rus82] Estelle Russek. Book review: [untitled]. Journal of the American Statistical Association, 77(377):223, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287811.

Rustagi:1989:BRB

[Rus89] Jagdish S. Rustagi. Book review: [untitled]. Journal of the American Statistical Association, 84(407):847, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289695.

Rao:1985:ISS

[RW85] J. N. K. Rao and C. F. J. Wu. Inference from stratified samples: second-order analysis of three methods for nonlinear statistics. Journal of the American Statistical Association, 80(391): 620–630, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288478.

Rao:1988:RIC

[RW88] J. N. K. Rao and C. F. J. Wu. Resampling inference with complex survey data. Journal of the American Statistical Association, 83 (401):231–241, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288945.

Ryan:1988:BRB

[Rya88] David L. Ryan. Book review: [untitled]. Journal of the American Statistical Association, 83(402):573–574, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288896.

Sadooghi-Alvandi:1986:CSS

[SA86] Mohammad Sadooghi-Alvandi. The choice of subsample size in two-stage sampling. Journal of the American Statistical Association, 81(394):555–558, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289249.

Samaniego:1980:MLE

[Sam80] Francisco J. Samaniego. Maximum likelihood estimation for binomially distributed signals in discrete noise. Journal of the American Statistical Association, 75(369):117–121, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287398.

Samuels:1981:MSR

[Sam81] Stephen M. Samuels. Minimax stopping rules when the underlying distribution is uniform. Journal of the American Statistical Association, 76(373):188–197, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287066.

Samarov:1985:BIR

[Sam85a] Alexander M. Samarov. Bounded-influence regression via local minimax mean squared error. Journal of the American Statistical Association, 80(392):1032–1040, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288571.

Sampson:1985:BRB

[Sam85b] Paul D. Sampson. Book review: [untitled]. Journal of the American Statistical Association, 80(390):487–488, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287933.

Samuels:1985:BCP

[Sam85c] Stephen M. Samuels. A best-choice problem with linear travel cost. Journal of the American Statistical Association, 80(390): 461–464, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287915.

Sampson:1988:BRB

[Sam88] Allan R. Sampson. Book review: [untitled]. Journal of the American Statistical Association, 83(402):577–578, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288902.

Samuelson:1989:BRB

[Sam89] Robert J. Samuelson. Book review: [untitled]. Journal of the American Statistical Association, 84(406):631, June 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289982.

Sandberg:1989:BRB

[San89] Valdimar C. Sandberg. Book review: [untitled]. Journal of the American Statistical Association, 84(406):614–615, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289955.

Sarndal:1981:ESR

[Sär81] Carl Erik Särndal. An empirical study of the ratio estimator and estimators of its variance: Comment. Journal of the American Statistical Association, 76(373):81–82, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287045.

Sarndal:1984:DCV

[Sär84] Carl Erik Särndal. Design-consistent versus model-dependent estimation for small domains. Journal of the American Statistical Association, 79(387):624–631, September 1984. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288409.

Sarkar:1987:CBE

[Sar87] Sanat K. Sarkar. A continuous bivariate exponential distribution. Journal of the American Statistical Association, 82(398):667–675, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289480.

Sarkar:1989:QIO

[Sar89] Sanat K. Sarkar. Quasi-independence in ordinal triangular contingency tables. Journal of the American Statistical Association, 84(406):592–597, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289947.

Saunders:1983:BRBa

[Sau83a] Ian W. Saunders. Book review: [untitled]. Journal of the American Statistical Association, 78(383):737, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288157.

Saunders:1983:BRBb

[Sau83b] Ian W. Saunders. Book review: [untitled]. Journal of the American Statistical Association, 78(383):745, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288174.

Saunders:1985:BRB

[Sau85] Ian W. Saunders. Book review: [untitled]. Journal of the American Statistical Association, 80(389):250, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288101.

Savage:1980:BRB

[Sav80] I. Richard Savage. Book review: [untitled]. Journal of the American Statistical Association, 75(370):477–478, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287494.

Savage:1981:BRB

[Sav81] I. Richard Savage. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1011, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287606.

Savage:1985:HSP

[Sav85] I. Richard Savage. Hard-soft problems. Journal of the American Statistical Association, 80(389):1–7, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288029.

Savage:1986:BRB

[Sav86] I. Richard Savage. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1112–1114, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289093.

Sayre:1980:DAE

[Say80] James W. Sayre. The distributions of the actual error rates in linear discriminant analysis. Journal of the American Statistical Association, 75(369):201–205, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287411.

Salvas-Bronsard:1980:BRB

[SB80a] Lise Salvas-Bronsard. Book review: [untitled]. Journal of the American Statistical Association, 75(369):244–245, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287427.

Simkowitz:1980:ASD

[SB80b] Michael A. Simkowitz and William L. Beedles. Asymmetric stable distributed security returns. Journal of the American Statistical Association, 75(370):306–312, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287449.

Sihota:1981:ASC

[SB81] S. S. Sihota and K. S. Banerjee. On the algebraic structures in the construction of confounding plans in mixed factorial designs on the lines of White and Hultquist. Journal of the American Statistical Association, 76(376):996–1001, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287601.

Spitzer:1983:SSP

[SB83] John J. Spitzer and Richard T. Baillie. Small-sample properties of predictions from the regression model with autoregressive errors. Journal of the American Statistical Association, 78(382):258–263, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288621.

Shapiro:1987:ACS

[SB87] Alexander Shapiro and Michael W. Browne. Analysis of covariance structures under elliptical distributions. Journal of the American Statistical Association, 82(400):1092–1097, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schaeffer:1989:RBM

[SB89] Nora Cate Schaeffer and Norman M. Bradburn. Respondent behavior in magnitude estimation. Journal of the American Statistical Association, 84(406):402–413, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289923.

Smith:1980:CSRa

[SC80a] Gary Smith and Frank Campbell. A critique of some ridge regression methods. Journal of the American Statistical Association, 75(369):74–81, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287386.

Smith:1980:CSRb

[SC80b] Gary Smith and Frank Campbell. A critique of some ridge regression methods: Rejoinder. Journal of the American Statistical Association, 75(369):100–103, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287395.

Storer:1985:DCR

[SC85] Barry E. Storer and John Crowley. A diagnostic for Cox regression and general conditional likelihoods. Journal of the American Statistical Association, 80(389):139–147, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288064.

Stern:1989:MEL

[SC89] Hal Stern and Thomas M. Cover. Maximum entropy and the lottery. Journal of the American Statistical Association, 84(408): 980–985, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290073.

Scarr:1987:BRB

[Sca87] Harry A. Scarr. Book review: [untitled]. Journal of the American Statistical Association, 82(398):679–680, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289485.

Schoenstadt:1980:CLR

[Sch80a] Arthur L. Schoenstadt. A comparison of the Lieberman–Ross and Man–Grubbs methods. Journal of the American Statistical Association, 75(369):212–216, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287413.

Schott:1980:BRB

[Sch80b] Brian Schott. Book review: [untitled]. Journal of the American Statistical Association, 75(370):475–476, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287492.

Scharf:1981:BRB

[Sch81a] Louis L. Scharf. Book review: [untitled]. Journal of the American Statistical Association, 76(374):494, June 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287864.

Schlicht:1981:SAP

[Sch81b] Ekkehart Schlicht. A seasonal adjustment principle and a seasonal adjustment method derived from this principle. Journal of the American Statistical Association, 76(374):374–378, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287838. See

corrigendum [Sch82b].

Schuenemeyer:1981:EVR

[Sch81c] John H. Schuenemeyer. Estimating volumes of remaining fossil fuel resources: a critical review: Comment. Journal of the American Statistical Association, 76(375):554–558, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287511.

Schervish:1982:BRB

[Sch82a] Mark J. Schervish. Book review: [untitled]. Journal of the American Statistical Association, 77(378):489, June 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287273.

Schlicht:1982:CSA

[Sch82b] Ekkehart Schlicht. Corrigendum: “A seasonal adjustment principle and a seasonal adjustment method derived from this principle” [J. Amer. Statist. Assoc. 76 (1981), no. 374, 374–378, MR 82h:62162]. Journal of the American Statistical Association, 77 (380):954, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287363. See [Sch81b].

Schmidt:1982:BRBa

[Sch82c] Peter Schmidt. Book review: [untitled]. Journal of the American Statistical Association, 77(377):217, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287802.

Schmidt:1982:BRBb

[Sch82d] Peter Schmidt. Book review: [untitled]. Journal of the American Statistical Association, 77(378):494, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287280.

Schervish:1983:BRB

[Sch83a] Mark J. Schervish. Book review: [untitled]. Journal of the American Statistical Association, 78(383):743–744, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288169.

Schervish:1983:UOI

[Sch83b] Mark J. Schervish. User-oriented inference. Journal of the American Statistical Association, 78(383):611–615, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288127.

Schmitt:1983:BRB

[Sch83c] Neal Schmitt. Book review: [untitled]. Journal of the American Statistical Association, 78(384):999, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288234.

Schwager:1983:RPS

[Sch83d] Steven J. Schwager. Run probabilities in sequences of Markovdependent trials. Journal of the American Statistical Association, 78(381):168–175, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287125.

Schervish:1984:BRB

[Sch84a] Mark J. Schervish. Book review: Prior Information in Linear Models, by Helge Toutenburg. Journal of the American Statistical Association, 79(385):235–236, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288377.

Schmidt:1984:BRB

[Sch84b] Peter Schmidt. Book review: Limited-Dependent and Qualitative Variables in Econometrics, by G. S. Maddala. Journal of the American Statistical Association, 79(387):738, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288445.

Schmoyer:1984:SCM

[Sch84c] Richard L. Schmoyer. Sigmoidally constrained maximum likelihood estimation in quantal bioassay. Journal of the American Statistical Association, 79(386):448–453, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288289.

Schnute:1984:LMN

[Sch84d] Jon Schnute. Linear mixtures: a new approach to bivariate trend lines. Journal of the American Statistical Association, 79(385): 1–8, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288325.

Schuster:1984:CPL

[Sch84e] Eugene F. Schuster. Classification of probability laws by tail behavior. Journal of the American Statistical Association, 79(388): 936–939, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288727.

Scheffer:1985:BRB

[Sch85a] C. L. Scheffer. Book review: [untitled]. Journal of the American Statistical Association, 80(391):788, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288524.

Schenker:1985:QAB

[Sch85b] Nathaniel Schenker. Qualms about bootstrap confidence intervals. Journal of the American Statistical Association, 80(390):360–361, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287897.

Schervish:1985:OSG

[Sch85c] Mark J. Schervish. A one-sided goodness-of-fit test for a multinomial population: Comment. Journal of the American Statistical Association, 80(391):562–563, September 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288470.

Schervish:1985:SCP

[Sch85d] Mark J. Schervish. Self-calibrating priors do not exist: Comment. Journal of the American Statistical Association, 80(390):341–342, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287893.

Schruben:1985:BRB

[Sch85e] Lee W. Schruben. Book review: [untitled]. Journal of the American Statistical Association, 80(390):494, June 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287950.

Schafer:1986:CIM

[Sch86a] Daniel W. Schafer. Combining information on measurement error in the errors-in-variables model. Journal of the American Statistical Association, 81(393):181–185, March 1986. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287987.

Schilling:1986:MTS

[Sch86b] Mark F. Schilling. Multivariate two-sample tests based on nearest neighbors. Journal of the American Statistical Association, 81 (395):799–806, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289012.

Schoenfeld:1986:BRB

[Sch86c] David Schoenfeld. Book review: [untitled]. Journal of the American Statistical Association, 81(394):572, June 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289259.

Schoenfeld:1986:CBN

[Sch86d] David A. Schoenfeld. Confidence bounds for normal means under order restrictions, with application to dose-response curves, toxicology experiments, and low-dose extrapolation. Journal of the American Statistical Association, 81(393):186–195, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287988.

Schaible:1987:BRB

[Sch87a] Wesley L. Schaible. Book review: [untitled]. Journal of the American Statistical Association, 82(397):351–352, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289187.

Schenker:1987:BBC

[Sch87b] Nathaniel Schenker. Better bootstrap confidence intervals: Comment. Journal of the American Statistical Association, 82(397): 192–194, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289150.

Schneiderman:1987:BRB

[Sch87c] Marvin A. Schneiderman. Book review: [untitled]. Journal of the American Statistical Association, 82(398):695–697, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289505.

Schoen:1987:BRB

[Sch87d] Robert Schoen. Book review: [untitled]. Journal of the American Statistical Association, 82(399):961, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288837.

Schrodt:1987:BRB

[Sch87e] Philip A. Schrodt. Book review: [untitled]. Journal of the American Statistical Association, 82(399):961–962, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288838.

Schucany:1987:BBC

[Sch87f] William R. Schucany. Better bootstrap confidence intervals: Comment. Journal of the American Statistical Association, 82(397): 196–197, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289152.

Schwarz:1987:LAV

[Sch87g] Gideon Schwarz. Least-absolute-value prediction lines in closed form. Journal of the American Statistical Association, 82(400): 1150–1152, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Schwarz:1987:MPL

[Sch87h] Gideon Schwarz. A minimax property of linear regression. Journal of the American Statistical Association, 82(397):220, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289157.

Schervish:1988:APC

[Sch88a] Mark J. Schervish. Applications of parallel computation to statistical inference. Journal of the American Statistical Association, 83(404):976–983, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290123.

Schmoyer:1988:LIN

[Sch88b] Richard L. Schmoyer. Linear interpolation with a nonparametric accelerated failure-time model. Journal of the American Statistical Association, 83(402):441–449, June 1988. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288860.

Schneiderman:1988:BRB

[Sch88c] M. A. Schneiderman. Book review: Correction: Book Reviews. Journal of the American Statistical Association, 83(404):1232, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290188.

Schwager:1988:BRB

[Sch88d] Steven J. Schwager. Book review: [untitled]. Journal of the American Statistical Association, 83(401):274, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288962.

Schervish:1989:BRB

[Sch89a] Mark J. Schervish. Book review: [untitled]. Journal of the American Statistical Association, 84(405):333–334, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289886.

Schmidley:1989:BRB

[Sch89b] A. Dianne Schmidley. Book review: [untitled]. Journal of the American Statistical Association, 84(407):849, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289699.

Schoenberg:1989:BRB

[Sch89c] Ronald Schoenberg. Book review: [untitled]. Journal of the American Statistical Association, 84(407):841, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289684.

Scott:1980:DEB

[Sco80a] David W. Scott. Density estimation and bump-hunting by the penalized likelihood method exemplified by scattering and meteorite data: Comment. Journal of the American Statistical Association, 75(369):61–62, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287380.

Scott:1980:UEB

[Sco80b] Elizabeth L. Scott. Using empirical Bayes techniques in the law school validity studies: Comment. Journal of the American Statistical Association, 75(372):821–823, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287166.

Scott:1982:BRB

[Sco82] Mark Scott. Book review: [untitled]. Journal of the American Statistical Association, 77(380):948, December 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287352.

Scott:1983:BRB

[Sco83] Mark Scott. Book review: [untitled]. Journal of the American Statistical Association, 78(383):736, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288155.

Scott:1985:FPT

[Sco85] David W. Scott. Frequency polygons: Theory and application. Journal of the American Statistical Association, 80(390):348–354, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287895.

Scott:1988:HFA

[Sco88] David W. Scott. How far are automatically chosen regression smoothing parameters from their optimum?: Comment. Journal of the American Statistical Association, 83(401):96–98, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288923.

Said:1985:HTA

[SD85a] Said E. Said and David A. Dickey. Hypothesis testing in ARIMA(p, 1, q) models. Journal of the American Statistical Association, 80(390):369–374, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287899.

Said:1985:HTM

[SD85b] Said E. Said and David A. Dickey. Hypothesis testing in arima(p, 1, q) models. Journal of the American Statistical Association, 80(390):369–374, ???? 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Scheers:1988:CRR

[SD88] N. J. Scheers and C. Mitchell Dayton. Covariate randomized response models. Journal of the American Statistical Association, 83(404):969–974, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290121.

Seaman:1982:BRB

[Sea82] Jeff Seaman. Book review: [untitled]. Journal of the American Statistical Association, 77(378):497, June 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287283.

Seaks:1986:BRB

[Sea86] Terry G. Seaks. Book review: [untitled]. Journal of the American Statistical Association, 81(395):867–868, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289038.

Searle:1989:BRB

[Sea89] Shayle R. Searle. Book review: [untitled]. Journal of the American Statistical Association, 84(407):847, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289696.

Sedransk:1984:BRB

[Sed84] N. Sedransk. Book review: [untitled]. Journal of the American Statistical Association, 79(386):475–476, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288303.

Sedransk:1985:EPC

[Sed85] J. Sedransk. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):126–127, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288060.

Sedransk:1986:I

[Sed86] Joseph Sedransk. [introduction]. Journal of the American Statistical Association, 81(394):337, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289221.

Seely:1980:SRE

[See80] Justus F. Seely. Some remarks on exact confidence intervals for positive linear combinations of variance components. Journal of the American Statistical Association, 75(370):372–374, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287461.

Selen:1986:AEC

[Sel86] Jan Selén. Adjusting for errors in classification and measurement in the analysis of partly and purely categorical data. Journal of the American Statistical Association, 81(393):75–81, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287969.

Sen:1980:BRB

[Sen80] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 75(370):473, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287486.

Sen:1981:BRBa

[Sen81a] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 76(373):210, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287079.

Sen:1981:BRBb

[Sen81b] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 76(374):494–495, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287865.

Sen:1981:BRBc

[Sen81c] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 76(375):752, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287559.

Sen:1982:BRB

[Sen82] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 77(377):221, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287807.

Sen:1983:BRBa

[Sen83a] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 78(381):212, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287151.

Sen:1983:BRBb

[Sen83b] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 78(383):732, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288147.

Sen:1983:BRBc

[Sen83c] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 78(384):989–990, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288220.

Sen:1984:BRBa

[Sen84a] P. K. Sen. Book review: Aspects of Multivariate Statistical Theory, by Robb J. Muirhead. Journal of the American Statistical Association, 79(385):231–232, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288369.

Sen:1984:BRBb

[Sen84b] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 79(386):474, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288301.

Sen:1984:BRBc

[Sen84c] P. K. Sen. Book review: [untitled]. Journal of the American Statistical Association, 79(387):732–733, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288433.

Sen:1986:GCP

[Sen86] Pranab Kumar Sen. The Gini coefficient and poverty indexes: some reconciliations. Journal of the American Statistical Association, 81(396):1050–1057, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289082.

Serow:1986:BRB

[Ser86] William J. Serow. Book review: [untitled]. Journal of the American Statistical Association, 81(393):261, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288017.

Scott:1981:CMC

[SF81a] David W. Scott and Lynette E. Factor. Corrigenda: Monte Carlo study of three data-based nonparametric probability density estimators. Journal of the American Statistical Association, 76(376): 1022, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287630.

Scott:1981:MCS

[SF81b] David W. Scott and Lynette E. Factor. Monte Carlo study of three data-based nonparametric probability density estimators. Journal of the American Statistical Association, 76(373):9–15, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287033.

Spruill:1982:ECC

[SG82] Nancy L. Spruill and Joseph L. Gastwirth. On the estimation of the correlation coefficient from grouped data. Journal of the American Statistical Association, 77(379):614–620, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287724.

Sebenius:1983:DBI

[SG83] James K. Sebenius and John Geanakoplos. Don’t bet on it: contingent agreements with asymmetric information. Journal of the American Statistical Association, 78(382):424–426, June 1983.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288651.

Sallas:1981:BLR

[SH81] William M. Sallas and David A. Harville. Best linear recursive estimation for mixed linear models. Journal of the American Statistical Association, 76(376):860–869, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287581.

Scott:1982:ETS

[SH82] A. J. Scott and D. Holt. The effect of two-stage sampling on ordinary least squares methods. Journal of the American Statistical Association, 77(380):848–854, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287317.

Schwertman:1986:SDM

[SH86] Neil C. Schwertman and Lance K. Heilbrun. A successive differences method for growth curves with missing data and random observation times. Journal of the American Statistical Association, 81(396):912–916, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289059.

Simkin:1987:IPA

[SH87] David Simkin and Reid Hastie. An information-processing analysis of graph perception. Journal of the American Statistical Association, 82(398):454–465, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289447.

Sarndal:1989:SDE

[SH89] Carl-Erik Särndal and Michael A. Hidiroglou. Small domain estimation: a conditional analysis. Journal of the American Statistical Association, 84(405):266–275, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289873.

Shaffer:1981:AVM

[Sha81a] Juliet Popper Shaffer. The analysis of variance mixed model with allocated observations: Application to repeated measurement designs. Journal of the American Statistical Association, 76(375):

607–611, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287518.

Shaffer:1981:CIC

[Sha81b] Juliet Popper Shaffer. Complexity: An interpretability criterion for multiple comparisons. Journal of the American Statistical Association, 76(374):395–401, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287841.

Shaughnessy:1981:MRD

[Sha81c] Peter W. Shaughnessy. Multiple runs distributions: Recurrences and critical values. Journal of the American Statistical Association, 76(375):732–736, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287540.

Shafer:1982:LP

[Sha82a] Glenn Shafer. Lindley’s paradox. Journal of the American Statistical Association, 77(378):325–334, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287244.

Shafer:1982:LPR

[Sha82b] Glenn Shafer. Lindley’s paradox: Rejoinder. Journal of the American Statistical Association, 77(378):349–351, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287251.

Shafer:1982:LPD

[Sha82c] Glenn Shafer. Lindley’s paradox. With discussion and with a reply by the author. Journal of the American Statistical Association, 77(378):325–351, ???? 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Shapiro:1983:SAO

[Sha83] C. P. Shapiro. Sequential allocation and optimal stopping in Bayesian simultaneous estimation. Journal of the American Statistical Association, 78(382):396–402, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288646.

Shafer:1984:BRBb

[Sha84a] Glenn R. Shafer. Book review: Expected Utility Hypotheses and the Allais Paradox, by Maurice Allais; Ole Hagen. Journal of the American Statistical Association, 79(385):224–225, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288363.

Shafer:1984:BRBa

[Sha84b] Glenn R. Shafer. Book review: Judgment Under Uncertainty: Heuristics and Biases, by Daniel Kahneman; Paul Slovic; Amos Tversky. Journal of the American Statistical Association, 79(385): 223–224, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288362.

Shapiro:1985:ASE

[Sha85] C. Page Shapiro. Allocation schemes for estimating the product of positive parameters. Journal of the American Statistical Association, 80(390):449–454, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287913.

Shaffer:1986:MSR

[Sha86a] Juliet Popper Shaffer. Modified sequentially rejective multiple test procedures. Journal of the American Statistical Association, 81 (395):826–831, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289016.

Shapiro:1986:ATO

[Sha86b] Alexander Shapiro. Asymptotic theory of overparameterized structural models. Journal of the American Statistical Association, 81(393):142–149, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287981.

Shao:1989:ECA

[Sha89a] Jun Shao. The efficiency and consistency of approximations to the jackknife variance estimators. Journal of the American Statistical Association, 84(405):114–119, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289853.

Shao:1989:MCA

[Sha89b] Jun Shao. Monte Carlo approximations in Bayesian decision theory. Journal of the American Statistical Association, 84(407):727– 732, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289654.

Shea:1980:BRB

[She80a] Gerald Shea. Book review: [untitled]. Journal of the American Statistical Association, 75(369):242, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287423.

Shelley:1980:BRB

[She80b] Mack Shelley. Book review: [untitled]. Journal of the American Statistical Association, 75(371):748–750, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287684.

Sherman:1981:BRB

[She81] Malcolm J. Sherman. Book review: [untitled]. Journal of the American Statistical Association, 76(375):745–746, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287549.

Shelley:1982:BRB

[She82a] M. C. Shelley. Book review: [untitled]. Journal of the American Statistical Association, 77(379):692–694, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287755.

Sherman:1982:BRBa

[She82b] Malcolm J. Sherman. Book review: [untitled]. Journal of the American Statistical Association, 77(378):491–492, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287276.

Sherman:1982:BRBb

[She82c] Malcolm J. Sherman. Book review: [untitled]. Journal of the American Statistical Association, 77(380):948–949, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287353.

Shelley:1983:BRB

[She83a] Mack Shelley. Book review: [untitled]. Journal of the American Statistical Association, 78(382):504, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288684.

Sherman:1983:BRB

[She83b] Malcolm J. Sherman. Book review: [untitled]. Journal of the American Statistical Association, 78(384):994–995, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288227.

Shelley:1984:BRB

[She84a] Mack Shelley. Book review: Content Analysis: An Introduction to its Methodology, by Klaus Krippendorff. Journal of the American Statistical Association, 79(385):240, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288384.

Sherman:1984:BRB

[She84b] Malcolm J. Sherman. Book review: [untitled]. Journal of the American Statistical Association, 79(388):957, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288753.

Sheatsley:1985:BRB

[She85] Paul B. Sheatsley. Book review: [untitled]. Journal of the American Statistical Association, 80(391):773–774, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288503.

Sherman:1989:BRB

[She89] Malcolm J. Sherman. Book review: [untitled]. Journal of the American Statistical Association, 84(405):348–349, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289908.

Shinozaki:1980:EMN

[Shi80] Nobuo Shinozaki. Estimation of a multivariate normal mean with a class of quadratic loss function. Journal of the American Statistical Association, 75(372):973–976, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287190.

Shibata:1984:BRB

[Shi84a] Hirofumi Shibata. Book review: [untitled]. Journal of the American Statistical Association, 79(387):739–740, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288448.

Shiller:1984:SPN

[Shi84b] Robert J. Shiller. Smoothness priors and nonlinear regression. Journal of the American Statistical Association, 79(387):609–615, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288407.

Shiraishi:1985:AAA

[Shi85] Taka-Aki Shiraishi. An asymptotic acceptance of aligned rank tests under alternatives of contaminated distributions in a randomized-blocks design. Journal of the American Statistical Association, 80(391):748–752, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288495.

Shreve:1984:BRB

[Shr84] Steven Shreve. Book review: [untitled]. Journal of the American Statistical Association, 79(388):955, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288748.

Shrout:1987:BRB

[Shr87] Patrick E. Shrout. Book review: [untitled]. Journal of the American Statistical Association, 82(397):361, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289201.

Skinner:1986:ESD

[SHS86] C. J. Skinner, D. J. Holmes, and T. M. F. Smith. The effect of sample design on principal component analysis. Journal of the American Statistical Association, 81(395):789–798, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289011.

Spurrier:1985:ESC

[SI85] John D. Spurrier and Steven P. Isham. Exact simultaneous confidence intervals for pairwise comparisons of three normal means. Journal of the American Statistical Association, 80(390):438–442, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287911.

Siegel:1980:TPT

[Sie80] Andrew F. Siegel. Testing for periodicity in a time series. Journal of the American Statistical Association, 75(370):345–348, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287456.

Siegel:1983:LML

[Sie83] Andrew F. Siegel. Low median and least absolute residual analysis of two-way tables. Journal of the American Statistical Association, 78(382):371–374, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288642.

Sikkel:1985:MME

[Sik85] Dirk Sikkel. Models for memory effects. Journal of the American Statistical Association, 80(392):835–841, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288541.

Silverman:1980:DEB

[Sil80] B. W. Silverman. Density estimation and bump-hunting by the penalized likelihood method exemplified by scattering and meteorite data: Comment. Journal of the American Statistical Association, 75(369):67–68, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287382.

Silverman:1984:FEC

[Sil84] B. W. Silverman. A fast and efficient cross-validation method for smoothing parameter choice in spline regression. Journal of the American Statistical Association, 79(387):584–589, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288404.

Sills:1986:BRB

[Sil86] David L. Sills. Book review: [untitled]. Journal of the American Statistical Association, 81(393):257, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288007.

Simon:1980:BRB

[Sim80] Gary Simon. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1036, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287207.

Simon:1985:BRBc

[Sim85a] Gary Simon. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1078, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288599.

Simon:1985:BRBa

[Sim85b] Gary A. Simon. Book review: [untitled]. Journal of the American Statistical Association, 80(389):240–241, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288086.

Simon:1985:BRBb

[Sim85c] Gary A. Simon. Book review: [untitled]. Journal of the American Statistical Association, 80(389):241–242, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288087.

Simonoff:1985:IGF

[Sim85d] Jeffrey S. Simonoff. An improved goodness-of-fit statistic for sparse multinomials. Journal of the American Statistical Association, 80(391):671–677, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288483.

Simonoff:1986:JBG

[Sim86] Jeffrey S. Simonoff. Jackknifing and bootstrapping goodness-of-fit statistics in sparse multinomials. Journal of the American Statistical Association, 81(396):1005–1011, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289075.

Simon:1987:BRB

[Sim87a] Gary Simon. Book review: [untitled]. Journal of the American Statistical Association, 82(397):348–349, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289184.

Simonoff:1987:CPS

[Sim87b] Jeffrey S. Simonoff. Comment on: “Performance of some resistant rules for outlier labeling” [J. Amer. Statist. Assoc. 81 (1986), no. 396, 991–999] by D. C. Hoaglin, B. Iglewicz and J. W. Tukey. With reply by Hoaglin, Iglewicz and Tukey. Journal of the American Statistical Association, 82(398):703–704, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [HIT86].

Simpson:1987:MHD

[Sim87c] Douglas G. Simpson. Minimum Hellinger distance estimation for the analysis of count data. Journal of the American Statistical Association, 82(399):802–807, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288789.

Simpson:1989:HDT

[Sim89] Douglas G. Simpson. Hellinger deviance tests: efficiency, breakdown points, and examples. Journal of the American Statistical Association, 84(405):107–113, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289852.

Singh:1980:BRB

[Sin80] Bahadur Singh. Book review: [untitled]. Journal of the American Statistical Association, 75(369):248–249, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287434.

Sincich:1982:BRB

[Sin82] Terry L. Sincich. Book review: [untitled]. Journal of the American Statistical Association, 77(380):939–940, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287339.

Singh:1986:SAO

[Sin86] N. Singh. A simple and asymptotically optimal test for the equality of normal populations. A pragmatic approach to one-way classification. Journal of the American Statistical Association, 81 (395):703–704, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289000.

Singpurwalla:1988:IPB

[Sin88] Nozer D. Singpurwalla. An interactive PC-based procedure for reliability assessment incorporating expert opinion and survival data. Journal of the American Statistical Association, 83(401): 43–51, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288917.

Singer:1989:WFS

[Sin89a] Burton Singer. The world fertility survey: An appraisal of methodology: Comment. Journal of the American Statistical Association, 84(407):771–772, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289662.

Singer:1989:BRB

[Sin89b] Eleanor Singer. Book review: [untitled]. Journal of the American Statistical Association, 84(405):344–345, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289902.

Singh:1989:BRB

[Sin89c] M. P. Singh. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1099–1100, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290102.

Sirken:1986:BRB

[Sir86] Monroe G. Sirken. Book review: [untitled]. Journal of the American Statistical Association, 81(395):871, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289043.

Schluchter:1989:LLA

[SJ89] Mark D. Schluchter and Kirby L. Jackson. Log-linear analysis of censored survival data with partially observed covariates. Journal of the American Statistical Association, 84(405):42–52, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289845.

Stephenson:1984:BRB

[SK84] W. Robert Stephenson and Oscar Kempthorne. Book review: An Introduction to Contemporary Statistics, by Lambert H. Koopmans. Journal of the American Statistical Association, 79(385): 228–229, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288366.

Skarpness:1983:BRB

[Ska83] Bradley Skarpness. Book review: [untitled]. Journal of the American Statistical Association, 78(381):208–209, March 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287143.

Skalski:1989:BRB

[Ska89] John R. Skalski. Book review: [untitled]. Journal of the American Statistical Association, 84(407):850–851, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289702.

Skillings:1987:BRB

[Ski87] John H. Skillings. Book review: [untitled]. Journal of the American Statistical Association, 82(399):952–953, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288822.

Skovgaard:1988:BRB

[Sko88] Ib M. Skovgaard. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1218, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290168.

Schmeiser:1980:SMG

[SL80] Bruce W. Schmeiser and Ram Lal. Squeeze methods for generating gamma variates. Journal of the American Statistical Association, 75(371):679–682, September 1980. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287668.

Sathe:1981:BCC

[SL81] Y. S. Sathe and S. R. Lingras. Bounds for the confidence coefficients of outer and inner confidence intervals for quantile intervals. Journal of the American Statistical Association, 76(374):473–475, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287853.

Self:1987:APM

[SL87] Steven G. Self and Kung-Yee Liang. Asymptotic properties of maximum likelihood estimators and likelihood ratio tests under nonstandard conditions. Journal of the American Statistical Association, 82(398):605–610, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289471.

Slottje:1987:BRB

[Slo87] Daniel J. Slottje. Book review: [untitled]. Journal of the American Statistical Association, 82(398):683, June 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289490.

Slotnick:1988:BRB

[Slo88] H. B. Slotnick. Book review: [untitled]. Journal of the American Statistical Association, 83(402):578, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288903.

Sly:1988:BRB

[Sly88] David F. Sly. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1225, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290178.

Subbaiah:1981:MAS

[SM81] Perla Subbaiah and Govind S. Mudholkar. On a multivariate analog of Studentized range test. Journal of the American Statistical Association, 76(375):725–728, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287538.

Saunders:1983:MLE

[SM83a] Sam C. Saunders and Janet M. Myhre. Maximum likelihood estimation for two-parameter decreasing hazard rate distributions using censored data. Journal of the American Statistical Association, 78(383):664–673, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288136.

Schmittlein:1983:PFR

[SM83b] David C. Schmittlein and Donald G. Morrison. Prediction of future random events with the condensed negative binomial distribution. Journal of the American Statistical Association, 78(382): 449–456, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288656.

Saunders:1984:BCM

[SM84a] Sam C. Saunders and Janet M. Myhre. On the behavior of certain maximum likelihood estimators from large, randomly censored samples. Journal of the American Statistical Association, 79(386):294–301, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288263.

Smith:1984:AEM

[SM84b] David W. Smith and Leigh W. Murray. An alternative to Eisenhart’s model II and mixed model in the case of negative variance estimates. Journal of the American Statistical Association, 79 (385):145–151, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288349.

Smith:1984:CPE

[SM84c] Stanley K. Smith and Marylou Mandell. A comparison of population estimation methods: Housing unit versus component II, ratio correlation, and administrative records. Journal of the American Statistical Association, 79(386):282–289, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288261.

Smith:1980:BRB

[Smi80] Daniel Scott Smith. Book review: [untitled]. Journal of the American Statistical Association, 75(371):748, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287683.

Smith:1981:ESR

[Smi81] T. M. F. Smith. An empirical study of the ratio estimator and estimators of its variance: Comment. Journal of the American Statistical Association, 76(373):83, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287046.

Smith:1983:BMC

[Smi83a] Adrian F. M. Smith. Bayes methods for combining the results of cancer studies in humans and other species: Comment. Journal of the American Statistical Association, 78(382):310–311, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288633.

Smith:1983:EMD

[Smi83b] T. M. Fred Smith. An evaluation of model-dependent and probability-sampling inferences in sample surveys: Comment. Journal of the American Statistical Association, 78(384):801–802, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288186.

Smith:1984:BRB

[Smi84] Woolcott Smith. Book review: [untitled]. Journal of the American Statistical Association, 79(386):476–477, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288306.

Smith:1985:BRBb

[Smi85a] Philip J. Smith. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1079–1080, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288602.

Smith:1985:BRBa

[Smi85b] Wray Smith. Book review: [untitled]. Journal of the American Statistical Association, 80(390):481, June 1985. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287922.

Smith:1986:REH

[Smi86] Stanley K. Smith. A review and evaluation of the housing unit method of population estimation. Journal of the American Statistical Association, 81(394):287–296, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289216.

Smith:1988:FRM

[Smi88] Richard L. Smith. Forecasting records by maximum likelihood. Journal of the American Statistical Association, 83(402):331–338, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288847.

Smith:1989:BRB

[Smi89a] Marjo V. Smith. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1101, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290104.

Smith:1989:TPS

[Smi89b] Philip J. Smith. Is two-phase sampling really better for estimating age composition? Journal of the American Statistical Association, 84(408):916–921, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290066.

Smith:1989:TME

[Smi89c] Stanley K. Smith. Toward a methodology for estimating temporary residents. Journal of the American Statistical Association, 84(406):430–436, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289926.

Smouse:1982:BEF

[Smo82] Evan P. Smouse. Bayesian estimation of a finite population total using auxiliary information in the presence of nonresponse. Journal of the American Statistical Association, 77(377):97– 102, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287774.

Smouse:1984:NBL

[Smo84] Evan P. Smouse. A note on Bayesian least squares inference for finite population models. Journal of the American Statistical Association, 79(386):390–392, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288280.

Snee:1982:NTW

[Sne82] Ronald D. Snee. Nonadditivity in a two-way classification: Is it interaction or nonhomogeneous variance? Journal of the American Statistical Association, 77(379):515–519, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287704.

Snell:1986:BRB

[Sne86] J. Laurie Snell. Book review: [untitled]. Journal of the American Statistical Association, 81(395):857–858, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289022.

Shafer:1983:AVA

[SO83] Glenn Shafer and Ingram Olkin. Adjusting P values to account for selection over dichotomies. Journal of the American Statistical Association, 78(383):674–678, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288137.

Sokal:1981:BRB

[Sok81] Robert R. Sokal. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1015–1016, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287618.

Sokal:1982:BRB

[Sok82] Robert R. Sokal. Book review: [untitled]. Journal of the American Statistical Association, 77(377):221–222, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287808.

Sokal:1983:BRB

[Sok83] Robert R. Sokal. Book review: [untitled]. Journal of the American Statistical Association, 78(382):508, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288692.

Sokal:1984:BRBa

[Sok84a] Robert R. Sokal. Book review: [untitled]. Journal of the American Statistical Association, 79(387):741, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288450.

Sokal:1984:BRBb

[Sok84b] Robert R. Sokal. Book review: [untitled]. Journal of the American Statistical Association, 79(387):741, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288451.

Solo:1984:ODA

[Sol84] Victor Solo. The order of differencing in ARIMA models. Journal of the American Statistical Association, 79(388):916–921, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288724.

Solo:1985:BRB

[Sol85] Victor Solo. Book review: [untitled]. Journal of the American Statistical Association, 80(391):784, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288517.

Solon:1987:BRB

[Sol87] Gary Solon. Book review: [untitled]. Journal of the American Statistical Association, 82(397):355–356, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289194.

Solo:1988:BRB

[Sol88] Victor Solo. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1222–1223, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290174.

Soms:1980:RBT

[Som80] Andrew P. Soms. Rational bounds for the t-tail area. Journal of the American Statistical Association, 75(370):438–440, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287473.

Soric:1989:SDE

[Sor89] Branko Soric. Statistical “discoveries” and effect-size estimation. Journal of the American Statistical Association, 84(406):608–610, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289950.

Spar:1988:BRB

[Spa88] Edward J. Spar. Book review: [untitled]. Journal of the American Statistical Association, 83(401):283–284, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288978.

Spencer:1985:ODQ

[Spe85a] Bruce D. Spencer. Optimal data quality. Journal of the American Statistical Association, 80(391):564–573, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288471.

Spencer:1985:SAE

[Spe85b] Bruce D. Spencer. Statistical aspects of equitable apportionment. Journal of the American Statistical Association, 80(392):815–822, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288538.

Speckman:1986:BRB

[Spe86] Paul Speckman. Book review: [untitled]. Journal of the American Statistical Association, 81(394):572, June 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289260.

Spitzer:1982:FEA

[Spi82] John J. Spitzer. A fast and efficient algorithm for the estimation of parameters in models with the Box-and-Cox transformation. Journal of the American Statistical Association, 77(380): 760–766, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287303.

Spitzer:1983:BRB

[Spi83] Frank Spitzer. Book review: [untitled]. Journal of the American Statistical Association, 78(382):500, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288674.

Spiegelman:1986:BRB

[Spi86] Clifford Spiegelman. Book review: [untitled]. Journal of the American Statistical Association, 81(395):871–872, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289044.

Spliid:1983:FEM

[Spl83] Henrik Spliid. A fast estimation method for the vector autoregressive moving average model with exogenous variables. Journal of the American Statistical Association, 78(384):843–849, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288194.

Sposito:1982:URE

[Spo82] V. A. Sposito. On unbiased L_{p} regression estimators. Journal of the American Statistical Association, 77(379):652–653, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287729.

Sprott:1983:EPC

[Spr83] David A. Sprott. Estimating the parameters of a convolution by maximum likelihood. Journal of the American Statistical Association, 78(382):457–460, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288657.

Spulber:1987:BRB

[Spu87a] Daniel F. Spulber. Book review: [untitled]. Journal of the American Statistical Association, 82(397):357, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289196.

Spurrier:1987:BRB

[Spu87b] John D. Spurrier. Book review: [untitled]. Journal of the American Statistical Association, 82(399):947, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288812.

Spurrier:1988:GST

[Spu88] John D. Spurrier. Generalizations of Steel’s treatments-versus-control multivariate sign test. Journal of the American Statistical Association, 83(402):471–476, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288864.

Swan:1986:ACW

[SR86] Shanna H. Swan and James M. Robins. An analysis of contaminated well water and health effects in Woburn, Massachusetts: Comment. Journal of the American Statistical Association, 81 (395):604–609, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288986.

Smith:1981:AMD

[SRMS81] Paul J. Smith, Donald S. Rae, Ronald W. Manderscheid, and Sam Silbergeld. Approximating the moments and distribution of the likelihood ratio statistic for multinomial goodness of fit. Journal of the American Statistical Association, 76(375):737–740, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287541.

Santner:1980:SSC

[SS80a] Thomas J. Santner and Mark K. Snell. Small-sample confidence intervals for p_{1} − p_{2} and p_{1}/p_{2} in 2 × 2 contingency tables. Journal of the American Statistical Association, 75(370):386–394, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287464.

Schmeiser:1980:ARM

[SS80b] Bruce W. Schmeiser and Mohamed A. Shalaby. Acceptance/rejection methods for beta variate generation. Journal of the American Statistical Association, 75(371):673–678, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287667.

Sethuraman:1982:THD

[SS82a] J. Sethuraman and Nozer D. Singpurwalla. Testing of hypotheses for distributions in accelerated life tests. Journal of the American Statistical Association, 77(377):204–208, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287790.

Smith:1982:BOE

[SS82b] Philip J. Smith and J. Sedransk. Bayesian optimization of the estimation of the age composition of a fish population. Journal of the American Statistical Association, 77(380):707–713, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287295.

Stem:1984:TIM

[SS84] Donald E. Stem, Jr. and R. Kirk Steinhorst. Telephone interview and mail questionnaire applications of the randomized response model. Journal of the American Statistical Association, 79(387): 555–564, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288401.

Sampson:1985:MSI

[SS85] Paul D. Sampson and Andrew F. Siegel. The measure of “size” independent of “shape” for multivariate lognormal populations. Journal of the American Statistical Association, 80(392):910–914, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288552.

Simon:1986:DPM

[SS86] Gary A. Simon and Jeffrey S. Simonoff. Diagnostic plots for missing data in least squares regression. Journal of the American Statistical Association, 81(394):501–509, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289241.

Scholz:1987:SAD

[SS87a] F. W. Scholz and M. A. Stephens. K-sample Anderson–Darling tests. Journal of the American Statistical Association, 82(399): 918–924, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288805.

Souvaine:1987:TSE

[SS87b] Diane L. Souvaine and J. Michael Steele. Time- and space-efficient algorithms for least median of squares regression. Journal of the American Statistical Association, 82(399):794–801, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288788.

Samanta:1988:SWT

[SS88a] M. Samanta and C. J. Schwarz. The Shapiro–Wilk test for exponentiality based on censored data. Journal of the American Statistical Association, 83(402):528–531, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288873.

Sinclair:1988:ADF

[SS88b] C. D. Sinclair and B. D. Spurr. Approximations to the distribution function of the Anderson–Darling test statistic. Journal of the American Statistical Association, 83(404):1190–1191, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2290155.

Stokes:1988:CRS

[SS88c] S. Lynne Stokes and Thomas W. Sager. Characterization of a ranked-set sample with application to estimating distribution functions. Journal of the American Statistical Association, 83 (402):374–381, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288852.

Shapiro:1985:RSSa

[SSdM85a] M. B. Shapiro, S. J. Schein, and F. M. de Monasterio. Regularity and structure of the spatial pattern of blue cones of macaque retina. Journal of the American Statistical Association, 80(392): 803–812, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288535.

Shapiro:1985:RSSb

[SSdM85b] M. B. Shapiro, S. J. Schein, and F. M. de Monasterio. Regularity and structure of the spatial pattern of blue cones of macaque retina: Rejoinder. Journal of the American Statistical Association, 80(392):814, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288537.

Stoffer:1988:WFA

[SSR^{+}88] David S. Stoffer, Mark S. Scher, Gale A. Richardson, Nancy L. Day, and Patricia A. Coble. A Walsh–Fourier analysis of the effects of moderate maternal alcohol consumption on neonatal sleepstate cycling. Journal of the American Statistical Association, 83 (404):954–963, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290119.

Sedransk:1987:ER

[SST87] J. Sedransk, Paul Switzer, and Judith M. Tanur. Editor’s report for 1986. Journal of the American Statistical Association, 82(398): 369–370, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289438.

Sedransk:1988:ER

[SST88] J. Sedransk, Paul Switzer, and Judith M. Tanur. Editors’ report for 1987. Journal of the American Statistical Association, 83(402): 287–288, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288842.

Sedransk:1989:ER

[SST89] J. Sedransk, Paul Switzer, and Judith M. Tanur. Editors’ report for 1988. Journal of the American Statistical Association, 84 (406):353, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289917.

Salemi:1982:ENL

[ST82a] Michael K. Salemi and George E. Tauchen. Estimation of nonlinear learning models. Journal of the American Statistical Association, 77(380):725–731, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287297.

Shaked:1982:EPM

[ST82b] Moshe Shaked and Lanh Tat Tran. Estimating parameters from mixed samples. Journal of the American Statistical Association, 77(377):196–203, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287789.

Scott:1987:BUC

[ST87] David W. Scott and George R. Terrell. Biased and unbiased crossvalidation in density estimation. Journal of the American Statistical Association, 82(400):1131–1146, ???? 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stafford:1985:BRB

[Sta85] Frank Stafford. Book review: [untitled]. Journal of the American Statistical Association, 80(389):256, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288109.

Stasny:1986:EGF

[Sta86] Elizabeth A. Stasny. Estimating gross flows using panel data with nonresponse: An example from the Canadian Labour Force Survey. Journal of the American Statistical Association, 81(393): 42–47, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287966.

Stallings:1989:BRB

[Sta89a] William M. Stallings. Book review: [untitled]. Journal of the American Statistical Association, 84(405):348, March 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289907.

Staniswalis:1989:KER

[Sta89b] Joan G. Staniswalis. The kernel estimate of a regression function in likelihood-based models. Journal of the American Statistical Association, 84(405):276–283, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289874.

Staniswalis:1989:LBS

[Sta89c] Joan G. Staniswalis. Local bandwidth selection for kernel estimates. Journal of the American Statistical Association, 84(405): 284–288, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289875.

Stephenson:1981:GCO

[Ste81] W. Robert Stephenson. A general class of one-sample nonparametric test statistics based on subsamples. Journal of the American Statistical Association, 76(376):960–966, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287596.

Stephenson:1982:BRB

[Ste82] W. Robert Stephenson. Book review: [untitled]. Journal of the American Statistical Association, 77(377):215, March 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287798.

Stefanski:1987:BRB

[Ste87a] Leonard A. Stefanski. Book review: [untitled]. Journal of the American Statistical Association, 82(399):949–950, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288817.

Stein:1987:MNQ

[Ste87b] Michael L. Stein. Minimum norm quadratic estimation of spatial variograms. Journal of the American Statistical Association, 82 (399):765–772, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288785.

Steinhorst:1988:BRB

[Ste88] R. Kirk Steinhorst. Book review: [untitled]. Journal of the American Statistical Association, 83(401):271, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288958.

Stephenson:1989:BRB

[Ste89] W. Robert Stephenson. Book review: [untitled]. Journal of the American Statistical Association, 84(407):836–837, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289677.

Stiner:1982:BRB

[Sti82] Frederic M. Stiner, Jr. Book review: [untitled]. Journal of the American Statistical Association, 77(378):493–494, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287279.

Stidham:1983:BRB

[Sti83a] Shaler Stidham, Jr. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1004, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288244.

Stigler:1983:PEB

[Sti83b] Stephen M. Stigler. Parametric empirical Bayes inference: Theory and applications: Comment. Journal of the American Statistical Association, 78(381):62–63, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287104.

Stiner:1983:BRB

[Sti83c] Frederic M. Stiner, Jr. Book review: [untitled]. Journal of the American Statistical Association, 78(383):742, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288166.

Stigler:1984:BRB

[Sti84] Stephen M. Stigler. Book review: [untitled]. Journal of the American Statistical Association, 79(388):947–948, December 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288736.

Stine:1985:BPI

[Sti85] Robert A. Stine. Bootstrap prediction intervals for regression. Journal of the American Statistical Association, 80(392):1026– 1031, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288570.

Stigler:1986:JCP

[Sti86] Stephen M. Stigler. John Craig and the probability of history: from the death of Christ to the birth of Laplace. Journal of the American Statistical Association, 81(396):879–887, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289056.

Stine:1987:EPA

[Sti87] Robert A. Stine. Estimating properties of autoregressive forecasts. Journal of the American Statistical Association, 82(400):

1072–1078, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stigler:1988:LBO

[Sti88] Stephen M. Stigler. A look backward on the occasion of the centenary of JASA. Journal of the American Statistical Association, 83(403):583–587, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Stokes:1980:ICC

[Sto80a] S. Lynne Stokes. Inferences on the correlation coefficient in bivariate normal populations from ranked set samples. Journal of the American Statistical Association, 75(372):989–995, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287193.

Stolwijk:1980:ISM

[Sto80b] Jan A. J. Stolwijk. Information sources for modeling the national economy: Comment. Journal of the American Statistical Association, 75(371):569–572, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287646.

Stoker:1982:UCS

[Sto82] Thomas M. Stoker. The use of cross-section data to characterize macrofunctions. Journal of the American Statistical Association, 77(378):369–380, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287254.

Stoline:1983:HMS

[Sto83a] Michael R. Stoline. The hunter method of simultaneous inference and its recommended use for applications having large known correlation structures. Journal of the American Statistical Association, 78(382):366–370, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288641.

Stoto:1983:APP

[Sto83b] Michael A. Stoto. The accuracy of population projections. Journal of the American Statistical Association, 78(381):13–20, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287094.

Stoker:1985:ASC

[Sto85a] Thomas M. Stoker. Aggregation, structural change, and crosssection estimation. Journal of the American Statistical Association, 80(391):720–729, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288491.

Stoto:1985:BRBb

[Sto85b] Michael Stoto. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1076, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288594.

Stoto:1985:BRBa

[Sto85c] Michael A. Stoto. Book review: [untitled]. Journal of the American Statistical Association, 80(389):245–246, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288095.

Stoffer:1986:EIS

[Sto86] David S. Stoffer. Estimation and identification of space-time ARMAX models in the presence of missing data. Journal of the American Statistical Association, 81(395):762–772, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289008.

Stock:1988:ECT

[Sto88] James H. Stock. Estimating continuous-time processes subject to time deformation: An application to postwar U.S. GNP. Journal of the American Statistical Association, 83(401):77–85, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288921.

Stock:1989:NPA

[Sto89] James H. Stock. Nonparametric policy analysis. Journal of the American Statistical Association, 84(406):567–575, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289944.

Stuart:1980:CTD

[Stu80] Michael Stuart. Components of χ^{2} for testing distributional assumptions against certain restricted alternatives. Journal of the American Statistical Association, 75(371):625–633, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287659.

Studden:1982:SRT

[Stu82] W. J. Studden. Some robust-type D-optimal designs in polynomial regression. Journal of the American Statistical Association, 77(380):916–921, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287327.

Stuetzle:1987:PW

[Stu87] Werner Stuetzle. Plot windows. Journal of the American Statistical Association, 82(398):466–475, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289448.

Stukel:1988:GLM

[Stu88] Thérèse A. Stukel. Generalized logistic models. Journal of the American Statistical Association, 83(402):426–431, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288858.

Sudderth:1983:BRB

[Sud83] William D. Sudderth. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1005, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288245.

Sudman:1987:BRBa

[Sud87a] Seymour Sudman. Book review: [untitled]. Journal of the American Statistical Association, 82(397):355, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289193.

Sudman:1987:BRBb

[Sud87b] Seymour Sudman. Book review: [untitled]. Journal of the American Statistical Association, 82(398):686–687, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289494.

Sudman:1989:BRB

[Sud89] Seymour Sudman. Book review: [untitled]. Journal of the American Statistical Association, 84(406):632–633, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289985.

Sugihara:1982:DCM

[Sug82] George Sugihara. Diversity as a concept and its measurement: Comment. Journal of the American Statistical Association, 77 (379):564–565, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287711.

Sukhatme:1987:BRB

[Suk87] Shashikala Sukhatme. Book review: [untitled]. Journal of the American Statistical Association, 82(399):953, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288823.

Susarla:1986:BRB

[Sus86] V. Susarla. Book review: [untitled]. Journal of the American Statistical Association, 81(393):264, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288021.

Sutherland:1985:BRB

[Sut85] Michael Sutherland. Book review: [untitled]. Journal of the American Statistical Association, 80(390):488–489, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287934.

Suzuki:1985:NEL

[Suz85] Kazuyuki Suzuki. Nonparametric estimation of lifetime distributions from a record of failures and follow-ups. Journal of the American Statistical Association, 80(389):68–72, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288041.

Shapiro:1980:BSE

[SW80a] C. P. Shapiro and Robert L. Wardrop. Bayesian sequential estimation for one-parameter exponential families. Journal of the American Statistical Association, 75(372):984–988, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287192.

Shapiro:1980:SBS

[SW80b] C. P. Shapiro and Robert L. Wardrop. Simultaneous Bayesian sequential estimation. Journal of the American Statistical Association, 75(370):359–365, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287459.

Spurrier:1980:TPE

[SW80c] John D. Spurrier and L. J. Wei. A test of the parameter of the exponential distribution in the Type I censoring case. Journal of the American Statistical Association, 75(370):405–409, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287467.

Scott:1981:ADR

[SW81] Alastair Scott and Chien-Fu Wu. On the asymptotic distribution of ratio and regression estimators. Journal of the American Statistical Association, 76(373):98–102, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287051.

Slud:1982:TSR

[SW82] Eric Slud and L. J. Wei. Two-sample repeated significance tests based on the modified Wilcoxon statistic. Journal of the American Statistical Association, 77(380):862–868, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287319.

Singpurwalla:1983:KEF

[SW83] Nozer D. Singpurwalla and Man Yuen Wong. Kernel estimators of the failure-rate function and density estimation: an analogy. Journal of the American Statistical Association, 78(382):478–481, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288663.

Srivastava:1986:LRT

[SW86] M. S. Srivastava and K. J. Worsley. Likelihood ratio tests for a change in the multivariate normal mean. Journal of the American Statistical Association, 81(393):199–204, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287990.

Silverman:1987:NEB

[SW87] B. W. Silverman and J. T. Wood. The nonparametric estimation of branching curves. Journal of the American Statistical Association, 82(398):551–558, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289464.

Stock:1988:TCT

[SW88] James H. Stock and Mark W. Watson. Testing for common trends. Journal of the American Statistical Association, 83(404):1097– 1107, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290142.

Sampson:1989:EMD

[SW89a] Allan R. Sampson and Lyn R. Whitaker. Estimation of multivariate distributions under stochastic ordering. Journal of the American Statistical Association, 84(406):541–548, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289940.

Skovlund:1989:ETD

[SW89b] Eva Skovlund and Lars Walloe. Estimation of treatment difference following a sequential clinical trial. Journal of the American Statistical Association, 84(407):823–828, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289673.

Swartzman:1981:BRB

[Swa81] Gordon Swartzman. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1013, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287612.

Swartzman:1984:BRB

[Swa84] Gordon Swartzman. Book review: [untitled]. Journal of the American Statistical Association, 79(386):477, June 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288307.

Sweeting:1984:AIL

[Swe84] Trevor J. Sweeting. Approximate inference in location-scale regression models. Journal of the American Statistical Association, 79(388):847–852, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288714.

Switzer:1983:BRB

[Swi83] Paul Switzer. Book review: [untitled]. Journal of the American Statistical Association, 78(382):510, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288695.

Switzer:1985:BRB

[Swi85] Paul Switzer. Book review: [untitled]. Journal of the American Statistical Association, 80(389):243, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288091.

Sirvanci:1984:EWP

[SY84] Mete Sirvanci and Grace Yang. Estimation of the Weibull parameters under Type I censoring. Journal of the American Statistical Association, 79(385):183–187, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288354.

Sylwester:1981:BRB

[Syl81] David L. Sylwester. Book review: [untitled]. Journal of the American Statistical Association, 76(376):1013, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287611.

Steele:1980:OSS

[SZ80] J. Michael Steele and James Zidek. Optimal strategies for second guessers. Journal of the American Statistical Association, 75(371): 596–601, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287655.

Stein:1987:PES

[SZJ87] Gillian Z. Stein, Walter Zucchini, and June M. Juritz. Parameter estimation for the Sichel distribution and its multivariate extension. Journal of the American Statistical Association, 82(399):

938–944, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288808.

Thompson:1989:ERT

[TA89] G. L. Thompson and L. P. Ammann. Efficacies of rank-transform statistics in two-way models with no interaction. Journal of the American Statistical Association, 84(405):325–330, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289881.

Tadikamalla:1980:RSE

[Tad80] Pandu R. Tadikamalla. Random sampling from the exponential power distribution. Journal of the American Statistical Association, 75(371):683–686, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287669.

Taeuber:1981:BRB

[Tae81] Conrad Taeuber. Book review: [untitled]. Journal of the American Statistical Association, 76(375):743–744, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287547.

Tai:1989:ABD

[Tai89] J. J. Tai. Application of Bayesian decision procedure to the inference of genetic linkage. Journal of the American Statistical Association, 84(407):669–673, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289647.

Takada:1981:ISE

[Tak81] Yoshikazu Takada. Invariant sequential estimation of the exponential mean life from the ordered observations. Journal of the American Statistical Association, 76(374):470–472, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287852.

Takemura:1983:TAA

[Tak83] Akimichi Takemura. Tensor analysis of ANOVA decomposition. Journal of the American Statistical Association, 78(384):894–900, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288201.

Tamhane:1981:RRT

[Tam81] Ajit C. Tamhane. Randomized response techniques for multiple sensitive attributes. Journal of the American Statistical Association, 76(376):916–923, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287588.

Tamhane:1985:SSP

[Tam85] Ajit C. Tamhane. Some sequential procedures for selecting the better Bernoulli treatment by using a matched samples design. Journal of the American Statistical Association, 80(390):455–460, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287914.

Tam:1988:SRR

[Tam88a] S. M. Tam. Some results on robust estimation in finite population sampling. Journal of the American Statistical Association, 83 (401):242–248, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288946.

Tamaki:1988:BAB

[Tam88b] Mitsushi Tamaki. A Bayesian approach to the best-choice problem. Journal of the American Statistical Association, 83(404): 1129–1133, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290145.

Tanaka:1983:ETF

[Tan83] Katsuto Tanaka. Estimation for transients in the frequency domain. Journal of the American Statistical Association, 78(383): 718–724, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288142.

Tarter:1980:DEB

[Tar80] Michael E. Tarter. Density estimation and bump-hunting by the penalized likelihood method exemplified by scattering and meteorite data: Comment. Journal of the American Statistical Association, 75(369):63–66, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287381.

Tarter:1986:MIS

[Tar86] Michael E. Tarter. Mean integrated squared error sampling. Journal of the American Statistical Association, 81(393):234– 242, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287996.

Tardif:1987:EOR

[Tar87] Serge Tardif. Efficiency and optimality results for tests based on weighted rankings. Journal of the American Statistical Association, 82(398):637–644, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289476.

Tatsuoka:1988:BRB

[Tat88] Maurice M. Tatsuoka. Book review: [untitled]. Journal of the American Statistical Association, 83(402):566–567, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288886.

Tavare:1984:BRB

[Tav84] Simon Tavare. Book review: [untitled]. Journal of the American Statistical Association, 79(387):743–744, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288456.

Tavare:1985:BRB

[Tav85] Simon Tavare. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1071–1072, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288587.

Taylor:1980:BRBa

[Tay80a] Amy K. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 75(370):479, June 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287496.

Taylor:1980:BRBb

[Tay80b] Ronald N. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1043, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287222.

Taylor:1984:BRB

[Tay84] William F. Taylor. Book review: Introduction to Data Analysis and Statistical Inference, by Carl N. Morris; John E. Rolph. Journal of the American Statistical Association, 79(385):242, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288386.

Taylor:1985:MLS

[Tay85] Jeremy M. G. Taylor. Measures of location of skew distributions obtained through Box–Cox transformations. Journal of the American Statistical Association, 80(390):427–432, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287909.

Taylor:1986:RMA

[Tay86] Jeremy M. G. Taylor. The retransformed mean after a fitted power transformation. Journal of the American Statistical Association, 81(393):114–118, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287976.

Taylor:1987:BRBa

[Tay87a] Lester D. Taylor. Book review: [untitled]. Journal of the American Statistical Association, 82(398):683–684, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289491.

Taylor:1987:BRBb

[Tay87b] Robert Lee Taylor. Book review: [untitled]. Journal of the American Statistical Association, 82(398):694, June 1987. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289503.

Taylor:1988:CGL

[Tay88] Jeremy M. G. Taylor. The cost of generalizing logistic regression. Journal of the American Statistical Association, 83(404):

1078–1083, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2290139.

Tiao:1981:MMT

[TB81] G. C. Tiao and G. E. P. Box. Modeling multiple time series with applications. Journal of the American Statistical Association, 76 (376):802–816, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287575.

Tryfos:1985:FR

[TB85] Peter Tryfos and Russell Blackmore. Forecasting records. Journal of the American Statistical Association, 80(389):46–50, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288039.

Tamura:1986:MHD

[TB86] Roy N. Tamura and Dennis D. Boos. Minimum Hellinger distance estimation for multivariate location and covariance. Journal of the American Statistical Association, 81(393):223–229, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287994.

Thompson:1989:MTS

[TBLM89] Patrick A. Thompson, William R. Bell, John F. Long, and Robert B. Miller. Multivariate time series projections of parameterized age-specific fertility rates. Journal of the American Statistical Association, 84(407):689–699, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289650.

Tanur:1989:BRB

[TCD^{+}89] Judith M. Tanur, George Casella, Richard Dykstra, Mark P. Finster, Donald P. Gaver, Jr., Joel Greenhouse, Gudmund R. Iversen, guillermina Jasso, Jan Kmenta, S. James Press, Seymour Sudman, Luke Tierney, Jessica Utts, Katherine K. Wallman, Stanley Wasserman, and Mark Watson. Book review: [untitled]. Journal of the American Statistical Association, 84(407):830–834, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289675.

Tella:1986:BRB

[Tel86] Dorothy M. Tella. Book review: [untitled]. Journal of the American Statistical Association, 81(394):579, June 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289274.

Tarone:1980:RCT

[TG80] Robert E. Tarone and John J. Gart. On the robustness of combined tests for trends in proportions. Journal of the American Statistical Association, 75(369):110–116, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287397.

Tang:1989:DGS

[TGG89] Dei-In Tang, Clare Gnecco, and Nancy L. Geller. Design of group sequential clinical trials with multiple endpoints. Journal of the American Statistical Association, 84(407):776–779, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289665.

Tosch:1980:BFM

[TH80] Thomas J. Tosch and Paul T. Holmes. A bivariate failure model. Journal of the American Statistical Association, 75(370):415–417, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287469.

Tu:1982:DAB

[TH82] Ching Tsao Tu and Chien Pai Han. Discriminant analysis based on binary and continuous variables. Journal of the American Statistical Association, 77(378):447–454, June 1982. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287266.

Tibshirani:1987:LLE

[TH87] Robert Tibshirani and Trevor Hastie. Local likelihood estimation. Journal of the American Statistical Association, 82(398):559–567, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289465.

Thaler:1984:NEH

[Tha84] H. T. Thaler. Nonparametric estimation of the hazard ratio. Journal of the American Statistical Association, 79(386):290–293, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288262.

Thaler:1985:SMP

[Tha85] H. T. Thaler. A statistical model for positron emission tomography: Comment. Journal of the American Statistical Association, 80(389):33, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288036.

Thisted:1980:CSR

[Thi80] Ronald A. Thisted. A critique of some ridge regression methods: Comment. Journal of the American Statistical Association, 75 (369):81–86, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287387.

Thisted:1989:BRB

[Thi89] Ronald A. Thisted. Book review: [untitled]. Journal of the American Statistical Association, 84(406):614, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289954.

Thompson:1981:ESR

[Tho81] M. E. Thompson. An empirical study of the ratio estimator and estimators of its variance: Comment. Journal of the American Statistical Association, 76(373):84–85, March 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287047.

Thomas:1982:BRB

[Tho82a] George Thomas. Book review: [untitled]. Journal of the American Statistical Association, 77(379):696, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287759.

Thompson:1982:BRBa

[Tho82b] M. E. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 77(377):212–213, March 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287794.

Thompson:1982:BRBb

[Tho82c] Velma Montoya Thompson. Book review: [untitled]. Journal of the American Statistical Association, 77(377):219–220, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287804.

Thompson:1983:BRBb

[Tho83a] W. A. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1001, December 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288237.

Thompson:1983:BRBa

[Tho83b] W. A. Thompson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 78(382):507–508, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288691.

Thorp:1983:BRB

[Tho83c] Edward O. Thorp. Book review: [untitled]. Journal of the American Statistical Association, 78(384):985, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288215.

Thompson:1984:BRBc

[Tho84a] M. E. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 79(386):480, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288312.

Thompson:1984:BRBe

[Tho84b] M. E. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 79(387):744–745, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288458.

Thompson:1984:BRBa

[Tho84c] W. A. Thompson, Jr. Book review: Statistics for Research, by Shirley Dowdy; Stanley Wearden. Journal of the American Statistical Association, 79(385):242, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288387.

Thompson:1984:BRBb

[Tho84d] W. A. Thompson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 79(386):478, June 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288309.

Thompson:1984:BRBd

[Tho84e] W. A. Thompson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 79(387):736, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288441.

Thompson:1985:BRBe

[Tho85a] Steven K. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 80(391):789, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288527.

Thompson:1985:BRBa

[Tho85b] W. A. Thompson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 80(390):491, June 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287941.

Thompson:1985:BRBb

[Tho85c] W. A. Thompson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 80(390):492, June 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287943.

Thompson:1985:BRBc

[Tho85d] W. A. Thompson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 80(390):492, June 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287944.

Thompson:1985:BRBd

[Tho85e] W. A. Thompson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 80(391):788, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288525.

Thompson:1985:BRBf

[Tho85f] W. A. Thompson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1066–1067, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288581.

Thornberry:1985:BRB

[Tho85g] Owen T. Thornberry, Jr. Book review: [untitled]. Journal of the American Statistical Association, 80(390):483–484, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287927.

Thompson:1986:BRBb

[Tho86a] M. E. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 81(396):1122–1123, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289103.

Thompson:1986:BRBa

[Tho86b] W. A. Thompson, Jr. Book review: [untitled]. Journal of the American Statistical Association, 81(395):861–862, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289030.

Thompson:1987:BRB

[Tho87] James R. Thompson. Book review: [untitled]. Journal of the American Statistical Association, 82(397):344, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289176.

Thomas:1989:BRB

[Tho89] William Thomas. Book review: [untitled]. Journal of the American Statistical Association, 84(406):616–617, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289957.

Thursby:1982:BRB

[Thu82] Jerry G. Thursby. Book review: [untitled]. Journal of the American Statistical Association, 77(380):943, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287345.

Thursby:1985:RAS

[Thu85] Jerry G. Thursby. The relationship among the specification tests of Hausman, Ramsey, and Chow. Journal of the American Statistical Association, 80(392):926–928, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288555.

Thursby:1987:BRB

[Thu87] Jerry G. Thursby. Book review: [untitled]. Journal of the American Statistical Association, 82(399):957, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288830.

Thursby:1988:BRB

[Thu88] Jerry G. Thursby. Book review: [untitled]. Journal of the American Statistical Association, 83(402):575, June 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288899.

Tibshirani:1988:ETR

[Tib88] Robert Tibshirani. Estimating transformations for regression via additivity and variance stabilization. Journal of the American Statistical Association, 83(402):394–405, June 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288855.

Tietjen:1981:BRBa

[Tie81a] Gary Tietjen. Book review: [untitled]. Journal of the American Statistical Association, 76(373):210, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287078.

Tietjen:1981:BRBb

[Tie81b] Gary L. Tietjen. Book review: [untitled]. Journal of the American Statistical Association, 76(374):497, June 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287871.

Tierney:1982:BRB

[Tie82] Luke Tierney. Book review: [untitled]. Journal of the American Statistical Association, 77(379):690–691, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287750.

Tierney:1984:BRB

[Tie84] Luke Tierney. Book review: [untitled]. Journal of the American Statistical Association, 79(388):955, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288749.

Tierney:1985:BRB

[Tie85] Luke Tierney. Book review: [untitled]. Journal of the American Statistical Association, 80(389):246, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288096.

Titterington:1985:BRB

[Tit85] D. M. Titterington. Book review: [untitled]. Journal of the American Statistical Association, 80(391):768–771, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288499.

Tsui:1988:NAT

[TJW88] Kwok-Leung Tsui, Nicholas P. Jewell, and C.-F. J. Wu. A nonparametric approach to the truncated regression problem. Journal of the American Statistical Association, 83(403):785–792, ???? 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tierney:1986:AAP

[TK86] Luke Tierney and Joseph B. Kadane. Accurate approximations for posterior moments and marginal densities. Journal of the American Statistical Association, 81(393):82–86, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287970.

Tierney:1989:FEL

[TKK89] Luke Tierney, Robert E. Kass, and Joseph B. Kadane. Fully exponential Laplace approximations to expectations and variances of nonpositive functions. Journal of the American Statistical Association, 84(407):710–716, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289652.

Thall:1988:ARE

[TL88] Peter F. Thall and John M. Lachin. Analysis of recurrent events: Nonparametric methods for random-interval count data. Journal of the American Statistical Association, 83(402):339–347, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288848.

Tolley:1982:BRB

[Tol82] H. Dennis Tolley. Book review: [untitled]. Journal of the American Statistical Association, 77(379):689, September 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287747.

Tomberlin:1988:PAF

[Tom88] Thomas J. Tomberlin. Predicting accident frequencies for drivers classified by two factors. Journal of the American Statistical Association, 83(402):309–321, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2288845.

Tortora:1982:BRB

[Tor82] Robert D. Tortora. Book review: [untitled]. Journal of the American Statistical Association, 77(378):496–497, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287282.

Torrez:1983:BRB

[Tor83] William C. Torrez. Book review: [untitled]. Journal of the American Statistical Association, 78(383):732–733, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288148.

Thomas:1987:SSC

[TR87] D. Roland Thomas and J. N. K. Rao. Small-sample comparisons of level and power for simple goodness-of-fit statistics under cluster sampling. Journal of the American Statistical Association, 82 (398):630–636, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289475.

Trader:1983:BRB

[Tra83] Ramona L. Trader. Book review: [untitled]. Journal of the American Statistical Association, 78(384):999–1000, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288235.

Tritchler:1984:AEL

[Tri84a] David Tritchler. An algorithm for exact logistic regression. Journal of the American Statistical Association, 79(387):709–711, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288420.

Tritchler:1984:IPT

[Tri84b] David Tritchler. On inverting permutation tests. Journal of the American Statistical Association, 79(385):200–207, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288357.

Terrell:1985:OND

[TS85] George R. Terrell and David W. Scott. Oversmoothed nonparametric density estimates. Journal of the American Statistical Association, 80(389):209–214, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288074.

Tsay:1984:RMT

[Tsa84] Ruey S. Tsay. Regression models with time series errors. Journal of the American Statistical Association, 79(385):118–124, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288345.

Tsay:1986:TSM

[Tsa86] Ruey S. Tsay. Time series model specification in the presence of outliers. Journal of the American Statistical Association, 81 (393):132–141, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287980.

Tsay:1987:CHT

[Tsa87] Ruey S. Tsay. Conditional heteroscedastic time series models. Journal of the American Statistical Association, 82(398):590–604, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289470.

Tsai:1989:BRB

[Tsa89a] Chih-Ling Tsai. Book review: [untitled]. Journal of the American Statistical Association, 84(405):336–337, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289891.

Tsay:1989:TMT

[Tsa89b] Ruey S. Tsay. Testing and modeling threshold autoregressive processes. Journal of the American Statistical Association, 84(405): 231–240, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289868.

Tsiatis:1982:RST

[Tsi82] Anastasios A. Tsiatis. Repeated significance testing for a general class of statistics used in censored survival analysis. Journal of the American Statistical Association, 77(380):855–861, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287318.

Tsui:1984:RCZ

[Tsu84] Kam-Wah Tsui. Robustness of Clevenson-Zidek-type estimators. Journal of the American Statistical Association, 79(385): 152–157, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288350.

Tsutakawa:1985:BRB

[Tsu85] Robert K. Tsutakawa. Book review: [untitled]. Journal of the American Statistical Association, 80(391):778, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288509.

Tsui:1986:FDR

[Tsu86] Kam-Wah Tsui. Further developments on the robustness of Clevenson-Zidek-type means estimators. Journal of the American Statistical Association, 81(393):176–180, March 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287986.

Tsutakawa:1988:MMA

[Tsu88] Robert K. Tsutakawa. Mixed model for analyzing geographic variability in mortality rates. Journal of the American Statistical Association, 83(401):37–42, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288916.

Tsay:1984:CEA

[TT84] Ruey S. Tsay and George C. Tiao. Consistent estimates of autoregressive parameters and extended sample autocorrelation function for stationary and nonstationary ARMA models. Journal of the American Statistical Association, 79(385):84–96, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288340.

Thomas:1986:CBP

[TT86] Dana L. Thomas and David R. Thomas. Confidence bands for percentiles in the linear regression model. Journal of the American Statistical Association, 81(395):705–708, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289001.

Tufte:1985:BRB

[Tuf85] Edward Tufte. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1074, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288591.

Tukey:1985:CEP

[Tuk85a] John W. Tukey. Comments on Estimating the population in a census year: 1980 and beyond. Journal of the American Statistical Association, 80(389):127–128, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tukey:1985:EPC

[Tuk85b] John W. Tukey. Estimating the population in a census year: 1980 and beyond: Comment. Journal of the American Statistical Association, 80(389):127–128, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288061.

Tukey:1985:RLR

[Tuk85c] John W. Tukey. The resistant line and related regression methods: Comment: The variance of slopes of lines fitted by groups: An analysis of the Johnston and Velleman Monte Carlo results. Journal of the American Statistical Association, 80(392):1055–1059, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288573.

Tuma:1982:EHH

[Tum82] Nancy Brandon Tuma. Estimating historical heights: Comment. Journal of the American Statistical Association, 77(378):298–301, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287236.

Turner:1983:MCT

[Tur83] D. W. Turner. A method for comparing two hierarchical clusterings: Comment. Journal of the American Statistical Association, 78(383):583, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288122.

Turgeon:1987:BRB

[Tur87] Livia R. Turgeon. Book review: [untitled]. Journal of the American Statistical Association, 82(397):358–359, March 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289198.

Turnbull:1989:BRB

[Tur89a] Craig D. Turnbull. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1093, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290090.

Turner:1989:WFS

[Tur89b] Anthony G. Turner. The world fertility survey: An appraisal of methodology: Comment. Journal of the American Statistical Association, 84(407):768–769, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289660.

Tukey:1985:VSL

[TV85] John W. Tukey and P. Velleman. The variance of slopes of lines fitted by groups: An analysis of the Johnstone and Velleman Monte Carlo results. comment on “the resistant line and related regression methods,” by I. Johnstone and P. Velleman. Journal of the American Statistical Association, 80(392):1055–1059, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Tveite:1984:BRB

[Tve84] Michael Tveite. Book review: [untitled]. Journal of the American Statistical Association, 79(388):963–964, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288768.

Tan:1980:ANN

[TW80] W. Y. Tan and S. P. Wong. On approximating the null and nonnull distributions of the F ratio in unbalanced random-effect models from nonnormal universes. Journal of the American Statistical Association, 75(371):655–662, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287664.

Thisted:1981:PMT

[TW81] Ronald A. Thisted and William E. Wecker. Predicting a multitude of time series. Journal of the American Statistical Association, 76 (375):516–523, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287504.

Tanner:1983:RSH

[TW83] Martin A. Tanner and Wing Hung Wong. A reanalysis of the standford heart transplant data: Comment. Journal of the American Statistical Association, 78(382):286–287, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288626.

Tanner:1984:DBN

[TW84a] Martin Abba Tanner and Wing Hung Wong. Data-based nonparametric estimation of the hazard function with applications to model diagnostis and exploratory analysis. Journal of the American Statistical Association, 79(385):174–182, March 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288353.

Tukey:1984:DAS

[TW84b] Paul A. Tukey and Kenneth W. Wachter. A demographic analogy for shareowner accounts. Journal of the American Statistical Association, 79(387):525–530, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288399.

Tanner:1987:CPDb

[TW87a] Martin A. Tanner and Wing Hung Wong. The calculation of posterior distributions by data augmentation: Rejoinder. Journal of the American Statistical Association, 82(398):548–550, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289463.

Tanner:1987:CPDa

[TW87b] Martin A. Tanner and Wing Hung Wong. The calculation of posterior distributions by data augmentation. With discussion and with a reply by the authors. Journal of the American Statistical Association, 82(398):528–540, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289457.

Tsui:1989:GVS

[TW89] Kam-Wah Tsui and Samaradasa Weerahandi. Generalized pvalues in significance testing of hypotheses in the presence of nuisance parameters. Journal of the American Statistical Association, 84(406):602–607, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289949.

Tanner:1985:MAA

[TY85] Martin A. Tanner and Michael A. Young. Modeling agreement among raters. Journal of the American Statistical Association, 80 (389):175–180, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288068.

Tishler:1981:NML

[TZ81] Asher Tishler and Israel Zang. A new maximum likelihood algorithm for piecewise regression. Journal of the American Statistical Association, 76(376):980–987, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287599.

Utts:1980:RCT

[UH80] Jessica M. Utts and Thomas P. Hettmansperger. A robust class of tests and estimates for multivariate location. Journal of the American Statistical Association, 75(372):939–946, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287185.

Ullah:1988:BRB

[Ull88] Aman Ullah. Book review: [untitled]. Journal of the American Statistical Association, 83(402):575, June 1988. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288898.

Umbach:1980:CGB

[Umb80] Dale Umbach. On characterizing generalized Bayes estimators for the natural parameter. Journal of the American Statistical Association, 75(371):729–733, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287678.

Utts:1987:BRB

[Utt87] Jessica Utts. Book review: [untitled]. Journal of the American Statistical Association, 82(397):364–365, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289209.

Uusipaikka:1983:ECB

[Uus83] Esa Uusipaikka. Exact confidence bands for linear regression over intervals. Journal of the American Statistical Association, 78 (383):638–644, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288132.

Uusipaikka:1985:ESC

[Uus85] Esa Uusipaikka. Exact simultaneous confidence intervals for multiple comparisons among three or four mean values. Journal of the American Statistical Association, 80(389):196–201, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288071.

Valliant:1985:NPT

[Val85] Richard Valliant. Nonlinear prediction theory and the estimation of proportions in a finite population. Journal of the American Statistical Association, 80(391):631–641, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288479.

Valliant:1987:CPS

[Val87a] Richard Valliant. Conditional properties of some estimators in stratified sampling. Journal of the American Statistical Association, 82(398):509–519, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289454.

Valliant:1987:GVF

[Val87b] Richard Valliant. Generalized variance functions in stratified twostage sampling. Journal of the American Statistical Association, 82(398):499–508, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289453.

Campenhout:1980:ARB

[Van80a] Jan M. Van Campenhout. The arbitrary relation between probability of error and measurement subset. Journal of the American Statistical Association, 75(369):104–109, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287396.

VanNess:1980:BRUa

[Van80b] John W. Van Ness. Book review: [untitled]. Journal of the American Statistical Association, 75(370):472, June 1980. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287484.

VanNess:1980:BRUb

[Van80c] John W. Van Ness. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1043–1044, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287223.

VanNostrand:1980:CSR

[Van80d] R. Craig Van Nostrand. A critique of some ridge regression methods: Comment. Journal of the American Statistical Association, 75(369):92–94, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287389.

VanDerGaag:1981:BRU

[Van81] Jacques Van Der Gaag. Book review: [untitled]. Journal of the American Statistical Association, 76(375):748–749, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287553.

Ness:1983:MCT

[Van83] John W. Van Ness. A method for comparing two hierarchical clusterings: Comment. Journal of the American Statistical Association, 78(383):576–579, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288119.

VanBowen:1986:BRU

[Van86a] J. Van Bowen, Jr. Book review: [untitled]. Journal of the American Statistical Association, 81(393):258–259, March 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288011.

VanRyzin:1986:BRUa

[Van86b] John Van Ryzin. Book review: [untitled]. Journal of the American Statistical Association, 81(393):253, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288003.

VanRyzin:1986:BRUb

[Van86c] John Van Ryzin. Book review: [untitled]. Journal of the American Statistical Association, 81(394):575–576, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289265.

VanRyzin:1986:BRUc

[Van86d] John Van Ryzin. Book review: [untitled]. Journal of the American Statistical Association, 81(394):577, June 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289269.

Vardeman:1980:BRB

[Var80] Stephen Vardeman. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1037, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287210.

Vardeman:1982:BRBa

[Var82a] Stephen Vardeman. Book review: [untitled]. Journal of the American Statistical Association, 77(379):691, September 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287752.

Vardeman:1982:BRBb

[Var82b] Stephen Vardeman. Book review: [untitled]. Journal of the American Statistical Association, 77(380):940, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287340.

Varadhan:1985:BRB

[Var85] S. R. S. Varadhan. Book review: [untitled]. Journal of the American Statistical Association, 80(390):490–491, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287939.

Vardeman:1986:BRB

[Var86] Stephen B. Vardeman. Book review: [untitled]. Journal of the American Statistical Association, 81(393):259, March 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288012.

Vardeman:1987:TPN

[Var87] Stephen B. Vardeman. Testing a point null hypothesis: The irreconcilability of P values and evidence: Comment. Journal of the American Statistical Association, 82(397):130–131, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289136.

Vardi:1988:SMI

[Var88] Yehuda Vardi. Statistical models for intercepted data. Journal of the American Statistical Association, 83(401):183–197, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288940.

Vardi:1989:BRB

[Var89] Yehuda Vardi. Book review: [untitled]. Journal of the American Statistical Association, 84(407):837–838, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289678.

Vaughn:1982:BRB

[Vau82] Richard C. Vaughn. Book review: [untitled]. Journal of the American Statistical Association, 77(377):216–217, March 1982. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287800.

vanHouwelingen:1987:MEM

[vd87] J. C. van Houwelingen and L. de Vries. Minimax estimation of the mixing proportion of two known distributions. Journal of the American Statistical Association, 82(397):300–304, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289167.

Velleman:1980:DCR

[Vel80] Paul F. Velleman. Definition and comparison of robust nonlinear data smoothing algorithms. Journal of the American Statistical Association, 75(371):609–615, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287657.

Velleman:1981:BRBb

[Vel81a] Paul Velleman. Book review: [untitled]. Journal of the American Statistical Association, 76(375):752–753, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287560.

Velleman:1981:BRBa

[Vel81b] Paul F. Velleman. Book review: [untitled]. Journal of the American Statistical Association, 76(374):492, June 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287858.

Verdinelli:1989:BRB

[Ver89a] Isabella Verdinelli. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1103–1104, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290109.

Verducci:1989:BRB

[Ver89b] Joseph S. Verducci. Book review: [untitled]. Journal of the American Statistical Association, 84(407):847–848, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289697.

Vinod:1980:CSR

[Vin80] H. D. Vinod. A critique of some ridge regression methods: Comment. Journal of the American Statistical Association, 75(369): 97–98, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287393.

Vere-Jones:1983:BRB

[VJ83] D. Vere-Jones. Book review: [untitled]. Journal of the American Statistical Association, 78(383):745, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288173.

Verrill:1988:TLS

[VJ88] Steve Verrill and Richard A. Johnson. Tables and large-sample distribution theory for censored-data correlation statistics for testing normality. Journal of the American Statistical Association, 83 (404):1192–1197, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290156.

Vardeman:1983:CSD

[VM83] Stephen Vardeman and Glen Meeden. Calibration, sufficiency, and domination considerations for Bayesian probability assessors. Journal of the American Statistical Association, 78(384):808–816, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288189.

Voorn:1981:CVT

[Voo81] Wim J. Voorn. A class of variate transformations causing unbounded likelihood. Journal of the American Statistical Association, 76(375):709–712, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287535.

Voss:1986:FOD

[Vos86] Daniel T. Voss. First-order deletion designs and the construction of efficient nearly orthogonal factorial designs in small blocks. Journal of the American Statistical Association, 81(395):813–818, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289014.

Vredevoe:1980:BRB

[Vre80] Donna L. Vredevoe. Book review: [untitled]. Journal of the American Statistical Association, 75(369):246, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287431.

Vardi:1985:SMPb

[VSK85a] Y. Vardi, L. A. Shepp, and L. Kaufman. A statistical model for positron emission tomography: Rejoinder. Journal of the American Statistical Association, 80(389):34–37, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288037.

Vardi:1985:SMPa

[VSK85b] Y. Vardi, L. A. Shepp, and L. Kaufman. A statistical model for positron emission tomography. With discussion. Journal of the American Statistical Association, 80(389):8–20, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288030.

Villalobos:1987:ICM

[VW87] Miguel Villalobos and Grace Wahba. Inequality-constrained multivariate smoothing splines with application to the estimation of posterior probabilities. Journal of the American Statistical Association, 82(397):239–248, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2289160.

Velleman:1980:CRC

[VY80] Paul F. Velleman and M. Agelia Ypelaar. Constructing regressions with controlled features: a method of probing regression performance. Journal of the American Statistical Association, 75(372): 839–844, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287170.

Wecker:1983:SEA

[WA83] William E. Wecker and Craig F. Ansley. The signal extraction approach to nonlinear regression and spline smoothing. Journal of the American Statistical Association, 78(381):81–89, March 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287113.

Williams:1989:ADD

[WABA89] Christopher J. Williams, Wyatt W. Anderson, Celeste J. Brown, and Jonathan Arnold. An analysis of density-dependent viability selection. Journal of the American Statistical Association, 84 (407):662–668, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289646.

Wahba:1980:ASL

[Wah80] Grace Wahba. Automatic smoothing of the log periodogram. Journal of the American Statistical Association, 75(369):122– 132, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287399.

Wainer:1981:RPM

[Wai81] Howard Wainer. Representing points in many dimensions by trees and castles: Comment. Journal of the American Statistical Association, 76(374):272–275, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287823.

Wainer:1983:BRB

[Wai83] Howard Wainer. Book review: [untitled]. Journal of the American Statistical Association, 78(382):492–496, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288666.

Waksberg:1985:BRB

[Wak85] Joseph Waksberg. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1077, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288597.

Wallenstein:1980:DSO

[Wal80a] Sylvan Wallenstein. Distributions of some one-sided k-sample Smirnov-type statistics. Journal of the American Statistical Association, 75(370):441–446, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287474.

Wallgren:1980:DPT

[Wal80b] Cheryl M. Wallgren. The distribution of the product of two correlated t variates. Journal of the American Statistical Association, 75(372):996–1000, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287194.

Wallis:1980:SRGa

[Wal80c] W. Allen Wallis. The Statistical Research Group, 1942–1945. Journal of the American Statistical Association, 75(370):320–330, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287451.

Wallis:1980:SRGb

[Wal80d] W. Allen Wallis. The Statistical Research Group, 1942–1945: Rejoinder. Journal of the American Statistical Association, 75(370): 334–335, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287454.

Wallis:1981:MDR

[Wal81a] Kenneth F. Wallis. Modeling demographic relationships: An analysis of forecast functions for Australian births: Comment. Journal of the American Statistical Association, 76(376):798–799, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287573.

Wallsten:1981:BRB

[Wal81b] Thomas S. Wallsten. Book review: [untitled]. Journal of the American Statistical Association, 76(374):500, June 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287876.

Wallace:1983:MCT

[Wal83] David L. Wallace. A method for comparing two hierarchical clusterings: Comment. Journal of the American Statistical Association, 78(383):569–576, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288118.

Waldman:1984:BRB

[Wal84a] Donald M. Waldman. Book review: [untitled]. Journal of the American Statistical Association, 79(386):472–473, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288299.

Walter:1984:RSS

[Wal84b] S. D. Walter. Required sample size for categorical matching. Journal of the American Statistical Association, 79(387):662–667, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288414.

Wallenstein:1986:BRB

[Wal86a] Sylvan Wallenstein. Book review: [untitled]. Journal of the American Statistical Association, 81(394):576–577, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289267.

Walton:1986:NOC

[Wal86b] Gerald S. Walton. The number of observed classes from a multiple hypergeometric distribution. Journal of the American Statistical Association, 81(393):169–171, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287984.

Wang:1980:SEM

[Wan80a] Y. H. Wang. Sequential estimation of the mean of a multinormal population. Journal of the American Statistical Association, 75 (372):977–983, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287191.

Wansbeek:1980:RIC

[Wan80b] Tom Wansbeek. A regression interpretation of the computation of MINQUE variance component estimates. Journal of the American Statistical Association, 75(370):375–376, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287462.

Wang:1989:SMR

[Wan89] Mei-Cheng Wang. A semiparametric model for randomly truncated data. Journal of the American Statistical Association, 84 (407):742–748, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289656.

Warren:1980:BRB

[War80] Richard D. Warren. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1043, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287221.

Wasserman:1980:ASN

[Was80a] Stanley Wasserman. Analyzing social networks as stochastic processes. Journal of the American Statistical Association, 75(370): 280–294, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287447.

Wasserman:1980:BRB

[Was80b] Stanley Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 75(370):470–471, June 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287482.

Wasserman:1985:BRB

[Was85] Stanley Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 80(389):237–238, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288084.

Wasserman:1989:BRB

[Was89] Stanley Wasserman. Book review: [untitled]. Journal of the American Statistical Association, 84(405):347, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289905.

Watson:1980:BRB

[Wat80] G. S. Watson. Book review: [untitled]. Journal of the American Statistical Association, 75(369):245–246, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287429.

Watson:1984:BRB

[Wat84a] G. S. Watson. Book review: [untitled]. Journal of the American Statistical Association, 79(388):947, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288735.

Watts:1984:BRB

[Wat84b] Donald G. Watts. Book review: [untitled]. Journal of the American Statistical Association, 79(387):737, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288443.

Watson:1985:BRB

[Wat85a] Mark W. Watson. Book review: [untitled]. Journal of the American Statistical Association, 80(391):783–784, September 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288516.

Watts:1985:BRB

[Wat85b] Harold Watts. Book review: [untitled]. Journal of the American Statistical Association, 80(390):493–494, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287948.

Watson:1987:UMB

[Wat87] Mark W. Watson. Uncertainty in model-based seasonal adjustment procedures and construction of minimax filters. Journal of the American Statistical Association, 82(398):395–408, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289441.

Watson:1988:BRB

[Wat88] Mark W. Watson. Book review: [untitled]. Journal of the American Statistical Association, 83(404):1212, December 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290160.

Watson:1989:SAS

[Wat89] Mark W. Watson. Sensitivity analysis of seasonal adjustments: Empirical case studies: Comment. Journal of the American Statistical Association, 84(405):27–28, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289841.

Wallenstein:1981:APT

[WB81] Sylvan Wallenstein and Agnes Berger. On the asymptotic power of tests for comparing K correlated proportions. Journal of the American Statistical Association, 76(373):114–118, March 1981.

CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287055.

Wachter:1985:PPF

[WB85] Kenneth W. Wachter and Richard A. Becker. Are productive people to be found? robust analysis of sparse two-way tables. Journal of the American Statistical Association, 80(390):266–276, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287882.

Weber:1989:BRB

[Web89] Elke U. Weber. Book review: [untitled]. Journal of the American Statistical Association, 84(405):349–350, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289909.

Wecker:1981:ATS

[Wec81] William E. Wecker. Asymmetric time series. Journal of the American Statistical Association, 76(373):16–21, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287034.

Wecker:1982:CAT

[Wec82] William E. Wecker. Corrigenda: Asymmetric time series. Journal of the American Statistical Association, 77(380):954, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287360.

Wegge:1980:BRB

[Weg80] Leon Wegge. Book review: [untitled]. Journal of the American Statistical Association, 75(369):239–241, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287420.

Wegman:1989:BRB

[Weg89] Edward J. Wegman. Book review: [untitled]. Journal of the American Statistical Association, 84(405):342, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289898.

Wehrung:1980:BRB

[Weh80] Donald Wehrung. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1033, December 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287203.

Wei:1980:GGG

[Wei80a] L. J. Wei. A generalized gehan and Gilbert test for paired observations that are subject to arbitrary right censorship. Journal of the American Statistical Association, 75(371):634–637, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287660.

Weisberg:1980:SLS

[Wei80b] Sanford Weisberg. Some large-sample tests for nonnormality in the linear regression model: Comment. Journal of the American Statistical Association, 75(369):28–31, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287374.

Weisberg:1980:UEB

[Wei80c] Sanford Weisberg. Using empirical Bayes techniques in the law school validity studies: Comment. Journal of the American Statistical Association, 75(372):824–825, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287167.

Wei:1981:ACE

[Wei81a] L. J. Wei. Asymptotic conservativeness and efficiency of Kruskal-Wallis test for K dependent samples. Journal of the American Statistical Association, 76(376):1006–1009, December 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287603.

Weiner:1981:BRB

[Wei81b] Howard J. Weiner. Book review: [untitled]. Journal of the American Statistical Association, 76(374):499–500, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287875.

Wei:1982:MLD

[Wei82] William W. S. Wei. Measurement of linear dependence and feedback between multiple time series: Comment. Journal of the American Statistical Association, 77(378):316–319, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287241.

Wei:1983:TII

[Wei83a] L. J. Wei. Tests for interchangeability with incomplete paired observations. Journal of the American Statistical Association, 78 (383):725–729, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288143.

Weisberg:1983:BRB

[Wei83b] Herbert I. Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 78(383):739, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288161.

Wei:1984:TGF

[Wei84a] L. J. Wei. Testing goodness of fit for proportional hazards model with censored observations. Journal of the American Statistical Association, 79(387):649–652, September 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288412.

Weisberg:1984:BRB

[Wei84b] Sanford Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 79(386):471–472, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288297.

Weisberg:1985:BRB

[Wei85] Sanford Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1081, December 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288606.

Weisberg:1986:LMA

[Wei86] Sanford Weisberg. A linear model approach to backcalculation of fish length. Journal of the American Statistical Association, 81 (396):922–929, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289061.

Weisberg:1989:BRB

[Wei89] Herbert I. Weisberg. Book review: [untitled]. Journal of the American Statistical Association, 84(406):620–621, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289963.

Welch:1984:BRB

[Wel84] Finis R. Welch. Book review: [untitled]. Journal of the American Statistical Association, 79(387):740–741, September 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288449.

Well:1986:BRB

[Wel86] Arnold D. Well. Book review: [untitled]. Journal of the American Statistical Association, 81(393):260, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288014.

Wellner:1989:BRB

[Wel89] Jon A. Wellner. Book review: [untitled]. Journal of the American Statistical Association, 84(406):627, June 1989. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289976.

Wermuth:1980:LRE

[Wer80] Nanny Wermuth. Linear recursive equations, covariance selection, and path analysis. Journal of the American Statistical Association, 75(372):963–972, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287189.

Westfall:1987:CVC

[Wes87a] Peter H. Westfall. A comparison of variance component estimates for arbitrary underlying distributions. Journal of the American Statistical Association, 82(399):866–874, September 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288798.

Westfall:1987:CMT

[Wes87b] Peter H. Westfall. Computable MINQUE-type estimates of variance components. Journal of the American Statistical Association, 82(398):586–589, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2289469.

Wolfenson:1982:BLD

[WF82] Marco Wolfenson and Terrence L. Fine. Bayes-like decision making with upper and lower probabilities. Journal of the American Statistical Association, 77(377):80–88, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287772.

Woodward:1981:RBA

[WG81] Wayne A. Woodward and H. L. Gray. On the relationship between the S array and the Box–Jenkins method of ARMA model identification. Journal of the American Statistical Association, 76 (375):579–587, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287515.

Wei:1983:NES

[WG83] L. J. Wei and M. H. Gail. Nonparametric estimation for a scalechange with censored observations. Journal of the American Statistical Association, 78(382):382–388, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288644.

Welch:1988:RPT

[WG88] William J. Welch and Leopoldo G. Gutierrez. Robust permutation tests for matched-pairs designs. Journal of the American Statistical Association, 83(402):450–455, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288861.

West:1986:MAB

[WH86] Mike West and P. Jeff Harrison. Monitoring and adaptation in Bayesian forecasting models. Journal of the American Statistical Association, 81(395):741–750, September 1986. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2289006.

Wu:1988:ETS

[WHH88] C. F. J. Wu, D. Holt, and D. J. Holmes. The effect of two-stage sampling on the F statistic. Journal of the American Statistical Association, 83(401):150–159, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288934.

White:1981:CDM

[Whi81a] Halbert White. Consequences and detection of misspecified nonlinear regression models. Journal of the American Statistical Association, 76(374):419–433, June 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2287845.

Whittemore:1981:CSS

[Whi81b] Alice S. Whittemore. Correction: “Sample size for logistic regression”. Journal of the American Statistical Association, 76(376): 1022–1023, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287632. See [Whi81c].

Whittemore:1981:SSL

[Whi81c] Alice S. Whittemore. Sample size for logistic regression with small response probability. Journal of the American Statistical Association, 76(373):27–32, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor. org/stable/2287036. See correction [Whi81b].

White:1983:BRB

[Whi83a] Marvin White. Book review: [untitled]. Journal of the American Statistical Association, 78(383):733–734, September 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288150.

Whitt:1983:BRB

[Whi83b] Ward Whitt. Book review: [untitled]. Journal of the American Statistical Association, 78(382):501–502, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288679.

Whitt:1986:BRB

[Whi86a] Ward Whitt. Book review: [untitled]. Journal of the American Statistical Association, 81(393):266, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288025.

Whittemore:1986:ACW

[Whi86b] Alice S. Whittemore. An analysis of contaminated well water and health effects in Woburn, Massachusetts: Comment. Journal of the American Statistical Association, 81(395):609–610, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288987.

White:1989:SAR

[Whi89] Halbert White. Some asymptotic results for learning in single hidden layer feedforward network models. Journal of the American Statistical Association, 84(408):1003–1013, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290076. See comment

[Fab94]. See correction [Whi92].

White:1992:SAR

[Whi92] Halbert White. “Some asymptotic results for learning in single hidden-layer feedforward network models” [J. Amer. Statist. Assoc. 84 (1989), no. 408, 1003–1013, MR 92e:62119]. Journal of the American Statistical Association, 87(420):1252, ???? 1992. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). See [Whi89, Fab94].

West:1985:DGLb

[WHM85a] Mike West, P. Jeff Harrison, and Helio S. Migon. Dynamic generalized linear models and Bayesian forecasting: Rejoinder. Journal of the American Statistical Association, 80(389):96–97, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288050.

West:1985:DGLa

[WHM85b] Mike West, P. Jeff Harrison, and Helio S. Migon. Dynamic generalized linear models and Bayesian forecasting. With discussion. Journal of the American Statistical Association, 80(389): 73–83, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288042.

Wichern:1980:BRB

[Wic80] Dean W. Wichern. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1036–1037, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287208.

Wichern:1981:BRB

[Wic81] Dean W. Wichern. Book review: [untitled]. Journal of the American Statistical Association, 76(374):496, June 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287869.

Widhelm:1983:BRBa

[Wid83a] William B. Widhelm. Book review: [untitled]. Journal of the American Statistical Association, 78(383):743, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288168.

Widhelm:1983:BRBb

[Wid83b] William B. Widhelm. Book review: [untitled]. Journal of the American Statistical Association, 78(384):1005, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288246.

Wijsman:1983:BRB

[Wij83] Robert A. Wijsman. Book review: [untitled]. Journal of the American Statistical Association, 78(382):501, June 1983. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288678.

Wijsman:1985:UIR

[Wij85] Robert A. Wijsman. A useful inequality on ratios of integrals, with application to maximum likelihood estimation. Journal of the American Statistical Association, 80(390):472–475, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2287917.

Wijsman:1986:BRB

[Wij86] Robert A. Wijsman. Book review: [untitled]. Journal of the American Statistical Association, 81(395):860–861, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289029.

Willemain:1980:EPM

[Wil80] Thomas R. Willemain. Estimating the population median by nomination sampling. Journal of the American Statistical Association, 75(372):908–911, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287180.

Wilcox:1984:SBP

[Wil84] Rand R. Wilcox. Selecting the best population, provided it is better than a standard: the unequal variance case. Journal of the American Statistical Association, 79(388):887–891, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288720.

Williams:1985:BRB

[Wil85a] R. J. Williams. Book review: [untitled]. Journal of the American Statistical Association, 80(391):787–788, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288523.

Wilson:1985:BRB

[Wil85b] Franklin D. Wilson. Book review: [untitled]. Journal of the American Statistical Association, 80(392):1075–1076, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288593.

Wilson:1986:BRB

[Wil86] Berry Wilson. Book review: [untitled]. Journal of the American Statistical Association, 81(395):866, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289036.

Williams:1988:BRB

[Wil88a] George W. Williams. Book review: [untitled]. Journal of the American Statistical Association, 83(402):576–577, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288901.

Willmot:1988:POF

[Wil88b] Gordon E. Willmot. Parameter orthogonality for a family of discrete distributions. Journal of the American Statistical Association, 83(402):517–521, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288871.

Winkler:1980:BRB

[Win80] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 75(369):251–252, March 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287436.

Winkler:1983:BRB

[Win83] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 78(382):497, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288669.

Winkler:1984:BRBa

[Win84a] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 79(388):963, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288766.

Winkler:1984:BRBb

[Win84b] Robert L. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 79(388):963, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288767.

Winkler:1985:DGL

[Win85] Robert L. Winkler. Dynamic generalized linear models and Bayesian forecasting: Comment. Journal of the American Statistical Association, 80(389):95, March 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288049.

Winkler:1989:BRB

[Win89] William E. Winkler. Book review: [untitled]. Journal of the American Statistical Association, 84(407):850, September 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289701.

Wiorkowski:1981:EVRa

[Wio81a] John J. Wiorkowski. Estimating volumes of remaining fossil fuel resources: a critical review. Journal of the American Statistical Association, 76(375):534–548, September 1981. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2287507.

Wiorkowski:1981:EVRb

[Wio81b] John J. Wiorkowski. Estimating volumes of remaining fossil fuel resources: a critical review: Rejoinder. Journal of the American Statistical Association, 76(375):558–559, September 1981. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287512.

Witmer:1984:BRB

[Wit84] Jeffrey A. Witmer. Book review: [untitled]. Journal of the American Statistical Association, 79(388):952, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288742.

Wittkowski:1988:FTS

[Wit88] Knut M. Wittkowski. Friedman-type statistics and consistent multiple comparisons for unbalanced designs with missing data. Journal of the American Statistical Association, 83(404):1163–1170, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2290150.

Whittemore:1988:ARC

[WK88] Alice S. Whittemore and Joseph B. Keller. Approximations for regression with covariate measurement error. Journal of the American Statistical Association, 83(404):1057–1066, December 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2290136.

Wei:1984:TSA

[WL84] L. J. Wei and J. M. Lachin. Two-sample asymptotically distribution-free tests for incomplete multivariate observations. Journal of the American Statistical Association, 79(387):653–661, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288413.

Waternaux:1989:MAL

[WLW89a] Christine Waternaux, Nan M. Laird, and James H. Ware. Methods for analysis of longitudinal data: Blood-lead concentrations and cognitive development. Journal of the American Statistical Association, 84(405):33–41, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289844.

Wei:1989:RAM

[WLW89b] L. J. Wei, D. Y. Lin, and L. Weissfeld. Regression analysis of multivariate incomplete failure time data by modeling marginal distributions. Journal of the American Statistical Association, 84 (408):1065–1073, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2290084.

White:1980:SLS

[WM80] Halbert White and Glenn M. MacDonald. Some large-sample tests for nonnormality in the linear regression model. Journal of the American Statistical Association, 75(369):16–28, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2287373. See

correction [WM81].

White:1981:CSL

[WM81] Halbert White and Glenn M. MacDonald. Correction: “Some large-sample tests for nonnormality in the linear regression model” [J. Amer. Statist. Assoc. 75 (1980), no. 369, 16–31, MR 81e:62077]. Journal of the American Statistical Association, 76 (376):1022, December 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287631. See [WM80].

Weber:1982:PDW

[WM82] Jean E. Weber and David E. Monarchi. Performance of the Durbin–Watson test and WLS estimation when the disturbance term includes serial dependence in addition to first-order autocorrelation. Journal of the American Statistical Association, 77 (377):117–128, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287777.

Wong:1985:HLR

[WM85] George Y. Wong and William M. Mason. The hierarchical logistic regression model for multilevel analysis. Journal of the American Statistical Association, 80(391):513–524, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288464.

Whitmore:1985:UGC

[WMH85] Roy W. Whitmore, Robert E. Mason, and Tyler D. Hartwell. Use of geographically classified telephone directory lists in multi-mode surveys. Journal of the American Statistical Association, 80(392): 842–844, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288542.

Wolins:1980:BRB

[Wol80] Leroy Wolins. Book review: [untitled]. Journal of the American Statistical Association, 75(371):759–760, September 1980. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2287694.

Wolins:1982:BRB

[Wol82] Leroy Wolins. Book review: [untitled]. Journal of the American Statistical Association, 77(377):220, March 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287805.

Wolins:1983:BRB

[Wol83] Leroy Wolins. Book review: [untitled]. Journal of the American Statistical Association, 78(382):497, June 1983. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288670.

Wolter:1984:ISE

[Wol84] Kirk M. Wolter. An investigation of some estimators of variance for systematic sampling. Journal of the American Statistical Association, 79(388):781–790, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288708.

Wolins:1985:BRB

[Wol85] Leroy Wolins. Book review: [untitled]. Journal of the American Statistical Association, 80(390):484–485, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287929.

Wolter:1986:SCE

[Wol86] Kirk M. Wolter. Some coverage error models for census data. Journal of the American Statistical Association, 81(394):338–346, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289222.

Wolak:1987:ETM

[Wol87a] Frank A. Wolak. An exact test for multiple inequality and equality constraints in the linear regression model. Journal of the American Statistical Association, 82(399):782–793, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288787.

Wollan:1987:BRB

[Wol87b] Peter C. Wollan. Book review: [untitled]. Journal of the American Statistical Association, 82(397):345, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289178.

Wolfe:1989:BRB

[Wol89] Robert A. Wolfe. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1098, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290099.

Wong:1982:RRA

[Won82a] George Y. Wong. Round robin analysis of variance via maximum likelihood. Journal of the American Statistical Association, 77 (380):714–724, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287296.

Wong:1982:HCM

[Won82b] M. Anthony Wong. A hybrid clustering method for identifying high-density clusters. Journal of the American Statistical Association, 77(380):841–847, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287316.

Wong:1983:MCT

[Won83a] M. Anthony Wong. A method for comparing two hierarchical clusterings: Comment. Journal of the American Statistical Association, 78(383):582–583, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288121.

Wong:1983:NML

[Won83b] Wing Hung Wong. A note on the modified likelihood for density estimation. Journal of the American Statistical Association, 78 (382):461–463, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288658.

Wong:1984:ROR

[Won84] George Y. Wong. Rank-one round Robin designs. Journal of the American Statistical Association, 79(385):136–144, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2288348.

Wong:1987:BMD

[Won87] George Y. Wong. Bayesian models for directed graphs. Journal of the American Statistical Association, 82(397):140–148, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289140.

Wonnacott:1989:BRB

[Won89] T. H. Wonnacott. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1093, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290091.

Woodroofe:1987:BRB

[Woo87] Michael Woodroofe. Book review: [untitled]. Journal of the American Statistical Association, 82(398):693–694, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289501.

Wei:1985:DFM

[WP85] L. J. Wei and D. Pee. Distribution-free methods of estimating location difference with censored paired data. Journal of the American Statistical Association, 80(390):405–410, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287905.

Woodward:1984:CMD

[WPSL84] Wayne A. Woodward, William C. Parr, William R. Schucany, and Hildegard Lindsey. A comparison of minimum distance and maximum likelihood estimation of a mixture proportion. Journal of the American Statistical Association, 79(387):590–598, September 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288405.

Warrack:1984:LRT

[WR84] Giles Warrack and Tim Robertson. A likelihood ratio test regarding two nested but oblique order-restricted hypotheses. Journal of the American Statistical Association, 79(388):881–886, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288719.

Wragge:1983:BRB

[Wra83] John H. Wragge. Book review: [untitled]. Journal of the American Statistical Association, 78(383):742, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288165.

Wretman:1981:ESR

[Wre81] Jan Wretman. An empirical study of the ratio estimator and estimators of its variance: Comment. Journal of the American Statistical Association, 76(373):86, March 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287048.

Wright:1983:FPS

[Wri83] Roger L. Wright. Finite population sampling with multivariate auxiliary information. Journal of the American Statistical Association, 78(384):879–884, December 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288199.

Whitaker:1989:ERS

[WS89] Lyn R. Whitaker and Francisco J. Samaniego. Estimating the reliability of systems subject to imperfect repair. Journal of the American Statistical Association, 84(405):301–309, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289878.

Wachter:1982:EHHa

[WT82a] Kenneth W. Wachter and James Trussell. Estimating historical heights. Journal of the American Statistical Association, 77(378): 279–293, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287231.

Wachter:1982:EHHb

[WT82b] Kenneth W. Wachter and James Trussell. Estimating historical heights: Rejoinder. Journal of the American Statistical Association, 77(378):301–303, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2287237.

Woltman:1980:CTM

[WTB80] Henry F. Woltman, Anthony G. Turner, and John M. Bushery. A comparison of three mixed-mode interviewing procedures in the nationa crime survey. Journal of the American Statistical Association, 75(371):534–543, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287642.

Wu:1985:ESD

[Wu85] C. F. Jeff Wu. Efficient sequential designs with binary data. Journal of the American Statistical Association, 80(392):974–984, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288563.

Wegman:1983:SS

[WW83] Edward J. Wegman and Ian W. Wright. Splines in statistics. Journal of the American Statistical Association, 78(382):351–365, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288640.

Wang:1987:SBD

[WW87a] Yuchung J. Wang and George Y. Wong. Stochastic blockmodels for directed graphs. Journal of the American Statistical Association, 82(397):8–19, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289119.

Wittes:1987:PMH

[WW87b] Janet Wittes and Sylvan Wallenstein. The power of the Mantel– Haenszel test. Journal of the American Statistical Association, 82 (400):1104–1109, ???? 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Westfall:1989:PVA

[WY89] Peter H. Westfall and S. Stanley Young. p value adjustments for multiple tests in multivariate binomial models. Journal of the American Statistical Association, 84(407):780–786, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289666.

Yakowitz:1985:NDE

[Yak85] Sidney J. Yakowitz. Nonparametric density estimation, prediction, and regression for Markov sequences. Journal of the American Statistical Association, 80(389):215–221, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288075.

Yang:1981:LFC

[Yan81] Shie-Shien Yang. Linear functions of concomitants of order statistics with application to nonparametric estimation of a regression function. Journal of the American Statistical Association, 76(375): 658–662, September 1981. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287526.

Yang:1983:BRB

[Yan83] Mark C. K. Yang. Book review: [untitled]. Journal of the American Statistical Association, 78(382):510–511, June 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288697.

Yang:1984:SAM

[Yan84] Shie-Shien Yang. Sample analogs to multidimensional coverages with applications to prediction. Journal of the American Statistical Association, 79(388):804–806, December 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288710.

Yang:1985:SNE

[Yan85] Shie-Shien Yang. A smooth nonparametric estimator of a quantile function. Journal of the American Statistical Association, 80(392): 1004–1011, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288567.

Yandell:1989:BRB

[Yan89] Brian S. Yandell. Book review: [untitled]. Journal of the American Statistical Association, 84(405):338, March 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289893.

Yao:1987:BRB

[Yao87] Yi-Ching Yao. Book review: [untitled]. Journal of the American Statistical Association, 82(397):345–346, March 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289179.

Yashchin:1987:BRB

[Yas87] Emmanuel Yashchin. Book review: [untitled]. Journal of the American Statistical Association, 82(398):691, June 1987. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289499.

Yeh:1985:NTF

[YBN85] Ching Ming Yeh, Ralph A. Bradley, and William I. Notz. Nearly trend-free block designs. Journal of the American Statistical Association, 80(392):985–992, December 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288564.

Yeo:1984:STA

[YD84] W. B. Yeo and H. A. David. Selection through an associated characteristic, with applications to the random effects model. Journal of the American Statistical Association, 79(386):399–405, June 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288282.

Yang:1985:DFC

[YM85] Mark C. K. Yang and Tse Chin Mo. Distribution-free confidence bounds for P (X_{1} + X_{2} + ⋯ + X_{k} < z). Journal of the American Statistical Association, 80(389):227–230, March 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288077. See correction

[YM86].

Yang:1986:CDF

[YM86] Mark C. K. Yang and T. C. Mo. Correction: “Distribution-free confidence bounds for P (X_{1} + X_{2} + ⋯ + X_{k} < z)” [J. Amer.

Statist. Assoc. 80 (1985), no. 389, 227–230, MR 87b:62061]. Journal of the American Statistical Association, 81(396):1132, December 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2289113. See

[YM85].

Youkeles:1983:BRB

[You83] Lee Youkeles. Book review: [untitled]. Journal of the American Statistical Association, 78(383):736, September 1983. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2288156.

Yohai:1988:HBP

[YZ88] Víctor J. Yohai and Ruben H. Zamar. High breakdown-point estimates of regression by means of the minimization of an efficient scale. Journal of the American Statistical Association, 83(402): 406–413, June 1988. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2288856.

Zacks:1986:ESP

[Zac86] S. Zacks. Estimating the scale parameter of an exponential distribution from a sample of time-censored rth-order statistics. Journal of the American Statistical Association, 81(393):205– 209, March 1986. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287991.

Zacks:1989:BRBa

[Zac89a] Shelemyahu Zacks. Book review: [untitled]. Journal of the American Statistical Association, 84(407):841–842, September 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2289685.

Zacks:1989:BRBb

[Zac89b] Shelemyahu Zacks. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1104, December 1989. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2290110.

Zahl:1984:BRBa

[Zah84a] Samuel Zahl. Book review: Basic Statistics for Business and Economics, by Paul G. Hoel; Raymond J. Jessen. Journal of the American Statistical Association, 79(385):243, March 1984. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288389.

Zahl:1984:BRBb

[Zah84b] Samuel Zahl. Book review: Ecological Diversity in Theory and Practice, by G. P. Patil; Frederick T. Grassle; Woollcott Smith; Charles Taillie. Journal of the American Statistical Association, 79(385):243, March 1984. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2288390.

Zahn:1989:BRB

[Zah89] Douglas A. Zahn. Book review: [untitled]. Journal of the American Statistical Association, 84(406):615–616, June 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289956.

Zalmai:1982:BRB

[Zal82] G. J. Zalmai. Book review: [untitled]. Journal of the American Statistical Association, 77(380):942–943, December 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287344.

Zeger:1986:RAC

[ZB86] Scott L. Zeger and Ron Brookmeyer. Regression analysis with censored autocorrelated data. Journal of the American Statistical Association, 81(395):722–729, September 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289003.

Zelterman:1988:HTA

[ZC88] Daniel Zelterman and Chan Fu Chen. Homogeneity tests against central-mixture alternatives. Journal of the American Statistical Association, 83(401):179–182, March 1988. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2288939.

Zeger:1985:EOS

[Zeg85] Scott L. Zeger. Exploring an ozone spatial time series in the frequency domain. Journal of the American Statistical Association, 80(390):323–331, June 1985. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2287889.

Zehnwirth:1988:GKF

[Zeh88] Ben Zehnwirth. A generalization of the Kalman filter for models with state-dependent observation variance. Journal of the American Statistical Association, 83(401):164–167, March 1988. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288936.

Zeleny:1980:BRB

[Zel80a] Milan Zeleny. Book review: [untitled]. Journal of the American Statistical Association, 75(372):1035, December 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287205.

Zellner:1980:ISM

[Zel80b] Arnold Zellner. Information sources for modeling the national economy: Comment. Journal of the American Statistical Association, 75(371):567–569, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287645.

Zellner:1982:MLD

[Zel82] Arnold Zellner. Measurement of linear dependence and feedback between multiple time series: Comment. Journal of the American Statistical Association, 77(378):313–314, June 1982. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287239.

Zelterman:1984:BRB

[Zel84] Daniel Zelterman. Book review: [untitled]. Journal of the American Statistical Association, 79(386):469, June 1984. CODEN JST-NAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//www.jstor.org/stable/2288294.

Zelterman:1985:BRB

[Zel85] Daniel Zelterman. Book review: [untitled]. Journal of the American Statistical Association, 80(391):790, September 1985. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2288528.

Zellner:1986:BEP

[Zel86a] Arnold Zellner. Bayesian estimation and prediction using asymmetric loss functions. Journal of the American Statistical Association, 81(394):446–451, June 1986. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289234.

Zelterman:1986:BRB

[Zel86b] Daniel Zelterman. Book review: [untitled]. Journal of the American Statistical Association, 81(395):875, September 1986. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2289051.

Zelen:1987:BRB

[Zel87a] Marvin Zelen. Book review: [untitled]. Journal of the American Statistical Association, 82(398):685–686, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2289493.

Zelterman:1987:GFT

[Zel87b] Daniel Zelterman. Goodness-of-fit tests for large sparse multinomial distributions. Journal of the American Statistical Association, 82(398):624–629, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2289474.

Zeger:1987:FDM

[ZH87] Scott L. Zeger and Clifford M. Hurvich. A frequency-domain median time series. Journal of the American Statistical Association, 82(399):832–835, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2288793.

Ziebart:1987:EAA

[Zie87] David A. Ziebart. The effects of annual accounting data on stock returns and trading activity: a causal model study. Journal of the American Statistical Association, 82(399):733–738, September 1987. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2288777.

Zilberman:1980:BRB

[Zil80] David Zilberman. Book review: [untitled]. Journal of the American Statistical Association, 75(369):245, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287428.

Zinger:1980:VEP

[Zin80] A. Zinger. Variance estimation in partially systematic sampling. Journal of the American Statistical Association, 75(369):206– 211, March 1980. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2287412.

Zinn:1989:BRB

[Zin89] Joel Zinn. Book review: [untitled]. Journal of the American Statistical Association, 84(408):1105–1106, December 1989. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2290112.

Zerbe:1980:AGC

[ZJ80] Gary O. Zerbe and Richard H. Jones. On application of growth curve techniques to time series data. Journal of the American Statistical Association, 75(371):507–509, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287638.